### v2

# Itô’s formula for flows of conditional measures on semimartingales

Abstract.
Motivated by recent development of mean-field systems with common noise, this paper establishes Itô’s formula for flows of conditional probability measures under a common filtration associated with general semimartingales. This generalizes existing works on flows of conditional measures on Itô processes and flows of deterministic measure on general semimartingales. The key technical components involve constructing conditional independent copies and establishing the equivalence between stochastic integrals with respect to the conditional law of semimartingales and the conditional expectation of stochastic integrals with respect to copies of semimartingales. Itô’s formula is then established for cylindrical functions through conditional independent copies, and extended to the general case through function approximations.
1. Introduction
Itô’s formula
The classical Itô’s formula for semimartingales is generally viewed as the stochastic counterpart of the chain rule in calculus. It stands as one of the key cornerstones of stochastic analysis [29], establishing an intrinsic connection between partial differential equations (PDEs) and diffusion processes. This formula plays a critical role in the theory of stochastic control.
With the recent development of theory of mean-field games and mean-field controls [6, 8, 11], Itô’s formula has been extended to the flows of probability measures. Initially applied to diffusions [6], this extension was further developed for (discontinuous) semimartingales [24]. The generalized Itô’s formula has been instrumental for deriving the master equations in mean-field games and the Bellman dynamic programming equation for McKean–Vlasov control problems, as discussed in [11] and the references therein.
One of the most recent advances in mean-field theory have been concerning mean-field controls and games with common noise (see for instance [1, 2, 3, 5, 9, 10, 13, 15, 18, 23, 27, 30, 31, 32, 33, 35, 40]) or with stochastic processes involving jumps ([7, 22, 25, 26, 28, 36]). In either case, Itô’s formula has been developed respectively in [11] and in [24]. In order to analyze rigorously the McKean–Vlasov dynamics with both semimartingales and idiosyncratic and common noises, it is necessary to establish the Itô’s formula for flows of conditional laws with semimartingales, which does not seem to exist to the best of our knowledge. This is the primary focus of this paper.
Our work
This paper establishes Itô’s formula (Theorem 2.7) for flow of conditional laws $\mu_{t}=$Law$(X_{t}|{\mathcal{G}}_{t})$ driven by (possible discontinuous) semimartingales $\{X_{t}\}_{t\in[0,T]}$ with a general form of common noise represented by a sub filtration ${\mathbb{G}}=({\mathcal{G}}_{t})_{t\in[0,T]}$ and ${\mathcal{G}}={\mathcal{G}}_{T}$.
The primary task to analyze the Itô integral $\int_{0}^{t}\eta_{s}{\mathrm{d}}X_{s}^{\mathcal{G}}$ with $\{X_{t}^{\mathcal{G}}:={\mathbb{E}}[X_{t}|{\mathcal{G}}]\}_{t\in[0,T]}$ over some appropriate adapted process $\{\eta_{t}\}_{t\in[0,T]}$. This is crucial as the integral of the conditional expectation such as $\int_{0}^{t}\eta_{s}{\mathrm{d}}X^{\mathcal{G}}_{s}$ is not necessarily equivalent to the conditional expectation of the integral of the form ${\mathbb{E}}\big{[}\int_{0}^{t}\eta_{s}{\mathrm{d}}X_{s}|{\mathcal{G}}\big{]}$ (See discussions and examples in Section 3). Our analysis is built on several key technical components: constructing conditional independent copies of general stochastic processes (Theorem 3.1), establishing the equivalence between the conditional expectation of stochastic integrals with respect to seminmartingales and stochastic integrals with respect to the conditional law of semimartingales, and building the equivalence between the quadratic variation of the conditional expectation of a semimartingale and the conditional expectation of the quadratic variation of its two conditional independent copies (Theorem 3.3).
With these key technical components, Itô’s formula for flows of conditional laws on semimartingales is established (Theorem 2.7): it is first derived for cylindrical functions (i.e., smooth mean-field functions with integrable forms 4.1)), which are then shown to be dense in the desired function space; finally, localization argument along with proper form of dominating convergence theorem finishes the task.
Related works
Conditional independent copied has been treated informally and in certain forms in the existing literature, for instance, when defining a semimartingale by stochastic differential equations involving both idiosyncratic and common noise in [11]. In this paper, we have generalized the construction of conditional independent copies for a broad class of semimartingales and with an arbitrary sub $\sigma$-algebra. This concept may be of independent theoretical interest, besides its crucial role in our derivation of Itô’s formula for flow of conditional laws on semimartingales.
There are several methods for deriving Itô’s formula and its variants, including the time discretization approach in [6] and [36] for mean-field jump diffusions, the density approach in [8] using the Fokker–Planck equation, and the particle approximation approach [11, 14], [19, 38] to approximate flows of measures by flows of empirical measures. The cylindrical function approach has proved appropriate and powerful for analysis of general semimartingales. They are initially explored in Fleming–Viot processes [21], later adopted in the analysis of polynomial diffusions [17], and most recently used in Itô’s formula for flow of measures with (discontinuous) semimartingales [24].
Meanwhile, linear derivatives on the space of probability measures are known to be appropriate for characterizing the infinitesimal changes in the functional of controlled McKean–Vlasov processes when jumps are added; this has been noted previously in [7] and further explored in [24] and [39].
Notation
Throughout the paper, we will adopt the following notations, unless otherwise specified.
  • ${\mathbb{N}}^{+}$ denotes the the set of all possible natural numbers, and we fix $d\in{\mathbb{N}}^{+}$.
  • ${\mathcal{P}}({\mathbb{R}}^{d})$ denotes the space of probability measures on ${\mathbb{R}}^{d}$ with the topology of weak convergence. ${\mathcal{P}}_{p}({\mathbb{R}}^{d})$ for $p\in[1,\infty)$ denotes the probability measures with finite $p$-th moment, equipped with the Wasserstein-$p$ metric.
  • $C^{k}({\mathbb{R}}^{d})$ is the set of all $k$-th differentiable functions on ${\mathbb{R}}^{d}$ with continuous derivatives up to the $k$-th order, and $C_{b}^{k}({\mathbb{R}}^{d})\subset C^{k}({\mathbb{R}}^{d})$ is the set of all $k$-th differentiable functions on ${\mathbb{R}}^{d}$ with bounded and continuous derivatives up to the $k$-th order. And we use the convention $C({\mathbb{R}}^{d})=C^{0}({\mathbb{R}}^{d})$ and $C_{b}({\mathbb{R}}^{d})=C^{0}_{b}({\mathbb{R}}^{d})$.
  • For a random variable $X$, $\|X\|_{L^{p}}$ represents its $p$-th moment, i.e., $\|X\|_{L^{p}}={\mathbb{E}}[|X|^{p}]^{\frac{1}{p}}$.
  • For vectors $\bm{a}=(a_{i})_{i=1,..,d},\bm{b}=(b_{i})_{i=1,..,d}\in{\mathbb{R}}^{d}$, we denote $\bm{a}\cdot\bm{b}=\sum_{i=1}^{d}a_{i}b_{i}$. For matrices $\bm{C}=(C_{ij})_{i,j=1,..,d},\bm{D}=(D_{ij})_{i,j=1,..,d}\in{\mathbb{R}}^{d\times d}$, we denote $\bm{C}:\bm{D}=\sum_{i,j=1}^{d}C_{ij}D_{ij}$.
  • For $\mu\in{\mathcal{P}}({\mathbb{R}}^{d})$ and $\varphi:{\mathbb{R}}^{d}\to{\mathbb{R}}$ such that $\int_{{\mathbb{R}}^{d}}|\varphi(x)|{\mathrm{d}}\mu(x)<\infty$, we set
    $\langle\mu,\varphi\rangle:=\int_{{\mathbb{R}}^{d}}\varphi(x){\mathrm{d}}\mu(x).$
2. Assumptions and main result
Given two ${\mathbb{R}}^{d}$-valued ${\mathbb{F}}$- semimartingales $\bm{X}=(\bm{X}_{t})_{t\in[0,T]}$ and $\bm{Y}=(\bm{Y}_{t})_{t\in[0,T]}$ for $T>0$ and $d\in\mathbb{N}^{+}$ on a completed filtered probability space $(\Omega,{\mathcal{F}},{\mathbb{F}}=({\mathcal{F}}_{t})_{t\in[0,T]},{\mathbb{P}})$, and $\mu=\big{(}\mu_{t}:=\text{Law}(\bm{X}_{t}|{\mathcal{G}}_{t})\big{)}_{t\in[0,T]}$ with a sub filtration ${\mathbb{G}}=({\mathcal{G}}_{t})_{t\in[0,T]}\subset{\mathbb{F}}$. Our focus is to develop an Itô’s formula for the functional $\Phi(\mu_{t},\bm{Y}_{t})$ with the function $\Phi:{\mathcal{P}}_{p}({\mathbb{R}}^{d})\times{\mathbb{R}}^{d}\to{\mathbb{R}}$, under suitable regularity conditions.
2.1. Assumptions
Throughout the paper, we will make the following assumptions on the semimartingales $\bm{X}$ and $\bm{Y}$, the filtrations ${\mathbb{G}}$ and ${\mathbb{F}}$, and the function $\Phi$. These are standard assumptions to ensure that stochastic integrals are properly defined (see [37]) with the presence of common noise (see [34]) and for flow of measures on semimartingales (see [24, 16]).
Assumptions on semimartingales.
Assumption 2.1 (Semimartingale).
  • We assume $\bm{X},\bm{Y}\in\mathcal{H}^{p}$ for $1\leq p\leq\infty$ [37, Section V.2], with
    $\big{\|}\bm{X}\big{\|}_{{\mathcal{H}}^{p}}:=\inf_{\bm{M},\bm{V}}\bigg{\|}|\bm{X}_{0}|+\sqrt{[\bm{M},\bm{M}]_{T}}+\int_{0}^{T}\big{|}{\mathrm{d}}\bm{V}_{s}\big{|}\bigg{\|}_{L^{p}}<\infty,$ (2.1)
    $\big{\|}\bm{Y}\big{\|}_{{\mathcal{H}}^{p}}:=\inf_{\bm{N},\bm{U}}\bigg{\|}|\bm{Y}_{0}|+\sqrt{[\bm{N},\bm{N}]_{T}}+\int_{0}^{T}\big{|}{\mathrm{d}}\bm{U}_{s}\big{|}\bigg{\|}_{L^{p}}<\infty,$
    where the infimum is taken over all possible decompositions
    $\bm{X}_{t}=\bm{X}_{0}+\bm{M}_{t}+\bm{V}_{t}\text{ and }\bm{Y}_{t}=\bm{Y}_{0}+\bm{N}_{t}+\bm{U}_{t}.$
    Here $\bm{V}=(\bm{V}_{t})_{t\in[0,T]}$ and $\bm{U}=(\bm{U}_{t})_{t\in[0,T]}$ are adapted càdlàg processes of finite variation with $\bm{V}_{0}=\bm{U}_{0}=\bm{0}$, and $\bm{M}=(\bm{M}_{t})_{t\in[0,T]}$ and $\bm{N}=(\bm{N}_{t})_{t\in[0,T]}$ are ${\mathbb{F}}$-local martingales such that $\bm{M}_{0}=\bm{N}_{0}=\bm{0}$.
  • We assume also
    ${\mathbb{E}}\bigg{[}\Big{(}\sum_{0\leq s\leq T}|\Delta\bm{X}_{s}|\Big{)}^{p}\bigg{]}+{\mathbb{E}}\bigg{[}\Big{(}\sum_{0\leq s\leq T}|\Delta\bm{Y}_{s}|\Big{)}^{p}\bigg{]}<\infty.$ (2.2)
Under these assumptions, a proper form of dominated convergence theorem has been established ([37, Page 273, Lemma]) and the following two propositions hold.
Proposition 2.2 ([37, Section V, Theorem 2]).
For any $1\leq p\leq\infty$, consider a norm on the space of ${\mathbb{R}}$-valued adapted càdlàg processes (or the space of ${\mathbb{R}}$-valued adapted càglàd processes) defined by
$\|H\|_{{\mathcal{S}}^{p}}:=\Big{\|}\sup_{0\leq t\leq T}|H_{t}|\Big{\|}_{L^{p}}.$
Then, for $1\leq p<\infty$, there exists a constant $c_{p}$ depending only on $p$ such that
$\|H\|_{{\mathcal{S}}^{p}}\leq c_{p}\|H\|_{{\mathcal{H}}^{p}},$
for any ${\mathbb{R}}$-valued semimartingale $H$.
Proposition 2.3 (Emery inequality, [37, Section V, Theorem 3]).
Let $Z$ be an ${\mathbb{R}}$-valued semimartingale, $H$ is an ${\mathbb{R}}$-valued adapted càglàd process, and $\frac{1}{p}+\frac{1}{q}=\frac{1}{r}$ ($1\leq p,q\leq\infty)$, then
$\Big{\|}\int_{0}^{\cdot}H_{s}{\mathrm{d}}Z_{s}\Big{\|}_{{\mathcal{H}}^{r}}\leq\|H\|_{{\mathcal{S}}^{p}}\|Z\|_{{\mathcal{H}}^{q}}.$
Remark 2.4.
Propositions 2.2 and 2.3 also hold for ${\mathbb{R}}^{d}$-valued semimartingales with appropriate forms in $d\in{\mathbb{N}}^{+}$.
These propositions are important for the subsequent analysis, for instance in the localization argument for establishing Itô’s lemma for the flow of conditional measure on semimartinagles, as will be clear in Section 4.2.
Assumptions on filtration.
Assumption 2.5 (Filtration).
Sub-filtration ${\mathbb{G}}\subset{\mathbb{F}}$ satisfies the compatibility assumption or the conditional independence condition. That is, for each $0\leq t\leq T$, ${\mathcal{F}}_{t}$ and ${\mathcal{G}}_{T}$ are independent, given ${\mathcal{G}}_{t}$, written as ${\mathcal{F}}_{t}\perp\!\!\!\perp{\mathcal{G}}_{T}\big{|}{\mathcal{G}}_{t}.$
In the context of mean field games, this assumption ensures a weak closure of adapted processes, as first noted in [12, Lemma 3.11], and then stated more explicitly in [4, Theorem 5.4].
One can deduce from is assumption that $\text{Law}(\bm{X}_{t}|{\mathcal{G}}_{t})=\text{Law}(\bm{X}_{t}|{\mathcal{G}}_{T})$, for any ${\mathbb{F}}$-adapted process $\bm{X}=(\bm{X}_{t})_{t\in[0,T]}$ almost surely for each $t\leq T$, according to a special form of the (H) hypothesis by freezing ${\mathcal{G}}_{u}={\mathcal{G}}_{T}$ when $u\geq T$ ([20]). This allows us to define a cádlág version of $\mu_{t}:=\text{Law}(\bm{X}_{t}|{\mathcal{G}}_{t})$ by taking $\text{Law}(\bm{X}_{t}|{\mathcal{G}}_{T})$, which will be assumed throughout the paper.
Assumptions on function $\Phi$. We will adopt the the linear derivative adapted from [16], which has shown to be appropriate for the analyzing the flow of measures on semimartingales [24, 16].
Definition 2.6 ($C^{2,2}({\mathcal{P}}_{p}({\mathbb{R}}^{d})\times{\mathbb{R}}^{d})$ functions).
For $p\geq 2$, a function $\Phi(\mu,\bm{y})$ is a $C^{2,2}({\mathcal{P}}_{p}({\mathbb{R}}^{d})\times{\mathbb{R}}^{d})$ if for any $\mu\in\mathcal{P}_{p}({\mathbb{R}}^{d})$ and $\bm{y}\in{\mathbb{R}}^{d}$, there exists a continuous mapping $(\mu,\bm{y},\bm{x}_{1})\to\frac{\delta\Phi}{\delta\mu}(\mu,\bm{y},\bm{x}_{1})$, and there exists a continuous mapping $(\mu,\bm{y},\bm{x}_{1},\bm{x}_{2})\to\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}(\mu,\bm{y},\bm{x}_{1},\bm{x}_{2})$ that is symmetric in its two arguments and with the following properties,
  • continuously differentiable: $\nabla_{\bm{y}}\Phi(\mu,\bm{y})$, $\nabla_{\bm{y}}^{2}\Phi(\mu,\bm{y})$, $\frac{\delta\Phi}{\delta\mu}(\mu,\bm{y},\bm{x}_{1})$, $\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}(\mu,\bm{y},\bm{x}_{1})$, $\nabla_{\bm{y}}\frac{\delta\Phi}{\delta\mu}(\mu,\bm{y},\bm{x}_{1})$ $\nabla^{2}_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}(\mu,\bm{y},\bm{x}_{1})$, $\nabla_{\bm{x}_{1}}\nabla_{\bm{y}}\frac{\delta\Phi}{\delta\mu}(\mu,\bm{y},\bm{x}_{1})$, $\nabla_{\bm{x}_{1}}\nabla_{\bm{x}_{2}}\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}(\mu,\bm{y},\bm{x}_{1},\bm{x}_{2})$ all exist and are continuous functions for all $\bm{y},\bm{x},\bm{x}^{\prime}\in{\mathbb{R}}^{d}$, $\mu\in{\mathcal{P}}_{p}({\mathbb{R}}^{d})$,
  • uniform polynomial-growth : there exists a constant $c>0$ such that for all $\bm{x}_{1},\bm{x}_{2},\bm{y}\in{\mathbb{R}}^{d}$, $\mu\in{\mathcal{P}}({\mathbb{R}}^{d})$, we have
    $\Big{|}\nabla_{\bm{y}}\Phi\Big{|}\leq c\big{(}1+|\bm{y}|^{p-1}\big{)},\quad\Big{|}\nabla^{2}_{\bm{y}}\Phi\Big{|}$ $\leq c\big{(}1+|\bm{y}|^{p-2}\big{)},$
    $\bigg{|}\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\bigg{|}+\bigg{|}\nabla_{\bm{y}}\frac{\delta\Phi}{\delta\mu}\bigg{|}$ $\leq c\big{(}1+|\bm{x}_{1}|^{p-1}+|\bm{y}|^{p-1}\big{)},$
    $\bigg{|}\nabla^{2}_{\bm{x}}\frac{\delta\Phi}{\delta\mu}\bigg{|}+\bigg{|}\nabla_{\bm{x}_{1}\bm{y}}\frac{\delta\Phi}{\delta\mu}\bigg{|}$ $\leq c\big{(}1+|\bm{x}_{1}|^{p-2}+|\bm{y}|^{{p-2}}\big{)},$
    $\bigg{|}\nabla_{\bm{x}_{1},\bm{x}_{2}}\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\bigg{|}$ $\leq c\big{(}1+|\bm{x}_{1}|^{p-2}+|\bm{x}_{2}|^{p-2}+|\bm{y}|^{p-2}\big{)},$
  • fundamental theorem of calculus: for everything $\mu,\nu\in{\mathcal{P}}_{2}({\mathbb{R}}^{d})$, we have
    $\Phi(\mu,\bm{y})-$ $\Phi(\nu,\bm{y})=\int_{0}^{1}\int_{{\mathbb{R}}^{d}}\frac{\delta\Phi}{\delta\mu}\big{(}\lambda\mu+(1-\lambda)\nu,\bm{y},\bm{x}_{1}\big{)}(\mu-\nu)({\mathrm{d}}\bm{x}_{1}){\mathrm{d}}\lambda,$
    $\Phi(\bm{x}_{0},\mu)-$ $\Phi(\bm{x}_{0},\nu)=\int_{0}^{1}\int_{{\mathbb{R}}^{d}}\frac{\delta\Phi}{\delta\mu}\big{(}\nu,\bm{y},\bm{x}_{1}\big{)}(\mu-\nu)({\mathrm{d}}\bm{x}_{1}){\mathrm{d}}\lambda$
    $+\int_{0}^{1}\int_{0}^{t}\int_{{\mathbb{R}}^{d}}\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}s\mu+(1-s)\nu,\bm{y},\bm{x}_{1},\bm{x}_{2}\big{)}(\mu-\nu)({\mathrm{d}}\bm{x}_{1})(\mu-\nu)({\mathrm{d}}\bm{x}_{2}){\mathrm{d}}s{\mathrm{d}}t.$
Such $\frac{\delta\Phi}{\delta\mu}$ and $\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}$ are called (a version of) the linear derivative and the second derivative of $\Phi$.
2.2. Main result
Theorem 2.7 (Itô’s formula for flows of conditional measure on semimartingales).
Given a completed filtered probability space $(\Omega,{\mathcal{F}},{\mathbb{F}}=({\mathcal{F}}_{t})_{t\in[0,T]},{\mathbb{P}})$, which supports two ${\mathbb{R}}^{d}$-valued ${\mathbb{F}}$-semimartingales $\bm{X}=(\bm{X}_{t})_{t\in[0,T]}$ and $\bm{Y}=(\bm{Y}_{t})_{t\in[0,T]}$ satisfying Assumption 2.1. Let $\mu_{t}=$Law$(X_{t}|{\mathcal{G}}_{t})$, with the subfiltration ${\mathbb{G}}=({\mathcal{G}}_{t})_{t\in[0,T]}\subset{\mathbb{F}}$ satisfying Assumption 2.5. Then, for $p\geq 2$ and any $\Phi\in C^{2,2}(\mathcal{P}_{p}({\mathbb{R}}^{d})\times{\mathbb{R}}^{d})$, we have
$\Phi(\mu_{t},\bm{Y}_{t})-\Phi(\mu_{0},\bm{Y}_{0})$ (2.3)
$=$ $\;\overline{\mathbb{E}}\bigg{[}\int_{0+}^{t}\bigg{(}\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\cdot{\mathrm{d}}(\bm{X}^{\prime})^{c}_{s}+\frac{1}{2}\nabla^{2}_{\bm{x}_{1}\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}:{\mathrm{d}}[\bm{X}^{\prime},\bm{X}^{\prime}]^{c}_{s}$
$\qquad\quad+\frac{1}{2}\nabla^{2}_{\bm{x}_{1}\bm{x}_{2}}\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-},\bm{X}^{\prime\prime}_{s-}\big{)}:{\mathrm{d}}[\bm{X}^{\prime},\bm{X}^{\prime\prime}]^{c}_{t}+\nabla^{2}_{\bm{x}_{1}\bm{y}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}:{\mathrm{d}}[\bm{X}^{\prime},\bm{Y}]^{c}_{s}\bigg{)}$
$\quad\quad+\sum_{0<s\leq t}\bigg{(}\frac{1}{2}\bigg{(}\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s},\bm{X}^{\prime\prime}_{s}\big{)}-\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-},\bm{X}^{\prime\prime}_{s}\big{)}$
$\qquad\qquad\qquad-\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s},\bm{X}^{\prime\prime}_{s-}\big{)}+\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-},\bm{X}^{\prime\prime}_{s-}\big{)}\bigg{)}$
$\qquad\qquad\qquad+\bigg{(}\nabla_{\bm{y}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s}\big{)}-\nabla_{\bm{y}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\bigg{)}\cdot\Delta\bm{Y}_{s}$
$\quad\qquad\qquad+\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s}\big{)}-\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\bigg{)}\mathbbm{1}_{\{\mu_{s}=\mu_{s-}\}}\bigg{|}{\mathcal{F}}\bigg{]}$
$+\int_{0+}^{t}\nabla_{\bm{y}}\Phi(\mu_{s-},\bm{Y}_{s-})\cdot{\mathrm{d}}(\bm{Y})^{c}_{s}+\frac{1}{2}\int_{0+}^{t}\nabla_{\bm{y}\bm{y}}^{2}\Phi(\mu_{s-},\bm{Y}_{s-}):{\mathrm{d}}[\bm{Y},\bm{Y}]^{c}_{s}$
$+\sum_{0<s\leq t}\Big{(}\Phi(\mu_{s},\bm{Y}_{s})-\Phi(\mu_{s-},\bm{Y}_{s-})\Big{)},$
for all $t\in[0,T]$, where $\bm{X}^{\prime},\bm{X}^{\prime\prime}$ are the conditional independent copies of $\bm{X}$ given the sub $\sigma$-algebra ${\mathcal{G}}_{T}$ defined in an enlarged probability space $(\overline{\Omega},\overline{{\mathcal{F}}},\overline{{\mathbb{P}}})$ and $\overline{{\mathbb{E}}}[\cdot|{\mathcal{F}}]$ is the conditional expectation given the extension ${\mathcal{F}}\subset\overline{{\mathcal{F}}}$ in the enlarged probability space.
Remark 2.8.
To ensure that the enlarged space $(\overline{\Omega},\overline{{\mathcal{F}}},\overline{{\mathbb{F}}},\overline{{\mathbb{P}}})$ is properly defined such that $\mu_{t},\bm{X}_{t},\bm{Y}_{t}$ and ${\mathcal{F}}$ are naturally extended in this enlarged space, one needs the notion of independent copies of stochastic processes, as will be developed in the next section and discussed after Corollary 3.2.
3. Conditional independent copy
3.1. Why conditional independent copy?
One of the key components in Theorem 2.7 is the concept of the conditional independent copy of stochastic processes. Here in this section, we illustrate through several examples its properties in the context of flow of conditional laws for semimartingales and its role in the derivation of Itô’s formula.
Let us first recall a mean-field game with common noise where $X_{t}$ is of the form
${\mathrm{d}}X_{t}=b_{t}{\mathrm{d}}t+\sigma_{t}{\mathrm{d}}W_{t}+\sigma^{0}_{t}{\mathrm{d}}W^{0}_{t},$
where $W,W^{0}$ are two independent Brownian motions and ${\mathbb{G}}=\{{\mathcal{G}}_{t}\}_{t\in[0,T]}$ is the sub filtration generated by $W^{0}$ and ${\mathcal{G}}={\mathcal{G}}_{T}=\sigma(W^{0}_{u},u\leq T)$. By the classical Itô’s formula, the dynamic of $g(X_{t})$ is given by
${\mathrm{d}}g(X_{t})=b_{t}g^{\prime}(X_{t})+\frac{1}{2}(\sigma_{t}^{2}+(\sigma^{0}_{t})^{2})g^{\prime\prime}(X_{t}){\mathrm{d}}t+\sigma_{t}g^{\prime}(X_{t}){\mathrm{d}}W_{t}+\sigma^{0}_{t}g^{\prime}(X_{t}){\mathrm{d}}W^{0}_{t},$
and consequently the dynamic of $Z_{t}={\mathbb{E}}[g(X_{t})|{\mathcal{G}}]$ is given by
${\mathrm{d}}Z_{t}={\mathbb{E}}\bigg{[}b_{t}g^{\prime}(X_{t})+\frac{1}{2}(\sigma_{t}^{2}+(\sigma^{0}_{t})^{2})g^{\prime\prime}(X_{t})\Big{|}{\mathcal{G}}\bigg{]}{\mathrm{d}}t+{\mathbb{E}}\big{[}\sigma^{0}_{t}g^{\prime}(X_{t})|{\mathcal{G}}\big{]}{\mathrm{d}}W^{0}_{t}.$ (3.1)
The above equation follows essentially from the fact that
${\mathbb{E}}\bigg{[}\int_{0}^{t}\eta_{s}{\mathrm{d}}W_{s}\Big{|}{\mathcal{G}}\bigg{]}=0,\qquad{\mathbb{E}}\bigg{[}\int_{0}^{t}\eta_{s}{\mathrm{d}}W^{0}_{s}\Big{|}{\mathcal{G}}\bigg{]}=\int_{0}^{t}{\mathbb{E}}[\eta_{s}|{\mathcal{G}}]{\mathrm{d}}W^{0}_{s},\text{ for all }t\in[0,T]$ (3.2)
see [34, Lemma B.1] for details. Indeed, from (3.1) one can calculate the quadratic variation of $Z_{t}$ to derive the Itô’s formula for $f(Z_{t})$:
${\mathrm{d}}\langle Z\rangle_{t}={\mathbb{E}}\big{[}\sigma^{0}_{t}g^{\prime}(X_{t})|{\mathcal{G}}\big{]}^{2}{\mathrm{d}}t={\mathbb{E}}[\sigma^{0}_{t}g^{\prime}(X_{t})\widetilde{\sigma}^{0}_{t}g^{\prime}(\widetilde{X}_{t})|{\mathcal{G}}]{\mathrm{d}}t,$
where $\widetilde{\sigma}_{t}^{0},\widetilde{X}_{t}$ are conditional respective independent copy of $\sigma^{0}_{t}$ and $X_{t}$ given ${\mathcal{G}}$, which is formally introduced in [11]. Under these expressions, one can derive Itô’s formula for $f(Z_{t})$, which depends on the precise expression of
$\int_{0}^{t}f^{\prime}(Z_{s}){\mathrm{d}}Z_{s}=$ $\int_{0}^{t}f^{\prime}(Y_{s})\bigg{(}{\mathbb{E}}\Big{[}b_{s}g^{\prime}(X_{s})+\frac{1}{2}(\sigma_{s}^{2}+(\sigma^{0}_{s})^{2})g^{\prime\prime}(X_{s})\Big{|}{\mathcal{G}}\Big{]}{\mathrm{d}}s+{\mathbb{E}}\big{[}\sigma^{0}_{t}g^{\prime}(X_{s})|{\mathcal{G}}\big{]}{\mathrm{d}}W^{0}_{s}\bigg{)}$
$=\int_{0}^{t}\bigg{(}{\mathbb{E}}$ $\Big{[}f^{\prime}(Z_{s})\Big{(}b_{s}g^{\prime}(X_{s})+\frac{1}{2}(\sigma_{s}^{2}+(\sigma^{0}_{s})^{2})g^{\prime\prime}(X_{s})\Big{)}\Big{|}{\mathcal{G}}\Big{]}{\mathrm{d}}s+{\mathbb{E}}\big{[}f^{\prime}(Z_{s})\sigma^{0}_{t}g^{\prime}(X_{s})|{\mathcal{G}}\big{]}{\mathrm{d}}W^{0}_{s}\bigg{)}$
and
$\int_{0}^{t}f^{\prime\prime}(Z_{s}){\mathrm{d}}\langle Z\rangle_{s}$ $=\int_{0}^{t}f^{\prime\prime}(Z_{s})\cdot\frac{1}{2}{\mathbb{E}}\big{[}\sigma^{0}_{s}g^{\prime}(X_{s})\widetilde{\sigma}^{0}_{s}g^{\prime}(\widetilde{X}_{s})\big{|}{\mathcal{G}}_{T}\big{]}{\mathrm{d}}s$
$=\frac{1}{2}\int_{0}^{t}{\mathbb{E}}\big{[}f^{\prime\prime}(Z_{s})\sigma^{0}_{s}g^{\prime}(X_{s})\widetilde{\sigma}^{0}_{s}g^{\prime}(\widetilde{X}_{s})\big{|}{\mathcal{G}}_{T}\big{]}{\mathrm{d}}s.$
Now, the issue arises when one considers a general continuous semimartingale $X$ and a general sub $\sigma$-algebra ${\mathcal{G}}$, where the Itô’s formula takes the form:
${\mathrm{d}}g(X_{t})=g^{\prime}(X_{t}){\mathrm{d}}X_{t}+\frac{1}{2}g^{\prime\prime}(X_{t}){\mathrm{d}}\langle X\rangle_{t}.$ (3.3)
In this case, in order to characterize the dynamics of $Z_{t}={\mathbb{E}}[g(X_{t})|{\mathcal{G}}]$, one needs an appropriate expression for ${\mathrm{d}}Z_{t}$, ${\mathrm{d}}\langle Z\rangle_{t}$ and more importantly for
$\int_{0}^{t}f^{\prime}(Z_{s}){\mathrm{d}}Z_{s},\text{ and }\int_{0}^{t}f^{\prime\prime}(Z_{s}){\mathrm{d}}\langle Z\rangle_{s}.$ (3.4)
One naive guess of the extension of (3.2) to this general case is
${\mathbb{E}}\bigg{[}\int_{0}^{t}\eta_{s}{\mathrm{d}}X_{s}\Big{|}{\mathcal{G}}\bigg{]}=\int_{0}^{t}{\mathbb{E}}[\eta_{s}|{\mathcal{G}}]{\mathrm{d}}X^{\mathcal{G}}_{s},$ (3.5)
where $X^{\mathcal{G}}_{t}:={\mathbb{E}}[X_{t}|{\mathcal{G}}]$.
Unfortunately this does not hold in general. For instance, take $X_{t}$ from (3.1) with $b_{t}=0$. By (3.2), the left hand side of (3.5) becomes
${\mathbb{E}}\bigg{[}\int_{0}^{t}\eta_{s}{\mathrm{d}}X_{s}\Big{|}{\mathcal{G}}\bigg{]}={\mathbb{E}}\bigg{[}\int_{0}^{t}\eta_{s}\big{(}\sigma_{s}{\mathrm{d}}W_{s}+\sigma^{0}_{s}{\mathrm{d}}W^{0}_{s}\big{)}\Big{|}{\mathcal{G}}\bigg{]}=\int_{0}^{t}{\mathbb{E}}\big{[}\eta_{s}\sigma^{0}_{s}\big{|}{\mathcal{G}}\big{]}{\mathrm{d}}W^{0}_{s},$
while the right hand side is
$\int_{0}^{t}{\mathbb{E}}[\eta_{s}|{\mathcal{G}}]{\mathrm{d}}X^{\mathcal{G}}_{s}=\int_{0}^{t}{\mathbb{E}}[\eta_{s}|{\mathcal{G}}]{\mathbb{E}}[\sigma^{0}_{s}|{\mathcal{G}}]{\mathrm{d}}W^{0}_{s},$
and (3.5) holds if and only if $\eta_{s}$ and $\sigma_{s}^{0}$ are conditionally uncorrelated of ${\mathcal{G}}$.
This is exactly why the concept of conditional independent copies of stochastic processes: instead of (3.2), let us consider the expression of the following form
$\int_{0}^{t}f^{\prime}(Z_{s}){\mathrm{d}}Z_{s}$ $=\int_{0}^{t}f^{\prime}(Z_{s}){\mathrm{d}}{\mathbb{E}}\big{[}g(X_{s})|{\mathcal{G}}\big{]}=\int_{0}^{t}f^{\prime}(Z_{s}){\mathrm{d}}{\mathbb{E}}\big{[}g(X^{(1)}_{s})|{\mathcal{F}}\big{]}={\mathbb{E}}\bigg{[}\int_{0}^{t}f^{\prime}(Z_{s}){\mathrm{d}}g(X^{(1)}_{s})\Big{|}{\mathcal{F}}\bigg{]}$
$={\mathbb{E}}\bigg{[}\int_{0}^{t}f^{\prime}(Z_{s})\Big{(}g^{\prime}(X^{(1)}_{s}){\mathrm{d}}X^{(1)}_{s}+\frac{1}{2}g^{\prime\prime}(X^{(1)}_{s}){\mathrm{d}}\langle X^{(1)}\rangle_{s}\Big{)}\Big{|}{\mathcal{F}}\bigg{]},$
and
$\int_{0}^{t}f^{\prime\prime}(Z_{s}){\mathrm{d}}\langle Z\rangle_{s}$ $=\int_{0}^{t}f^{\prime}(Z_{s}){\mathrm{d}}\langle{\mathbb{E}}\big{[}g(X_{\cdot})|{\mathcal{G}}\big{]}\rangle_{s}={\mathbb{E}}\bigg{[}\int_{0}^{t}f^{\prime}(Y_{s}){\mathrm{d}}\langle g(X^{(1)}_{\cdot}),g(X^{(2)}_{\cdot})\rangle_{s}|{\mathcal{F}}\bigg{]}$
$={\mathbb{E}}\bigg{[}\int_{0}^{t}f^{\prime\prime}(Z_{s})g^{\prime}(X^{(1)}_{s})g^{\prime}(X^{(2)}_{s}){\mathrm{d}}\langle X^{(1)},X^{(2)}\rangle_{s}\Big{|}{\mathcal{F}}\bigg{]},$
where $X^{(1)}$, $X^{(2)}$ are two conditional independent copies of $X$ give ${\mathcal{G}}$. With these expressions, one can analyze (3.4) for a general semimartingale $X$, as will be detailed in (2c) and (2d) of Theorem 3.3.
3.2. Examples
To get some intuition for this notion of conditional independent copies of stochastic processes, let us see some examples when ${\mathcal{F}}$ and the sub $\sigma$-algebra ${\mathcal{G}}$ have some special structure.
Example 1. Let $A,B$ are two independent random variables with ${\mathcal{F}}=\sigma(A,B)$ and ${\mathcal{G}}=\sigma(A)$. Then one can enlarge the probability space to include $A^{(i)}$ such that $A^{(i)}$ are independent and with the same distribution of $A$, and are independent of $B$. Then for any $X=f(A,B)$, the conditional independent copies of $X$ are $X^{(i)}=f(A^{(i)},B)$.
Example 2. Let $X$ be the unique strong solution of the SDE
${\mathrm{d}}X_{t}=b(X_{t}){\mathrm{d}}t+\sigma(X_{t}){\mathrm{d}}W_{t}+\sigma^{(0)}(X_{t}){\mathrm{d}}W^{(0)}_{t},$
where $W,W^{(0)}$ are two independent Brownian motions, and $b$, $\sigma$, and $\sigma^{0}$ satisfy appropriate continuity conditions. Let ${\mathcal{F}}_{t}=\sigma\big{(}\{W_{s},W^{(0}_{s}\}_{s\leq t}\big{)}$ and sub $\sigma$-algebra ${\mathcal{G}}=\sigma(W^{(0)})$, then one can enlarge the probability space to include independent Brownian motions $W^{(1)}$ and $X^{(1)}$, with $X^{(1)}$ the unique strong solution of
${\mathrm{d}}X^{(1)}_{t}=b(X^{(1)}_{t}){\mathrm{d}}t+\sigma(X^{(1)}_{t}){\mathrm{d}}W^{(1)}_{t}+\sigma^{(0)}(X^{(1)}_{t}){\mathrm{d}}W^{(0)}_{t},$
and $X^{(1)}$ is the conditional independent copies of $X$ given ${\mathcal{G}}$.
There are more examples of conditional independent copies of stochastic processes in the literature of mean-field games. For instance, when defining a semimartingale by stochastic differential equations involving both idiosyncratic and common noise, conditional copies are formally formulated in [11] as $\Omega=\Omega_{0}\times\widetilde{\Omega}_{1}$, where the original probability space follows the structure $\Omega=\Omega_{0}\times\Omega_{1}$, and $\widetilde{\Omega}_{1}$ represents the independent copy of the probability space $\Omega_{1}$. [11, Theorem 4.14, Volume II]
In the next section, we will provide the construction of conditional independent copies for the general case of a semimartingale $\bm{X}$ and an arbitrary sub $\sigma$-algebra ${\mathcal{G}}$ which is not necessarily associated with any probability space. As emphasized earlier, this construction is necessary for establishing Itô’s formula with the flow of conditional measures on semimartigales.
3.3. Construction of conditional independent copies: general cases
Given a probability space $(\Omega,{\mathcal{F}},{\mathbb{P}})$, let us consider the extended probability space $\overline{\Omega}=\Omega^{n+1}$ defined by
$\overline{\Omega}=\Omega^{n+1}=\{(\omega_{0},\omega_{1},\cdots,\omega_{n})|\omega_{0},\omega_{1},\cdots,\omega_{n}\in\Omega\},$ (3.6)
and the extended $\sigma$-algebra $\overline{\mathcal{F}}$
¯$\overline{\mathcal{F}}=\sigma\{A_{0}\times A_{1}\times\cdots\times A_{n}:A_{0},A_{1},\cdots,A_{n}\in{\mathcal{F}}\}.$ (3.7)
To ensure that these copies are conditional independent, let us define the probability measure $\overline{\mathbb{P}}$ as follows: for $A_{0},A_{1},\cdots,A_{n}\in{\mathcal{F}}$, define
¯$\overline{\mathbb{P}}\big{(}A_{0}\times A_{1}\times\cdots\times A_{n}\big{)}:={\mathbb{E}}\bigg{[}\mathbbm{1}_{A_{0}}\prod_{i=1}^{n}{\mathbb{P}}(A_{i}|{\mathcal{G}})\bigg{]}.$ (3.8)
This measure is properly defined since the set $\{A_{0}\times A_{1}\times\cdots\times A_{n}\}$ is a $\pi$-system. Finally, define ~$\widetilde{\mathcal{F}},\widetilde{\mathcal{G}}$ for the extension of ${\mathcal{F}}$ and ${\mathcal{G}}$ into this enlarged probability space as
~:=$\widetilde{\mathcal{F}}:=\{A\times\Omega^{n}:A\in{\mathcal{F}}\},\quad\widetilde{\mathcal{G}}:=\{A\times\Omega^{n}:A\in{\mathcal{G}}\}.$ (3.9)
Then we have:
Theorem 3.1.
Given a probability space $(\Omega,{\mathcal{F}},{\mathbb{P}})$ and a sub $\sigma$-algebra ${\mathcal{G}}\subset{\mathcal{F}}$. Then there exists uniquely an enlarged probability space $(\overline{\Omega},\overline{\mathcal{F}},{\mathbb{P}})$ satisfying (3.6), (3.7) and (3.8); and for any random variable $R$ in $(\Omega,{\mathbb{F}},{\mathbb{P}})$, there exist an $\widetilde{R}$ and $n$ conditional independent copies $\{R^{(i)}\}_{i=1,2,..n}$ by
$\widetilde{R}(\omega_{0},\omega_{1},\cdots,\omega_{n})=R(\omega_{0}),\;R^{(i)}(\omega_{0},\omega_{1},\cdots,\omega_{n})=R(\omega_{i}),$ (3.10)
for $\overline{\omega}=(\omega_{0},,\omega_{1},\cdots,\omega_{n})\in\overline{\Omega}$ and $\widetilde{R}$, $\{R^{(i)}\}_{i=1,2,..n}$ satisfying
  • $\widetilde{R},\big{\{}R^{(i)}\big{\}}_{i=1,2,..n}$ are independent given $\widetilde{\mathcal{G}}$,
  • $\widetilde{R},\big{\{}R^{(i)}\big{\}}_{i=1,2,..n}$ are independent given $\widetilde{\mathcal{F}}$,
  • Law$\big{(}R\big{|}{\mathcal{G}}\big{)}(\omega_{0})=$Law$\big{(}\widetilde{R}\big{|}\widetilde{\mathcal{G}}\big{)}(\overline{\omega})=$Law$\big{(}R^{(i)}\big{|}{\widetilde{\mathcal{G}}}\big{)}(\overline{\omega})=$Law$\big{(}R^{(i)}\big{|}{\widetilde{\mathcal{F}}}\big{)}(\overline{\omega})$, for all $i=1,2,..,n$, a.s.
with $\overline{\omega}=(\omega_{0},\cdots,\omega_{n})\in\overline{\Omega}$ and $\widetilde{\mathcal{F}}$, $\widetilde{\mathcal{G}}$ defined by (3.9).
Note that when ${\mathcal{G}}$ is a trivial $\sigma$-algebra $\{\emptyset,\Omega\}$, then the above theorem reduces to the standard construction of independent copies. Its proof is given in Section 5.1.
Now, in the context of semimartingales, there is a corresponding version of the conditional independent copy where the probability space is a filtered space and sub $\sigma$-algebra ${\mathcal{G}}={\mathcal{G}}_{T}$ comes from the sub filtration ${\mathbb{G}}=({\mathcal{G}}_{t})_{t\in[0,T]}$. Now define $\overline{\mathbb{F}}:=(\overline{\mathcal{F}}_{t})_{t\in(0,T)}$ by
¯$\overline{\mathcal{F}}_{t}=\sigma\{A_{0}\times A_{1}\times\cdots\times A_{n}:A_{0},A_{1},\cdots,A_{n}\in{\mathcal{F}}_{t}\},$ (3.11)
and further define $\widetilde{\mathbb{F}}=(\widetilde{\mathcal{F}}_{t})_{t\in(0,T)}$ and $\widetilde{\mathbb{G}}=(\widetilde{\mathcal{G}}_{t})_{t\in(0,T)}$ respectively for the extension of ${\mathbb{F}}$ and ${\mathbb{G}}$ into this enlarged probability space as
~$\widetilde{\mathcal{F}}_{t}:=\{A\times\Omega^{n}:A\in{\mathcal{F}}_{t}\},\quad\widetilde{\mathcal{G}}_{t}:=\{A\times\Omega^{n}:A\in{\mathcal{G}}_{t}\}.$ (3.12)
With this construction, we have
Corollary 3.2.
Given two ${\mathbb{R}}^{d}$-valued ${\mathbb{F}}$-semimartingales $\bm{X}=(\bm{X}_{t})_{t\in[0,T]}$ and $\bm{Y}=(\bm{X}_{t})_{t\in[0,T]}$ on a filtered probability space $(\Omega,{\mathcal{F}},{\mathbb{F}}=({\mathcal{F}}_{t})_{t\in[0,T]},{\mathbb{P}})$, and a sub $\sigma$-algebra ${\mathcal{G}}\subset{\mathcal{F}}$, there exists uniquely an enlarged probability space $(\overline{\Omega},\overline{\mathcal{F}},\overline{\mathbb{F}}=({\mathcal{F}}_{t})_{t\in[0,T]},\overline{\mathbb{P}})$ satisfying (3.6), (3.7), (3.8) and (3.11). Moreover, one can define two conditional independent copies $(\bm{X}^{\prime},\bm{Y}^{\prime})$, $(\bm{X}^{\prime\prime},\bm{Y}^{\prime\prime})$ of the random processes $(\bm{X},\bm{Y})$ by
$(\bm{X}^{\prime},\bm{Y}^{\prime})(\omega_{0},\omega_{1},\cdots,\omega_{n})=(\bm{X},\bm{Y})(\omega_{1}),\;(\bm{X}^{\prime\prime},\bm{Y}^{\prime\prime})(\omega_{0},\omega_{1},\cdots,\omega_{n})=(\bm{X},\bm{Y})(\omega_{2}),$
and $({\bm{X}},{\bm{Y}})$, $(\bm{X}^{\prime},\bm{Y}^{\prime})$ and $(\bm{X}^{\prime\prime},\bm{Y}^{\prime\prime})$ satisfying
$\begin{cases}\text{Law}\big{(}{\bm{X}},{\bm{Y}}\big{|}{\mathcal{G}}\big{)}=\text{Law}\big{(}\bm{X}^{\prime},\bm{Y}^{\prime}\big{|}{\mathcal{G}}\big{)}=\text{Law}\big{(}\bm{X}^{\prime},\bm{Y}^{\prime}\big{|}{\mathcal{F}}\big{)}\\ \qquad=\text{Law}\big{(}\bm{X}^{\prime\prime},\bm{Y}^{\prime\prime}\big{|}{\mathcal{G}}\big{)}=\text{Law}\big{(}\bm{X}^{\prime\prime},\bm{Y}^{\prime\prime}\big{|}{\mathcal{F}}\big{)}\text{ a.s.}\\ ({\bm{X}},{\bm{Y}}),(\bm{X}^{\prime},\bm{Y}^{\prime}),(\bm{X}^{\prime\prime},\bm{Y}^{\prime\prime})\text{ are independent given }{\mathcal{G}},\text{and are also independent given }{\mathcal{F}}.\end{cases}$ (3.13)
Furthermore, if the sub $\sigma$-algebra ${\mathcal{G}}={\mathcal{G}}_{T}$ from a sub filtration ${\mathbb{G}}=({\mathcal{G}}_{t})_{t\in[0,T]}\subset{\mathbb{F}}$ satisfies the compatibility conditions as in Assumption 2.5, then this compatibility condition holds for the natural extension of $\overline{\mathbb{F}}$ and $\overline{\mathbb{G}}$ defined by (3.12) in the extended probability space. That is, for each $0\leq t\leq T$, ~$\widetilde{\mathcal{F}}_{t}$ and $\widetilde{\mathcal{G}}_{T}$ are independent given $\overline{\mathcal{G}}_{t}$, written as ~$\widetilde{\mathcal{F}}_{t}\perp\!\!\!\perp\widetilde{\mathcal{G}}_{T}\big{|}\widetilde{\mathcal{G}}_{t}.$
Note that Theorem 3.1 allows us to use the notation $(\bm{X},\bm{Y}),{\mathcal{F}}$, and ${\mathcal{G}}$ to respectively represent $(\widetilde{\bm{X}},\widetilde{\bm{Y}}),\widetilde{\mathcal{F}}$, and $\widetilde{\mathcal{G}}$ without any ambiguity, as is adopted in the above corollary. Furthermore, Corollary 3.2 enables simplifying the notation $\widetilde{\mathbb{F}}$ and $\widetilde{\mathbb{G}}$ respectively by ${\mathbb{F}}$ and ${\mathbb{G}}$ without any ambiguity. Moreover, if we define the conditional law in the enlarged probability space as $\widetilde{\mu}_{t}:=\text{Law}\big{(}\widetilde{\bm{X}}\big{|}\widetilde{\mathcal{G}}_{t}\big{)}$, then by the compatibility condition and Theorem 3.3, we have
$\widetilde{\mu}_{t}(\omega_{0},\omega_{1},\cdots,\omega_{n})=\text{Law}(\widetilde{\bm{X}_{t}}\big{|}\widetilde{\mathcal{G}}_{T})(\omega_{0},\omega_{1},\cdots,\omega_{n})=\text{Law}(\bm{X}_{t}\big{|}\widetilde{\mathcal{G}}_{T})(\omega_{0})=\mu_{t}(\omega_{0}),$
which in turn allows for using $\mu_{t}$ for $\widetilde{\mu}_{t}$ without any ambiguity.
3.4. Properties of the conditional independent copies of semimartingales
Next, we study properties of these conditional independent copies of semimartingales and explore the relation between the conditional expectation of their stochastic integrals and stochastic integrals with respect to the conditional law of semimartingales. For the sake of simplifying the notations, we will present the results in the one-dimensional case and their corresponding multi-dimensional cases can be adapted accordingly.
Theorem 3.3.
Give two ${\mathbb{R}}$-valued ${\mathbb{F}}$-semimartingales $\{X_{t}\}_{t\in[0,T]}$ and $\{Y_{t}\}_{t\in[0,T]}$ on a filtered probability space $(\Omega,{\mathcal{F}},{\mathbb{F}},{\mathbb{P}})$. Let $(X^{\prime},Y^{\prime})$ and $(X^{\prime\prime},Y^{\prime\prime})$ be their conditional independent copies defined in Corollary 3.2 in the enlarged filtered probability space $(\overline{\Omega},\overline{\mathcal{F}},\overline{\mathbb{F}}=({\mathcal{F}}_{t})_{t\in[0,T]},\overline{\mathbb{P}})$ satisfying (3.6), (3.8) and (3.11). Then,
  1. (1)
    $X$, $Y$, $X^{\prime}$, $Y^{\prime}$, $X^{\prime\prime}$, $Y^{\prime\prime}$ are $\overline{\mathbb{F}}$-semimartingales.
  2. (2)
    Suppose ${\mathbb{E}}[|X_{t}|]<\infty$ and ${\mathbb{E}}[|Y_{t}|]<\infty$ for all $t\in[0,T]$, and define
    $X^{\mathcal{G}}_{t}:=\overline{\mathbb{E}}\big{[}X_{t}\big{|}{\mathcal{G}}_{T}\big{]}=\overline{\mathbb{E}}\big{[}X_{t}\big{|}{\mathcal{G}}_{t}\big{]},\;\;Y^{\mathcal{G}}_{t}:=\overline{\mathbb{E}}\big{[}Y_{t}\big{|}{\mathcal{G}}_{T}\big{]}=\overline{\mathbb{E}}\big{[}Y_{t}\big{|}{\mathcal{G}}_{t}\big{]}.$
    1. (a)
      If $\{X_{t}\}_{t\in[0,T]}$ is a finite variation process and ${\mathbb{E}}\big{[}\big{(}\int_{0}^{T}|{\mathrm{d}}X_{s}|\big{)}^{p}\big{]}<\infty,$ for $p\geq 1$, then $\{X^{\mathcal{G}}_{t}\}_{t\in[0,T]}$ is also a finite variation process and
      $\overline{\mathbb{E}}\bigg{[}\bigg{(}\int_{0}^{T}\big{|}{\mathrm{d}}X^{\mathcal{G}}_{s}\big{|}\bigg{)}^{p}\bigg{]}<\infty.$
    2. (b)
      $\{X^{\mathcal{G}}_{t}\}_{t\in[0,T]}$ and $\{Y^{\mathcal{G}}_{t}\}_{t\in[0,T]}$ are semimartingales with respect to $\overline{\mathbb{F}}$, ${\mathbb{F}}$ and ${\mathbb{G}}$.
    3. (c)
      If $\{Z_{t}\}_{t\in[0,T]}$ is an ${\mathbb{R}}$-valued ${\mathbb{F}}$-adapted càdlàg process in the filtered probability space $(\Omega,{\mathcal{F}},{\mathbb{F}},{\mathbb{P}})$ and $\|X\|_{\mathcal{H}_{p}}+\|Z\|_{\mathcal{H}_{q}}<\infty$ with $\frac{1}{p}+\frac{1}{q}=1$, then
      $\int_{0}^{t}Z_{s-}{\mathrm{d}}X^{\mathcal{G}}_{s}=\overline{\mathbb{E}}\bigg{[}\int_{0}^{t}Z_{s-}{\mathrm{d}}X^{\prime}_{s}\bigg{|}{\mathcal{F}}\bigg{]}.$
    4. (d)
      If $\{Z_{t}\}_{t\in[0,T]}$ is an ${\mathbb{R}}$-valued ${\mathbb{F}}$-adapted càdlàg process in a filtered probability space $(\Omega,{\mathcal{F}},{\mathbb{F}},{\mathbb{P}})$ and $\|X\|_{\mathcal{H}_{p}}+\|Y\|_{\mathcal{H}_{q}}+\|Z\|_{\mathcal{H}_{r}}<\infty$ with $\frac{1}{p}+\frac{1}{q}+\frac{1}{r}=1$, then
      $\int_{0}^{t}Z_{s-}{\mathrm{d}}\big{[}X^{\mathcal{G}},Y^{\mathcal{G}}\big{]}_{s}=\overline{\mathbb{E}}\bigg{[}\int_{0}^{t}Z_{s-}{\mathrm{d}}\big{[}X^{\prime\prime},Y^{\prime}\big{]}_{s}\bigg{|}{\mathcal{F}}\bigg{]},$
      $\int_{0}^{t}Z_{s}{\mathrm{d}}\big{[}X^{\mathcal{G}},Y\big{]}_{s}=\overline{\mathbb{E}}\bigg{[}\int_{0}^{t}Z_{s}{\mathrm{d}}\big{[}X^{\prime},Y\big{]}_{s}\bigg{|}{\mathcal{F}}\bigg{]}.$
The first statement in Theorem 3.3 shows that semi-martingales remain semimartingales under the enlarged filtration $\overline{F}$. (2a) and (2b) suggest that under the conditional expectation operation, the finite variation property, the semimartingale property with respect to $\overline{\mathcal{F}}$ and ${\mathcal{F}}$ are preserved and the martingale property of ${\mathcal{G}}$ is unchanged because of the compatibility condition in Assumption 2.5. In particular, (2c) and (2d) are the key equations for the derivation of the Itô’s formula for cylindrical function; they ensure crucially the equivalence between the stochastic integral of the conditional expectation of a semimartingale and the conditional expectation of the Itô’s integral of its conditional independent copy, as well as the equivalence of the quadratic variation of the conditional expectation of a semimartingale and the conditional expectation of the quadratic variation of its two conditional independent copies.
The proof of the theorem is presented in Section 5.2.
4. Proof of Theorem 2.7
Having established the properties of conditional independent copy of semimartingales, we are now ready to establish Itô’s formula for flow of conditional laws on semimratingales, through several steps. First, we will derive the formula for cylindrical functions; we will then show that cylindrical functions are dense in $C^{2,2}({\mathbb{R}}^{d}\times\mathcal{P}_{p}({\mathbb{R}}^{d}))$. Finally, appropriate localization techniques and applications of appropriate dominance convergence theorem finish the proof.
4.1. Step 1: Ito’s formula for cylindrical functions.
First, let us recall the cylindrical functions (see for instance ([17]).
Definition 4.1 ($C^{2,2}$ cylindrical functions).
A function $\Phi:{\mathcal{P}}({\mathbb{R}}^{d})\times{\mathbb{R}}^{d}\to{\mathbb{R}}$ is a $C^{2,2}$ cylindrical function if
$\Phi(\mu,\bm{y})=f\big{(}\langle\mu,g^{(1)}\rangle,..\langle\mu,g^{(n)}\rangle,\bm{y}\big{)},$
where $f\in C^{2}({\mathbb{R}}^{n+d})$ and $g\in C^{2}_{b}({\mathbb{R}}^{d})$.
Next, we show that when restricted to the smaller set of functions $\Phi$, Theorem 2.7 holds. For semimartingales $\{\bm{X}_{t}\}_{t\in[0,T]}$ and $g^{(k)}\in C_{b}^{2}({\mathbb{R}}^{d})$, for $k=1,2,..,n$, by the classical Ito’s formula [37, Theorem 33, Chapter II], we have
$g^{(i)}(\bm{X}_{t})-g^{(i)}(\bm{X}_{0})=$ $\int_{0+}^{t}\nabla g^{(i)}(\bm{X}_{s-})\cdot{\mathrm{d}}\bm{X}^{c}_{s}+\frac{1}{2}\int_{0+}^{t}\nabla^{2}g^{(i)}(\bm{X}_{s-}):{\mathrm{d}}[\bm{X},\bm{X}]^{c}_{s}$
$\;\;+\sum_{0<s\leq t}\Big{\{}g^{(i)}(\bm{X}_{s})-g^{(i)}(\bm{X}_{s-})\Big{\}},\text{ for }t\in[0,T],$
and $\{g^{(i)}(\bm{X}_{t})\}_{t\in[0,T]}$ for $i=1,..,N$ are semimartingales. Taking conditional expectation given ${\mathcal{G}}_{T}$, and recall that $\mu_{t}=\text{Law}(\bm{X}_{t}|{\mathcal{G}}_{t})=\text{Law}(\bm{X}_{t}|{\mathcal{G}}_{T})$, we have
$\big{\langle}\mu_{t},g^{(i)}\big{\rangle}-\big{\langle}\mu_{0},g^{(i)}\big{\rangle}=$ $\;{\mathbb{E}}\bigg{[}\int_{0+}^{t}\nabla g^{(i)}(\bm{X}_{s-})\cdot{\mathrm{d}}\bm{X}^{c}_{s}+\frac{1}{2}\int_{0+}^{t}\nabla^{2}g^{(i)}(\bm{X}_{s-}):{\mathrm{d}}[\bm{X},\bm{X}]^{c}_{s}$ (4.1)
$\qquad+\sum_{0<s\leq t}\Big{\{}g^{(i)}(\bm{X}_{s})-g^{(i)}(\bm{X}_{s-})\Big{\}}\bigg{|}{\mathcal{G}}_{T}\bigg{]}.$
Define $\bm{Z}$ by
$\bm{Z}:=\Big{\{}\bm{Z}_{t}=\Big{(}\big{\langle}\mu_{t},g^{(1)}\big{\rangle},\big{\langle}\mu_{t},g^{(2)}\big{\rangle},..,\big{\langle}\mu_{t},g^{(n)}\big{\rangle}\Big{)}\Big{\}}_{t\in[0,T]},$
and by (2b) of Theorem 3.3, $\bm{Z}$ is an $\overline{\mathbb{F}}$-semimartingale. Using the localization argument (which is detailed in Step 2), let us assume that $\bm{Y}_{t}$ is bounded. Then applying Itô’s formula to $f\big{(}\bm{Z},\bm{Y}\big{)}$ yields
$f(\bm{Z}_{t},\bm{Y}_{t})-f(\bm{Z}_{0},\bm{Y}_{0})$ (4.2)
$=$ $\underbrace{\int_{0+}^{t}\nabla_{\bm{z}}f(\bm{Z}_{s-},\bm{Y}_{s-})\cdot{\mathrm{d}}\bm{Z}_{s}}_{J_{1}}+\frac{1}{2}\underbrace{\int_{0+}^{t}\nabla_{\bm{z}\bm{z}}^{2}f(\bm{Z}_{s-},\bm{Y}_{s-}):{\mathrm{d}}[\bm{Z},\bm{Z}]_{s}+\int_{0+}^{t}\nabla_{\bm{z}\bm{y}}^{2}f(\bm{Z}_{s-},\bm{Y}_{s-}):{\mathrm{d}}[\bm{Z},\bm{Y}]_{s}}_{J_{2}}$
-- $\underbrace{\sum_{0<s\leq t}\Big{\{}\nabla_{\bm{z}}f(\bm{Y}_{s-},\bm{Z}_{s-})\cdot\Delta\bm{Z}_{s}+\frac{1}{2}\nabla^{2}_{\bm{z}\bm{z}}f(\bm{Y}_{s-},\bm{Z}_{s-}):\big{(}\Delta\bm{Z}_{s}\Delta\bm{Z}_{s}^{\top}\big{)}+\nabla^{2}_{\bm{z}\bm{y}}f(\bm{Y}_{s-},\bm{Z}_{s-}):\big{(}\Delta\bm{Z}_{s}\Delta\bm{Y}_{s}^{\top}\big{)}\Big{\}}}_{J_{3}}$
$\quad+\int_{0+}^{t}\nabla_{\bm{y}}f(\bm{Z}_{s-},\bm{Y}_{s-})\cdot{\mathrm{d}}\bm{Y}^{c}_{s}+\frac{1}{2}\int_{0+}^{t}\nabla_{\bm{y}\bm{y}}^{2}f(\bm{Z}_{s-},\bm{Y}_{s-}):{\mathrm{d}}[\bm{Y},\bm{Y}]^{c}_{s}$
$\quad+\sum_{0<s\leq t}\Big{\{}f(\bm{Y}_{s},\bm{Z}_{s})-f(\bm{Y}_{s-},\bm{Z}_{s-})\Big{\}}.$
Let us compute $J_{1},J_{2},J_{3}$ separately. First, we note that $g^{(i)}$ are bounded, $\bm{Y}_{t}$ is bounded, $f\in C^{2}({\mathbb{R}}^{n+d})$, we have $\nabla_{\bm{z}}f(\bm{Z}_{s-},\bm{Y}_{s-})$, $\nabla^{2}_{\bm{z}\bm{y}}f(\bm{Z}_{s-},\bm{Y}_{s-})$ and $\nabla^{2}_{\bm{z}}f(\bm{Z}_{s-},\bm{Y}_{s-})$ are bounded and we confirm the conditions for applying (2c),(2d) of Theorem 3.3. Note that for the conditional copy $\bm{X}^{\prime}$, we have
$g^{(i)}(\bm{X}^{\prime}_{t})-g^{(i)}(\bm{X}^{\prime}_{0})=$ $\int_{0+}^{t}\nabla g^{(i)}(\bm{X}^{\prime}_{s-})\cdot{\mathrm{d}}(\bm{X}^{\prime})^{c}_{s}+\frac{1}{2}\int_{0+}^{t}\nabla^{2}g^{(i)}(\bm{X}_{s-}):{\mathrm{d}}[\bm{X}^{\prime},\bm{X}^{\prime}]^{c}_{s}$
$\;\;+\sum_{0<s\leq t}\Big{\{}g^{(i)}(\bm{X}^{\prime}_{s})-g^{(i)}(\bm{X}^{\prime}_{s-})\Big{\}},\text{ for }t\in[0,T],$
therefore, by (2c) of Theorem 3.3, ${\mathrm{d}}\bm{Z}_{s}={\mathrm{d}}\big{(}g^{(1)}(\bm{X}_{s}),\cdots,g^{(n)}(\bm{X}_{s})\big{)}^{\mathcal{G}})$, we have
$J_{1}=\overline{\mathbb{E}}\Bigg{[}\int_{0+}^{t}$ $\sum_{i=1}^{n}\partial_{z_{i}}f(\bm{Z}_{s-},\bm{Y}_{s-})\bigg{(}\nabla g^{(i)}(\bm{X}^{\prime}_{s-})\cdot{\mathrm{d}}(\bm{X}^{\prime})^{c}_{s}+\frac{1}{2}\nabla^{2}g^{(i)}(\bm{X}^{\prime}_{s-}):{\mathrm{d}}[\bm{X}^{\prime},\bm{X}^{\prime}]^{c}_{s}\bigg{)}$ (4.3)
$+\sum_{0<s\leq t}\sum_{i=1}^{n}\partial_{z_{i}}f(\bm{Z}_{s-},\bm{Y}_{s-})\Big{\{}g^{(i)}(\bm{X}^{\prime}_{s})-g^{(i)}(\bm{X}^{\prime}_{s-})\Big{\}}\bigg{|}{\mathcal{F}}\Bigg{]}$
For $J_{2}$, computing $[g^{(i)}(\bm{X}^{\prime}),g^{(j)}(\bm{X}^{\prime\prime})]_{t}$ for $1\leq i,j\leq n$, we have
[…]
$=\int_{0+}^{t}\Big{(}\nabla g^{(i)}(\bm{X}^{\prime}_{s-})^{\top}\nabla g^{(j)}(\bm{X}^{\prime\prime}_{s-})\Big{)}:{\mathrm{d}}[\bm{X}^{\prime},\bm{X}^{\prime\prime}]^{c}_{s}$
$\qquad+\sum_{0<s\leq t}\Big{\{}\big{(}g^{(i)}(\bm{X}^{\prime}_{s})-g^{(i)}(\bm{X}^{\prime}_{s-})\big{)}\big{(}g^{(j)}(\bm{X}^{\prime\prime}_{s})-g^{(j)}(\bm{X}^{\prime\prime}_{s-})\big{)}\Big{\}},$
and computing $[g^{(i)}(\bm{X}^{\prime}),Y^{(k)}]_{t}$ for $1\leq i\leq n$, $1\leq k\leq d$, we have
[…]
$=\int_{0+}^{t}\nabla g^{(i)}(\bm{X}^{\prime}_{s-})\cdot{\mathrm{d}}[\bm{X}^{\prime},Y^{(j)}]^{c}_{s}+\sum_{0<s\leq t}\Big{\{}\big{(}g^{(i)}(\bm{X}^{\prime}_{s})-g^{(i)}(\bm{X}^{\prime}_{s-})\big{)}\big{(}Y^{(k)}_{s}-Y^{(k)}_{s-}\big{)}\Big{\}}.$
Consequently, using (2d) of Theorem 3.3, we have
$J_{2}=\overline{\mathbb{E}}\Bigg{[}\int_{0+}^{t}$ $\bigg{(}\frac{1}{2}\sum_{i,j=1}^{n}\partial^{2}_{z_{i}z_{j}}f(\bm{Z}_{s-},\bm{Y}_{s-})\Big{(}\big{(}\nabla g^{(i)}(\bm{X}^{\prime}_{s-})\nabla g^{(j)}(\bm{X}^{\prime\prime}_{s-})^{\top}\big{)}:{\mathrm{d}}[\bm{X}^{\prime},\bm{X}^{\prime\prime}]^{c}_{s}\Big{)}$ (4.4)
$+\sum_{i=1}^{n}\sum_{k=1}^{d}\partial^{2}_{z_{i}y_{k}}f(\bm{Z}_{s-},\bm{Y}_{s-})\Big{(}\nabla g^{(i)}(\bm{X}^{\prime}_{s-})\cdot{\mathrm{d}}[\bm{X}^{\prime},Y^{(k)}]^{c}_{s}\Big{)}\bigg{)}$
$+\sum_{0<s\leq t}$ $\bigg{(}\frac{1}{2}\sum_{i,j=1}^{n}\partial^{2}_{z_{i}z_{j}}f(\bm{Z}_{s-},\bm{Y}_{s-})\Big{\{}\big{(}g^{(i)}(\bm{X}^{\prime}_{s})-g^{(i)}(\bm{X}^{\prime}_{s-})\big{)}\big{(}g^{(j)}(\bm{X}^{\prime}_{s})-g^{(j)}(\bm{X}^{\prime}_{s-})\big{)}\Big{\}}$
$+\sum_{i=1}^{n}\sum_{k=1}^{d}\partial^{2}_{z_{i}y_{k}}f(\bm{Z}_{s-},\bm{Y}_{s-})\Big{\{}\big{(}g^{(i)}(\bm{X}^{\prime}_{s})-g^{(i)}(\bm{X}^{\prime}_{s-})\big{)}\big{(}Y^{(k)}_{s}-Y^{(k)}_{s-}\big{)}\Big{\}}\bigg{)}\bigg{|}{\mathcal{F}}\Bigg{]}.$
Now, to deal with $J_{3}$, note that
$\Delta Z^{(i)}_{t}$ $=\big{\langle}\mu_{t},g^{(i)}\big{\rangle}-\lim_{s\nearrow t}\big{\langle}\mu_{s},g^{(i)}\big{\rangle}=\overline{\mathbb{E}}\big{[}g^{(i)}(\bm{X}_{t})\big{|}{\mathcal{G}}_{T}\big{]}-\lim_{s\nearrow t}\overline{\mathbb{E}}\big{[}g^{(i)}(\bm{X}_{s})\big{|}{\mathcal{G}}_{T}\big{]}$
$=\overline{\mathbb{E}}\big{[}g^{(i)}(\bm{X}_{t})\big{|}{\mathcal{G}}_{T}\big{]}-\overline{\mathbb{E}}\big{[}\lim_{s\nearrow t}g^{(i)}(\bm{X}_{s})\big{|}{\mathcal{G}}_{T}\big{]}=\overline{\mathbb{E}}\big{[}g^{(i)}(\bm{X}_{t})-g^{(i)}(\bm{X}_{t-})\big{|}{\mathcal{G}}_{T}\big{]}$
$=\overline{\mathbb{E}}\big{[}g^{(i)}(\bm{X}^{\prime}_{t})-g^{(i)}(\bm{X}^{\prime}_{t-})\big{|}{\mathcal{F}}\big{]},$
where the third equation holds by the conditional dominated convergence theorem. Therefore, we can rewrite $J_{3}$ as
$J_{3}=\sum_{0<s\leq t}\bigg{\{}\overline{\mathbb{E}}\bigg{[}$ $\frac{1}{2}\sum_{i,j=1}^{n}\partial^{2}_{z_{i}z_{j}}f(\bm{Y}_{s-},\bm{Z}_{s-})\big{(}g^{(i)}(\bm{X}^{\prime}_{s})-g^{(i)}(\bm{X}^{\prime}_{s-})\big{)}\big{(}g^{(j)}(\bm{X}^{\prime\prime}_{s})-g^{(j)}(\bm{X}^{\prime\prime}_{s-})\big{)}$ (4.5)
$+\sum_{i=1}^{n}\partial_{z_{i}}f(\bm{Y}_{s-},\bm{Z}_{s-})\big{(}g^{(i)}(\bm{X}^{\prime}_{s})-g^{(i)}(\bm{X}^{\prime}_{s-})\big{)}\bigg{|}{\mathcal{F}}\bigg{]}\bigg{\}}$
$+\sum_{0<s\leq t}\overline{\mathbb{E}}\bigg{[}$ $\sum_{i=1}^{n}\sum_{k=1}^{d}\partial^{2}_{z_{i}y_{k}}f(\bm{Y}_{s-},\bm{Z}_{s-})\big{(}g^{(i)}(\bm{X}^{\prime}_{s})-g^{(i)}(\bm{X}^{\prime}_{s-})\big{)}\big{(}Y^{(k)}_{s}-Y^{(k)}_{s-}\big{)}\bigg{|}{\mathcal{F}}\Bigg{]},$
since $\partial^{2}_{z_{i}z_{j}}f(\bm{Y}_{s-},\bm{Z}_{s-})$, $\partial_{z_{i}}f(\bm{Y}_{s-},\bm{Z}_{s-})$, $\partial^{2}_{z_{i}y_{k}}f(\bm{Y}_{s-},\bm{Z}_{s-})$, $Y^{(k)}_{s}$ and $Y^{(k)}_{s-}$ are ${\mathcal{F}}$-measurable random variables for $s\in(0,t]$, $1\leq i,j\leq n$ and $1\leq k\leq d$. Recall that $\Phi(\mu,\bm{y})=f\big{(}\langle\mu,g^{(1)}\rangle,..\langle\mu,g^{(n)}\rangle,\bm{y}\big{)}$, simple calculation of linear derivatives gives us a version of the derivatives by
$\nabla_{\bm{y}}\Phi(\mu,\bm{y})=\nabla_{\bm{y}}f\big{(}\langle\mu,g^{(1)}\rangle,..\langle\mu,g^{(n)},\bm{y}\rangle\big{)},\quad\nabla^{2}_{0}\Phi(\mu,\bm{y})=\nabla^{2}_{\bm{y}\bm{y}}f\big{(}\langle\mu,g^{(1)}\rangle,..\langle\mu,g^{(n)}\rangle,\bm{y}\big{)},$
$\frac{\delta\Phi}{\delta\mu}(\mu,\bm{y},\bm{x}_{1})=\sum_{i=1}^{n}\partial_{z_{i}}f\big{(}\langle\mu,g^{(1)}\rangle,..\langle\mu,g^{(n)}\rangle,\bm{y}\big{)}g^{(i)}(\bm{x}_{1}),$
$\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}(\mu,\bm{y},\bm{x}_{1},\bm{x}_{2})=\sum_{i,j=1}^{n}\partial_{z_{i}z_{j}}f\big{(}\langle\mu,g^{(1)}\rangle,..\langle\mu,g^{(n)}\rangle,\bm{y}\big{)}g^{(i)}(\bm{x}_{1})g^{(j)}(\bm{x}_{2}).$
Therefore, we can rewrite $J_{1}$ by
$\;\overline{\mathbb{E}}\bigg{[}\int_{0+}^{t}\bigg{(}\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\cdot{\mathrm{d}}(\bm{X}^{\prime})^{c}_{s}+\frac{1}{2}\nabla^{2}_{\bm{x}_{1}\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s},\bm{X}^{\prime}_{s}\big{)}:{\mathrm{d}}[\bm{X}^{\prime},\bm{X}^{\prime}]^{c}_{t}\bigg{)}$ (4.6)
$+\sum_{0<s\leq t}\bigg{(}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s}\big{)}-\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\bigg{)}\bigg{|}{\mathcal{F}}\bigg{]},$
$J_{2}$ by
$\overline{\mathbb{E}}\bigg{[}$ $\int_{0+}^{t}\bigg{(}\frac{1}{2}\nabla^{2}_{\bm{x}_{1}\bm{x}_{2}}\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{Y}_{s},\bm{X}^{\prime}_{s},\bm{X}^{\prime\prime}_{s}\big{)}:{\mathrm{d}}[\bm{X}^{\prime},\bm{X}^{\prime\prime}]^{c}_{t}+\nabla^{2}_{\bm{x}_{1}\bm{y}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}:{\mathrm{d}}[\bm{X}^{\prime},\bm{Y}]^{c}_{s}\bigg{)}$ (4.7)
$+\sum_{0<s\leq t}\bigg{(}\frac{1}{2}\bigg{(}\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s},\bm{X}^{\prime\prime}_{s}\big{)}-\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-},\bm{X}^{\prime\prime}_{s}\big{)}$
$\qquad\qquad\qquad-\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s},\bm{X}^{\prime\prime}_{s-}\big{)}+\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-},\bm{X}^{\prime\prime}_{s-}\big{)}\bigg{)}$
$\qquad\qquad\qquad+\bigg{(}\nabla_{\bm{y}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s}\big{)}-\nabla_{\bm{y}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\bigg{)}\cdot\Delta\bm{Y}_{s}\bigg{)}\bigg{|}{\mathcal{F}}\bigg{]},$
and $J_{3}$ by
$\sum_{0<s\leq t}\bigg{(}\frac{1}{2}\overline{\mathbb{E}}$ $\bigg{[}\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s},\bm{X}^{\prime\prime}_{s}\big{)}-\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-},\bm{X}^{\prime\prime}_{s}\big{)}$ (4.8)
$-\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s},\bm{X}^{\prime\prime}_{s-}\big{)}+\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-},\bm{X}^{\prime\prime}_{s-}\big{)}\bigg{|}{\mathcal{F}}\bigg{]}$
$+$ $\overline{\mathbb{E}}\bigg{[}\bigg{(}\nabla_{\bm{y}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s}\big{)}-\nabla_{\bm{y}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\bigg{)}\cdot\Delta\bm{Y}_{s}\bigg{|}{\mathcal{F}}\bigg{]}$
$+$ $\overline{\mathbb{E}}\bigg{[}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s}\big{)}-\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\bigg{|}{\mathcal{F}}\bigg{]}\bigg{)}.$
Using the argument at the end of [24, Section 3.2.1], one can cancel out redundant terms in $J_{2}$ and $J_{3}$ to get the first summation in the right hand side of (2.3). Therefore, combining (4.6), (4.7), and (4.8) and plugging them back into (4.2) complete the proof.
4.2. Step 2: localization argument.
We claim that for a sequence of ${\mathbb{F}}$ stopping time $\{\tau_{n}\}_{n\in{\mathbb{N}}}$ such that $\tau_{n}\to T$, in order to prove that (2.3) holds for $\Phi\in C^{2,2}({\mathcal{P}}_{p}({\mathbb{R}}^{d})\times{\mathbb{R}}^{d})$ with the process $(\mu,\bm{Y})=(\mu_{t\wedge\tau_{n}},\bm{Y}_{t\wedge\tau_{n}})_{t\in[0,T]}$, it suffices to prove that (2.3) holds for $\Phi$ with the truncated process $(\mu^{(\tau_{n})},\bm{Y}^{(\tau_{n})}):=(\mu_{t\wedge\tau_{n}},\bm{Y}_{t\wedge\tau_{n}})_{t\in[0,T]}$.
Indeed, suppose (2.3) holds for the truncated process, for $t<T$. Let $n\to\infty$, it is not hard to see that the left hand side of (2.3) converges and last three terms of the right hand side of (2.3) converge to the desired equation. Now, let us split the conditional expectation term of the right hand side of (2.3) into several terms. First, in order to prove that
$\overline{\mathbb{E}}\bigg{[}\int_{0+}^{t\wedge\tau_{n}}\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\cdot{\mathrm{d}}(\bm{X}^{\prime})^{c}_{s}\bigg{|}{\mathcal{F}}\bigg{]}\to\overline{\mathbb{E}}\bigg{[}\int_{0+}^{t}\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\cdot{\mathrm{d}}(\bm{X}^{\prime})^{c}_{s}\bigg{|}{\mathcal{F}}\bigg{]},$ (4.9)
almost surely as $n\to\infty$, we separate it by,
$\int_{0+}^{t\wedge\tau_{n}}$ $\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\cdot{\mathrm{d}}(\bm{X}^{\prime})^{c}_{s}=$ (4.10)
$\int_{0+}^{t\wedge\tau_{n}}\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\cdot{\mathrm{d}}\bm{X}^{\prime}_{s}-\sum_{0<s\leq t\wedge\tau_{n}}\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\cdot\Delta\bm{X}^{\prime}_{s}$
Note that the first term is uniformly bounded by
$\sup_{0\leq s\leq T}\bigg{|}\int_{0+}^{s}\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{u-},\bm{Y}_{u-},\bm{X}^{\prime}_{u-}\big{)}\cdot{\mathrm{d}}\bm{X}^{\prime}_{u}\bigg{|},$
and
$\overline{\mathbb{E}}\bigg{[}\sup_{0\leq s\leq T}\bigg{|}\int_{0+}^{s}$ $\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{u-},\bm{Y}_{u-},\bm{X}^{\prime}_{u-}\big{)}\cdot{\mathrm{d}}\bm{X}^{\prime}_{u}\bigg{|}\bigg{]}=\bigg{\|}\int_{0+}^{\cdot}\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{u-},\bm{Y}_{u-},\bm{X}^{\prime}_{u-}\big{)}\cdot{\mathrm{d}}\bm{X}^{\prime}_{u}\bigg{\|}_{{\mathcal{S}}_{1}}$
$\leq c_{1}\bigg{\|}\int_{0+}^{\cdot}\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{u-},\bm{Y}_{u-},\bm{X}^{\prime}_{u-}\big{)}\cdot{\mathrm{d}}\bm{X}^{\prime}_{u}\bigg{\|}_{{\mathcal{H}}_{1}}$
$\leq c_{1}\bigg{\|}\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{\cdot},\bm{Y}_{\cdot},\bm{X}^{\prime}_{\cdot}\big{)}\bigg{\|}_{{\mathcal{S}}_{p^{\prime}}}\big{\|}\bm{X}^{\prime}\big{\|}_{{\mathcal{H}}_{p}}\leq c_{1}C\big{\|}|\bm{X}|^{p-1}+|\bm{Y}|^{p-1}+1\big{\|}_{{\mathcal{S}}_{p^{\prime}}}\big{\|}\bm{X}\big{\|}_{{\mathcal{H}}_{p}}$
$\leq c_{p}c_{1}C\big{(}\big{\|}\bm{X}\big{\|}^{p}_{{\mathcal{H}}_{p}}+\big{\|}\bm{Y}\big{\|}^{p}_{{\mathcal{H}}_{p}}+1\big{)}\big{\|}\bm{X}\big{\|}_{{\mathcal{H}}_{p}}<\infty,$
where $p^{\prime}=\frac{p}{p-1}$, $c_{1},c_{p}$ are constants in Proposition 2.2 and $C$ is a generic constant that may vary line by line. The first inequality is due to Proposition 2.2, the second inequality holds because of Proposition 2.3, the third inequality holds by $\Phi\in C^{2,2}({\mathcal{P}}_{p}({\mathbb{R}}^{d})\times{\mathbb{R}}^{d})$, and the last one holds because $\frac{1}{p}+\frac{1}{p^{\prime}}=1$. Meanwhile, the second term of the right hand side of (4.10) is bounded by
$\sup_{0\leq s\leq T}\bigg{|}\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},$ $\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\bigg{|}\cdot\sum_{0<s\leq T}|\Delta\bm{X}^{\prime}_{s}|\leq C\sup_{0\leq s\leq T}\big{(}1+|\bm{X}_{s}|^{p-1}+|\bm{Y}_{s}|^{p-1}\big{)}\cdot\sum_{0<s\leq T}|\Delta\bm{X}^{\prime}_{s}|$ (4.11)
$\leq C\bigg{(}1+\sup_{0\leq s\leq T}|\bm{X}_{s}|^{p}+\sup_{0\leq s\leq T}|\bm{Y}_{s}|^{p}+\bigg{(}\sum_{0<s\leq T}|\Delta\bm{X}^{\prime}_{s}|\bigg{)}^{p}\bigg{)},$
where $C$ is a generic constant that may vary line by line, and the right hand side of the above inequality is integrable due to the fact that $\|\bm{X}\|_{{\mathcal{S}}_{p}}\leq c_{p}\|\bm{X}\|_{{\mathcal{H}}_{p}}<\infty$, $\|\bm{Y}\|_{{\mathcal{S}}_{p}}\leq c_{p}\|\bm{Y}\|_{{\mathcal{H}}_{p}}<\infty$ and the assumption on $\bm{X}$. Therefore, using the conditional dominated theorem, we conclude (4.9). Next, to prove that
$\overline{\mathbb{E}}\bigg{[}\int_{0+}^{t\wedge\tau_{n}}$ $\nabla^{2}_{\bm{x}_{1}\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}:{\mathrm{d}}[\bm{X}^{\prime},\bm{X}^{\prime}]^{c}_{s}\bigg{|}{\mathcal{F}}\bigg{]}$ (4.12)
$\to\overline{\mathbb{E}}\bigg{[}\int_{0+}^{t}\nabla^{2}_{\bm{x}_{1}\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}:{\mathrm{d}}[\bm{X}^{\prime},\bm{X}^{\prime}]^{c}_{s}\bigg{|}{\mathcal{F}}\bigg{]},$
almost surely as $n\to\infty$, we note that
$\int_{0+}^{t\wedge\tau_{n}}\nabla^{2}_{\bm{x}_{1}\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}:{\mathrm{d}}[\bm{X}^{\prime},\bm{X}^{\prime}]^{c}_{s}$
is bounded by
$\sup_{0\leq s\leq T}\bigg{|}\int_{0+}^{s}\nabla^{2}_{\bm{x}_{1}\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{u-},\bm{Y}_{u-},\bm{X}^{\prime}_{u-}\big{)}:{\mathrm{d}}[\bm{X}^{\prime},\bm{X}^{\prime}]^{c}_{u}\bigg{|}\leq\sup_{0\leq s\leq T}\bigg{|}\nabla^{2}_{\bm{x}_{1}\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\bigg{|}\cdot[\bm{X}^{\prime},\bm{X}^{\prime}]^{c}_{T}$
and
$\overline{\mathbb{E}}\bigg{[}\sup_{0\leq s\leq T}\bigg{|}\nabla^{2}_{\bm{x}_{1}\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}$ $\bigg{|}\cdot[\bm{X}^{\prime},\bm{X}^{\prime}]^{c}_{T}\bigg{]}\leq\bigg{\|}\nabla^{2}_{\bm{x}_{1}\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{\cdot},\bm{Y}_{\cdot},\bm{X}^{\prime}_{\cdot}\big{)}\bigg{\|}_{{\mathcal{S}}_{p^{\prime\prime}}}\big{\|}[\bm{X}^{\prime},\bm{X}^{\prime}]^{c}_{T}\big{\|}_{L_{p/2}}$
$\leq C\Big{\|}1+|\bm{X}^{\prime}|^{p-2}+|\bm{Y}|^{p-2}\Big{\|}_{{\mathcal{S}}_{p^{\prime\prime}}}\big{\|}[\bm{X}^{\prime},\bm{X}^{\prime}]^{c}_{T}\big{\|}_{L_{p/2}}$
$\leq c_{p^{\prime\prime}}C\Big{(}1+\|\bm{X}^{\prime}\|^{p}_{{\mathcal{H}}_{p}}+\|\bm{Y}\|^{p}_{{\mathcal{H}}_{p}}\Big{)}\big{\|}[\bm{X}^{\prime},\bm{X}^{\prime}]^{c}_{T}\big{\|}_{L_{p/2}}<\infty,$
where $c_{p^{\prime\prime}}$ is the constant in Proposition 2.2, $C$ is a generic constant that may vary from line to line and $p^{\prime\prime}=\frac{p-2}{2p}$. Thus, (4.12) holds due to the conditional dominated theorem, we conclude . Similar arguments holds also for the convergence of $\int_{0+}^{t\wedge\tau_{n}}\nabla^{2}_{\bm{x}_{1}\bm{x}_{2}}\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-},\bm{X}^{\prime\prime}_{s-}\big{)}:{\mathrm{d}}[\bm{X}^{\prime},\bm{X}^{\prime\prime}]_{s}$ and $\int_{0+}^{t\wedge\tau_{n}}\nabla^{2}_{\bm{x}_{1}\bm{y}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}:{\mathrm{d}}[\bm{X}^{\prime},\bm{Y}]_{s}$.
Next, let us deal with the convergence of
$\sum_{0<s\leq t\wedge\tau_{n}}\bigg{(}$ $\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s},\bm{X}^{\prime\prime}_{s}\big{)}-\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-},\bm{X}^{\prime\prime}_{s}\big{)}$ (4.13)
$-\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s},\bm{X}^{\prime\prime}_{s-}\big{)}+\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-},\bm{X}^{\prime\prime}_{s-}\big{)}\bigg{)}\mathbbm{1}_{\{\mu_{s}=\mu_{s-}\}}.$
With simple application of the fundamental theorem of calculus for the linear derivative, we have
$\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s},\bm{X}^{\prime\prime}_{s}\big{)}-\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-},\bm{X}^{\prime\prime}_{s}\big{)}$
$-\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s},\bm{X}^{\prime\prime}_{s-}\big{)}+\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-},\bm{X}^{\prime\prime}_{s-}\big{)}$
$=\int_{0}^{1}$ $\int_{0}^{1}\nabla^{2}_{\bm{x}_{1}\bm{x}_{2}}\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}+\lambda(\bm{X}^{\prime}_{s}-\bm{X}^{\prime}_{s-}),\bm{X}^{\prime\prime}_{s-}+\gamma(\bm{X}^{\prime\prime}_{s}-\bm{X}^{\prime\prime}_{s-})\big{)}:\Big{(}\big{(}\Delta\bm{X}^{\prime}_{s}\big{)}^{\top}\Delta\bm{X}^{\prime\prime}_{s}\Big{)}{\mathrm{d}}\lambda{\mathrm{d}}\gamma,$
and the right hand side of the above equation is bounded by
$C\cdot|\Delta\bm{X}^{\prime}_{s}|\cdot|\Delta\bm{X}^{\prime\prime}_{s}|\cdot\Big{(}1+|\bm{X}^{\prime}_{s}|^{p-2}+|\bm{X}^{\prime\prime}_{s}|^{p-2}\Big{)}.$
Therefore, (4.13) is bounded by
$\sup_{0\leq s\leq T}\Big{(}1+|\bm{X}^{\prime}_{s}|^{p-2}+|\bm{X}^{\prime\prime}_{s}|^{p-2}\Big{)}\big{(}\sum_{0<s\leq T}|\Delta\bm{X}^{\prime}_{s}|\cdot|\Delta\bm{X}^{\prime\prime}_{s}|\big{)}$
$\leq C\bigg{(}\sup_{0\leq t\leq T}\Big{(}1+|\bm{X}^{\prime}_{t}|^{p-2}+|\bm{X}^{\prime\prime}_{t}|^{p-2}\Big{)}^{\frac{p}{p-2}}+\Big{(}\sum_{0<s\leq T}|\Delta\bm{X}^{\prime}_{s}|^{2}+\sum_{0<s\leq T}|\Delta\bm{X}^{\prime\prime}_{s}|^{2}\Big{)}^{\frac{p}{2}}\bigg{)}$
$\leq C\bigg{(}\sup_{0\leq t\leq T}\Big{(}1+|\bm{X}^{\prime}_{t}|^{p}+|\bm{X}^{\prime\prime}_{t}|^{p}\Big{)}+\sum_{0<s\leq T}|\Delta\bm{X}^{\prime}_{s}|^{p}+\sum_{0<s\leq T}|\Delta\bm{X}^{\prime\prime}_{s}|^{p}\bigg{)},$
where $C$ is a generic constant that may vary from line by line. Therefore, the right hand side of the above inequality is integrable and the convergence of the conditional expectation of (4.13) is ensured by conditional dominated convergence theorem. Similar arguments can be applied to deduce the convergence of the conditional expectation of $\big{(}\nabla_{\bm{y}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s}\big{)}-\nabla_{\bm{y}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\big{)}\cdot\Delta\bm{Y}_{s}$ and $\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s}\big{)}-\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}$, s and thus we conclude that (2.3) holds for the original process $(\mu,\bm{Y})$ for $0\leq t<T$.
For the case of $t=T$, taking $n\to\infty$, left hand side equals $\Phi(\mu_{T-},\bm{Y}_{T-})-\Phi(\mu_{0},\bm{Y}_{0})$ while all terms at the right hand sides is evaluated at $T$ except the last term being
$\sum_{0<s<t}\Big{(}\Phi(\mu_{s},\bm{Y}_{s})-\Phi(\mu_{s-},\bm{Y}_{s-})\Big{)},$
adding $\Phi(\mu_{T},\bm{Y}_{T})-\Phi(\mu_{T-},\bm{Y}_{T-})$ to both sides, we finish the proof of the claim.
4.3. Step 3: cylindrical functions are dense in $C^{2,2}({\mathbb{R}}^{d}\times\mathcal{P}_{p}({\mathbb{R}}^{d}))$.
The following result slightly generalizes that in [16, Theorem 4.4, Theorem 4.10, Section 4] as it includes the convergences of the derivatives of the linear derivative.
Proposition 4.2.
Let $\Phi\in C^{2,2}({\mathcal{P}}_{p}({\mathbb{R}}^{d})\times{\mathbb{R}}^{d})$, then there exist a sequence of $C^{2,2}$ cylinder functions $\{f_{n}\}_{n\in{\mathbb{N}}^{+}}$ such that one has the point-wise convergence
$\bigg{(}f_{n},\nabla_{\bm{y}}f_{n}$ $,\nabla_{\bm{y}}^{2}f_{n},\frac{\delta f_{n}}{\delta\mu},\nabla_{\bm{x}_{1}}\frac{\delta f_{n}}{\delta\mu},\nabla^{2}_{\bm{x}_{1}}\frac{\delta f_{n}}{\delta\mu},\nabla_{\bm{x}_{1}\bm{y}}\frac{\delta f_{n}}{\delta\mu},\nabla_{\bm{x}_{1}\bm{x}_{2}}\frac{\delta^{2}f_{n}}{(\delta\mu)^{2}}\bigg{)}$ (4.14)
$\to\bigg{(}\Phi,\nabla_{\bm{y}}\Phi,\nabla_{\bm{y}}^{2}\Phi,\frac{\delta\Phi}{\delta\mu},\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu},\nabla^{2}_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu},\nabla_{\bm{x}_{1}\bm{y}}\frac{\delta\Phi}{\delta\mu},\nabla_{\bm{x}_{1}\bm{x}_{2}}\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\bigg{)}$
as $n\to\infty$, for all $\bm{x}_{1},\bm{x}_{2},\bm{y},\in{\mathbb{R}}^{d}$, $\mu\in{\mathcal{P}}_{p}({\mathbb{R}}^{d})$. Moreover, there exists a constant $c>0$ such that
$|\nabla_{\bm{y}}f_{n}|(\mu,\bm{y})\leq c(1+|\bm{y}|^{p-1}),\qquad|\nabla^{2}_{\bm{y}}f_{n}|(\mu,\bm{y})$ $\leq c(1+|\bm{y}|^{p-2}),$ (4.15)
$\bigg{|}\frac{\delta f_{n}}{\delta\mu}\bigg{|}\leq c\big{(}1+|\bm{x}_{1}|^{p}+|\bm{y}|^{p}\big{)},\qquad\bigg{|}\nabla_{\bm{x}_{1}}\frac{\delta f_{n}}{\delta\mu}\bigg{|}$ $\leq c\big{(}1+|\bm{x}_{1}|^{p-1}+|\bm{y}|^{p-1}\big{)},$
$\bigg{|}\nabla^{2}_{\bm{x}_{1}}\frac{\delta f_{n}}{\delta\mu}\bigg{|}\leq c_{K}\big{(}1+|\bm{x}_{1}|^{p-2}+|\bm{y}|^{p-2}\big{)},\qquad\bigg{|}\nabla^{2}_{\bm{x}_{1}\bm{y}}\frac{\delta f_{n}}{\delta\mu}\bigg{|}$ $\leq c\big{(}1+|\bm{x}_{1}|^{{p-2}}+|\bm{y}|^{p-2}\big{)},$
$\bigg{|}\nabla^{2}_{\bm{x}_{1}\bm{x}_{2}}\frac{\delta^{2}f_{n}}{(\delta\mu)^{2}}\bigg{|}$ $\leq c\big{(}1+|\bm{x}_{1}|^{p-2}+|\bm{x}_{2}|^{p-2}+|\bm{y}|^{p-2}\big{)},$
for all $\bm{x}_{1},\bm{x}_{2},\bm{y}\in{\mathbb{R}}^{d}$, $\mu\in{\mathcal{P}}_{p}({\mathbb{R}}^{d})$.
The proof is similar to that of [16, Section 4]. The main step of their proof consists of constructing an operator $T_{n}:C_{b}({\mathbb{R}}^{d})\to C_{b}({\mathbb{R}}^{d})$ and the corresponding adjoint $T_{n}^{*}:{\mathcal{P}}_{p}({\mathbb{R}}^{d})\to{\mathcal{P}}_{p}({\mathbb{R}}^{d})$ corresponding square $T_{n}^{\otimes 2}:C_{b}({\mathbb{R}}^{2d})\to C_{b}({\mathbb{R}}^{2d})$ such that $T_{n}\varphi_{n}\to\varphi$, when $\varphi_{n}\to\varphi$ and $T_{n}^{*}\mu\to\mu$. Define $f_{n}(\mu,\bm{y}):=\Phi(T_{n}^{*}\mu,\bm{y})$, then the key identities to be established are
$\frac{\delta f_{n}}{\delta\mu}(\mu,\bm{y},\bm{x}_{1})=T_{n}\Big{(}\frac{\delta\Phi}{\delta\mu}(T_{n}^{*}\mu,\bm{y},\cdot)\Big{)}(\bm{x}_{1}),\quad\frac{\delta^{2}f_{n}}{(\delta\mu)^{2}}(\mu,\bm{y},\bm{x}_{1},\bm{x}_{2})=T_{n}^{\otimes 2}\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}(T_{n}^{*}\mu,\bm{y},\cdot,\cdot)(\bm{x}_{1},\bm{x}_{2}).$
for all $\bm{x}_{1},\bm{x}_{2},\bm{y}\in{\mathbb{R}}^{d}$ and $\mu\in{\mathcal{P}}_{p}({\mathbb{R}}^{d})$. To this end, it suffices to note that the construction of $T_{n}$ ensures
$T_{n}\nabla_{\bm{x}_{1}}\varphi(\bm{x}_{1})=\nabla_{\bm{x}_{1}}T_{n}\varphi(\bm{x}_{1}),\quad T_{n}^{\otimes 2}\nabla^{2}_{\bm{x}_{1}\bm{x}_{2}}\psi(\bm{x}_{1},\bm{x}_{2})=\nabla^{2}_{\bm{x}_{1}\bm{x}_{2}}T_{n}\psi(\bm{x}_{1},\bm{x}_{2}),$
for all $\varphi\in C^{1}_{b}({\mathbb{R}}^{d})$ and $\psi\in C^{2}_{b}({\mathbb{R}}^{2d}).$
4.4. Final step
Define $\tau_{k}:=\inf\{t\in[0,T]:|\bm{Y}_{t}|>k\}$, then by Step 1, we know that Itô’s formula (2.3) holds for the truncated process $(\mu^{(\tau_{n})},\bm{Y}^{(\tau_{n})}):=(\mu_{t\wedge\tau_{n}},\bm{Y}_{t\wedge\tau_{n}})_{t\in[0,T]}$ with the sequences of cylindrical functions $\{f_{n}\}_{n\in{\mathbb{N}}^{+}}$. By Step 2, the localization argument shows that Itô’s formula (2.3) holds for the sequences of cylindrical functions $\{f_{n}\}_{n\in{\mathbb{N}}^{+}}$.
In order to pass the limit from $f_{n}$ to $\Phi$, let us consider the first term in the right hand side of (2.3) and the remaining terms follow similarly. For the first term, we need to prove that there exists a subsequence $\{n_{k}\}_{k\in{\mathbb{N}}^{+}}$ such that
${\mathbb{E}}\bigg{[}\int_{0+}^{t}\nabla_{\bm{x}_{1}}\frac{\delta f_{n_{k}}}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\cdot{\mathrm{d}}(\bm{X}^{\prime})^{c}_{s}\Big{|}{\mathcal{F}}\bigg{]}\to{\mathbb{E}}\bigg{[}\int_{0+}^{t}\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\cdot{\mathrm{d}}(\bm{X}^{\prime})^{c}_{s}\Big{|}{\mathcal{F}}\bigg{]},$ (4.16)
as $k\to\infty.$ To this end, we first prove that
$\int_{0+}^{t}\nabla_{\bm{x}_{1}}\frac{\delta f_{n}}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\cdot{\mathrm{d}}(\bm{X}^{\prime})^{c}_{s}\overset{L^{1}}{\to}\int_{0+}^{t}\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\cdot{\mathrm{d}}(\bm{X}^{\prime})^{c}_{s},$ (4.17)
as $n\to\infty$. By Proposition 4.2, we have
$\bigg{|}|\nabla_{\bm{x}_{1}}\frac{\delta f_{n}}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\bigg{|}\leq c\big{(}1+|\bm{X}_{s-}|^{p-1}+|\bm{Y}_{s-}|^{p-1}\big{)},$
where the ${\mathcal{S}}_{p^{\prime}}$ norm of the right hand side is finite with $p^{\prime}=\frac{p}{p-1}$. Now (4.17) follows from $\bm{X}\in{\mathcal{H}}^{p}$ and the dominated convergence theorem of stochastic integral ([37, Page 273, Lemma]). Next, since
${\mathbb{E}}\bigg{[}\Big{|}$ ${\mathbb{E}}\Big{[}\int_{0+}^{t}\nabla_{\bm{x}_{1}}\frac{\delta f_{n}}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\cdot{\mathrm{d}}(\bm{X}^{\prime})^{c}_{s}\Big{|}{\mathcal{F}}\Big{]}-{\mathbb{E}}\Big{[}\int_{0+}^{t}\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\cdot{\mathrm{d}}(\bm{X}^{\prime})^{c}_{s}\Big{|}{\mathcal{F}}\Big{]}\Big{|}\bigg{]}$
$\leq{\mathbb{E}}\bigg{[}\Big{|}\int_{0+}^{t}\nabla_{\bm{x}_{1}}\frac{\delta f_{n}}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\cdot{\mathrm{d}}(\bm{X}^{\prime})^{c}_{s}-\int_{0+}^{t}\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\cdot{\mathrm{d}}(\bm{X}^{\prime})^{c}_{s}\Big{|}\bigg{]}\to 0,$
and $L_{1}$ convergences implies convergence in subsequence, (4.16) holds. The remaining terms follow similarly due to the uniform bound (4.15), and thus the proof of Theorem 2.7 is complete.
5. Proof of Section 3
5.1. Construction of the conditional independent copy
This subsection is devoted to the proof of the precise construction of conditional independent copy as in Theorem 3.1. For ease of exposition, we will show how to enlarge the probability space with two independent copies, and any finite number of copies can be easily adapted.
Recall the definition in Section 3, let us consider the probability space $\overline{\Omega}$ and the $\sigma$-algrebra $\overline{\mathcal{F}}$ defined by
$\overline{\Omega}=\Omega^{3}=\{(\omega_{0},\omega_{1},\omega_{2})|\omega_{0},\omega_{1},\omega_{2}\in\Omega\},\qquad\overline{\mathcal{F}}=\sigma\{A_{0}\times A_{1}\times A_{2}:A_{0},A_{1},A_{2}\in{\mathcal{F}}\}$ (5.1)
and probability measure $\overline{\mathbb{P}}$ is defined as follows: for $A_{0},A_{1},A_{2}\in{\mathcal{F}}$,
¯$\overline{\mathbb{P}}\big{(}A_{0}\times A_{1}\times A_{2}\big{)}:={\mathbb{E}}\Big{[}\mathbbm{1}_{A_{0}}{\mathbb{P}}(A_{1}|{\mathcal{G}}){\mathbb{P}}(A_{2}|{\mathcal{G}})\Big{]},$ (5.2)
Since the set $\{A_{0}\times A_{1}\times A_{2}:A_{0},A_{1},A_{2}\in{\mathcal{F}}\}$ is a $\pi$-system, the measure on $(\overline{\Omega},\overline{\mathcal{F}},\overline{\mathbb{P}})$ is properly defined. Next, use the notation ~$\widetilde{\mathcal{F}},\widetilde{\mathcal{G}}$ for the nature extension into this enlarged probability space as
~:=$\widetilde{\mathcal{F}}:=\{A\times\Omega\times\Omega:A\in{\mathcal{F}}\},\quad\widetilde{\mathcal{G}}:=\{A\times\Omega\times\Omega:A\in{\mathcal{G}}\},$ (5.3)
and define ${\mathcal{F}}^{\prime}$, ${\mathcal{F}}^{\prime\prime}$ as
:=${\mathcal{F}}^{\prime}:=\{\Omega\times A\times\Omega:A\in{\mathcal{F}}\},\quad{\mathcal{F}}^{\prime\prime}:=\{\Omega\times\Omega\times A:A\in{\mathcal{F}}\}.$ (5.4)
With these constructions, we have the following propositions.
Proposition 5.1.
Given $(\overline{\Omega},\overline{\mathcal{F}},\overline{\mathbb{P}})$ defined by (5.1) and (5.2) and the sub-algebras ~$\widetilde{\mathcal{F}},\widetilde{\mathcal{G}}$, ${\mathcal{F}}^{\prime}$, and ${\mathcal{F}}^{\prime\prime}$ defined by (5.3) and (5.4). Suppose $A,B\in{\mathcal{F}}$, let $\widetilde{A}=A\times\Omega\times\Omega$, $A^{\prime}=\Omega\times A\times\Omega$, $A^{\prime\prime}=\Omega\times\Omega\times A$, $B^{\prime}=\Omega\times B\times\Omega$ and $C^{\prime\prime}=\Omega\times\Omega\times C$. Then
  1. (1)
    For $\overline{\omega}=(\omega_{0},\omega_{1},\omega_{2})\in\overline{\Omega}$
    ${\mathbb{P}}(A|\mathcal{G})(\omega_{0})=\overline{\mathbb{P}}(\widetilde{A}|\widetilde{\mathcal{G}})(\omega_{0},\omega_{1},\omega_{2}),\text{ a.s. under }\overline{\mathbb{P}}.$
  2. (2)
    Given $\widetilde{\mathcal{G}}$, then $\widetilde{A},A^{\prime},$ and $A^{\prime\prime}$ have the same conditional law. That is,
    ¯$\overline{\mathbb{P}}(\widetilde{A}|\widetilde{\mathcal{G}})=\overline{\mathbb{P}}(A^{\prime}|\widetilde{\mathcal{G}})=\overline{\mathbb{P}}(A^{\prime\prime}|\widetilde{\mathcal{G}}),\text{ a.s. under }\overline{\mathbb{P}}.$
    Moreover, this conditional probability ¯$\overline{\mathbb{P}}(\widetilde{A}|\widetilde{\mathcal{G}})$ coincides with the conditional probability of $A^{\prime}$ and $A^{\prime\prime}$ given $\widetilde{\mathcal{F}}$, that is
    ¯$\overline{\mathbb{P}}(\widetilde{A}|\widetilde{\mathcal{G}})=\overline{\mathbb{P}}(A^{\prime}|\widetilde{\mathcal{F}})=\overline{\mathbb{P}}(A^{\prime\prime}|\widetilde{\mathcal{F}}),\text{ a.s. under }\overline{\mathbb{P}}.$
  3. (3)
    Given $\widetilde{\mathcal{G}}$, then $\widetilde{\mathcal{F}},{\mathcal{F}}^{\prime},$ and ${\mathcal{F}}^{\prime\prime}$ are conditionally independent. That is,
    ¯$\overline{\mathbb{P}}\big{(}\widetilde{A}\cap B^{\prime}\cap C^{\prime\prime}|\widetilde{\mathcal{G}}\big{)}=\overline{\mathbb{P}}\big{(}\widetilde{A}|\widetilde{\mathcal{G}}\big{)}\overline{\mathbb{P}}\big{(}B^{\prime}|\widetilde{\mathcal{G}}\big{)}\overline{\mathbb{P}}\big{(}C^{\prime\prime}|\widetilde{\mathcal{G}}\big{)},\text{ a.s. under }\overline{\mathbb{P}}.$
    Moreover, given $\widetilde{\mathcal{F}}$, then ${\mathcal{F}}^{\prime}$ and ${\mathcal{F}}^{\prime\prime}$ are conditionally independent, That is
    ¯$\overline{\mathbb{P}}\big{(}B^{\prime}\cap C^{\prime\prime}|\widetilde{\mathcal{F}}\big{)}=\overline{\mathbb{P}}\big{(}B^{\prime}|\widetilde{\mathcal{F}}\big{)}\overline{\mathbb{P}}\big{(}C^{\prime\prime}|\widetilde{\mathcal{F}}\big{)},\text{ a.s. under }\overline{\mathbb{P}}.$
Proof.
To prove (1), note that ${\mathbb{P}}[A|{\mathcal{G}}]$ is ${\mathcal{G}}$ measurable in $(\Omega,{\mathcal{F}},{\mathbb{F}},{\mathbb{P}})$, then ${\mathbb{P}}[A|{\mathcal{G}}](\omega_{1})$ as a random variable in $(\overline{\Omega},\overline{\mathcal{F}},\overline{\mathbb{F}},\overline{\mathbb{P}})$ is $\widetilde{\mathcal{G}}$ measurable . For any $\widetilde{G}\in\widetilde{\mathcal{G}}$, there exists a $G\in{\mathcal{G}}$ such that $\widetilde{G}=G\times\Omega\times\Omega\in\widetilde{\mathcal{G}}$. Recall the definition of $\overline{{\mathbb{P}}}$ at (5.2) and the property of conditional expectation, we have
¯$\overline{\mathbb{P}}\big{[}\widetilde{A}\cap\widetilde{G}\big{]}=\overline{\mathbb{P}}\big{[}(A\cap G)\times\Omega\times\Omega\big{]}={\mathbb{P}}[A\cap G]={\mathbb{E}}\Big{[}{\mathbb{P}}[A|{\mathcal{G}}]\mathbbm{1}_{G}\Big{]}=\int_{\Omega}{\mathbb{P}}[A|{\mathcal{G}}](\omega)\mathbbm{1}_{G}(\omega){\mathbb{P}}({\mathrm{d}}\omega).$
Again by (5.2), we obtain ¯$\overline{\mathbb{P}}({\mathrm{d}}\omega_{1}\times\Omega\times\Omega)={\mathbb{P}}({\mathrm{d}}\omega_{1})$, therefore,
¯$\overline{\mathbb{P}}\big{[}\widetilde{A}\cap\widetilde{G}\big{]}$ $=\int_{\Omega^{3}}{\mathbb{P}}[A|{\mathcal{G}}](\omega_{0})\mathbbm{1}_{B}(\omega_{0})\overline{\mathbb{P}}({\mathrm{d}}\omega_{0}{\mathrm{d}}\omega_{1}{\mathrm{d}}\omega_{2})$ (5.5)
$=\int_{\Omega^{3}}{\mathbb{P}}[A|{\mathcal{G}}](\omega_{0})\mathbbm{1}_{\widetilde{G}}(\omega_{0},\omega_{1},\omega_{2})\overline{\mathbb{P}}({\mathrm{d}}\omega_{0}{\mathrm{d}}\omega_{1}{\mathrm{d}}\omega_{2})=\overline{\mathbb{E}}\Big{[}{\mathbb{P}}[A|{\mathcal{G}}]\mathbbm{1}_{\widetilde{G}}\Big{]},$
which leads to ${\mathbb{P}}[A|{\mathcal{G}}](\omega_{0})=\overline{\mathbb{P}}[\widetilde{A}|\widetilde{\mathcal{G}}](\omega_{0},\omega_{1},\omega_{2})$ a.s. under $\overline{P}$.
To prove (2), for any $\widetilde{G}\in\widetilde{\mathcal{G}}$, there exists a $G\in{\mathcal{G}}$ such that $\widetilde{G}=G\times\Omega\times\Omega\in\widetilde{\mathcal{G}}$. Then,
¯$\overline{\mathbb{P}}\big{[}A^{\prime}\cap\widetilde{G}\big{]}=\overline{\mathbb{P}}\big{[}G\times A\times\Omega\big{]}={\mathbb{E}}\big{[}\mathbbm{1}_{G}{\mathbb{P}}(A|{\mathcal{G}})\big{]}=\overline{\mathbb{E}}\Big{[}{\mathbb{P}}[A|{\mathcal{G}}]\mathbbm{1}_{\widetilde{G}}\Big{]}=\overline{\mathbb{E}}\Big{[}\overline{\mathbb{P}}[\widetilde{A}|\widetilde{\mathcal{G}}]\mathbbm{1}_{\widetilde{G}}\Big{]},$
where the second equation holds because of (5.2), the third equation holds due to the same calculation as (5.5) and the last equation holds by (1). This leads to ¯$\overline{\mathbb{P}}(A^{\prime}|\widetilde{\mathcal{G}})=\overline{\mathbb{P}}(\widetilde{A}|\widetilde{\mathcal{G}})$ and similarly, we have ¯$\overline{\mathbb{P}}(A^{\prime\prime}|\widetilde{\mathcal{G}})=\overline{\mathbb{P}}(\widetilde{A}|\widetilde{\mathcal{G}})$.
As for the second statement of (2), note that ¯$\overline{\mathbb{P}}(A^{\prime}|\widetilde{\mathcal{G}})$ is $\widetilde{\mathcal{G}}$ measurable, thus $\widetilde{\mathcal{F}}$ measurable. For any $\widetilde{F}\in\widetilde{\mathcal{F}}$, there exists a $F\in{\mathcal{F}}$ such that $\widetilde{F}=F\times\Omega\times\Omega\in\widetilde{\mathcal{F}}$. Then,
¯$\overline{\mathbb{P}}(A^{\prime}\cap\widetilde{F})=\overline{\mathbb{P}}\big{[}F\times A\times\Omega\big{]}={\mathbb{E}}\big{[}\mathbbm{1}_{F}{\mathbb{P}}(A|{\mathcal{G}})\big{]}=\overline{\mathbb{E}}\Big{[}{\mathbb{P}}[A|{\mathcal{G}}]\mathbbm{1}_{\widetilde{F}}\Big{]}=\overline{\mathbb{E}}\Big{[}\overline{\mathbb{P}}[\widetilde{A}|\widetilde{\mathcal{G}}]\mathbbm{1}_{\widetilde{F}}\Big{]},$
where the second equation holds because of (5.2), the third equation holds due to the same calculation as (5.5) and the last equation holds by (1). This leads to ¯$\overline{\mathbb{P}}(A^{\prime}|\widetilde{\mathcal{F}})=\overline{\mathbb{P}}(\widetilde{A}|\widetilde{\mathcal{G}})$ and similarly, we have ¯$\overline{\mathbb{P}}(A^{\prime\prime}|\widetilde{\mathcal{F}})=\overline{\mathbb{P}}(\widetilde{A}|\widetilde{\mathcal{G}})$.
To prove (3), for any $\widetilde{G}\in\widetilde{\mathcal{G}}$, there exists a $G\in{\mathcal{G}}$ such that $\widetilde{G}=G\times\Omega\times\Omega\in\widetilde{\mathcal{G}}$. Then,
¯$\overline{{\mathbb{P}}}(\widetilde{A}\cap B^{\prime}\cap C^{\prime\prime}\cap\widetilde{G})$ =¯$=\overline{{\mathbb{P}}}((A\cap G)\times B\times C)={\mathbb{E}}\Big{[}\mathbbm{1}_{A\cap G}{\mathbb{P}}(B|{\mathcal{G}}){\mathbb{P}}(C|{\mathcal{G}})\Big{]}$
$={\mathbb{E}}\Big{[}{\mathbb{E}}[\mathbbm{1}_{A\cap G}|{\mathcal{G}}]{\mathbb{P}}(B|{\mathcal{G}}){\mathbb{P}}(C|{\mathcal{G}})\Big{]}={\mathbb{E}}\Big{[}\mathbbm{1}_{G}{\mathbb{P}}(A|{\mathcal{G}}){\mathbb{P}}(B|{\mathcal{G}}){\mathbb{P}}(C|{\mathcal{G}})\Big{]}$
Using the same calculation as in (5.5) and note that ${\mathbb{P}}(A|{\mathcal{G}})=\overline{\mathbb{P}}(\widetilde{A}|\widetilde{\mathcal{G}})$, ${\mathbb{P}}(B|{\mathcal{G}})=\overline{\mathbb{P}}(B^{\prime}|\widetilde{\mathcal{G}})$ and ${\mathbb{P}}(C|{\mathcal{G}})=\overline{\mathbb{P}}(C^{\prime\prime}|\widetilde{\mathcal{G}})$, we have
¯$\overline{{\mathbb{P}}}(\widetilde{A}\cap B^{\prime}\cap C^{\prime\prime}\cap\widetilde{G})$ $=\overline{\mathbb{E}}\Big{[}\mathbbm{1}_{\widetilde{G}}{\mathbb{P}}(A|{\mathcal{G}}){\mathbb{P}}(B|{\mathcal{G}}){\mathbb{P}}(C|{\mathcal{G}})\Big{]}=\overline{\mathbb{E}}\Big{[}\mathbbm{1}_{\widetilde{G}}\overline{\mathbb{P}}(\widetilde{A}|\widetilde{\mathcal{G}})\overline{\mathbb{P}}(B^{\prime}|\widetilde{\mathcal{G}})\overline{\mathbb{P}}(C^{\prime\prime}|\widetilde{\mathcal{G}})\Big{]}.$
This leads to
¯$\overline{\mathbb{P}}\big{(}\widetilde{A}\cap B^{\prime}\cap C^{\prime\prime}|\widetilde{\mathcal{G}}\big{)}=\overline{\mathbb{P}}(\widetilde{A}|\widetilde{\mathcal{G}})\overline{\mathbb{P}}(B^{\prime}|\widetilde{\mathcal{G}})\overline{\mathbb{P}}(C^{\prime\prime}|\widetilde{\mathcal{G}}),\text{ a.s. under }\overline{\mathbb{P}}.$
As for the second statement of (3), for any $\widetilde{F}\in\widetilde{\mathcal{F}}$, there exists a $F\in{\mathcal{F}}$ such that $\widetilde{F}=F\times\Omega\times\Omega\in\widetilde{\mathcal{F}}$. Then,
¯$\overline{\mathbb{P}}(B^{\prime}\cap C^{\prime\prime}\cap\widetilde{F})$ =¯$=\overline{\mathbb{P}}(F\times B\times C)={\mathbb{E}}\big{[}\mathbbm{1}_{F}{\mathbb{P}}(B|{\mathcal{G}}){\mathbb{P}}(C|{\mathcal{G}})\big{]}$
$=\overline{\mathbb{E}}\big{[}\mathbbm{1}_{\widetilde{F}}{\mathbb{P}}(B|{\mathcal{G}}){\mathbb{P}}(C|{\mathcal{G}})\big{]}=\overline{\mathbb{E}}\big{[}\mathbbm{1}_{\widetilde{F}}\overline{\mathbb{P}}(B^{\prime}|\widetilde{\mathcal{F}})\overline{\mathbb{P}}(C^{\prime\prime}|\overline{\mathcal{F}})\big{]},$
where the second equation holds because of (5.2), the third equation holds due to the same calculation as (5.5) and the last equation holds by (1) and (2). Therefore,
¯$\overline{\mathbb{P}}\big{(}B^{\prime}\cap C^{\prime\prime}|\widetilde{\mathcal{F}}\big{)}=\overline{\mathbb{P}}\big{(}B^{\prime}|\widetilde{\mathcal{F}}\big{)}\overline{\mathbb{P}}\big{(}C^{\prime\prime}|\widetilde{\mathcal{F}}\big{)},\text{ a.s. under }\overline{\mathbb{P}},$
which completes the whole proof. ∎
Proof of Theorem 3.1. For any random variable $R$ in $(\Omega,{\mathcal{F}},{\mathbb{P}})$, recall that the natural extension $\widetilde{R}$ and the two conditional independent copies $R^{\prime},R^{\prime\prime}$ of $R$ are defined by
$\widetilde{R}(\omega_{0},\omega_{1},\omega_{2})=R(\omega_{0}),\;R^{\prime}(\omega_{0},\omega_{1},\omega_{2})=R(\omega_{1}),\;R^{\prime\prime}(\omega_{0},\omega_{1},\omega_{2})=R(\omega_{2}).$
Then
$\text{Law}\big{(}R\big{|}{\mathcal{G}}\big{)}(\omega_{0})=\text{Law}\big{(}\widetilde{R}\big{|}\widetilde{\mathcal{G}}\big{)}(\overline{\omega})=\text{Law}\big{(}R^{\prime}\big{|}{\widetilde{\mathcal{G}}}\big{)}(\overline{\omega})=\text{Law}\big{(}R^{\prime\prime}\big{|}{\widetilde{\mathcal{F}}}\big{)}(\overline{\omega})\text{ a.s. under }\overline{\mathbb{P}},$
[…]
$R,R^{\prime}\text{ and }R^{\prime\prime}\text{ are independent given }\widetilde{\mathcal{G}}\text{ and }R,R^{\prime}\text{ and }R^{\prime\prime}\text{ are independent given }{\mathcal{F}},$
which completes the proof of Theorem 3.1.∎
Proof of Corollary 3.2. It suffices to show that extension of the sub-filtration $\widetilde{\mathbb{G}}\subset\widetilde{\mathbb{F}}$ satisfies the compatibility assumption or the conditional independence condition, that is
~${\widetilde{\mathcal{F}}}_{t}\perp\!\!\!\perp{\widetilde{\mathcal{G}}}_{T}\big{|}{\widetilde{\mathcal{G}}}_{t},\text{ for all }0\leq t\leq T.$
For any $\widetilde{F}_{t}\in\widetilde{\mathcal{F}}_{t}$, $\widetilde{G}_{t}\in{\widetilde{\mathcal{G}}}_{t}$ and $\widetilde{G}_{T}\in{\widetilde{\mathcal{G}}}_{T}$, there exists $F_{t}\in{\mathcal{F}}_{t}$, $G_{t}\in{\mathcal{G}}_{t}$ and $G_{T}\in{\mathcal{G}}_{T}$ such that $\widetilde{F}_{t}=F_{t}\times\Omega\times\Omega$, $\widetilde{G}_{t}=G_{t}\times\Omega\times\Omega$ and $\widetilde{G}_{T}=G_{T}\times\Omega\times\Omega$. Then,
¯$\overline{{\mathbb{P}}}(\widetilde{F}_{t}\cap\widetilde{G}_{t}\cap\widetilde{G}_{T})$ =¯$=\overline{\mathbb{P}}((F_{t}\cap G_{t}\cap G_{T})\times\Omega\times\Omega)={\mathbb{P}}(F_{t}\cap G_{t}\cap G_{T})$
$={\mathbb{E}}\big{[}{\mathbb{P}}(F_{t}\cap G_{T}|{\mathcal{G}}_{t})\mathbbm{1}_{G_{t}}\big{]}={\mathbb{E}}\big{[}{\mathbb{P}}(F_{t}|{\mathcal{G}}_{t}){\mathbb{P}}(G_{T}|{\mathcal{G}}_{t})\mathbbm{1}_{G_{t}}\big{]}$
$=\overline{\mathbb{E}}\big{[}{\mathbb{P}}(F_{t}|{\mathcal{G}}_{t}){\mathbb{P}}(G_{T}|{\mathcal{G}}_{t})\mathbbm{1}_{\widetilde{G}_{t}}\big{]}=\overline{\mathbb{E}}\big{[}\overline{\mathbb{P}}(\widetilde{F}_{t}|\widetilde{\mathcal{G}}_{t})\overline{\mathbb{P}}(\widetilde{G}_{T}|\widetilde{\mathcal{G}}_{t})\mathbbm{1}_{\widetilde{G}_{t}}\big{]},$
where the fourth equation holds since ${\mathcal{F}}_{t}\perp\!\!\!\perp{\mathcal{G}}_{T}|{\mathcal{G}}_{t}$, the fifth equation holds because of the same calculation as (5.5) and the last equation holds because of (2) in Proposition 5.1. Therefore,
¯$\overline{\mathbb{P}}(\widetilde{F}_{t}\cap\widetilde{G}_{T}|\widetilde{\mathcal{G}}_{t})=\overline{\mathbb{P}}(\widetilde{F}_{t}|\widetilde{\mathcal{G}}_{t})\overline{\mathbb{P}}(\widetilde{G}_{T}|\widetilde{\mathcal{G}}_{t}),$
and thus completes the proof.∎
5.2. Proof of Theorem 3.3
For (1), due to the decomposablity (see [37, Theorem 1, Chapter 4]) let us assume that
$X_{t}=M_{t}+V_{t},$
where $M_{t}$ is an ${\mathbb{F}}$-local martingale and $V_{t}$ is finite variation process. To prove that $X$ is a $\overline{\mathbb{F}}$-semimartingale, it suffices to show that $M_{t}$ is a $\overline{\mathbb{F}}$-local martingale. Let $\{\tau_{k}\}_{k\in{\mathbb{N}}^{+}}$ be a sequence of ${\mathbb{F}}$-stopping times such that $M^{\tau_{k}}:=(M_{t\wedge\tau_{k}})_{t\in[0,T]}$ is an ${\mathbb{F}}$-martingale for all $k\in{\mathbb{N}}^{+}$ and as $k\to\infty$, we have $\tau_{k}\to\infty$. We claim that
$\overline{\mathbb{E}}[M_{t\wedge\tau_{k}}|\overline{\mathcal{F}}_{s}]=M_{s\wedge\tau_{k}}.$
Recall that
¯$\overline{\mathcal{F}}_{s}=\sigma\{A_{0}\times A_{1}\times A_{2}:A_{0},A_{1},A_{2}\in{\mathcal{F}}_{s}\},$
it suffices to show that
$\overline{\mathbb{E}}\big{[}M_{t\wedge\tau_{k}}\mathbbm{1}_{A_{0}\times A_{1}\times A_{2}}\big{]}=\overline{\mathbb{E}}\big{[}M_{s\wedge\tau_{k}}\mathbbm{1}_{A_{0}\times A_{1}\times A_{2}}\big{]}$ (5.6)
for all $A_{0},A_{1},A_{2}\in{\mathcal{F}}_{s}$. Note that $M_{t\wedge\tau_{k}}\mathbbm{1}_{A_{0}\times\Omega\times\Omega}$, $\mathbbm{1}_{\Omega\times A_{1}\times\Omega}$ and $\mathbbm{1}_{\Omega\times\Omega\times A_{2}}$ are conditional independent of given ${\mathcal{G}}_{T}$ by (3) in Proposition 5.1, we have
$\overline{\mathbb{E}}\big{[}M_{t\wedge\tau_{k}}\mathbbm{1}_{A_{0}\times A_{1}\times A_{2}}\big{]}=\overline{\mathbb{E}}\Big{[}\overline{\mathbb{E}}\big{[}M_{t\wedge\tau_{k}}\mathbbm{1}_{A_{0}\times\Omega\times\Omega}\big{|}{\mathcal{G}}_{T}\big{]}\overline{\mathbb{E}}\big{[}\mathbbm{1}_{\Omega\times A_{1}\times\Omega}\big{|}{\mathcal{G}}_{T}\big{]}\overline{\mathbb{E}}\big{[}\mathbbm{1}_{\Omega\times\Omega\times A_{2}}\big{|}{\mathcal{G}}_{T}\big{]}\Big{]}.$ (5.7)
By (2) in Proposition 5.1 and ${\mathcal{F}}_{s}\perp\!\!\!\perp{\mathcal{G}}_{T}|{\mathcal{G}}_{s}$, we know that
$\overline{\mathbb{E}}\big{[}\mathbbm{1}_{\Omega\times A_{1}\times\Omega}\big{|}{\mathcal{G}}_{T}\big{]}=\overline{\mathbb{P}}(A_{1}|{\mathcal{G}}_{T})=\overline{\mathbb{P}}(A_{1}|{\mathcal{G}}_{s}),\text{ and similarly }\overline{\mathbb{E}}\big{[}\mathbbm{1}_{\Omega\times\Omega\times A_{2}}\big{|}{\mathcal{G}}_{T}\big{]}=\overline{\mathbb{P}}(A_{2}|{\mathcal{G}}_{s}).$ (5.8)
Plugging back in (5.7), we have
$\overline{\mathbb{E}}\big{[}M_{t\wedge\tau_{k}}\mathbbm{1}_{A_{0}\times A_{1}\times A_{2}}\big{]}$ $=\overline{\mathbb{E}}\Big{[}\overline{\mathbb{E}}\big{[}M_{t\wedge\tau_{k}}\mathbbm{1}_{A_{0}\times\Omega\times\Omega}\big{|}{\mathcal{G}}_{T}\big{]}\overline{\mathbb{P}}\big{(}A_{1}|{\mathcal{G}}_{s}\big{)}\overline{\mathbb{P}}\big{(}A_{2}|{\mathcal{G}}_{s}\big{)}\Big{]}$ (5.9)
$=\overline{\mathbb{E}}\Big{[}\overline{\mathbb{E}}\big{[}M_{t\wedge\tau_{k}}\mathbbm{1}_{A_{0}\times\Omega\times\Omega}\big{|}{\mathcal{G}}_{s}\big{]}\overline{\mathbb{P}}\big{(}A_{1}|{\mathcal{G}}_{s}\big{)}\overline{\mathbb{P}}\big{(}A_{2}|{\mathcal{G}}_{s}\big{)}\Big{]}$
where the second inequality holds by the tower rule. Since $M^{\tau_{k}}$ is a ${\mathcal{F}}$-martingale, we get
$\overline{\mathbb{E}}\big{[}M_{t\wedge\tau_{k}}\mathbbm{1}_{A_{0}\times\Omega\times\Omega}\big{|}{\mathcal{G}}_{s}\big{]}=\overline{\mathbb{E}}\big{[}\mathbbm{1}_{A_{0}\times\Omega\times\Omega}{\mathbb{E}}[M_{t\wedge\tau_{k}}|{\mathcal{F}}_{s}]\big{|}{\mathcal{G}}_{s}\big{]}=\overline{\mathbb{E}}\big{[}M_{s\wedge\tau_{k}}\mathbbm{1}_{A_{0}\times\Omega\times\Omega}\big{|}{\mathcal{G}}_{s}\big{]}.$
Therefore,
$\overline{\mathbb{E}}\big{[}M_{t\wedge\tau_{k}}\mathbbm{1}_{A_{0}\times A_{1}\times A_{2}}\big{]}$ $=\overline{\mathbb{E}}\Big{[}\overline{\mathbb{E}}\big{[}M_{s\wedge\tau_{k}}\mathbbm{1}_{A_{0}\times\Omega\times\Omega}\big{|}{\mathcal{G}}_{s}\big{]}\overline{\mathbb{P}}\big{(}A_{1}|{\mathcal{G}}_{s}\big{)}\overline{\mathbb{P}}\big{(}A_{2}|{\mathcal{G}}_{s}\big{)}\Big{]}.$
Repeating the calculation as in (5.7), (5.8) and (5.9) we achieve
$\overline{\mathbb{E}}\big{[}M_{s\wedge\tau_{k}}\mathbbm{1}_{A_{0}\times A_{1}\times A_{2}}\big{]}$ $=\overline{\mathbb{E}}\Big{[}\overline{\mathbb{E}}\big{[}M_{s\wedge\tau_{k}}\mathbbm{1}_{A_{0}\times\Omega\times\Omega}\big{|}{\mathcal{G}}_{s}\big{]}\overline{\mathbb{P}}\big{(}A_{1}|{\mathcal{G}}_{s}\big{)}\overline{\mathbb{P}}\big{(}A_{2}|{\mathcal{G}}_{s}\big{)}\Big{]},$
which leads to (5.6).
To show that $X^{\prime}$ is an $\overline{\mathbb{F}}$-semimartingale, by the construction of the extended probability space, we can define $M^{\prime},V^{\prime}$, and $\tau_{k}^{\prime}$ to be respectively the conditional independent copy of $M,V$, and $\tau_{k}$. Then it suffices to show that it suffices to show that
${\mathbb{E}}\big{[}M^{\prime}_{t\wedge\tau^{\prime}_{k}}\mathbbm{1}_{A_{0}\times A_{1}\times A_{2}}\big{]}={\mathbb{E}}\big{[}M^{\prime}_{s\wedge\tau^{\prime}_{k}}\mathbbm{1}_{A_{0}\times A_{1}\times A_{2}}\big{]}.$
Similarly as the procedure of showing (5.6), we have
$\overline{\mathbb{E}}\big{[}M^{\prime}_{t\wedge\tau_{k}^{\prime}}\mathbbm{1}_{A_{0}\times A_{1}\times A_{2}}\big{]}=\overline{\mathbb{E}}\Big{[}\overline{\mathbb{E}}\big{[}M^{\prime}_{t\wedge\tau^{\prime}_{k}}\mathbbm{1}_{\Omega\times A_{2}\times\Omega}\big{|}{\mathcal{G}}_{T}\big{]}\overline{\mathbb{P}}(A_{0}|{\mathcal{G}}_{s})\overline{\mathbb{P}}(A_{2}|{\mathcal{G}}_{s})\Big{]}$
$=\overline{\mathbb{E}}\Big{[}\overline{\mathbb{E}}\big{[}M_{t\wedge\tau_{k}}\mathbbm{1}_{A_{1}\times\Omega\times\Omega}\big{|}{\mathcal{G}}_{T}\big{]}\overline{\mathbb{P}}(A_{0}|{\mathcal{G}}_{s})\overline{\mathbb{P}}(A_{2}|{\mathcal{G}}_{s})\Big{]}=\overline{\mathbb{E}}\Big{[}\overline{\mathbb{E}}\big{[}M_{t\wedge\tau_{k}}\mathbbm{1}_{A_{1}\times\Omega\times\Omega}\big{|}{\mathcal{G}}_{s}\big{]}\overline{\mathbb{P}}(A_{0}|{\mathcal{G}}_{s})\overline{\mathbb{P}}(A_{2}|{\mathcal{G}}_{s})\Big{]}$
$=\overline{\mathbb{E}}\Big{[}\overline{\mathbb{E}}\big{[}M_{s\wedge\tau_{k}}\mathbbm{1}_{A_{1}\times\Omega\times\Omega}\big{|}{\mathcal{G}}_{s}\big{]}\overline{\mathbb{P}}(A_{0}|{\mathcal{G}}_{s})\overline{\mathbb{P}}(A_{2}|{\mathcal{G}}_{s})\Big{]}=\overline{\mathbb{E}}\Big{[}\overline{\mathbb{E}}\big{[}M_{s\wedge\tau_{k}}\mathbbm{1}_{A_{1}\times\Omega\times\Omega}\big{|}{\mathcal{G}}_{T}\big{]}\overline{\mathbb{P}}(A_{0}|{\mathcal{G}}_{s})\overline{\mathbb{P}}(A_{2}|{\mathcal{G}}_{s})\Big{]}$
$=\overline{\mathbb{E}}\Big{[}\overline{\mathbb{E}}\big{[}M^{\prime}_{s\wedge\tau^{\prime}_{k}}\mathbbm{1}_{\Omega\times A_{2}\times\Omega}\big{|}{\mathcal{G}}_{T}\big{]}\overline{\mathbb{P}}(A_{0}|{\mathcal{G}}_{s})\overline{\mathbb{P}}(A_{2}|{\mathcal{G}}_{s})\Big{]}={\mathbb{E}}\big{[}M^{\prime}_{s\wedge\tau^{\prime}_{k}}\mathbbm{1}_{A_{0}\times A_{1}\times A_{2}}\big{]},$
where the first and last equation mimic the argument in (5.7), (5.8) and (5.9), the second and the sixth equation hold since (2) in Proposition 5.1, the third and the fifth equation is because of the tower rule and the fourth equation follows from the martingale property. This completes the proof of (1).
For (2a), it suffices to prove that
$\overline{\mathbb{E}}\bigg{[}\bigg{(}\sup_{\sigma_{m}}\sum_{i=1}^{k_{m}}\big{|}X^{\mathcal{G}}_{t_{i+1}^{m}}-X^{\mathcal{G}}_{t_{i}^{m}}\big{|}\bigg{)}^{p}\bigg{]}<\infty,$
where the supreme is taken over all partition $\sigma_{m}:\{0=t_{0}^{m}\leq t_{1}^{m}\leq\cdots\leq t_{k_{m}}^{m}=T\}$ of $[0,T]$. Since
$\sum_{i=1}^{k_{m}}\big{|}X^{\mathcal{G}}_{t_{i+1}^{m}}-X^{\mathcal{G}}_{t_{i}^{m}}\big{|}=\sum_{i=1}^{k_{m}}\Big{|}\overline{\mathbb{E}}\big{[}X_{t_{i+1}^{m}}-X_{t_{i}^{m}}\big{|}{\mathcal{G}}_{T}\big{]}\Big{|}\leq\overline{\mathbb{E}}\bigg{[}\sum_{i=1}^{k_{m}}\big{|}X_{t_{i+1}^{m}}-X_{t_{i}^{m}}\big{|}\Big{|}{\mathcal{G}}_{T}\bigg{]},$
we have,
$\overline{\mathbb{E}}$ $\bigg{[}\bigg{(}\sup_{\sigma_{m}}\sum_{i=1}^{k_{m}}\big{|}X^{\mathcal{G}}_{t_{i+1}^{m}}-X^{\mathcal{G}}_{t_{i}^{m}}\big{|}\bigg{)}^{p}\bigg{]}$
$=\overline{\mathbb{E}}\bigg{[}\sup_{\sigma_{m}}\bigg{(}\sum_{i=1}^{k_{m}}\big{|}X^{\mathcal{G}}_{t_{i+1}^{m}}-X^{\mathcal{G}}_{t_{i}^{m}}\big{|}\bigg{)}^{p}\bigg{]}\leq\overline{\mathbb{E}}\bigg{[}\sup_{\sigma_{m}}\bigg{(}\overline{\mathbb{E}}\bigg{[}\sum_{i=1}^{k_{m}}\big{|}X_{t_{i+1}^{m}}-X_{t_{i}^{m}}\big{|}\Big{|}{\mathcal{G}}_{T}\bigg{]}\bigg{)}^{p}\bigg{]}$
$\leq\overline{\mathbb{E}}\bigg{[}\bigg{(}{\mathbb{E}}\bigg{[}\sup_{\sigma_{m}}\sum_{i=1}^{k_{m}}\big{|}X_{t_{i+1}^{m}}-X_{t_{i}^{m}}\big{|}\Big{|}{\mathcal{G}}_{T}\bigg{]}\bigg{)}^{p}\bigg{]}\leq\overline{\mathbb{E}}\bigg{[}\overline{\mathbb{E}}\bigg{[}\bigg{(}\sup_{\sigma_{m}}\sum_{i=1}^{k_{m}}\big{|}X_{t_{i+1}^{m}}-X_{t_{i}^{m}}\big{|}\bigg{)}^{p}\bigg{|}{\mathcal{G}}_{T}\bigg{]}\bigg{]}$
$=\overline{\mathbb{E}}\bigg{[}\bigg{(}\sup_{\sigma_{m}}\sum_{i=1}^{k_{m}}\big{|}X_{t_{i+1}^{m}}-X_{t_{i}^{m}}\big{|}\bigg{)}^{p}\bigg{]}=\overline{\mathbb{E}}\bigg{[}\bigg{(}\int_{0}^{T}\big{|}{\mathrm{d}}X_{s}\big{|}\bigg{)}^{p}\bigg{]}<\infty,$
which leads to (2a).
For (2b), due to (2a) and the decomposablity (see [37, Theorem 1, Chapter 4]), it suffices to show that if $X$ is a ${\mathbb{F}}$-local martingale, then $X^{\mathcal{G}}$ is a local martingale with respect to $\overline{\mathbb{F}}$, ${\mathbb{F}}$ and ${\mathbb{G}}$ and thanks to (1), it can be further reduced to prove that $X^{\mathcal{G}}$ is a local martingale with respect to ${\mathbb{F}}$ and ${\mathbb{G}}$. Let $\tau_{k}$ be the stopping time $\tau_{k}:=\inf\{t\in[0,T]:[X,X]_{t}>k\}$ and consider the truncated process $X^{\tau_{k}}:=\{X_{t\wedge\tau_{k}}\}_{t\in[0,T]}$. By Burkholder-Davis-Gundy inequality,
$\overline{\mathbb{E}}\big{[}(X^{\tau_{k}})^{*}_{t}\big{]}^{2}\leq\overline{\mathbb{E}}\big{[}\big{(}(X^{\tau_{k}}\big{)}^{*}_{t})^{2}\big{]}\leq C\overline{\mathbb{E}}\big{[}\big{[}X^{\tau_{k}},X^{\tau_{k}}\big{]}_{t}\big{]}\leq k,$ (5.10)
for some constant $C$ and $(X^{\tau_{k}})^{*}_{t}=\sup_{0\leq s\leq t}|X_{s\wedge\tau_{k}}|.$ Let $\{\widetilde{\tau}_{\ell}\}_{\ell\in{\mathbb{N}}^{+}}$ be the sequence of the ${\mathbb{F}}$-stopping times such that $\tau_{\ell}\to T$ and $X^{\tilde{\tau}_{\ell}}=\{X_{t\wedge\tilde{\tau}_{\ell}}\}_{t\in[0,T]}$ are ${\mathbb{F}}$-martingales for all $\ell\in{\mathbb{N}}^{+}$. Then $X^{\tilde{\tau}_{\ell}\wedge\tau_{k}}=\{X_{t\wedge\tilde{\tau}_{\ell}\wedge\tau_{k}}\}_{t\in[0,T]}$ are ${\mathbb{F}}$-martingales for all $\ell\in{\mathbb{N}}^{+}$. Using conditional dominate convergence theorem, $|X_{t\wedge\tau_{k}\wedge\tilde{\tau}_{\ell}}|\leq(X^{\tau_{k}})^{*}_{t}$ and (5.10), we have
$X_{s\wedge\tau_{k}}=\lim_{\ell\to\infty}X_{s\wedge\tau_{k}\wedge\tilde{\tau}_{\ell}}=\lim_{\ell\to\infty}\overline{\mathbb{E}}\big{[}X_{t\wedge\tau_{k}\wedge\tilde{\tau}_{\ell}}|{\mathcal{F}}_{s}\big{]}=\overline{\mathbb{E}}\Big{[}\lim_{\ell\to\infty}X_{t\wedge\tau_{k}\wedge\tilde{\tau}_{\ell}}|{\mathcal{F}}_{s}\Big{]}=\overline{\mathbb{E}}\big{[}X_{t\wedge\tau_{k}}|{\mathcal{F}}_{s}\big{]}.$
Combining with (5.10), we have that $X^{\tau_{k}}=\{X_{t\wedge\tau_{k}}\}_{t\in[0,T]}$ is indeed a square integrable ${\mathbb{F}}$-martingale.
Now we claim that $(X^{\tau_{k}})^{\mathcal{G}}$ is a ${\mathbb{G}}$-martingale, and using the property of H hyphothesis, and the compatiability condition of ${\mathbb{F}}$ and ${\mathbb{G}}$ (i.e. ${\mathcal{F}}_{t}\perp\!\!\!\perp{\mathcal{G}}_{T}|{\mathcal{G}}_{t}$ for all $t\in[0,T]$), it is also an ${\mathbb{F}}$-martingale (see [20]). The integrability holds trivially since $X_{t}$ is integrable. And the martingale property holds since
$\overline{\mathbb{E}}\Big{[}X_{t\wedge\tau_{k}}^{\mathcal{G}}\big{|}{\mathcal{G}}_{s}\Big{]}$ $=\overline{\mathbb{E}}\Big{[}\overline{\mathbb{E}}\big{[}X_{t\wedge\tau_{k}}|{\mathcal{G}}_{T}\big{]}\big{|}{\mathcal{G}}_{s}\Big{]}=\overline{\mathbb{E}}\big{[}X_{t\wedge\tau_{k}}|{\mathcal{G}}_{s}\big{]}=\overline{\mathbb{E}}\Big{[}\overline{\mathbb{E}}\big{[}X_{t\wedge\tau_{k}}|{\mathcal{F}}_{s}\big{]}\big{|}{\mathcal{G}}_{s}\Big{]}$
$=\overline{\mathbb{E}}\big{[}X_{s\wedge\tau_{k}}|{\mathcal{G}}_{s}\big{]}=\overline{\mathbb{E}}\big{[}X_{s\wedge\tau_{k}}|{\mathcal{G}}_{T}\big{]}=X_{s\wedge\tau_{k}}^{\mathcal{G}}.$
This completes the proof of the claim that $X^{\mathcal{G}}$ is an ${\mathbb{G}}$-martingale and thus leads to(2b).
For (2c), let $\sigma_{m}:0=t_{0}^{m}\leq t_{1}^{m}\leq...\leq t_{k_{m}}^{m}=t$ be a sequence of partitions tend to identity (the definition of a sequence tending to identity can be found in the [37, Page 64, defintion] with a slight modification of the terminal time), such that
$\int_{0}^{t}Z_{s-}{\mathrm{d}}X^{\mathcal{G}}_{s}=\lim_{m\to\infty}\sum_{k=0}^{k_{m}-1}Z_{t_{k}^{m}}(X^{\mathcal{G}}_{t_{k+1}^{m}}-X^{\mathcal{G}}_{t_{k}^{m}}),\;\;\int_{0}^{t}Z_{s-}{\mathrm{d}}X^{\prime}_{s}=\lim_{m\to\infty}\sum_{k=0}^{k_{m}-1}Z_{t_{k}^{m}}(X^{\prime}_{t_{k+1}^{m}}-X^{\prime}_{t_{k}^{m}}),$
where the convergence is understood as almost surely convergence. Such partition can be found thanks to [37, Theorem 21, Chapter II]. Then
$\int_{0}^{t}Z_{s-}{\mathrm{d}}X^{\mathcal{G}}_{s}$ $=\lim_{m\to\infty}\sum_{k=0}^{k_{m}-1}Z_{t_{k}^{m}}(X^{\mathcal{G}}_{t_{k+1}^{m}}-X^{\mathcal{G}}_{t_{k}^{m}})=\lim_{m\to\infty}\sum_{k=0}^{k_{m}-1}Z_{t_{k}^{m}}{\mathbb{E}}\Big{[}X_{t_{k+1}^{m}}-X_{t_{k}^{m}}\Big{|}{\mathcal{G}}_{T}\Big{]}$
$=\lim_{m\to\infty}\sum_{k=0}^{k_{m}-1}Z_{t_{k}^{m}}\overline{\mathbb{E}}\Big{[}X^{\prime}_{t_{k+1}^{m}}-X^{\prime}_{t_{k}^{m}}\Big{|}{\mathcal{F}}\Big{]}=\lim_{m\to\infty}\sum_{k=0}^{k_{m}-1}{\mathbb{E}}\Big{[}Z_{t_{k}^{m}}(X^{\prime}_{t_{k+1}^{m}}-X^{\prime}_{t_{k}^{m}})\Big{|}{\mathcal{F}}\Big{]}.$
Define ${Z}^{m}$ as the process
$Z^{m}_{t}=Z_{0}\mathbbm{1}_{\{0\}}(t)+\sum_{k=0}^{k_{m}-1}Z_{t_{k}^{m}}\mathbbm{1}_{(t_{k}^{m},t_{k+1}^{m}]}(t),$
then $\int_{0}^{t}Z_{s-}{\mathrm{d}}X^{\mathcal{G}}_{s}=\lim_{m\to\infty}{\mathbb{E}}\big{[}\int_{0}^{t}Z_{s}^{m}{\mathrm{d}}X^{\prime}_{s}\Big{|}{\mathcal{F}}\big{]},$ and
$\sum_{k=0}^{k_{m}-1}\overline{\mathbb{E}}\Big{[}Z_{t_{k}^{m}}\big{(}X^{\prime}_{t_{k+1}^{m}}-X^{\prime}_{t_{k}^{m}}\big{)}\Big{]}-\overline{\mathbb{E}}\bigg{[}\int_{0}^{t}Z_{s-}{\mathrm{d}}X^{\prime}_{s}\bigg{]}=\overline{\mathbb{E}}\bigg{[}\int_{0}^{t}(Z_{s}^{m}-Z_{s-}){\mathrm{d}}X^{\prime}_{s}\bigg{]}.$
Recall that $X^{\prime}$ has the decomposition $X^{\prime}=M^{\prime}+V^{\prime}$ where $M^{\prime}$ is a local martingale and $V^{\prime}$ being a adapted finite variation process and
$\bigg{\|}\sqrt{[M^{\prime},M^{\prime}]_{T}}+\int_{0}^{T}\big{|}{\mathrm{d}}V^{\prime}_{s}\big{|}\bigg{\|}_{L^{p}}<\infty.$
Therefore,
$\overline{\mathbb{E}}\bigg{[}\bigg{|}\int_{0}^{t}(Z_{s}^{m}-Z^{\prime\prime}_{s-}){\mathrm{d}}X^{\prime}_{s}\bigg{|}\bigg{]}\leq\overline{\mathbb{E}}\bigg{[}\bigg{|}\int_{0}^{t}(Z_{s}^{m}-Z_{s-}){\mathrm{d}}M^{\prime}_{s}\bigg{|}\bigg{]}+\overline{\mathbb{E}}\bigg{[}\bigg{|}\int_{0}^{t}(Z_{s}^{m}-Z_{s-}){\mathrm{d}}V^{\prime}_{s}\bigg{|}\bigg{]}.$ (5.11)
For the first term in (5.11), we use Burkholder-Davis-Gundy inequality so that
$\overline{\mathbb{E}}\bigg{[}\bigg{|}\int_{0}^{t}(Z_{s}^{m}-Z_{s-}){\mathrm{d}}M^{\prime}_{s}\bigg{|}\bigg{]}$ $\leq\overline{\mathbb{E}}\bigg{[}\sup_{0\leq s\leq t}\bigg{|}\int_{0}^{s}(Z_{u}^{m}-Z_{u-}){\mathrm{d}}M^{\prime}_{u}\bigg{|}^{p}\bigg{]}$
$\leq c_{1}\overline{\mathbb{E}}\bigg{[}\bigg{(}\int_{0}^{t}(Z_{u}^{m}-Z_{u-})^{2}{\mathrm{d}}[M^{\prime},M^{\prime}]_{s}\bigg{)}^{p/2}\bigg{]}$
for some constant $c_{1}$. Note that $\int_{0}^{t}(Z_{u}^{m}-Z_{u-})^{2}{\mathrm{d}}[M^{\prime},M^{\prime}]_{s}\leq 4\sup_{0\leq s\leq t}(Z_{s})^{2}[M^{\prime},M^{\prime}]_{t}$, and
$\overline{\mathbb{E}}\Big{[}\Big{(}\sup_{0\leq s\leq t}(Z_{s})^{2}[M^{\prime},M^{\prime}]_{t}\Big{)}^{1/2}\Big{]}\leq\bigg{(}\overline{\mathbb{E}}\Big{[}\sup_{0\leq s\leq t}(Z_{s})^{q}\Big{]}\overline{\mathbb{E}}\big{[}[M^{\prime},M^{\prime}]^{p/2}_{t}\big{]}\bigg{)}^{1/2}\leq C\sqrt{\|Z\|^{q}_{{\mathcal{H}}^{q}}\overline{\mathbb{E}}\big{[}[M^{\prime},M^{\prime}]_{t}^{p/2}\big{]}}<\infty$
Since $Z$ is a càdlàg function, we know that $Z^{m}_{s}\to Z_{s-},a.s.$ and $|Z^{m}_{s}-Z_{s}|\leq\sup_{0\leq u\leq s}|Z_{u}|\in S_{q}$, and we can use the dominated convergence theorem (the classical one and the one in ${\mathcal{H}}_{p}$, see [37, Page 273, Lemma]) to get
$\overline{\mathbb{E}}\bigg{[}\int_{0}^{t}(Z_{s}^{m}-Z_{s-}){\mathrm{d}}M^{\prime}_{s}\bigg{]}\to 0.$
For the second term in (5.11), we have
$\overline{\mathbb{E}}\bigg{[}\bigg{|}\int_{0}^{t}(Z_{s}^{m}-Z_{s-}){\mathrm{d}}V^{\prime}_{s}\bigg{|}\bigg{]}\leq\overline{\mathbb{E}}\bigg{[}\int_{0}^{t}|Z_{s}^{m}-Z_{s-}|\big{|}{\mathrm{d}}V^{\prime}_{s}\big{|}\bigg{]}.$
Noting that $\int_{0}^{t}|Z_{s}^{m}-Z_{s-}|\big{|}{\mathrm{d}}V^{\prime}_{s}\big{|}\leq 2\sup_{0\leq s\leq t}|Z_{s}|\int_{0}^{t}|{\mathrm{d}}V_{t}|$, and similarly, we have
$\overline{\mathbb{E}}\bigg{[}\bigg{|}\int_{0}^{t}(Z_{s}^{m}-Z_{s-}){\mathrm{d}}V^{\prime}_{s}\bigg{|}\bigg{]}\to 0,$
and therefore,
$\lim_{m\to\infty}\overline{\mathbb{E}}\bigg{[}\bigg{|}\int_{0}^{t}Z_{s}^{m}{\mathrm{d}}X^{\prime}_{s}-\int_{0}^{t}Z_{s-}{\mathrm{d}}X^{\prime}_{s}\bigg{|}\bigg{]}=0.$
Now define
$R^{m}:=\overline{\mathbb{E}}\bigg{[}\int_{0}^{t}Z^{m}_{s-}{\mathrm{d}}X^{\prime}_{s}\Big{|}{\mathcal{F}}\bigg{]},\;R:=\overline{\mathbb{E}}\bigg{[}\int_{0}^{t}Z_{s-}{\mathrm{d}}X^{\prime}_{s}\Big{|}{\mathcal{F}}\bigg{]},$
for all $A\in{\mathcal{F}}$, we have
$\overline{\mathbb{E}}[(R-R_{m})\mathbbm{1}_{A}]=\overline{\mathbb{E}}\bigg{[}\bigg{(}\int_{0}^{t}Z_{s}^{m}{\mathrm{d}}X^{\prime}_{s}-\int_{0}^{t}Z_{s-}{\mathrm{d}}X^{\prime}_{s}\bigg{)}\mathbbm{1}_{A}\bigg{]}\leq\overline{\mathbb{E}}\bigg{[}\bigg{|}\int_{0}^{t}Z_{s}^{m}{\mathrm{d}}X^{\prime}_{s}-\int_{0}^{t}Z_{s-}{\mathrm{d}}X^{\prime}_{s}\bigg{|}\bigg{]}\to 0.$
Let $A=\{R-R_{m}>\epsilon\}$, we have
¯$\overline{\mathbb{P}}\big{(}R_{m}-R>\epsilon\big{)}\leq\frac{\overline{\mathbb{E}}[(R_{m}-R)\mathbbm{1}_{A}]}{\epsilon}\to 0,$
and similarly
¯$\overline{\mathbb{P}}\big{(}R_{m}-R<\epsilon\big{)}\to 0.$
Therefore $R_{m}\to R$ in probability. Meanwhile $R_{m}\to\int_{0}^{t}Z_{s-}{\mathrm{d}}X^{\mathcal{G}}_{s}$ almost surely. Therefore,
$\int_{0}^{t}Z_{s-}{\mathrm{d}}X^{\mathcal{G}}_{s}=\overline{\mathbb{E}}\bigg{[}\int_{0}^{t}Z^{m}_{s-}{\mathrm{d}}X^{\prime}_{s}\Big{|}{\mathcal{F}}\bigg{]}.$
For (2d), let us first prove that
$[X^{\mathcal{G}},Y^{\mathcal{G}}]_{t}=\overline{\mathbb{E}}\Big{[}[X^{\prime\prime},Y^{\prime}]_{t}\Big{|}{\mathcal{G}}\Big{]}.$
By the definition of the quadratic variation,
$\big{[}X^{\mathcal{G}},Y^{\mathcal{G}}\big{]}_{t}=X^{\mathcal{G}}_{t}Y^{\mathcal{G}}_{t}-\int_{0}^{t}X^{\mathcal{G}}_{s-}{\mathrm{d}}Y^{\mathcal{G}}_{s}-\int_{0}^{t}Y^{\mathcal{G}}_{s-}{\mathrm{d}}X^{\mathcal{G}}_{s},$
where the last two terms are well defined because of (2b). The first term $X^{\mathcal{G}}_{t}Y^{\mathcal{G}}_{t}$ can be calculated by
$X^{\mathcal{G}}_{t}Y^{\mathcal{G}}_{t}=\overline{\mathbb{E}}\big{[}X_{t}|{\mathcal{G}}_{T}\big{]}\overline{\mathbb{E}}\big{[}Y_{t}|{\mathcal{G}}_{T}\big{]}=\overline{\mathbb{E}}\big{[}X^{\prime\prime}_{t}|{\mathcal{F}}\big{]}\overline{\mathbb{E}}\big{[}Y^{\prime}_{t}|{\mathcal{F}}\big{]}=\overline{\mathbb{E}}\big{[}X^{\prime\prime}_{t}Y^{\prime}_{t}|{\mathcal{F}}\big{]},$
where the last equation holds due to (3) in Proposition 5.1. For the second term $\int_{0}^{t}X^{\mathcal{G}}_{s-}{\mathrm{d}}Y^{\mathcal{G}}_{s}$, let $\sigma_{m}:0=t_{0}^{m}\leq t_{1}^{m}\leq...\leq t_{k_{m}}^{m}=t$ be a sequence of partitions tend to identity in the sense of [37, Page 64, defintion], with a slight modification of the terminal time, such that
$\int_{0}^{t}X^{\mathcal{G}}_{s-}{\mathrm{d}}Y^{\mathcal{G}}_{s}=\lim_{m\to\infty}\sum_{k=0}^{k_{m}-1}X^{\mathcal{G}}_{t_{k}^{m}}(Y^{G}_{t_{k+1}^{m}}-Y^{\mathcal{G}}_{t_{k}^{m}}),\;\;\int_{0}^{t}X^{\prime\prime}_{s-}{\mathrm{d}}Y^{\prime}_{s}=\lim_{m\to\infty}\sum_{k=0}^{k_{m}-1}X^{\prime\prime}_{t_{k}^{m}}(Y^{\prime}_{t_{k+1}^{m}}-Y^{\prime}_{t_{k}^{m}}),$
where the convergence is understood as almost surely convergence. Therefore,
$\int_{0}^{t}X^{\mathcal{G}}_{s-}{\mathrm{d}}Y^{\mathcal{G}}_{s}$ $=\lim_{m\to\infty}\sum_{k=0}^{k_{m}-1}X^{\mathcal{G}}_{t_{k}^{m}}(Y^{\mathcal{G}}_{t_{k+1}^{m}}-Y^{\mathcal{G}}_{t_{k}^{m}})=\lim_{m\to\infty}\sum_{k=0}^{k_{m}-1}{\mathbb{E}}\big{[}X_{t_{k}^{m}}|{\mathcal{G}}_{T}\big{]}\overline{\mathbb{E}}\big{[}Y_{t_{k+1}^{m}}-Y_{t_{k}^{m}}|{\mathcal{G}}_{T}\big{]}$
$=\lim_{m\to\infty}\sum_{k=0}^{k_{m}-1}{\mathbb{E}}\big{[}X^{\prime\prime}_{t_{k}^{m}}|{\mathcal{F}}\big{]}{\mathbb{E}}\big{[}Y^{\prime}_{t_{k+1}^{m}}-Y^{\prime}_{t_{k}^{m}}|{\mathcal{F}}\big{]}=\lim_{m\to\infty}\sum_{k=0}^{k_{m}-1}\overline{\mathbb{E}}\big{[}X^{\prime\prime}_{t_{k}^{m}}\big{(}Y^{\prime}_{t_{k+1}^{m}}-Y^{\prime}_{t_{k}^{m}}\big{)}|{\mathcal{F}}\big{]},$
where the last equation holds due to (3) in Proposition 5.1. Similar as the proof of (2c), we can conclude that
$\int_{0}^{t}X^{\mathcal{G}}_{s-}{\mathrm{d}}Y^{\mathcal{G}}_{s}=\overline{\mathbb{E}}\bigg{[}\int_{0}^{t}X^{\prime\prime}_{s-}{\mathrm{d}}Y^{\prime}_{s}\bigg{|}{\mathcal{F}}\bigg{]},$
and similarly
$\int_{0}^{t}Y^{\mathcal{G}}_{s-}{\mathrm{d}}X^{\mathcal{G}}_{s}=\overline{\mathbb{E}}\bigg{[}\int_{0}^{t}Y^{\prime}_{s-}{\mathrm{d}}X^{\prime\prime}_{s}\bigg{|}{\mathcal{F}}\bigg{]}.$
Therefore
$[X^{\mathcal{G}},Y^{\mathcal{G}}]_{t}=\overline{\mathbb{E}}\Big{[}[X^{\prime\prime},Y^{\prime}]_{t}\Big{|}{\mathcal{F}}\Big{]}.$
Now let $\sigma_{m}:0=t_{0}^{m}\leq t_{1}^{m}\leq...\leq t_{k_{m}}^{m}=t$ being a sequence of partitions tend to identity in the sense of [37, Page 64, defintion] with a slight modification of the terminal time, such that
$\int_{0}^{t}Z_{s-}{\mathrm{d}}[X^{\mathcal{G}},Y^{\mathcal{G}}]_{s}=\lim_{m\to\infty}\sum_{k=0}^{k_{m}-1}Z_{t_{k}^{m}}\big{(}[X^{\mathcal{G}},Y^{\mathcal{G}}]_{t_{k+1}^{m}}-[X^{\mathcal{G}},Y^{\mathcal{G}}]_{t_{k}^{m}}\big{)},$
and
$\int_{0}^{t}Z_{s-}{\mathrm{d}}[X^{\prime\prime},Y^{\prime}]_{s}=\lim_{m\to\infty}\sum_{k=0}^{k_{m}-1}Z_{t_{k}^{m}}([X^{\prime\prime},Y^{\prime}]_{t_{k+1}^{m}}-[X^{\prime\prime},Y^{\prime}]_{t_{k}^{m}}),$
To repeat the same argument as before, we need to notice that $\big{|}\int_{0}^{t}(Z_{s}^{m}-Z_{s-}){\mathrm{d}}[X^{\prime\prime},Y^{\prime}]_{s}\big{|}\leq 2\sup_{0\leq s\leq t}|Z_{s}|[X^{\prime\prime},Y^{\prime}]_{t}$ and
[…]
where the first inequality is due to Cauchy inequality and Kunita-Watanabe Inequality and $C$ is a generic constant. Now repeating the arguments in (2c), we achieve
$\int_{0}^{t}Z_{s-}{\mathrm{d}}\big{[}X^{\mathcal{G}},Y^{\mathcal{G}}\big{]}_{t}=\overline{\mathbb{E}}\bigg{[}\int_{0}^{t}Z_{s-}{\mathrm{d}}\big{[}X^{\prime\prime},Y^{\prime}\big{]}_{s}\bigg{|}{\mathcal{F}}\bigg{]},$
and similarly, we can get
$\int_{0}^{t}Z_{s-}\big{[}X^{\mathcal{G}},Y\big{]}_{s}=\overline{\mathbb{E}}\bigg{[}\int_{0}^{t}Z_{s-}\big{[}X^{\prime},Y\big{]}_{s}\bigg{|}{\mathcal{F}}\bigg{]},$
which completes the proof.∎

### v3

@@
 Abstract.
-Motivated by recent development of mean-field systems with common noise, this paper establishes Itô’s formula for flows of conditional probability measures under a common filtration associated with general semimartingales.
-This generalizes existing works on flows of conditional measures on Itô processes and flows of deterministic measure on general semimartingales.
-The key technical components involve constructing conditional independent copies and establishing the equivalence between stochastic integrals with respect to the conditional law of semimartingales and the conditional expectation of stochastic integrals with respect to copies of semimartingales.
-Itô’s formula is then established for cylindrical functions through conditional independent copies, and extended to the general case through function approximations.
+Motivated by recent developments of mean-field systems with common noise, this paper establishes Itô’s formula for flows of conditional probability measures under a common filtration associated with general semimartingales.
+It generalizes existing works on flows of conditional measures on Itô processes and flows of deterministic measures on general semimartingales.
+It is derived via the cylindrical function approach.
+One key technical component involves constructing conditionally independent copies of stochastic processes, which allows for establishing the equivalence between stochastic integrals with respect to the conditional laws of semimartingales and the conditional expectation of stochastic integrals with respect to copies of semimartingales.
+As an application, a class of control problems for jump-diffusions with partial observations is analyzed.
 1. Introduction
 Itô’s formula
-The classical Itô’s formula for semimartingales is generally viewed as the stochastic counterpart of the chain rule in calculus.
-It stands as one of the key cornerstones of stochastic analysis [29], establishing an intrinsic connection between partial differential equations (PDEs) and diffusion processes.
-This formula plays a critical role in the theory of stochastic control.
-With the recent development of theory of mean-field games and mean-field controls [6, 8, 11], Itô’s formula has been extended to the flows of probability measures.
-Initially applied to diffusions [6], this extension was further developed for (discontinuous) semimartingales [24].
-The generalized Itô’s formula has been instrumental for deriving the master equations in mean-field games and the Bellman dynamic programming equation for McKean–Vlasov control problems, as discussed in [11] and the references therein.
-One of the most recent advances in mean-field theory have been concerning mean-field controls and games with common noise (see for instance [1, 2, 3, 5, 9, 10, 13, 15, 18, 23, 27, 30, 31, 32, 33, 35, 40]) or with stochastic processes involving jumps ([7, 22, 25, 26, 28, 36]).
-In either case, Itô’s formula has been developed respectively in [11] and in [24].
-In order to analyze rigorously the McKean–Vlasov dynamics with both semimartingales and idiosyncratic and common noises, it is necessary to establish the Itô’s formula for flows of conditional laws with semimartingales, which does not seem to exist to the best of our knowledge.
+The classical Itô’s formula for semimartingales, the stochastic equivalent of the chain rule in calculus, is a cornerstone of stochastic analysis [36].
+It intrinsically links partial differential equations (PDEs) with diffusion processes.
+Recent advancements in mean-field game and control theories [9, 11, 14] have led to its extension to flows of probability measures, initially for diffusions [9] and later for discontinuous semimartingales ([31, 53]).
+This generalized Itô’s formula is crucial for deriving master equations in mean-field games and Bellman dynamic programming equations for McKean–Vlasov control problems, see [14] and the references therein.
+Most recent studies in mean-field theory involve controls and games with common noise (see for instance [1, 2, 5, 7, 12, 13, 17, 19, 24, 30, 34, 40, 41, 44, 45, 48, 54]) or stochastic jump processes ([10, 28, 32, 33, 35, 49]).
+In either case, Itô’s formula has been developed respectively in ([9, 14]) and ([31, 53]).
+In order to analyze rigorously the McKean–Vlasov dynamics with both semimartingales and idiosyncratic and common noises, it is necessary to establish the Itô’s formula for flows of conditional laws with semimartingales, which does not seem to exist, to the best of our knowledge.
 This is the primary focus of this paper.
 Our work
-This paper establishes Itô’s formula (Theorem 2.7) for flow of conditional laws $\mu_{t}=$Law$(X_{t}|{\mathcal{G}}_{t})$ driven by (possible discontinuous) semimartingales $\{X_{t}\}_{t\in[0,T]}$ with a general form of common noise represented by a sub filtration ${\mathbb{G}}=({\mathcal{G}}_{t})_{t\in[0,T]}$ and ${\mathcal{G}}={\mathcal{G}}_{T}$.
-The primary task to analyze the Itô integral $\int_{0}^{t}\eta_{s}{\mathrm{d}}X_{s}^{\mathcal{G}}$ with $\{X_{t}^{\mathcal{G}}:={\mathbb{E}}[X_{t}|{\mathcal{G}}]\}_{t\in[0,T]}$ over some appropriate adapted process $\{\eta_{t}\}_{t\in[0,T]}$.
+[…]
+One key step is to analyze the Itô integral $\int_{0}^{t}\eta_{s}{\mathrm{d}}X_{s}^{\mathcal{G}}$ with $\{X_{t}^{\mathcal{G}}:={\mathbb{E}}[X_{t}|{\mathcal{G}}]\}_{t\in[0,T]}$ over some appropriate adapted process $\{\eta_{t}\}_{t\in[0,T]}$.
 This is crucial as the integral of the conditional expectation such as $\int_{0}^{t}\eta_{s}{\mathrm{d}}X^{\mathcal{G}}_{s}$ is not necessarily equivalent to the conditional expectation of the integral of the form ${\mathbb{E}}\big{[}\int_{0}^{t}\eta_{s}{\mathrm{d}}X_{s}|{\mathcal{G}}\big{]}$ (See discussions and examples in Section 3).
-Our analysis is built on several key technical components: constructing conditional independent copies of general stochastic processes (Theorem 3.1), establishing the equivalence between the conditional expectation of stochastic integrals with respect to seminmartingales and stochastic integrals with respect to the conditional law of semimartingales, and building the equivalence between the quadratic variation of the conditional expectation of a semimartingale and the conditional expectation of the quadratic variation of its two conditional independent copies (Theorem 3.3).
-With these key technical components, Itô’s formula for flows of conditional laws on semimartingales is established (Theorem 2.7): it is first derived for cylindrical functions (i.e., smooth mean-field functions with integrable forms 4.1)), which are then shown to be dense in the desired function space; finally, localization argument along with proper form of dominating convergence theorem finishes the task.
+Our analysis is built on two key technical components: one is to construct conditionally independent copies of general stochastic processes (Theorem 3.1), the other is to establish the equivalence between the conditional expectation of stochastic integrals with respect to semimartingales and stochastic integrals with respect to the conditional law of semimartingales, as well as the equivalence between the quadratic variation of the conditional expectation of a semimartingale and the conditional expectation of the quadratic variation of its two conditionally independent copies (Theorem 3.5).
+With these key technical components, Itô’s formula for flows of conditional laws on semimartingales is established (Theorem 2.11): the result is first established for cylindrical functions, defined as smooth mean-field functions with integrable representations (see Definition 4.1).
+Since these functions are dense in the desired function space, the general form is then obtained by a localization argument combined with suitable versions of the dominated convergence theorem.
+This Itô’s formula (Theorem 2.11) is applied to analyze control problems with partially observed jump-diffusions.
+An appropriate form of the verification theorem 6.2 is derived, which recovers the results in [4] as special cases.
 Related works
-Conditional independent copied has been treated informally and in certain forms in the existing literature, for instance, when defining a semimartingale by stochastic differential equations involving both idiosyncratic and common noise in [11].
-In this paper, we have generalized the construction of conditional independent copies for a broad class of semimartingales and with an arbitrary sub $\sigma$-algebra.
-This concept may be of independent theoretical interest, besides its crucial role in our derivation of Itô’s formula for flow of conditional laws on semimartingales.
-There are several methods for deriving Itô’s formula and its variants, including the time discretization approach in [6] and [36] for mean-field jump diffusions, the density approach in [8] using the Fokker–Planck equation, and the particle approximation approach [11, 14], [19, 38] to approximate flows of measures by flows of empirical measures.
+One of the key technical elements is the construction of conditionally independent copies of stochastic processes.
+This concept has been treated informally, for instance, when defining a semimartingale by stochastic differential equations involving both idiosyncratic and common noise in [14].
+Our construction of conditionally independent copies of general stochastic processes is inspired by the proofs of Yamada-Watanabe theorem (see for instance [39, 42]) and the theory of weak and strong solutions of stochastic differential equations (see for instance [37, 43]).
+Most existing methods specify the common information $\mathcal{G}$} as a stochastic input, and make use of the existence of the regular conditional distribution to construct an extended probability space.
+In contrast, we construct conditionally independent copies directly, rather than utilizing the regular conditional distribution.
+(See the discussion at the beginning of Section 5 for a detailed comparison).
+Our construction approach generalizes the construction of conditionally independent copies to a broad class of semimartingales and allows for an arbitrary sub $\sigma$-algebra.
+This result may be of independent theoretical interest, besides its essential role in our derivation of Itô’s formula for flow of conditional laws on semimartingales.
+A crucial assumption in our analysis is a condition on the filtration known as the compatibility condition.
+This condition appears in various contexts under different names, such as immersion [38], very good extensions [37], and natural extensions [50].
+In the theory of filtration enlargement, it is known as the H-hypothesis ([8, 26]), which plays a central role for characterizing semimartingale properties with filtration expansion or shrinkage.
+Following [43], we adopt the term compatibility to be aligned with recent literature on mean-field games ([6, 14, 15, 16, 46]).
+Finally, it is worth pointing out that there are several methods to derive Itô’s formula and its variants, including the time discretization approach in [9] and [49] for mean-field jump diffusions, the density approach in [11] using the Fokker–Planck equation, and the particle approximation approach of [14, 18], [25, 52] by flows of empirical measures.
 The cylindrical function approach has proved appropriate and powerful for analysis of general semimartingales.
-They are initially explored in Fleming–Viot processes [21], later adopted in the analysis of polynomial diffusions [17], and most recently used in Itô’s formula for flow of measures with (discontinuous) semimartingales [24].
-Meanwhile, linear derivatives on the space of probability measures are known to be appropriate for characterizing the infinitesimal changes in the functional of controlled McKean–Vlasov processes when jumps are added; this has been noted previously in [7] and further explored in [24] and [39].
+It is initially explored for Fleming–Viot processes [27], later adopted in the analysis of polynomial diffusions [23], and for Itô’s formulas on flow of measures with (discontinuous) semimartingales [31] and for non-anticipative maps of càdlàg rough paths [22].
 Notation
 Throughout the paper, we will adopt the following notations, unless otherwise specified.
-  • ${\mathbb{N}}^{+}$ denotes the the set of all possible natural numbers, and we fix $d\in{\mathbb{N}}^{+}$.
+  • We use the bold letter $\bm{x},\bm{X}$ to denote a (random) vector with multiple dimensions or a matrix depending on the context.
+$|\bm{X}|$ denotes the Euclidean norm of a vector $\bm{X}\in{\mathbb{R}}^{d}$, or the Frobenius norm of a matrix.
   • ${\mathcal{P}}({\mathbb{R}}^{d})$ denotes the space of probability measures on ${\mathbb{R}}^{d}$ with the topology of weak convergence.
 ${\mathcal{P}}_{p}({\mathbb{R}}^{d})$ for $p\in[1,\infty)$ denotes the probability measures with finite $p$-th moment, equipped with the Wasserstein-$p$ metric.
-  • $C^{k}({\mathbb{R}}^{d})$ is the set of all $k$-th differentiable functions on ${\mathbb{R}}^{d}$ with continuous derivatives up to the $k$-th order, and $C_{b}^{k}({\mathbb{R}}^{d})\subset C^{k}({\mathbb{R}}^{d})$ is the set of all $k$-th differentiable functions on ${\mathbb{R}}^{d}$ with bounded and continuous derivatives up to the $k$-th order.
-And we use the convention $C({\mathbb{R}}^{d})=C^{0}({\mathbb{R}}^{d})$ and $C_{b}({\mathbb{R}}^{d})=C^{0}_{b}({\mathbb{R}}^{d})$.
-  • For a random variable $X$, $\|X\|_{L^{p}}$ represents its $p$-th moment, i.e., $\|X\|_{L^{p}}={\mathbb{E}}[|X|^{p}]^{\frac{1}{p}}$.
-  • For vectors $\bm{a}=(a_{i})_{i=1,..,d},\bm{b}=(b_{i})_{i=1,..,d}\in{\mathbb{R}}^{d}$, we denote $\bm{a}\cdot\bm{b}=\sum_{i=1}^{d}a_{i}b_{i}$.
-For matrices $\bm{C}=(C_{ij})_{i,j=1,..,d},\bm{D}=(D_{ij})_{i,j=1,..,d}\in{\mathbb{R}}^{d\times d}$, we denote $\bm{C}:\bm{D}=\sum_{i,j=1}^{d}C_{ij}D_{ij}$.
-  • For $\mu\in{\mathcal{P}}({\mathbb{R}}^{d})$ and $\varphi:{\mathbb{R}}^{d}\to{\mathbb{R}}$ such that $\int_{{\mathbb{R}}^{d}}|\varphi(x)|{\mathrm{d}}\mu(x)<\infty$, we set
-    $\langle\mu,\varphi\rangle:=\int_{{\mathbb{R}}^{d}}\varphi(x){\mathrm{d}}\mu(x).$
+  • $C^{k}({\mathbb{R}}^{d})$ is the set of all $k$-th differentiable functions on ${\mathbb{R}}^{d}$ with continuous derivatives up to the $k$-th order, and $C_{b}^{k}({\mathbb{R}}^{d})\subset C^{k}({\mathbb{R}}^{d})$ is the set of all $k$-th differentiable functions on ${\mathbb{R}}^{d}$ with bounded and continuous derivatives up to the $k$-th order, with the convention $C({\mathbb{R}}^{d})=C^{0}({\mathbb{R}}^{d})$ and $C_{b}({\mathbb{R}}^{d})=C^{0}_{b}({\mathbb{R}}^{d})$.
+  • For a $d$-dimensional vector-valued random variable $\bm{X}$, $\|\bm{X}\|_{L^{p}}$ represents its $p$-th moment, i.e., $\|\bm{X}\|_{L^{p}}={\mathbb{E}}[|\bm{X}|^{p}]^{\frac{1}{p}}$.
+  • For vectors $\bm{a}=(a_{i})_{i=1,..,d},\bm{b}=(b_{i})_{i=1,..,d}\in{\mathbb{R}}^{d}$, denote $\bm{a}\cdot\bm{b}=\sum_{i=1}^{d}a_{i}b_{i}$.
+For matrices $\bm{C}=(C_{ij})_{i,j=1,..,d},\bm{D}=(D_{ij})_{i,j=1,..,d}\in{\mathbb{R}}^{d\times d}$, denote $\bm{C}:\bm{D}=\sum_{i,j=1}^{d}C_{ij}D_{ij}$.
+  • For $\mu\in{\mathcal{P}}({\mathbb{R}}^{d})$ and $\varphi:{\mathbb{R}}^{d}\to{\mathbb{R}}$ such that $\int_{{\mathbb{R}}^{d}}|\varphi(\bm{x})|{\mathrm{d}}\mu(\bm{x})<\infty$, set
+    $\langle\mu,\varphi\rangle:=\int_{{\mathbb{R}}^{d}}\varphi(\bm{x}){\mathrm{d}}\mu(\bm{x}).$
+  • For a càdlàg process $\bm{X}_{t}$, we adopt the standard notation:
+    $\bm{X}_{t-}:=\lim_{s\nearrow t}\bm{X}_{s},\quad\Delta\bm{X}_{t}:=\bm{X}_{t}-\bm{X}_{t-},$
+    and $\bm{X}^{c}$ to represent the continuous part of the process.
 2. Assumptions and main result
-Given two ${\mathbb{R}}^{d}$-valued ${\mathbb{F}}$- semimartingales $\bm{X}=(\bm{X}_{t})_{t\in[0,T]}$ and $\bm{Y}=(\bm{Y}_{t})_{t\in[0,T]}$ for $T>0$ and $d\in\mathbb{N}^{+}$ on a completed filtered probability space $(\Omega,{\mathcal{F}},{\mathbb{F}}=({\mathcal{F}}_{t})_{t\in[0,T]},{\mathbb{P}})$, and $\mu=\big{(}\mu_{t}:=\text{Law}(\bm{X}_{t}|{\mathcal{G}}_{t})\big{)}_{t\in[0,T]}$ with a sub filtration ${\mathbb{G}}=({\mathcal{G}}_{t})_{t\in[0,T]}\subset{\mathbb{F}}$.
-Our focus is to develop an Itô’s formula for the functional $\Phi(\mu_{t},\bm{Y}_{t})$ with the function $\Phi:{\mathcal{P}}_{p}({\mathbb{R}}^{d})\times{\mathbb{R}}^{d}\to{\mathbb{R}}$, under suitable regularity conditions.
-2.1. Assumptions
+Fix $T>0$ and $d\in\mathbb{N}^{+}$.
+Consider two ${\mathbb{R}}^{d}$-valued $\mathbb{F}$}-semimartingales $\bm{X}=(\bm{X}_{t})_{t\in[0,T]}$ and $\bm{Y}=(\bm{Y}_{t})_{t\in[0,T]}$ on a completed filtered probability space $(\Omega,{\mathcal{F}},{\mathbb{F}}=({\mathcal{F}}_{t})_{t\in[0,T]},{\mathbb{P}})$, and $\mu=\big{\{}\mu_{t}:=\text{Law}(\bm{X}_{t}|{\mathcal{G}}_{t})\big{\}}_{t\in[0,T]}$ for a sub-filtration ${\mathbb{G}}=({\mathcal{G}}_{t})_{t\in[0,T]}\subset{\mathbb{F}}$.
+Our focus is to establish an Itô’s formula for the functional $\Phi(\mu_{t},\bm{Y}_{t})$, with $\Phi:{\mathcal{P}}_{p}({\mathbb{R}}^{d})\times{\mathbb{R}}^{d}\to{\mathbb{R}}$ for some $p\in[2,\infty)$ under suitable regularity conditions.
+2.1. Assumptions and preliminary tools
 Throughout the paper, we will make the following assumptions on the semimartingales $\bm{X}$ and $\bm{Y}$, the filtrations $\mathbb{G}$} and $\mathbb{F}$}, and the function $\Phi$.
-These are standard assumptions to ensure that stochastic integrals are properly defined (see [37]) with the presence of common noise (see [34]) and for flow of measures on semimartingales (see [24, 16]).
-Assumptions on semimartingales.
-Assumption 2.1 (Semimartingale).
-  • We assume $\bm{X},\bm{Y}\in\mathcal{H}^{p}$ for $1\leq p\leq\infty$ [37, Section V.2], with
-    $\big{\|}\bm{X}\big{\|}_{{\mathcal{H}}^{p}}:=\inf_{\bm{M},\bm{V}}\bigg{\|}|\bm{X}_{0}|+\sqrt{[\bm{M},\bm{M}]_{T}}+\int_{0}^{T}\big{|}{\mathrm{d}}\bm{V}_{s}\big{|}\bigg{\|}_{L^{p}}<\infty,$ (2.1)
-    $\big{\|}\bm{Y}\big{\|}_{{\mathcal{H}}^{p}}:=\inf_{\bm{N},\bm{U}}\bigg{\|}|\bm{Y}_{0}|+\sqrt{[\bm{N},\bm{N}]_{T}}+\int_{0}^{T}\big{|}{\mathrm{d}}\bm{U}_{s}\big{|}\bigg{\|}_{L^{p}}<\infty,$
+These are standard assumptions to ensure that stochastic integrals are properly defined ([51]) with the presence of common noise ([47]), and for flow of measures on semimartingales ([21, 31]).
+First, we specify the assumptions on the semimartingales $\bm{X}$ and $\bm{Y}$.
+Assumption 2.1.
+For some $1\leq p\leq\infty$,
+  • $\big{\|}\bm{X}\big{\|}_{{\mathcal{H}}^{p}}=\inf_{\bm{M},\bm{V}}\bigg{\|}|\bm{X}_{0}|+\sqrt{[\bm{M},\bm{M}]_{T}}+\int_{0}^{T}\big{|}{\mathrm{d}}\bm{V}_{s}\big{|}\bigg{\|}_{L^{p}}<\infty,$ (2.1)
+    $\big{\|}\bm{Y}\big{\|}_{{\mathcal{H}}^{p}}=\inf_{\bm{N},\bm{U}}\bigg{\|}|\bm{Y}_{0}|+\sqrt{[\bm{N},\bm{N}]_{T}}+\int_{0}^{T}\big{|}{\mathrm{d}}\bm{U}_{s}\big{|}\bigg{\|}_{L^{p}}<\infty,$
     where the infimum is taken over all possible decompositions
-    $\bm{X}_{t}=\bm{X}_{0}+\bm{M}_{t}+\bm{V}_{t}\text{ and }\bm{Y}_{t}=\bm{Y}_{0}+\bm{N}_{t}+\bm{U}_{t}.$
-    Here $\bm{V}=(\bm{V}_{t})_{t\in[0,T]}$ and $\bm{U}=(\bm{U}_{t})_{t\in[0,T]}$ are adapted càdlàg processes of finite variation with $\bm{V}_{0}=\bm{U}_{0}=\bm{0}$, and $\bm{M}=(\bm{M}_{t})_{t\in[0,T]}$ and $\bm{N}=(\bm{N}_{t})_{t\in[0,T]}$ are ${\mathbb{F}}$-local martingales such that $\bm{M}_{0}=\bm{N}_{0}=\bm{0}$.
-  • We assume also
+    $\bm{X}_{t}=\bm{X}_{0}+\bm{M}_{t}+\bm{V}_{t}\text{ and }\bm{Y}_{t}=\bm{Y}_{0}+\bm{N}_{t}+\bm{U}_{t},$
+    where $\bm{V}=(\bm{V}_{t})_{t\in[0,T]}$ and $\bm{U}=(\bm{U}_{t})_{t\in[0,T]}$ are adapted càdlàg processes of finite variation with $\bm{V}_{0}=\bm{U}_{0}=\bm{0}$, and $\bm{M}=(\bm{M}_{t})_{t\in[0,T]}$ and $\bm{N}=(\bm{N}_{t})_{t\in[0,T]}$ are $\mathbb{F}$}-local martingales such that $\bm{M}_{0}=\bm{N}_{0}=\bm{0}$.
   • ${\mathbb{E}}\bigg{[}\Big{(}\sum_{0\leq s\leq T}|\Delta\bm{X}_{s}|\Big{)}^{p}\bigg{]}+{\mathbb{E}}\bigg{[}\Big{(}\sum_{0\leq s\leq T}|\Delta\bm{Y}_{s}|\Big{)}^{p}\bigg{]}<\infty.$ (2.2)
-Under these assumptions, a proper form of dominated convergence theorem has been established ([37, Page 273, Lemma]) and the following two propositions hold.
-Proposition 2.2 ([37, Section V, Theorem 2]).
-For any $1\leq p\leq\infty$, consider a norm on the space of ${\mathbb{R}}$-valued adapted càdlàg processes (or the space of ${\mathbb{R}}$-valued adapted càglàd processes) defined by
+There are two propositions that are necessary for the subsequent analysis, especially for the localization argument in establishing the Itô’s lemma for the flow of conditional measures on semimartingales, as will be clear in Section 4.2.
+Proposition 2.2 ([51, Section V, Theorem 2]).
+For any $1\leq p\leq\infty$, define a norm on the space $D$ of all $\mathbb{R}$}-valued adapted càdlàg processes, such that for $H=(H_{t})_{t\in[0,T]}\in D$,
 $\|H\|_{{\mathcal{S}}^{p}}:=\Big{\|}\sup_{0\leq t\leq T}|H_{t}|\Big{\|}_{L^{p}}.$
-Then, for $1\leq p<\infty$, there exists a constant $c_{p}$ depending only on $p$ such that
+Then, there exists a constant $c_{p}$ depending only on $p$ such that
 $\|H\|_{{\mathcal{S}}^{p}}\leq c_{p}\|H\|_{{\mathcal{H}}^{p}},$
 for any $\mathbb{R}$}-valued semimartingale $H$.
-Proposition 2.3 (Emery inequality, [37, Section V, Theorem 3]).
-Let $Z$ be an ${\mathbb{R}}$-valued semimartingale, $H$ is an ${\mathbb{R}}$-valued adapted càglàd process, and $\frac{1}{p}+\frac{1}{q}=\frac{1}{r}$ ($1\leq p,q\leq\infty)$, then
-$\Big{\|}\int_{0}^{\cdot}H_{s}{\mathrm{d}}Z_{s}\Big{\|}_{{\mathcal{H}}^{r}}\leq\|H\|_{{\mathcal{S}}^{p}}\|Z\|_{{\mathcal{H}}^{q}}.$
+Proposition 2.3 (Emery inequality, [51, Section V, Theorem 3]).
+Let $Z$ be an $\mathbb{R}$}-valued semimartingale, $H$ be an $\mathbb{R}$}-valued adapted càdlàg process, and $\frac{1}{p}+\frac{1}{q}=\frac{1}{r}$ ($1\leq p,q\leq\infty)$, then
+$\Big{\|}\int_{0}^{\cdot}H_{s-}{\mathrm{d}}Z_{s}\Big{\|}_{{\mathcal{H}}^{r}}\leq\|H\|_{{\mathcal{S}}^{p}}\|Z\|_{{\mathcal{H}}^{q}}.$
 Remark 2.4.
 Propositions 2.2 and 2.3 also hold for ${\mathbb{R}}^{d}$-valued semimartingales with appropriate forms in $d\in{\mathbb{N}}^{+}$.
-These propositions are important for the subsequent analysis, for instance in the localization argument for establishing Itô’s lemma for the flow of conditional measure on semimartinagles, as will be clear in Section 4.2.
-Assumptions on filtration.
-Assumption 2.5 (Filtration).
-Sub-filtration ${\mathbb{G}}\subset{\mathbb{F}}$ satisfies the compatibility assumption or the conditional independence condition.
-That is, for each $0\leq t\leq T$, ${\mathcal{F}}_{t}$ and ${\mathcal{G}}_{T}$ are independent, given ${\mathcal{G}}_{t}$, written as ${\mathcal{F}}_{t}\perp\!\!\!\perp{\mathcal{G}}_{T}\big{|}{\mathcal{G}}_{t}.$
-In the context of mean field games, this assumption ensures a weak closure of adapted processes, as first noted in [12, Lemma 3.11], and then stated more explicitly in [4, Theorem 5.4].
-One can deduce from is assumption that $\text{Law}(\bm{X}_{t}|{\mathcal{G}}_{t})=\text{Law}(\bm{X}_{t}|{\mathcal{G}}_{T})$, for any ${\mathbb{F}}$-adapted process $\bm{X}=(\bm{X}_{t})_{t\in[0,T]}$ almost surely for each $t\leq T$, according to a special form of the (H) hypothesis by freezing ${\mathcal{G}}_{u}={\mathcal{G}}_{T}$ when $u\geq T$ ([20]).
-This allows us to define a cádlág version of $\mu_{t}:=\text{Law}(\bm{X}_{t}|{\mathcal{G}}_{t})$ by taking $\text{Law}(\bm{X}_{t}|{\mathcal{G}}_{T})$, which will be assumed throughout the paper.
-Assumptions on function $\Phi$. We will adopt the the linear derivative adapted from [16], which has shown to be appropriate for the analyzing the flow of measures on semimartingales [24, 16].
-Definition 2.6 ($C^{2,2}({\mathcal{P}}_{p}({\mathbb{R}}^{d})\times{\mathbb{R}}^{d})$ functions).
-For $p\geq 2$, a function $\Phi(\mu,\bm{y})$ is a $C^{2,2}({\mathcal{P}}_{p}({\mathbb{R}}^{d})\times{\mathbb{R}}^{d})$ if for any $\mu\in\mathcal{P}_{p}({\mathbb{R}}^{d})$ and $\bm{y}\in{\mathbb{R}}^{d}$, there exists a continuous mapping $(\mu,\bm{y},\bm{x}_{1})\to\frac{\delta\Phi}{\delta\mu}(\mu,\bm{y},\bm{x}_{1})$, and there exists a continuous mapping $(\mu,\bm{y},\bm{x}_{1},\bm{x}_{2})\to\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}(\mu,\bm{y},\bm{x}_{1},\bm{x}_{2})$ that is symmetric in its two arguments and with the following properties,
-  • continuously differentiable: $\nabla_{\bm{y}}\Phi(\mu,\bm{y})$, $\nabla_{\bm{y}}^{2}\Phi(\mu,\bm{y})$, $\frac{\delta\Phi}{\delta\mu}(\mu,\bm{y},\bm{x}_{1})$, $\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}(\mu,\bm{y},\bm{x}_{1})$, $\nabla_{\bm{y}}\frac{\delta\Phi}{\delta\mu}(\mu,\bm{y},\bm{x}_{1})$ $\nabla^{2}_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}(\mu,\bm{y},\bm{x}_{1})$, $\nabla_{\bm{x}_{1}}\nabla_{\bm{y}}\frac{\delta\Phi}{\delta\mu}(\mu,\bm{y},\bm{x}_{1})$, $\nabla_{\bm{x}_{1}}\nabla_{\bm{x}_{2}}\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}(\mu,\bm{y},\bm{x}_{1},\bm{x}_{2})$ all exist and are continuous functions for all $\bm{y},\bm{x},\bm{x}^{\prime}\in{\mathbb{R}}^{d}$, $\mu\in{\mathcal{P}}_{p}({\mathbb{R}}^{d})$,
-  • uniform polynomial-growth : there exists a constant $c>0$ such that for all $\bm{x}_{1},\bm{x}_{2},\bm{y}\in{\mathbb{R}}^{d}$, $\mu\in{\mathcal{P}}({\mathbb{R}}^{d})$, we have
+Moreover, a proper form of dominated convergence theorem has been established as follows.
+Lemma 2.5.
+([51, Page 273, Lemma]) Let $p,q,r$ be given such that $\frac{1}{p}+\frac{1}{q}=\frac{1}{r}$ where $1<r<+\infty$.
+Let $Z$ be a semimartingale such that $\|Z\|_{\mathcal{H}_{q}}<+\infty$, and let $\{H^{n}=(H^{n}_{t})_{t\in[0,T]}\}_{n\in{\mathbb{N}}}$ and $Y=(Y_{t})_{t\in[0,T]}$ be semimartingales such that $\|H^{n}\|_{{\mathcal{S}}^{p}}<+\infty$ for all $n\in{\mathbb{N}}$, $\|Y\|_{{\mathcal{S}}^{p}}<+\infty$ and $|H^{n}_{t}|\leq Y_{t}$.
+Suppose $\lim_{n\to\infty}H^{n}_{t-}(\omega)=0$, for all $(t,\omega)$.
+Then
+$\lim_{n\to\infty}\bigg{\|}\int_{0}^{T}H^{n}_{s-}{\mathrm{d}}Z_{s}\bigg{\|}_{{\mathcal{H}}^{r}}=0.$
+Next, we impose the compatibility condition for filtrations, in the following sense.
+Assumption 2.6 (Compatibility).
+The sub-filtration ${\mathbb{G}}\subset{\mathbb{F}}$ satisfies the compatibility condition, if for each $0\leq t\leq T$, ${\mathcal{F}}_{t}$ and ${\mathcal{G}}_{T}$ are conditionally independent given ${\mathcal{G}}_{t}$, expressed as ${\mathcal{F}}_{t}\perp\!\!\!\perp{\mathcal{G}}_{T}\big{|}{\mathcal{G}}_{t}.$
+Remark 2.7.
+There are a number of alternative characterizations for the compatibility condition, such as the immersion or the H-hypothesis, in the literature of filtration enlargement and shrinkage (see for instance [8, 26, 38]).
+Given two filtrations ${\mathbb{G}}=({\mathcal{G}}_{t})_{t\in[0,T]}$ and ${\mathbb{F}}=({\mathcal{F}}_{t})_{t\in[0,T]}$ defined on the same space, with ${\mathbb{G}}\subset{\mathbb{F}}$, the following are known to be equivalent ([8, Theorem 3]):
+  • ${\mathcal{F}}_{t}\perp\!\!\!\perp{\mathcal{G}}_{T}|{\mathcal{G}}_{t}$, for all $0\leq t\leq T.$ (Compatibility condition).
+  • Every square integrable $\mathbb{G}$}-martingale is an $\mathbb{F}$}-martingale. (H-hypothesis).
+  • ${\mathbb{E}}[Z|{\mathcal{G}}_{t}]$ = ${\mathbb{E}}[Z|{\mathcal{F}}_{t}]$ a.s., for each $t\in[0,T]$ and each bounded ${\mathcal{F}}_{T}$-measurable random variable $Z$.
+In the context of mean-field games, this compatibility assumption ensures a weak closure of adapted processes, as first noted in [15, Lemma 3.11] and stated more explicitly in [6, Theorem 5.4].
+Moreover, as shown in the following lemma, this compatibility condition allows us to define a càdlàg version of $\mu_{t}$ by taking $\mu_{t}=\text{Law}(\bm{X}_{t}|{\mathcal{G}}_{T})$, which will be used throughout the paper.
+Lemma 2.8.
+Under Assumption 2.6, $\text{Law}(\bm{X}_{t}|{\mathcal{G}}_{T})=\text{Law}(\bm{X}_{t}|{\mathcal{G}}_{t})$ almost surely.
+Proof.
+It suffices to show that for any bounded ${\mathcal{G}}_{T}$- measurable random variable $A$, ${\mathbb{E}}\big{[}\varphi(\bm{X}_{t})A\big{]}={\mathbb{E}}\Big{[}{\mathbb{E}}\big{[}\varphi(\bm{X}_{t})|{\mathcal{G}}_{t}\big{]}A\Big{]}.$ This is guaranteed by ${\mathcal{F}}_{t}\perp\!\!\!\perp{\mathcal{G}}_{T}|{\mathcal{G}}_{t}$ and
+${\mathbb{E}}\big{[}\varphi(X_{t})A\big{]}$ $={\mathbb{E}}\Big{[}{\mathbb{E}}\big{[}\varphi(X_{t})A|{\mathcal{G}}_{t}\big{]}\Big{]}={\mathbb{E}}\Big{[}{\mathbb{E}}\big{[}\varphi(X_{t})|{\mathcal{G}}_{t}\big{]}{\mathbb{E}}\big{[}A|{\mathcal{G}}_{t}\big{]}\Big{]}$
+$={\mathbb{E}}\Big{[}{\mathbb{E}}\Big{[}{\mathbb{E}}\big{[}\varphi(X_{t})|{\mathcal{G}}_{t}\big{]}A\big{|}{\mathcal{G}}_{t}\Big{]}\Big{]}={\mathbb{E}}\Big{[}{\mathbb{E}}\big{[}\varphi(X_{t})|{\mathcal{G}}_{t}\big{]}A\Big{]}.$
+Finally, we will assume $\Phi\in C^{2,2}({\mathcal{P}}_{p}({\mathbb{R}}^{d})\times{\mathbb{R}}^{d})$ according to the following definition.
+Definition 2.9 ($C^{2,2}({\mathcal{P}}_{p}({\mathbb{R}}^{d})\times{\mathbb{R}}^{d})$).
+For $p\geq 2$, a function $\Phi(\mu,\bm{y})$ is $C^{2,2}({\mathcal{P}}_{p}({\mathbb{R}}^{d})\times{\mathbb{R}}^{d})$ if for any $\mu\in\mathcal{P}_{p}({\mathbb{R}}^{d})$ and $\bm{y}\in{\mathbb{R}}^{d}$, there exists a continuous mapping $(\mu,\bm{y},\bm{x}_{1})\to\frac{\delta\Phi}{\delta\mu}(\mu,\bm{y},\bm{x}_{1})$, as well as a continuous mapping $(\mu,\bm{y},\bm{x}_{1},\bm{x}_{2})\to\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}(\mu,\bm{y},\bm{x}_{1},\bm{x}_{2})$ that is symmetric in its two arguments $(\bm{x}_{1},\bm{x}_{2})$ and with the following properties:
+  • (Continuous differentiability) $\nabla_{\bm{y}}\Phi(\mu,\bm{y})$, $\nabla_{\bm{y}}^{2}\Phi(\mu,\bm{y})$, $\frac{\delta\Phi}{\delta\mu}(\mu,\bm{y},\bm{x}_{1})$, $\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}(\mu,\bm{y},\bm{x}_{1})$,
+    $\nabla_{\bm{y}}\frac{\delta\Phi}{\delta\mu}(\mu,\bm{y},\bm{x}_{1})$ $\nabla^{2}_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}(\mu,\bm{y},\bm{x}_{1})$, $\nabla_{\bm{x}_{1}}\nabla_{\bm{y}}\frac{\delta\Phi}{\delta\mu}(\mu,\bm{y},\bm{x}_{1})$, $\nabla_{\bm{x}_{1}}\nabla_{\bm{x}_{2}}\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}(\mu,\bm{y},\bm{x}_{1},\bm{x}_{2})$ all exist and are continuous for all $\bm{y},\bm{x}_{1},\bm{x}_{2}\in{\mathbb{R}}^{d}$, $\mu\in{\mathcal{P}}_{p}({\mathbb{R}}^{d})$.
+  • (Uniform polynomial-growth) there exists a constant $c>0$ such that for all $\bm{x}_{1},\bm{x}_{2},\bm{y}\in{\mathbb{R}}^{d}$, $\mu\in{\mathcal{P}}({\mathbb{R}}^{d})$,
     $\Big{|}\nabla_{\bm{y}}\Phi\Big{|}\leq c\big{(}1+|\bm{y}|^{p-1}\big{)},\quad\Big{|}\nabla^{2}_{\bm{y}}\Phi\Big{|}$ $\leq c\big{(}1+|\bm{y}|^{p-2}\big{)},$
     $\bigg{|}\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\bigg{|}+\bigg{|}\nabla_{\bm{y}}\frac{\delta\Phi}{\delta\mu}\bigg{|}$ $\leq c\big{(}1+|\bm{x}_{1}|^{p-1}+|\bm{y}|^{p-1}\big{)},$
-    $\bigg{|}\nabla^{2}_{\bm{x}}\frac{\delta\Phi}{\delta\mu}\bigg{|}+\bigg{|}\nabla_{\bm{x}_{1}\bm{y}}\frac{\delta\Phi}{\delta\mu}\bigg{|}$ $\leq c\big{(}1+|\bm{x}_{1}|^{p-2}+|\bm{y}|^{{p-2}}\big{)},$
-    $\bigg{|}\nabla_{\bm{x}_{1},\bm{x}_{2}}\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\bigg{|}$ $\leq c\big{(}1+|\bm{x}_{1}|^{p-2}+|\bm{x}_{2}|^{p-2}+|\bm{y}|^{p-2}\big{)},$
-  • fundamental theorem of calculus: for everything $\mu,\nu\in{\mathcal{P}}_{2}({\mathbb{R}}^{d})$, we have
+    $\bigg{|}\nabla^{2}_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\bigg{|}+\bigg{|}\nabla_{\bm{x}_{1}\bm{y}}\frac{\delta\Phi}{\delta\mu}\bigg{|}$ $\leq c\big{(}1+|\bm{x}_{1}|^{p-2}+|\bm{y}|^{{p-2}}\big{)},$
+    $\bigg{|}\nabla_{\bm{x}_{1},\bm{x}_{2}}\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\bigg{|}$ $\leq c\big{(}1+|\bm{x}_{1}|^{p-2}+|\bm{x}_{2}|^{p-2}+|\bm{y}|^{p-2}\big{)}.$
+  • (Fundamental theorem of calculus) for everything $\mu,\nu\in{\mathcal{P}}_{2}({\mathbb{R}}^{d})$,
     $\Phi(\mu,\bm{y})$- $\Phi(\nu,\bm{y})=\int_{0}^{1}\int_{{\mathbb{R}}^{d}}\frac{\delta\Phi}{\delta\mu}\big{(}\lambda\mu+(1-\lambda)\nu,\bm{y},\bm{x}_{1}\big{)}(\mu-\nu)({\mathrm{d}}\bm{x}_{1}){\mathrm{d}}\lambda,$
-    $\Phi(\bm{x}_{0},\mu)-$ $\Phi(\bm{x}_{0},\nu)=\int_{0}^{1}\int_{{\mathbb{R}}^{d}}\frac{\delta\Phi}{\delta\mu}\big{(}\nu,\bm{y},\bm{x}_{1}\big{)}(\mu-\nu)({\mathrm{d}}\bm{x}_{1}){\mathrm{d}}\lambda$
-    $+\int_{0}^{1}\int_{0}^{t}\int_{{\mathbb{R}}^{d}}\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}s\mu+(1-s)\nu,\bm{y},\bm{x}_{1},\bm{x}_{2}\big{)}(\mu-\nu)({\mathrm{d}}\bm{x}_{1})(\mu-\nu)({\mathrm{d}}\bm{x}_{2}){\mathrm{d}}s{\mathrm{d}}t.$
-Such $\frac{\delta\Phi}{\delta\mu}$ and $\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}$ are called (a version of) the linear derivative and the second derivative of $\Phi$.
+    $\Phi(\mu,\bm{y})$- $\Phi(\nu,\bm{y})-\int_{0}^{1}\int_{{\mathbb{R}}^{d}}\frac{\delta\Phi}{\delta\mu}\big{(}\nu,\bm{y},\bm{x}_{1}\big{)}(\mu-\nu)({\mathrm{d}}\bm{x}_{1})$
+    $=\int_{0}^{1}\int_{0}^{t}\int_{{\mathbb{R}}^{d}}\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}s\mu+(1-s)\nu,\bm{y},\bm{x}_{1},\bm{x}_{2}\big{)}(\mu-\nu)({\mathrm{d}}\bm{x}_{1})(\mu-\nu)({\mathrm{d}}\bm{x}_{2}){\mathrm{d}}s{\mathrm{d}}t.$
+Here $\frac{\delta\Phi}{\delta\mu}$ and $\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}$ are respectively (a version of) the first order and the second order linear derivatives of $\Phi$.
+Remark 2.10.
+Note that the linear derivative is in the sense of [21], which is properly defined up to an additive constant.
+This form of linear derivative is shown to be appropriate for analyzing the flow of measures on semimartingales [21, 31].
 2.2. Main result
+Now we are ready to state the main result.
 Theorem 2.11 (Itô’s formula for flows of conditional measure on semimartingales).
-Given a completed filtered probability space $(\Omega,{\mathcal{F}},{\mathbb{F}}=({\mathcal{F}}_{t})_{t\in[0,T]},{\mathbb{P}})$, which supports two ${\mathbb{R}}^{d}$-valued ${\mathbb{F}}$-semimartingales $\bm{X}=(\bm{X}_{t})_{t\in[0,T]}$ and $\bm{Y}=(\bm{Y}_{t})_{t\in[0,T]}$ satisfying Assumption 2.1.
-Let $\mu_{t}=$Law$(X_{t}|{\mathcal{G}}_{t})$, with the subfiltration ${\mathbb{G}}=({\mathcal{G}}_{t})_{t\in[0,T]}\subset{\mathbb{F}}$ satisfying Assumption 2.5.
-Then, for $p\geq 2$ and any $\Phi\in C^{2,2}(\mathcal{P}_{p}({\mathbb{R}}^{d})\times{\mathbb{R}}^{d})$, we have
+[…]
 $\Phi(\mu_{t},\bm{Y}_{t})-\Phi(\mu_{0},\bm{Y}_{0})$ (2.3)
 $=$ $\;\overline{{\mathbb{E}}}\bigg{[}\int_{0+}^{t}\bigg{(}\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\cdot{\mathrm{d}}(\bm{X}^{\prime})^{c}_{s}+\frac{1}{2}\nabla^{2}_{\bm{x}_{1}\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}:{\mathrm{d}}[\bm{X}^{\prime},\bm{X}^{\prime}]^{c}_{s}$
-$\qquad\quad+\frac{1}{2}\nabla^{2}_{\bm{x}_{1}\bm{x}_{2}}\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-},\bm{X}^{\prime\prime}_{s-}\big{)}:{\mathrm{d}}[\bm{X}^{\prime},\bm{X}^{\prime\prime}]^{c}_{t}+\nabla^{2}_{\bm{x}_{1}\bm{y}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}:{\mathrm{d}}[\bm{X}^{\prime},\bm{Y}]^{c}_{s}\bigg{)}$
+$\qquad\quad+\frac{1}{2}\nabla^{2}_{\bm{x}_{1}\bm{x}_{2}}\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-},\bm{X}^{\prime\prime}_{s-}\big{)}:{\mathrm{d}}[\bm{X}^{\prime},\bm{X}^{\prime\prime}]^{c}_{s}+\nabla^{2}_{\bm{x}_{1}\bm{y}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}:{\mathrm{d}}[\bm{X}^{\prime},\bm{Y}]^{c}_{s}\bigg{)}$
 $\quad\quad+\sum_{0<s\leq t}\bigg{(}\frac{1}{2}\bigg{(}\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s},\bm{X}^{\prime\prime}_{s}\big{)}-\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-},\bm{X}^{\prime\prime}_{s}\big{)}$
@@
 $+\sum_{0<s\leq t}\Big{(}\Phi(\mu_{s},\bm{Y}_{s})-\Phi(\mu_{s-},\bm{Y}_{s-})\Big{)},$
-for all $t\in[0,T]$, where $\bm{X}^{\prime},\bm{X}^{\prime\prime}$ are the conditional independent copies of $\bm{X}$ given the sub $\sigma$-algebra ${\mathcal{G}}_{T}$ defined in an enlarged probability space $(\overline{\Omega},\overline{{\mathcal{F}}},\overline{{\mathbb{P}}})$ and $\overline{{\mathbb{E}}}[\cdot|{\mathcal{F}}]$ is the conditional expectation given the extension ${\mathcal{F}}\subset\overline{{\mathcal{F}}}$ in the enlarged probability space.
+for all $t\in[0,T]$, where $\bm{X}^{\prime},\bm{X}^{\prime\prime}$ are two conditionally independent copies of $\bm{X}$ given the sub $\sigma$-algebra ${\mathcal{G}}_{T}$ defined in some enlarged probability space $(\overline{\Omega},\overline{{\mathcal{F}}},\overline{{\mathbb{P}}})$, and $\overline{{\mathbb{E}}}[\cdot|{\mathcal{F}}]$ is the conditional expectation given $\mathcal{F}$} in the enlarged probability space.
 Remark 2.12.
-To ensure that the enlarged space $(\overline{\Omega},\overline{{\mathcal{F}}},\overline{{\mathbb{F}}},\overline{{\mathbb{P}}})$ is properly defined such that $\mu_{t},\bm{X}_{t},\bm{Y}_{t}$ and ${\mathcal{F}}$ are naturally extended in this enlarged space, one needs the notion of independent copies of stochastic processes, as will be developed in the next section and discussed after Corollary 3.2.
-3. Conditional independent copy
-3.1. Why conditional independent copy?
-One of the key components in Theorem 2.7 is the concept of the conditional independent copy of stochastic processes.
-Here in this section, we illustrate through several examples its properties in the context of flow of conditional laws for semimartingales and its role in the derivation of Itô’s formula.
-Let us first recall a mean-field game with common noise where $X_{t}$ is of the form
+To ensure that the enlarged space $(\overline{\Omega},\overline{{\mathcal{F}}},\overline{{\mathbb{F}}},\overline{{\mathbb{P}}})$ is properly defined such that $\mu_{t},\bm{X}_{t},\bm{Y}_{t}$ and $\mathcal{F}$} are naturally extended to this enlarged space and $\bm{X}^{\prime}$$\bm{X}^{\prime\prime}$ are well-defined, one needs the precise notion of conditionally independent copies of stochastic processes.
+In addition, one needs to show the existence of these conditionally independent copies in Theorem 2.11.
+These will be studied in the next section and specified in Theorem 3.1 and Corollary 3.2.
+Remark 2.13.
+Theorem 2.11 includes the following Itô’s formulas as special cases:
+  • The Itô’s formula for mean-field games and control problems with common noise developed in [9, 14]: take $X_{t}$ as an Itô’s process driven by both idiosyncratic noise $W$ and common noise $W^{0}$, let the common filtration be the filtration generated by $W_{0}$.
+  • The Itô’s formula for flows of measures associated with possibly discontinuous processes without common noise studied in [31, 53]: set $X_{t}$ as a general semimartingale and choose $\mathcal{G}$} as the trivial $\sigma$-algebra.
+  • The Itô’s formula for partially observed diffusions as in [4]: let $X_{t}$ be an Itô’s process and letting $\mathcal{G}$} be the filtration generated by the observation process.
+For more details, see Section 6 for an extension of this problem to jump-diffusion processes.
+3. Conditionally Independent Copy of Stochastic Process
+3.1. Why conditionally independent copy?
+One of the key components in Theorem 2.11 is the notion of conditionally independent copies of stochastic processes.
+We now illustrate this concept through several examples in the context of flow of conditional laws for semimartingales and its role in the derivation of Itô’s formula.
+Let us first consider a real-valued stochastic process $X_{t}$ of the form
 ${\mathrm{d}}X_{t}=b_{t}{\mathrm{d}}t+\sigma_{t}{\mathrm{d}}W_{t}+\sigma^{0}_{t}{\mathrm{d}}W^{0}_{t},$
-where $W,W^{0}$ are two independent Brownian motions and ${\mathbb{G}}=\{{\mathcal{G}}_{t}\}_{t\in[0,T]}$ is the sub filtration generated by $W^{0}$ and ${\mathcal{G}}={\mathcal{G}}_{T}=\sigma(W^{0}_{u},u\leq T)$.
-By the classical Itô’s formula, the dynamic of $g(X_{t})$ is given by
+where $W,W^{0}$ are two independent real valued Brownian motions, ${\mathbb{G}}=\{{\mathcal{G}}_{t}\}_{t\in[0,T]}$ is the filtration generated by $W^{0}$, ${\mathcal{G}}={\mathcal{G}}_{T}=\sigma(W^{0}_{u},u\leq T)$, and $\sigma_{t}$ and $\sigma_{t}^{0}$ are appropriate functions to ensure the well-definedness of the above SDE.
+By the classical Itô’s formula, the dynamic of $g(X_{t})$ for a smooth function $g$ is given by
 ${\mathrm{d}}g(X_{t})=b_{t}g^{\prime}(X_{t})+\frac{1}{2}(\sigma_{t}^{2}+(\sigma^{0}_{t})^{2})g^{\prime\prime}(X_{t}){\mathrm{d}}t+\sigma_{t}g^{\prime}(X_{t}){\mathrm{d}}W_{t}+\sigma^{0}_{t}g^{\prime}(X_{t}){\mathrm{d}}W^{0}_{t},$
@@
 ${\mathrm{d}}Z_{t}={\mathbb{E}}\bigg{[}b_{t}g^{\prime}(X_{t})+\frac{1}{2}(\sigma_{t}^{2}+(\sigma^{0}_{t})^{2})g^{\prime\prime}(X_{t})\Big{|}{\mathcal{G}}\bigg{]}{\mathrm{d}}t+{\mathbb{E}}\big{[}\sigma^{0}_{t}g^{\prime}(X_{t})|{\mathcal{G}}\big{]}{\mathrm{d}}W^{0}_{t}.$ (3.1)
-The above equation follows essentially from the fact that
-${\mathbb{E}}\bigg{[}\int_{0}^{t}\eta_{s}{\mathrm{d}}W_{s}\Big{|}{\mathcal{G}}\bigg{]}=0,\qquad{\mathbb{E}}\bigg{[}\int_{0}^{t}\eta_{s}{\mathrm{d}}W^{0}_{s}\Big{|}{\mathcal{G}}\bigg{]}=\int_{0}^{t}{\mathbb{E}}[\eta_{s}|{\mathcal{G}}]{\mathrm{d}}W^{0}_{s},\text{ for all }t\in[0,T]$ (3.2)
-see [34, Lemma B.1] for details. Indeed, from (3.1) one can calculate the quadratic variation of $Z_{t}$ to derive the Itô’s formula for $f(Z_{t})$:
+The above equation follows essentially from the fact that for progressively measurable process $(\eta_{t})_{t\in[0,T]},$
+${\mathbb{E}}\bigg{[}\int_{0}^{t}\eta_{s}{\mathrm{d}}W_{s}\Big{|}{\mathcal{G}}\bigg{]}=0,\qquad{\mathbb{E}}\bigg{[}\int_{0}^{t}\eta_{s}{\mathrm{d}}W^{0}_{s}\Big{|}{\mathcal{G}}\bigg{]}=\int_{0}^{t}{\mathbb{E}}[\eta_{s}|{\mathcal{G}}]{\mathrm{d}}W^{0}_{s},\text{ for all }t\in[0,T].$ (3.2)
+(See [47, Lemma B.1] for more details).
+From (3.1) one can calculate the quadratic variation of $Z_{t}$:
 ${\mathrm{d}}\langle Z\rangle_{t}={\mathbb{E}}\big{[}\sigma^{0}_{t}g^{\prime}(X_{t})|{\mathcal{G}}\big{]}^{2}{\mathrm{d}}t={\mathbb{E}}[\sigma^{0}_{t}g^{\prime}(X_{t})\widetilde{\sigma}^{0}_{t}g^{\prime}(\widetilde{X}_{t})|{\mathcal{G}}]{\mathrm{d}}t,$
-where $\widetilde{\sigma}_{t}^{0},\widetilde{X}_{t}$ are conditional respective independent copy of $\sigma^{0}_{t}$ and $X_{t}$ given ${\mathcal{G}}$, which is formally introduced in [11].
-Under these expressions, one can derive Itô’s formula for $f(Z_{t})$, which depends on the precise expression of
-$\int_{0}^{t}f^{\prime}(Z_{s}){\mathrm{d}}Z_{s}=$ $\int_{0}^{t}f^{\prime}(Y_{s})\bigg{(}{\mathbb{E}}\Big{[}b_{s}g^{\prime}(X_{s})+\frac{1}{2}(\sigma_{s}^{2}+(\sigma^{0}_{s})^{2})g^{\prime\prime}(X_{s})\Big{|}{\mathcal{G}}\Big{]}{\mathrm{d}}s+{\mathbb{E}}\big{[}\sigma^{0}_{t}g^{\prime}(X_{s})|{\mathcal{G}}\big{]}{\mathrm{d}}W^{0}_{s}\bigg{)}$
+where $\widetilde{\sigma}_{t}^{0},\widetilde{X}_{t}$ are respectively conditionally independent copies of $\sigma^{0}_{t}$ and $X_{t}$ given $\mathcal{G}$}, as formally introduced in [14].
+Moreover, one can derive for a smooth function $f$ the Itô’s formula for $f(Z_{t})$, which depends on the precise expression of
+$\int_{0}^{t}f^{\prime}(Z_{s}){\mathrm{d}}Z_{s}=$ $\int_{0}^{t}f^{\prime}(Z_{s})\bigg{(}{\mathbb{E}}\Big{[}b_{s}g^{\prime}(X_{s})+\frac{1}{2}(\sigma_{s}^{2}+(\sigma^{0}_{s})^{2})g^{\prime\prime}(X_{s})\Big{|}{\mathcal{G}}\Big{]}{\mathrm{d}}s+{\mathbb{E}}\big{[}\sigma^{0}_{t}g^{\prime}(X_{s})|{\mathcal{G}}\big{]}{\mathrm{d}}W^{0}_{s}\bigg{)}$
 $=\int_{0}^{t}\bigg{(}{\mathbb{E}}$ $\Big{[}f^{\prime}(Z_{s})\Big{(}b_{s}g^{\prime}(X_{s})+\frac{1}{2}(\sigma_{s}^{2}+(\sigma^{0}_{s})^{2})g^{\prime\prime}(X_{s})\Big{)}\Big{|}{\mathcal{G}}\Big{]}{\mathrm{d}}s+{\mathbb{E}}\big{[}f^{\prime}(Z_{s})\sigma^{0}_{t}g^{\prime}(X_{s})|{\mathcal{G}}\big{]}{\mathrm{d}}W^{0}_{s}\bigg{)}$
@@
 $=\frac{1}{2}\int_{0}^{t}{\mathbb{E}}\big{[}f^{\prime\prime}(Z_{s})\sigma^{0}_{s}g^{\prime}(X_{s})\widetilde{\sigma}^{0}_{s}g^{\prime}(\widetilde{X}_{s})\big{|}{\mathcal{G}}_{T}\big{]}{\mathrm{d}}s.$
-Now, the issue arises when one considers a general continuous semimartingale $X$ and a general sub $\sigma$-algebra ${\mathcal{G}}$, where the Itô’s formula takes the form:
+However, there is an issue when one considers a general continuous semimartingale $X$ and a general sub $\sigma$-algebra $\mathcal{G}$}, where the Itô’s formula takes the form:
 ${\mathrm{d}}g(X_{t})=g^{\prime}(X_{t}){\mathrm{d}}X_{t}+\frac{1}{2}g^{\prime\prime}(X_{t}){\mathrm{d}}\langle X\rangle_{t}.$ (3.3)
-In this case, in order to characterize the dynamics of $Z_{t}={\mathbb{E}}[g(X_{t})|{\mathcal{G}}]$, one needs an appropriate expression for ${\mathrm{d}}Z_{t}$, ${\mathrm{d}}\langle Z\rangle_{t}$ and more importantly for
-$\int_{0}^{t}f^{\prime}(Z_{s}){\mathrm{d}}Z_{s},\text{ and }\int_{0}^{t}f^{\prime\prime}(Z_{s}){\mathrm{d}}\langle Z\rangle_{s}.$ (3.4)
+Indeed, in order to characterize the dynamics of $Z_{t}={\mathbb{E}}[g(X_{t})|{\mathcal{G}}]$, one needs an appropriate expression for ${\mathrm{d}}Z_{t}$, ${\mathrm{d}}\langle Z\rangle_{t}$, and more importantly for
+[…]
 One naive guess of the extension of (3.2) to this general case is
-${\mathbb{E}}\bigg{[}\int_{0}^{t}\eta_{s}{\mathrm{d}}X_{s}\Big{|}{\mathcal{G}}\bigg{]}=\int_{0}^{t}{\mathbb{E}}[\eta_{s}|{\mathcal{G}}]{\mathrm{d}}X^{\mathcal{G}}_{s},$ (3.5)
-where $X^{\mathcal{G}}_{t}:={\mathbb{E}}[X_{t}|{\mathcal{G}}]$.
-Unfortunately this does not hold in general. For instance, take $X_{t}$ from (3.1) with $b_{t}=0$. By (3.2), the left hand side of (3.5) becomes
+${\mathbb{E}}\bigg{[}\int_{0}^{t}\eta_{s}{\mathrm{d}}X_{s}\Big{|}{\mathcal{G}}\bigg{]}=\int_{0}^{t}{\mathbb{E}}[\eta_{s}|{\mathcal{G}}]{\mathrm{d}}{\mathbb{E}}[X_{s}|{\mathcal{G}}].$ (3.5)
+Unfortunately this guess does not hold in general. For instance, take $X_{t}$ from (3.1) with $b_{t}=0$. By (3.2), the left hand side of (3.5) becomes
 ${\mathbb{E}}\bigg{[}\int_{0}^{t}\eta_{s}{\mathrm{d}}X_{s}\Big{|}{\mathcal{G}}\bigg{]}={\mathbb{E}}\bigg{[}\int_{0}^{t}\eta_{s}\big{(}\sigma_{s}{\mathrm{d}}W_{s}+\sigma^{0}_{s}{\mathrm{d}}W^{0}_{s}\big{)}\Big{|}{\mathcal{G}}\bigg{]}=\int_{0}^{t}{\mathbb{E}}\big{[}\eta_{s}\sigma^{0}_{s}\big{|}{\mathcal{G}}\big{]}{\mathrm{d}}W^{0}_{s},$
@@
 $\int_{0}^{t}{\mathbb{E}}[\eta_{s}|{\mathcal{G}}]{\mathrm{d}}X^{\mathcal{G}}_{s}=\int_{0}^{t}{\mathbb{E}}[\eta_{s}|{\mathcal{G}}]{\mathbb{E}}[\sigma^{0}_{s}|{\mathcal{G}}]{\mathrm{d}}W^{0}_{s},$
-and (3.5) holds if and only if $\eta_{s}$ and $\sigma_{s}^{0}$ are conditionally uncorrelated of ${\mathcal{G}}$.
-This is exactly why the concept of conditional independent copies of stochastic processes: instead of (3.2), let us consider the expression of the following form
-$\int_{0}^{t}f^{\prime}(Z_{s}){\mathrm{d}}Z_{s}$ $=\int_{0}^{t}f^{\prime}(Z_{s}){\mathrm{d}}{\mathbb{E}}\big{[}g(X_{s})|{\mathcal{G}}\big{]}=\int_{0}^{t}f^{\prime}(Z_{s}){\mathrm{d}}{\mathbb{E}}\big{[}g(X^{(1)}_{s})|{\mathcal{F}}\big{]}={\mathbb{E}}\bigg{[}\int_{0}^{t}f^{\prime}(Z_{s}){\mathrm{d}}g(X^{(1)}_{s})\Big{|}{\mathcal{F}}\bigg{]}$
-$={\mathbb{E}}\bigg{[}\int_{0}^{t}f^{\prime}(Z_{s})\Big{(}g^{\prime}(X^{(1)}_{s}){\mathrm{d}}X^{(1)}_{s}+\frac{1}{2}g^{\prime\prime}(X^{(1)}_{s}){\mathrm{d}}\langle X^{(1)}\rangle_{s}\Big{)}\Big{|}{\mathcal{F}}\bigg{]},$
+and (3.5) holds if and only if $\eta_{s}$ and $\sigma_{s}^{0}$ are conditionally uncorrelated, given $\mathcal{G}$}.
+[…]
+$\int_{0}^{t}f^{\prime}(Z_{s}){\mathrm{d}}Z_{s}$ $=\int_{0}^{t}f^{\prime}(Z_{s}){\mathrm{d}}{\mathbb{E}}\big{[}g(X_{s})|{\mathcal{G}}\big{]}=\int_{0}^{t}f^{\prime}(Z_{s}){\mathrm{d}}\widetilde{\mathbb{E}}\big{[}g(X^{(1)}_{s})\big{]}=\widetilde{\mathbb{E}}\bigg{[}\int_{0}^{t}f^{\prime}(Z_{s}){\mathrm{d}}g(X^{(1)}_{s})\bigg{]}$
+$=\widetilde{\mathbb{E}}\bigg{[}\int_{0}^{t}f^{\prime}(Z_{s})\Big{(}g^{\prime}(X^{(1)}_{s}){\mathrm{d}}X^{(1)}_{s}+\frac{1}{2}g^{\prime\prime}(X^{(1)}_{s}){\mathrm{d}}\langle X^{(1)}\rangle_{s}\Big{)}\bigg{]},$
 and
-$\int_{0}^{t}f^{\prime\prime}(Z_{s}){\mathrm{d}}\langle Z\rangle_{s}$ $=\int_{0}^{t}f^{\prime}(Z_{s}){\mathrm{d}}\langle{\mathbb{E}}\big{[}g(X_{\cdot})|{\mathcal{G}}\big{]}\rangle_{s}={\mathbb{E}}\bigg{[}\int_{0}^{t}f^{\prime}(Y_{s}){\mathrm{d}}\langle g(X^{(1)}_{\cdot}),g(X^{(2)}_{\cdot})\rangle_{s}|{\mathcal{F}}\bigg{]}$
-$={\mathbb{E}}\bigg{[}\int_{0}^{t}f^{\prime\prime}(Z_{s})g^{\prime}(X^{(1)}_{s})g^{\prime}(X^{(2)}_{s}){\mathrm{d}}\langle X^{(1)},X^{(2)}\rangle_{s}\Big{|}{\mathcal{F}}\bigg{]},$
-where $X^{(1)}$, $X^{(2)}$ are two conditional independent copies of $X$ give ${\mathcal{G}}$.
-With these expressions, one can analyze (3.4) for a general semimartingale $X$, as will be detailed in (2c) and (2d) of Theorem 3.3.
+$\int_{0}^{t}f^{\prime\prime}(Z_{s}){\mathrm{d}}\langle Z\rangle_{s}$ $=\int_{0}^{t}f^{\prime}(Z_{s}){\mathrm{d}}\langle{\mathbb{E}}\big{[}g(X_{\cdot})|{\mathcal{G}}\big{]}\rangle_{s}=\widetilde{\mathbb{E}}\bigg{[}\int_{0}^{t}f^{\prime}(Z_{s}){\mathrm{d}}\langle g(X^{(1)}_{\cdot}),g(X^{(2)}_{\cdot})\rangle_{s}\bigg{]}$
+$=\widetilde{\mathbb{E}}\bigg{[}\int_{0}^{t}f^{\prime\prime}(Z_{s})g^{\prime}(X^{(1)}_{s})g^{\prime}(X^{(2)}_{s}){\mathrm{d}}\langle X^{(1)},X^{(2)}\rangle_{s}\bigg{]},$
+where $\widetilde{\mathbb{E}}$ denotes the expectation on $X^{(1)}$ and $X^{(2)}$ while keeping the information that is $\mathcal{G}$}-measurable.
+With these expressions, one can analyze (3.4) for a general semimartingale $X$, as will be detailed in (1) and (2) of Theorem 3.5.
+Next, we will provide the construction of conditionally independent copies for the general case of a semimartingale $\bm{X}$ and an arbitrary sub $\sigma$-algebra $\mathcal{G}$} which is not necessarily associated with any random variable.
+As emphasized earlier, this construction is necessary for establishing Itô’s formula with the flow of conditional measures on semimartigales.
+3.2. Construction of conditionally independent copies of stochastic process
+Given a probability space $(\Omega,{\mathcal{F}},{\mathbb{P}})$, consider the extended probability space $\overline{\Omega}=\Omega^{n+1}$ defined by
+$\overline{\Omega}=\Omega^{n+1}=\{(\omega_{0},\omega_{1},\cdots,\omega_{n})|\omega_{0},\omega_{1},\cdots,\omega_{n}\in\Omega\},$ (3.6)
+and the extended $\sigma$-algebra $\overline{{\mathcal{F}}}$ by
+$\overline{{\mathcal{F}}}=\sigma\{A_{0}\times A_{1}\times\cdots\times A_{n}:A_{0},A_{1},\cdots,A_{n}\in{\mathcal{F}}\}.$ (3.7)
+To ensure that these copies are conditionally independent, let us define the probability measure $\overline{{\mathbb{P}}}$ as follows: for $A_{0},A_{1},\cdots,A_{n}\in{\mathcal{F}}$,
+$\overline{{\mathbb{P}}}\big{(}A_{0}\times A_{1}\times\cdots\times A_{n}\big{)}:={\mathbb{E}}\bigg{[}\mathbbm{1}_{A_{0}}\prod_{i=1}^{n}{\mathbb{P}}(A_{i}|{\mathcal{G}})\bigg{]}.$ (3.8)
+This measure is well defined since the set $\{A_{0}\times A_{1}\times\cdots\times A_{n}\}$ is a $\pi$-system.
+Finally, we define the completion of the above extended probability space and still use $(\overline{\Omega},\overline{{\mathcal{F}}},\overline{{\mathbb{P}}})$ to denote the completed probability space.
+Further, define $\widetilde{\mathcal{F}},\widetilde{\mathcal{G}}$ for the extension of $\mathcal{F}$} and $\mathcal{G}$} into this enlarged probability space as
+[…]
+Then we have:
+Theorem 3.1.
+Given a probability space $(\Omega,{\mathcal{F}},{\mathbb{P}})$ and a sub $\sigma$-algebra $\mathcal{G}$$\subset{\mathcal{F}}$, there exists an enlarged completed probability space $(\overline{\Omega},\overline{{\mathcal{F}}},\overline{{\mathbb{P}}})$ satisfying (3.6), (3.7) and (3.8); and for any random variable $R$ taking values in a standard Borel space $(E,\mathcal{E})$, there exist an $\widetilde{R}$ and $n$ conditionally independent copies $\{R^{(i)}\}_{i=1,2,..n}$ given $\widetilde{\mathcal{G}}$ such that
+$\widetilde{R}(\omega_{0},\omega_{1},\cdots,\omega_{n})=R(\omega_{0}),\;R^{(i)}(\omega_{0},\omega_{1},\cdots,\omega_{n})=R(\omega_{i}),$ (3.10)
+with $\widetilde{R}$ and $\{R^{(i)}\}_{i=1,2,..n}$ satisfying the following:
+  • Given $\widetilde{\mathcal{G}}$, $\widetilde{R}$ and $\big{\{}R^{(i)}\big{\}}_{i=1,2,..n}$ are conditionally independent.
+  • Given $\widetilde{\mathcal{F}}$, $\widetilde{R}$ and $\big{\{}R^{(i)}\big{\}}_{i=1,2,..n}$ are conditionally independent.
+  • $\big{(}R\big{|}{\mathcal{G}}\big{)}(\omega_{0})=Law$$\big{(}\widetilde{R}\big{|}\widetilde{\mathcal{G}}\big{)}(\overline{\omega})=$$\big{(}R^{(i)}\big{|}{\widetilde{\mathcal{G}}}\big{)}(\overline{\omega})=Law$$\big{(}R^{(i)}\big{|}{\widetilde{\mathcal{F}}}\big{)}(\overline{\omega})$, for all $i=1,2,..,n$, a.s.,
+for $\overline{\omega}=(\omega_{0},,\omega_{1},\cdots,\omega_{n})\in\overline{\Omega}$.
+Note that when $\mathcal{G}$} is a trivial $\sigma$-algebra $\{\emptyset,\Omega\}$, then the above theorem reduces to the standard construction of independent copies of stochastic processes.
+3.3. Construction of independent copies of semimartingale
+In the context of semimartingales, we can similarly define a corresponding version of the conditionally independent copy as follows.
+Take the completed filtered probability space $(\Omega,{\mathcal{F}},{\mathbb{F}}=({\mathcal{F}}_{t})_{t\in[0,T]},{\mathbb{P}})$ with a sub-filtration ${\mathbb{G}}=({\mathcal{G}}_{t})_{t\in[0,T]}\subset{\mathbb{F}}$ satisfying ${\mathcal{F}}_{t}\perp\!\!\!\perp{\mathcal{G}}_{T}|{\mathcal{G}}_{t}$, define a complete extended probability space $(\overline{\Omega},\overline{{\mathcal{F}}},\overline{{\mathbb{P}}})$ satisfying (3.6), (3.7) and (3.8) by setting the sub $\sigma$-algebra ${\mathcal{G}}={\mathcal{G}}_{T}$.
+Define $\overline{{\mathbb{F}}}:=(\overline{{\mathcal{F}}}_{t})_{t\in(0,T)}$ by
+$\overline{{\mathcal{F}}}_{t}:=\text{completion of }\sigma\{A_{0}\times A_{1}\times\cdots\times A_{n}:A_{0},A_{1},\cdots,A_{n}\in{\mathcal{F}}_{t}\},$ (3.11)
+and define $\widetilde{\mathbb{F}}:=(\widetilde{\mathcal{F}}_{t})_{t\in(0,T)}$ and $\widetilde{\mathbb{G}}:=(\widetilde{\mathcal{G}}_{t})_{t\in(0,T)}$ respectively for the extension of $\mathbb{F}$} and $\mathbb{G}$} in this enlarged probability space as
+[…]
+With this construction, we have
+Corollary 3.2.
+[…]
+$(\bm{X}^{\prime},\bm{Y}^{\prime})(\omega_{0},\omega_{1},\cdots,\omega_{n})=(\bm{X},\bm{Y})(\omega_{1}),\;(\bm{X}^{\prime\prime},\bm{Y}^{\prime\prime})(\omega_{0},\omega_{1},\cdots,\omega_{n})=(\bm{X},\bm{Y})(\omega_{2}),$
+$(\widetilde{\bm{X}},\widetilde{\bm{Y}})(\omega_{0},\omega_{1},\cdots,\omega_{n})=(\bm{X},\bm{Y})(\omega_{0}),$
+[…]
+$\begin{cases}\text{Law}\big{(}\widetilde{\bm{X}},\widetilde{\bm{Y}}\big{|}\widetilde{\mathcal{G}}\big{)}=\text{Law}\big{(}\bm{X}^{\prime},\bm{Y}^{\prime}\big{|}\widetilde{\mathcal{G}}\big{)}=\text{Law}\big{(}\bm{X}^{\prime},\bm{Y}^{\prime}\big{|}\widetilde{\mathcal{F}}\big{)}\\ \qquad=\text{Law}\big{(}\bm{X}^{\prime\prime},\bm{Y}^{\prime\prime}\big{|}\widetilde{\mathcal{G}}\big{)}=\text{Law}\big{(}\bm{X}^{\prime\prime},\bm{Y}^{\prime\prime}\big{|}\widetilde{\mathcal{F}}\big{)}\text{ a.s.,}\\ (\widetilde{\bm{X}},\widetilde{\bm{Y}}),(\bm{X}^{\prime},\bm{Y}^{\prime}),(\bm{X}^{\prime\prime},\bm{Y}^{\prime\prime})\text{ are conditionally independent, given }{\mathcal{G}}\ or\ {\mathcal{F}}.\par\par\end{cases}$ (3.13)
+Furthermore, if the sub $\sigma$-algebra ${\mathcal{G}}={\mathcal{G}}_{T}$, and ${\mathcal{F}}_{t}\perp\!\!\!\perp{\mathcal{G}}_{T}\big{|}{\mathcal{G}}_{t},$ for all $0\leq t\leq T$, then the sub-filtration $\widetilde{\mathcal{G}}\subset\widetilde{\mathcal{F}}$ satisfies $\widetilde{\mathcal{F}}_{t}\perp\!\!\!\perp\widetilde{\mathcal{G}}_{T}\big{|}\widetilde{\mathcal{G}}_{t},$ for all $0\leq t\leq T$.
+[…]
+$\widetilde{\mu}_{t}(\omega_{0},\omega_{1},\cdots,\omega_{n})=\text{Law}(\widetilde{\bm{X}_{t}}\big{|}\widetilde{\mathcal{G}}_{T})(\omega_{0},\omega_{1},\cdots,\omega_{n})=\text{Law}(\bm{X}_{t}\big{|}\widetilde{\mathcal{G}}_{T})(\omega_{0})=\mu_{t}(\omega_{0}),$
+which in turn allows for using $\mu_{t}$ for $\widetilde{\mu}_{t}$ without any ambiguity. Therefore, from now on we sometimes will simply use the original notation to denote the natural extensions.
 3.4. Examples
-To get some intuition for this notion of conditional independent copies of stochastic processes, let us see some examples when ${\mathcal{F}}$ and the sub $\sigma$-algebra ${\mathcal{G}}$ have some special structure.
+To get some intuition for this notion of conditionally independent copies of stochastic processes, let us see some special examples of $\mathcal{F}$} and the sub $\sigma$-algebra $\mathcal{G}$}.
 Example 1.
-Let $A,B$ are two independent random variables with ${\mathcal{F}}=\sigma(A,B)$ and ${\mathcal{G}}=\sigma(A)$.
-Then one can enlarge the probability space to include $A^{(i)}$ such that $A^{(i)}$ are independent and with the same distribution of $A$, and are independent of $B$.
-Then for any $X=f(A,B)$, the conditional independent copies of $X$ are $X^{(i)}=f(A^{(i)},B)$.
-Example 2. Let $X$ be the unique strong solution of the SDE
+Let $A,B$ be two independent random variables with ${\mathcal{F}}=\sigma(A,B)$ and ${\mathcal{G}}=\sigma(B)$.
+One can enlarge the probability space $(\overline{\Omega},\overline{{\mathcal{F}}},\overline{{\mathbb{P}}})$ to include $\{(A^{(i)},B^{(i)})\}_{i=1,2,..,n}$ satisfying (3.6), (3.7),(3.8) and (3.10).
+Moreover,
+  1. (1)
+    $\{(A^{(i)},B^{(i)})\}_{i=1,2,..,n}$ have the same distribution as $(A,B)$.
+  2. (2)
+    $B^{(i)}=B$ almost surely.
+  3. (3)
+    $\overline{{\mathcal{F}}}=\sigma\big{(}A,B,\{A^{(i)},B^{(i)}\}_{i=1,2,..,n}\big{)}=\sigma\big{(}A,B,\{A^{(i)}\}_{i=1,2,..,n}\big{)}$.
+  4. (4)
+    $\{A^{(i)}\}_{i=1,2,..,n}$ and $A,B$ are independent.
+Consequently, for any $X=f(A,B)$, the conditionally independent copies of $X$ are $X^{(i)}=f(A^{(i)},B^{(i)})=f(A^{(i)},B),\ \ i=1,\cdots,n.$
+Example 2. Let $X$ satisfy the following stochastic differential equation
 ${\mathrm{d}}X_{t}=b(X_{t}){\mathrm{d}}t+\sigma(X_{t}){\mathrm{d}}W_{t}+\sigma^{(0)}(X_{t}){\mathrm{d}}W^{(0)}_{t}$,
 where $W,W^{(0)}$ are two independent Brownian motions, and $b$, $\sigma$, and $\sigma^{0}$ satisfy appropriate continuity conditions.
-Let ${\mathcal{F}}_{t}=\sigma\big{(}\{W_{s},W^{(0}_{s}\}_{s\leq t}\big{)}$ and sub $\sigma$-algebra ${\mathcal{G}}=\sigma(W^{(0)})$, then one can enlarge the probability space to include independent Brownian motions $W^{(1)}$ and $X^{(1)}$, with $X^{(1)}$ the unique strong solution of
-${\mathrm{d}}X^{(1)}_{t}=b(X^{(1)}_{t}){\mathrm{d}}t+\sigma(X^{(1)}_{t}){\mathrm{d}}W^{(1)}_{t}+\sigma^{(0)}(X^{(1)}_{t}){\mathrm{d}}W^{(0)}_{t},$
-and $X^{(1)}$ is the conditional independent copies of $X$ given ${\mathcal{G}}$.
-There are more examples of conditional independent copies of stochastic processes in the literature of mean-field games.
-For instance, when defining a semimartingale by stochastic differential equations involving both idiosyncratic and common noise, conditional copies are formally formulated in [11] as $\Omega=\Omega_{0}\times\widetilde{\Omega}_{1}$, where the original probability space follows the structure $\Omega=\Omega_{0}\times\Omega_{1}$, and $\widetilde{\Omega}_{1}$ represents the independent copy of the probability space $\Omega_{1}$.
-[11, Theorem 4.14, Volume II]
-In the next section, we will provide the construction of conditional independent copies for the general case of a semimartingale $\bm{X}$ and an arbitrary sub $\sigma$-algebra ${\mathcal{G}}$ which is not necessarily associated with any probability space.
-As emphasized earlier, this construction is necessary for establishing Itô’s formula with the flow of conditional measures on semimartigales.
-3.3. Construction of conditional independent copies: general cases
-Given a probability space $(\Omega,{\mathcal{F}},{\mathbb{P}})$, let us consider the extended probability space $\overline{\Omega}=\Omega^{n+1}$ defined by
-$\overline{\Omega}=\Omega^{n+1}=\{(\omega_{0},\omega_{1},\cdots,\omega_{n})|\omega_{0},\omega_{1},\cdots,\omega_{n}\in\Omega\},$ (3.6)
-and the extended $\sigma$-algebra $\overline{\mathcal{F}}$
-¯$\overline{\mathcal{F}}=\sigma\{A_{0}\times A_{1}\times\cdots\times A_{n}:A_{0},A_{1},\cdots,A_{n}\in{\mathcal{F}}\}.$ (3.7)
-To ensure that these copies are conditional independent, let us define the probability measure $\overline{\mathbb{P}}$ as follows: for $A_{0},A_{1},\cdots,A_{n}\in{\mathcal{F}}$, define
-¯$\overline{\mathbb{P}}\big{(}A_{0}\times A_{1}\times\cdots\times A_{n}\big{)}:={\mathbb{E}}\bigg{[}\mathbbm{1}_{A_{0}}\prod_{i=1}^{n}{\mathbb{P}}(A_{i}|{\mathcal{G}})\bigg{]}.$ (3.8)
-This measure is properly defined since the set $\{A_{0}\times A_{1}\times\cdots\times A_{n}\}$ is a $\pi$-system.
-Finally, define ~$\widetilde{\mathcal{F}},\widetilde{\mathcal{G}}$ for the extension of ${\mathcal{F}}$ and ${\mathcal{G}}$ into this enlarged probability space as
-~:=$\widetilde{\mathcal{F}}:=\{A\times\Omega^{n}:A\in{\mathcal{F}}\},\quad\widetilde{\mathcal{G}}:=\{A\times\Omega^{n}:A\in{\mathcal{G}}\}.$ (3.9)
-Then we have:
-Theorem 3.1.
-Given a probability space $(\Omega,{\mathcal{F}},{\mathbb{P}})$ and a sub $\sigma$-algebra ${\mathcal{G}}\subset{\mathcal{F}}$.
-Then there exists uniquely an enlarged probability space $(\overline{\Omega},\overline{\mathcal{F}},{\mathbb{P}})$ satisfying (3.6), (3.7) and (3.8); and for any random variable $R$ in $(\Omega,{\mathbb{F}},{\mathbb{P}})$, there exist an $\widetilde{R}$ and $n$ conditional independent copies $\{R^{(i)}\}_{i=1,2,..n}$ by
-$\widetilde{R}(\omega_{0},\omega_{1},\cdots,\omega_{n})=R(\omega_{0}),\;R^{(i)}(\omega_{0},\omega_{1},\cdots,\omega_{n})=R(\omega_{i}),$ (3.10)
-for $\overline{\omega}=(\omega_{0},,\omega_{1},\cdots,\omega_{n})\in\overline{\Omega}$ and $\widetilde{R}$, $\{R^{(i)}\}_{i=1,2,..n}$ satisfying
-  • $\widetilde{R},\big{\{}R^{(i)}\big{\}}_{i=1,2,..n}$ are independent given $\widetilde{\mathcal{G}}$,
-  • $\widetilde{R},\big{\{}R^{(i)}\big{\}}_{i=1,2,..n}$ are independent given $\widetilde{\mathcal{F}}$,
-  • Law$\big{(}R\big{|}{\mathcal{G}}\big{)}(\omega_{0})=$Law$\big{(}\widetilde{R}\big{|}\widetilde{\mathcal{G}}\big{)}(\overline{\omega})=$Law$\big{(}R^{(i)}\big{|}{\widetilde{\mathcal{G}}}\big{)}(\overline{\omega})=$Law$\big{(}R^{(i)}\big{|}{\widetilde{\mathcal{F}}}\big{)}(\overline{\omega})$, for all $i=1,2,..,n$, a.s.
-with $\overline{\omega}=(\omega_{0},\cdots,\omega_{n})\in\overline{\Omega}$ and $\widetilde{\mathcal{F}}$, $\widetilde{\mathcal{G}}$ defined by (3.9).
-Note that when ${\mathcal{G}}$ is a trivial $\sigma$-algebra $\{\emptyset,\Omega\}$, then the above theorem reduces to the standard construction of independent copies.
-Its proof is given in Section 5.1.
-Now, in the context of semimartingales, there is a corresponding version of the conditional independent copy where the probability space is a filtered space and sub $\sigma$-algebra ${\mathcal{G}}={\mathcal{G}}_{T}$ comes from the sub filtration ${\mathbb{G}}=({\mathcal{G}}_{t})_{t\in[0,T]}$.
-Now define $\overline{\mathbb{F}}:=(\overline{\mathcal{F}}_{t})_{t\in(0,T)}$ by
-¯$\overline{\mathcal{F}}_{t}=\sigma\{A_{0}\times A_{1}\times\cdots\times A_{n}:A_{0},A_{1},\cdots,A_{n}\in{\mathcal{F}}_{t}\},$ (3.11)
-and further define $\widetilde{\mathbb{F}}=(\widetilde{\mathcal{F}}_{t})_{t\in(0,T)}$ and $\widetilde{\mathbb{G}}=(\widetilde{\mathcal{G}}_{t})_{t\in(0,T)}$ respectively for the extension of ${\mathbb{F}}$ and ${\mathbb{G}}$ into this enlarged probability space as
-~$\widetilde{\mathcal{F}}_{t}:=\{A\times\Omega^{n}:A\in{\mathcal{F}}_{t}\},\quad\widetilde{\mathcal{G}}_{t}:=\{A\times\Omega^{n}:A\in{\mathcal{G}}_{t}\}.$ (3.12)
-With this construction, we have
-Corollary 3.2.
-Given two ${\mathbb{R}}^{d}$-valued ${\mathbb{F}}$-semimartingales $\bm{X}=(\bm{X}_{t})_{t\in[0,T]}$ and $\bm{Y}=(\bm{X}_{t})_{t\in[0,T]}$ on a filtered probability space $(\Omega,{\mathcal{F}},{\mathbb{F}}=({\mathcal{F}}_{t})_{t\in[0,T]},{\mathbb{P}})$, and a sub $\sigma$-algebra ${\mathcal{G}}\subset{\mathcal{F}}$, there exists uniquely an enlarged probability space $(\overline{\Omega},\overline{\mathcal{F}},\overline{\mathbb{F}}=({\mathcal{F}}_{t})_{t\in[0,T]},\overline{\mathbb{P}})$ satisfying (3.6), (3.7), (3.8) and (3.11).
-Moreover, one can define two conditional independent copies $(\bm{X}^{\prime},\bm{Y}^{\prime})$, $(\bm{X}^{\prime\prime},\bm{Y}^{\prime\prime})$ of the random processes $(\bm{X},\bm{Y})$ by
-$(\bm{X}^{\prime},\bm{Y}^{\prime})(\omega_{0},\omega_{1},\cdots,\omega_{n})=(\bm{X},\bm{Y})(\omega_{1}),\;(\bm{X}^{\prime\prime},\bm{Y}^{\prime\prime})(\omega_{0},\omega_{1},\cdots,\omega_{n})=(\bm{X},\bm{Y})(\omega_{2}),$
-and $({\bm{X}},{\bm{Y}})$, $(\bm{X}^{\prime},\bm{Y}^{\prime})$ and $(\bm{X}^{\prime\prime},\bm{Y}^{\prime\prime})$ satisfying
-$\begin{cases}\text{Law}\big{(}{\bm{X}},{\bm{Y}}\big{|}{\mathcal{G}}\big{)}=\text{Law}\big{(}\bm{X}^{\prime},\bm{Y}^{\prime}\big{|}{\mathcal{G}}\big{)}=\text{Law}\big{(}\bm{X}^{\prime},\bm{Y}^{\prime}\big{|}{\mathcal{F}}\big{)}\\ \qquad=\text{Law}\big{(}\bm{X}^{\prime\prime},\bm{Y}^{\prime\prime}\big{|}{\mathcal{G}}\big{)}=\text{Law}\big{(}\bm{X}^{\prime\prime},\bm{Y}^{\prime\prime}\big{|}{\mathcal{F}}\big{)}\text{ a.s.}\\ ({\bm{X}},{\bm{Y}}),(\bm{X}^{\prime},\bm{Y}^{\prime}),(\bm{X}^{\prime\prime},\bm{Y}^{\prime\prime})\text{ are independent given }{\mathcal{G}},\text{and are also independent given }{\mathcal{F}}.\end{cases}$ (3.13)
-Furthermore, if the sub $\sigma$-algebra ${\mathcal{G}}={\mathcal{G}}_{T}$ from a sub filtration ${\mathbb{G}}=({\mathcal{G}}_{t})_{t\in[0,T]}\subset{\mathbb{F}}$ satisfies the compatibility conditions as in Assumption 2.5, then this compatibility condition holds for the natural extension of $\overline{\mathbb{F}}$ and $\overline{\mathbb{G}}$ defined by (3.12) in the extended probability space.
-That is, for each $0\leq t\leq T$, ~$\widetilde{\mathcal{F}}_{t}$ and $\widetilde{\mathcal{G}}_{T}$ are independent given $\overline{\mathcal{G}}_{t}$, written as ~$\widetilde{\mathcal{F}}_{t}\perp\!\!\!\perp\widetilde{\mathcal{G}}_{T}\big{|}\widetilde{\mathcal{G}}_{t}.$
-Note that Theorem 3.1 allows us to use the notation $(\bm{X},\bm{Y}),{\mathcal{F}}$, and ${\mathcal{G}}$ to respectively represent $(\widetilde{\bm{X}},\widetilde{\bm{Y}}),\widetilde{\mathcal{F}}$, and $\widetilde{\mathcal{G}}$ without any ambiguity, as is adopted in the above corollary.
-Furthermore, Corollary 3.2 enables simplifying the notation $\widetilde{\mathbb{F}}$ and $\widetilde{\mathbb{G}}$ respectively by ${\mathbb{F}}$ and ${\mathbb{G}}$ without any ambiguity.
-Moreover, if we define the conditional law in the enlarged probability space as $\widetilde{\mu}_{t}:=\text{Law}\big{(}\widetilde{\bm{X}}\big{|}\widetilde{\mathcal{G}}_{t}\big{)}$, then by the compatibility condition and Theorem 3.3, we have
-$\widetilde{\mu}_{t}(\omega_{0},\omega_{1},\cdots,\omega_{n})=\text{Law}(\widetilde{\bm{X}_{t}}\big{|}\widetilde{\mathcal{G}}_{T})(\omega_{0},\omega_{1},\cdots,\omega_{n})=\text{Law}(\bm{X}_{t}\big{|}\widetilde{\mathcal{G}}_{T})(\omega_{0})=\mu_{t}(\omega_{0}),$
-which in turn allows for using $\mu_{t}$ for $\widetilde{\mu}_{t}$ without any ambiguity.
-3.4. Properties of the conditional independent copies of semimartingales
-Next, we study properties of these conditional independent copies of semimartingales and explore the relation between the conditional expectation of their stochastic integrals and stochastic integrals with respect to the conditional law of semimartingales.
-For the sake of simplifying the notations, we will present the results in the one-dimensional case and their corresponding multi-dimensional cases can be adapted accordingly.
+Let ${\mathcal{F}}_{t}=\sigma\big{(}\{W_{s},W^{(0)}_{s}\}_{s\leq t}\big{)}$ and the sub-filtration ${\mathcal{G}}_{t}=\sigma\big{(}\{W^{(0)}_{s}\}_{s\leq t}\big{)}$, then one can enlarge the probability space to include independent Brownian motions $\big{\{}(X^{\prime},W^{\prime},{W^{(0)}}^{\prime})\big{\}}_{i=1,2,..,N}$ satisfying (3.6), (3.7),(3.8), (3.10) and (3.11).
+Moreover,
+  1. (1)
+    $X^{\prime}$ satisfies the following stochastic differential equation
+    ${\mathrm{d}}X^{\prime}_{t}=b(X^{\prime}_{t}){\mathrm{d}}t+\sigma(X^{\prime}_{t}){\mathrm{d}}W^{\prime}_{t}+\sigma^{(0)}(X^{\prime}_{t}){\mathrm{d}}(W^{(0)}_{t})^{\prime}$.
+  2. (2)
+    $W^{(0)}=(W^{(0)})^{\prime}$ almost surely.
+  3. (3)
+    $\overline{{\mathcal{F}}}_{t}=\sigma\big{(}\{W_{s},W^{(0)}_{s},W^{\prime}_{s},(W^{(0}_{s})^{\prime}\}_{s\leq t}\big{)}=\sigma\big{(}\{W_{s},W^{(0}_{s},W^{\prime}_{s}\}_{s\leq t}\big{)}$
+  4. (4)
+    $W,W^{\prime},W^{(0)}$ are independent Brownian motions.
+Consequently, $X^{\prime}$ is one conditionally independent copy of $X$ given $\mathcal{G}$}.
+There are more examples of conditionally independent copies of stochastic processes in the literature of mean-field games.
+For instance, when defining a semimartingale by stochastic differential equations involving both idiosyncratic and common noise, conditionally independent copies are formally formulated in [14] as $\overline{\Omega}=\Omega_{0}\times\Omega_{1}\times\widetilde{\Omega}_{1}$, where the original probability space follows the structure $\Omega=\Omega_{0}\times\Omega_{1}$, and $\widetilde{\Omega}_{1}$ represents the independent copy of the probability space $\Omega_{1}$ ([14, Theorem 4.14, Volume II]).
+3.5. Properties of the conditionally independent copies of semimartingales
+Next, we study properties of these conditionally independent copies of semimartingales, and explore the relation between the conditional expectation of their stochastic integrals and stochastic integrals with respect to the conditional law of semimartingales.
+For notational simplicity, we will present the results for the one-dimensional case and their corresponding multi-dimensional cases can be obtained with appropriate adaptations.
+First, by the compatibility condition and the classical results regarding filtration enlargement, we have:
+Proposition 3.3.
+Given an $\mathbb{R}$}-valued $\mathbb{F}$}-semimartingales $\{X_{t}\}_{t\in[0,T]}$ on a filtered probability space $(\Omega,{\mathcal{F}},{\mathbb{F}},{\mathbb{P}})$.
+Let $X^{\prime}$ and $X^{\prime\prime}$ be its conditionally independent copies defined in Proposition 3.2 in the enlarged filtered probability space $(\overline{\Omega},\overline{{\mathcal{F}}},\overline{{\mathbb{F}}}=(\overline{{\mathcal{F}}}_{t})_{t\in[0,T]},\overline{{\mathbb{P}}})$ satisfying (3.6), (3.8) and (3.11).
+Then,
+  1. (1)
+    $X$, $^{\prime}, X′′X^{\prime\prime}$ are $\overline{{\mathbb{F}}}$-semimartingales.
+  2. (2)
+    Suppose ${\mathbb{E}}[|X_{t}|]<\infty$ for all $t\in[0,T]$, and define
+    $X^{\mathcal{G}}_{t}:=\overline{{\mathbb{E}}}\big{[}X_{t}\big{|}{\mathcal{G}}_{T}\big{]}=\overline{{\mathbb{E}}}\big{[}X_{t}\big{|}{\mathcal{G}}_{t}\big{]}.$
+    1. (a)
+      If $\{X_{t}\}_{t\in[0,T]}$ is a finite variation process and ${\mathbb{E}}\big{[}\big{(}\int_{0}^{T}|{\mathrm{d}}X_{s}|\big{)}^{p}\big{]}<\infty$ for $p\geq 1$, then $\{X^{\mathcal{G}}_{t}\}_{t\in[0,T]}$ is also a finite variation process, and
+      $\overline{{\mathbb{E}}}\bigg{[}\bigg{(}\int_{0}^{T}\big{|}{\mathrm{d}}X^{\mathcal{G}}_{s}\big{|}\bigg{)}^{p}\bigg{]}<\infty.$
+    2. (b)
+      $\{X^{\mathcal{G}}_{t}\}_{t\in[0,T]}$ is a semimartingale with respect to $\overline{{\mathbb{F}}}$, $\mathbb{F}$}, and $\mathbb{G}$}.
+Remark 3.4.
+Proposition 3.3 shows that semimartingales remain semimartingales under the enlarged filtration $\overline{{\mathbb{F}}}$.
+This is closely related to the so called $H^{\prime}$-hypothesis where any $\mathbb{F}$}-martingale is an $\overline{{\mathbb{F}}}$-semimartingale with the enlargement filtration from $\mathbb{F}$} to $\overline{{\mathbb{F}}}$ (see for instance [3, 29]).
+Here we have obtained a slightly stronger result that any $\mathbb{F}$}-semimartingale $X$ and its conditionally independent copy remain $\overline{{\mathbb{F}}}$-semimartingales; meanwhile the $\mathbb{G}$}-semimartingale $X^{\mathcal{G}}$ (the projection taken with respect to the $\sigma$-algebra $\mathcal{G}$} as in [8]) remains $\overline{{\mathbb{F}}}$-semimartingale.
+Despite the apparent similarity, we are unable to locate exactly the same statement, and will provide the proof for completeness.
 Theorem 3.5.
 Give two $\mathbb{R}$}-valued $\mathbb{F}$}-semimartingales $\{X_{t}\}_{t\in[0,T]}$ and $\{Y_{t}\}_{t\in[0,T]}$ on a filtered probability space $(\Omega,{\mathcal{F}},{\mathbb{F}},{\mathbb{P}})$.
-Let $(X^{\prime},Y^{\prime})$ and $(X^{\prime\prime},Y^{\prime\prime})$ be their conditional independent copies defined in Corollary 3.2 in the enlarged filtered probability space $(\overline{\Omega},\overline{\mathcal{F}},\overline{\mathbb{F}}=({\mathcal{F}}_{t})_{t\in[0,T]},\overline{\mathbb{P}})$ satisfying (3.6), (3.8) and (3.11).
-Then,
+Let $(X^{\prime},Y^{\prime})$ and $(X^{\prime\prime},Y^{\prime\prime})$ be their conditionally independent copies defined in Proposition 3.2 in the enlarged filtered probability space $(\overline{\Omega},\overline{{\mathcal{F}}},\overline{{\mathbb{F}}}=({\mathcal{F}}_{t})_{t\in[0,T]},\overline{{\mathbb{P}}})$ satisfying (3.6), (3.8) and (3.11).
+Suppose ${\mathbb{E}}[|X_{t}|]<\infty$ and ${\mathbb{E}}[|Y_{t}|]<\infty$ for all $t\in[0,T]$.
   1. (1)
-    $X$, $Y$, $X^{\prime}$, $Y^{\prime}$, $X^{\prime\prime}$, $Y^{\prime\prime}$ are $\overline{\mathbb{F}}$-semimartingales.
+    If $\{Z_{t}\}_{t\in[0,T]}$ is an $\mathbb{R}$}-valued $\mathbb{F}$}-adapted càdlàg process in the filtered probability space $(\Omega,{\mathcal{F}},{\mathbb{F}},{\mathbb{P}})$, and $\|X\|_{\mathcal{H}_{p}}+\|Z\|_{\mathcal{H}_{q}}<\infty$ with $\frac{1}{p}+\frac{1}{q}=1$, then
+    $\int_{0}^{t}Z_{s-}{\mathrm{d}}X^{\mathcal{G}}_{s}=\overline{{\mathbb{E}}}\bigg{[}\int_{0}^{t}Z_{s-}{\mathrm{d}}X^{\prime}_{s}\bigg{|}{\mathcal{F}}\bigg{]}.$
   2. (2)
-    Suppose ${\mathbb{E}}[|X_{t}|]<\infty$ and ${\mathbb{E}}[|Y_{t}|]<\infty$ for all $t\in[0,T]$, and define
-    $X^{\mathcal{G}}_{t}:=\overline{\mathbb{E}}\big{[}X_{t}\big{|}{\mathcal{G}}_{T}\big{]}=\overline{\mathbb{E}}\big{[}X_{t}\big{|}{\mathcal{G}}_{t}\big{]},\;\;Y^{\mathcal{G}}_{t}:=\overline{\mathbb{E}}\big{[}Y_{t}\big{|}{\mathcal{G}}_{T}\big{]}=\overline{\mathbb{E}}\big{[}Y_{t}\big{|}{\mathcal{G}}_{t}\big{]}.$
-    1. (a)
-      If $\{X_{t}\}_{t\in[0,T]}$ is a finite variation process and ${\mathbb{E}}\big{[}\big{(}\int_{0}^{T}|{\mathrm{d}}X_{s}|\big{)}^{p}\big{]}<\infty,$ for $p\geq 1$, then $\{X^{\mathcal{G}}_{t}\}_{t\in[0,T]}$ is also a finite variation process and
-      $\overline{\mathbb{E}}\bigg{[}\bigg{(}\int_{0}^{T}\big{|}{\mathrm{d}}X^{\mathcal{G}}_{s}\big{|}\bigg{)}^{p}\bigg{]}<\infty.$
-    2. (b)
-      $\{X^{\mathcal{G}}_{t}\}_{t\in[0,T]}$ and $\{Y^{\mathcal{G}}_{t}\}_{t\in[0,T]}$ are semimartingales with respect to $\overline{\mathbb{F}}$, ${\mathbb{F}}$ and ${\mathbb{G}}$.
-    3. (c)
-      If $\{Z_{t}\}_{t\in[0,T]}$ is an ${\mathbb{R}}$-valued ${\mathbb{F}}$-adapted càdlàg process in the filtered probability space $(\Omega,{\mathcal{F}},{\mathbb{F}},{\mathbb{P}})$ and $\|X\|_{\mathcal{H}_{p}}+\|Z\|_{\mathcal{H}_{q}}<\infty$ with $\frac{1}{p}+\frac{1}{q}=1$, then
-      $\int_{0}^{t}Z_{s-}{\mathrm{d}}X^{\mathcal{G}}_{s}=\overline{\mathbb{E}}\bigg{[}\int_{0}^{t}Z_{s-}{\mathrm{d}}X^{\prime}_{s}\bigg{|}{\mathcal{F}}\bigg{]}.$
-    4. (d)
     If $\{Z_{t}\}_{t\in[0,T]}$ is an $\mathbb{R}$}-valued $\mathbb{F}$}-adapted càdlàg process in a filtered probability space $(\Omega,{\mathcal{F}},{\mathbb{F}},{\mathbb{P}})$ and $\|X\|_{\mathcal{H}_{p}}+\|Y\|_{\mathcal{H}_{q}}+\|Z\|_{\mathcal{H}_{r}}<\infty$ with $\frac{1}{p}+\frac{1}{q}+\frac{1}{r}=1$, then
@@
     $\int_{0}^{t}Z_{s}{\mathrm{d}}\big{[}X^{\mathcal{G}},Y\big{]}_{s}=\overline{{\mathbb{E}}}\bigg{[}\int_{0}^{t}Z_{s}{\mathrm{d}}\big{[}X^{\prime},Y\big{]}_{s}\bigg{|}{\mathcal{F}}\bigg{]}.$
-The first statement in Theorem 3.3 shows that semi-martingales remain semimartingales under the enlarged filtration $\overline{F}$.
-(2a) and (2b) suggest that under the conditional expectation operation, the finite variation property, the semimartingale property with respect to $\overline{\mathcal{F}}$ and ${\mathcal{F}}$ are preserved and the martingale property of ${\mathcal{G}}$ is unchanged because of the compatibility condition in Assumption 2.5.
-In particular, (2c) and (2d) are the key equations for the derivation of the Itô’s formula for cylindrical function; they ensure crucially the equivalence between the stochastic integral of the conditional expectation of a semimartingale and the conditional expectation of the Itô’s integral of its conditional independent copy, as well as the equivalence of the quadratic variation of the conditional expectation of a semimartingale and the conditional expectation of the quadratic variation of its two conditional independent copies.
-The proof of the theorem is presented in Section 5.2.
-4. Proof of Theorem 2.7
-Having established the properties of conditional independent copy of semimartingales, we are now ready to establish Itô’s formula for flow of conditional laws on semimratingales, through several steps.
-First, we will derive the formula for cylindrical functions; we will then show that cylindrical functions are dense in $C^{2,2}({\mathbb{R}}^{d}\times\mathcal{P}_{p}({\mathbb{R}}^{d}))$.
-Finally, appropriate localization techniques and applications of appropriate dominance convergence theorem finish the proof.
+Remark 3.6.
+Statements (1) and (2) in Theorem 3.5 are the key equations for deriving the Itô’s formula (Theorem 2.11.
+They ensure crucially a) the equivalence between the stochastic integral of the conditional expectation of a semimartingale and the conditional expectation of the Itô’s integral of its conditionally independent copy, and b) the equivalence between the quadratic variation of the conditional expectation of a semimartingale and the conditional expectation of the quadratic variation of its two conditionally independent copies.
+The proof of Theorem 3.5 is presented in Section 5.3.
+4. Proof of Main Result (Theorem 2.11)
+Having established necessary properties of conditionally independent copies of semimartingales, we are now ready to prove Theorem 2.11, the Itô’s formula for flow of conditional laws on semimartingales, through several steps.
+First, we will derive the formula for cylindrical functions; we will then show that cylindrical functions are dense in $C^{2,2}({\mathbb{R}}^{d}\times\mathcal{P}_{p}({\mathbb{R}}^{d}))$; finally, with appropriate localization technique and dominated convergence theorem, we will finish the proof.
 4.1. Step 1: Ito’s formula for cylindrical functions.
@@
 $\Phi(\mu,\bm{y})=f\big{(}\langle\mu,g^{(1)}\rangle,..\langle\mu,g^{(n)}\rangle,\bm{y}\big{)},$
-where $f\in C^{2}({\mathbb{R}}^{n+d})$ and $g\in C^{2}_{b}({\mathbb{R}}^{d})$.
-Next, we show that when restricted to the smaller set of functions $\Phi$, Theorem 2.7 holds.
-For semimartingales $\{\bm{X}_{t}\}_{t\in[0,T]}$ and $g^{(k)}\in C_{b}^{2}({\mathbb{R}}^{d})$, for $k=1,2,..,n$, by the classical Ito’s formula [37, Theorem 33, Chapter II], we have
+where $f\in C^{2}({\mathbb{R}}^{n+d})$ and $g=(g^{(1)},..,g^{(n)})$, with $g^{(i)}\in C^{2}_{b}({\mathbb{R}}^{d})$.
+We will first show that for smooth cylindrical functions $\Phi$, Theorem 2.11 holds.
+Indeed, for semimartingales $\{\bm{X}_{t}\}_{t\in[0,T]}$ and $g^{(i)}\in C_{b}^{2}({\mathbb{R}}^{d})$ with $i=1,2,..,n$, by the classical Itô’s formula [51, Theorem 33, Chapter II],
 $g^{(i)}(\bm{X}_{t})-g^{(i)}(\bm{X}_{0})=$ $\int_{0+}^{t}\nabla g^{(i)}(\bm{X}_{s-})\cdot{\mathrm{d}}\bm{X}^{c}_{s}+\frac{1}{2}\int_{0+}^{t}\nabla^{2}g^{(i)}(\bm{X}_{s-}):{\mathrm{d}}[\bm{X},\bm{X}]^{c}_{s}$
@@
 $\bm{Z}:=\Big{\{}\bm{Z}_{t}=\Big{(}\big{\langle}\mu_{t},g^{(1)}\big{\rangle},\big{\langle}\mu_{t},g^{(2)}\big{\rangle},..,\big{\langle}\mu_{t},g^{(n)}\big{\rangle}\Big{)}\Big{\}}_{t\in[0,T]},$
-and by (2b) of Theorem 3.3, $\bm{Z}$ is an $\overline{\mathbb{F}}$-semimartingale.
-Using the localization argument (which is detailed in Step 2), let us assume that $\bm{Y}_{t}$ is bounded.
-Then applying Itô’s formula to $f\big{(}\bm{Z},\bm{Y}\big{)}$ yields
+then by (2b) of Theorem 3.5, $\bm{Z}$ is an $\overline{{\mathbb{F}}}$-semimartingale.
+Let us for the moment assume that $\bm{Y}_{t}$ is bounded. Then applying Itô’s formula to $f\big{(}\bm{Z},\bm{Y}\big{)}$ yields
 $f(\bm{Z}_{t},\bm{Y}_{t})-f(\bm{Z}_{0},\bm{Y}_{0})$ (4.2)
@@
 $\quad+\sum_{0<s\leq t}\Big{\{}f(\bm{Y}_{s},\bm{Z}_{s})-f(\bm{Y}_{s-},\bm{Z}_{s-})\Big{\}}.$
-Let us compute $J_{1},J_{2},J_{3}$ separately.
-First, we note that $g^{(i)}$ are bounded, $\bm{Y}_{t}$ is bounded, $f\in C^{2}({\mathbb{R}}^{n+d})$, we have $\nabla_{\bm{z}}f(\bm{Z}_{s-},\bm{Y}_{s-})$, $\nabla^{2}_{\bm{z}\bm{y}}f(\bm{Z}_{s-},\bm{Y}_{s-})$ and $\nabla^{2}_{\bm{z}}f(\bm{Z}_{s-},\bm{Y}_{s-})$ are bounded and we confirm the conditions for applying (2c),(2d) of Theorem 3.3.
-Note that for the conditional copy $\bm{X}^{\prime}$, we have
+We will compute $J_{1},J_{2},$ and $J_{3}$ separately.
+First, note that both $g^{(i)}$ and $\bm{Y}_{t}$ are bounded, and $f\in C^{2}({\mathbb{R}}^{n+d})$, therefore $\nabla_{\bm{z}}f(\bm{Z}_{s-},\bm{Y}_{s-})$, $\nabla^{2}_{\bm{z}\bm{y}}f(\bm{Z}_{s-},\bm{Y}_{s-})$, and $\nabla^{2}_{\bm{z}}f(\bm{Z}_{s-},\bm{Y}_{s-})$ are bounded, and the conditions for applying (1) and (2) of Theorem 3.5 are satisfied.
+For the conditionally independent copy $\bm{X}^{\prime}$, we have
 $g^{(i)}(\bm{X}^{\prime}_{t})-g^{(i)}(\bm{X}^{\prime}_{0})=$ $\int_{0+}^{t}\nabla g^{(i)}(\bm{X}^{\prime}_{s-})\cdot{\mathrm{d}}(\bm{X}^{\prime})^{c}_{s}+\frac{1}{2}\int_{0+}^{t}\nabla^{2}g^{(i)}(\bm{X}_{s-}):{\mathrm{d}}[\bm{X}^{\prime},\bm{X}^{\prime}]^{c}_{s}$
-$\;\;+\sum_{0<s\leq t}\Big{\{}g^{(i)}(\bm{X}^{\prime}_{s})-g^{(i)}(\bm{X}^{\prime}_{s-})\Big{\}},\text{ for }t\in[0,T],$
-therefore, by (2c) of Theorem 3.3, ${\mathrm{d}}\bm{Z}_{s}={\mathrm{d}}\big{(}g^{(1)}(\bm{X}_{s}),\cdots,g^{(n)}(\bm{X}_{s})\big{)}^{\mathcal{G}})$, we have
+$\;\;+\sum_{0<s\leq t}\Big{\{}g^{(i)}(\bm{X}^{\prime}_{s})-g^{(i)}(\bm{X}^{\prime}_{s-})\Big{\}},\text{ for }t\in[0,T].$
+Therefore, by (1) of Theorem 3.5, ${\mathrm{d}}\bm{Z}_{s}={\mathrm{d}}\big{(}{\mathbb{E}}[g^{(1)}(\bm{X}_{s})|{\mathcal{G}}],\cdots,{\mathbb{E}}[g^{(n)}(\bm{X}_{s})|{\mathcal{G}}]\big{)}$, and
 $J_{1}=\overline{{\mathbb{E}}}\Bigg{[}\int_{0+}^{t}$ $\sum_{i=1}^{n}\partial_{z_{i}}f(\bm{Z}_{s-},\bm{Y}_{s-})\bigg{(}\nabla g^{(i)}(\bm{X}^{\prime}_{s-})\cdot{\mathrm{d}}(\bm{X}^{\prime})^{c}_{s}+\frac{1}{2}\nabla^{2}g^{(i)}(\bm{X}^{\prime}_{s-}):{\mathrm{d}}[\bm{X}^{\prime},\bm{X}^{\prime}]^{c}_{s}\bigg{)}$ (4.3)
-$+\sum_{0<s\leq t}\sum_{i=1}^{n}\partial_{z_{i}}f(\bm{Z}_{s-},\bm{Y}_{s-})\Big{\{}g^{(i)}(\bm{X}^{\prime}_{s})-g^{(i)}(\bm{X}^{\prime}_{s-})\Big{\}}\bigg{|}{\mathcal{F}}\Bigg{]}$
-For $J_{2}$, computing $[g^{(i)}(\bm{X}^{\prime}),g^{(j)}(\bm{X}^{\prime\prime})]_{t}$ for $1\leq i,j\leq n$, we have
-[…]
+$+\sum_{0<s\leq t}\sum_{i=1}^{n}\partial_{z_{i}}f(\bm{Z}_{s-},\bm{Y}_{s-})\Big{\{}g^{(i)}(\bm{X}^{\prime}_{s})-g^{(i)}(\bm{X}^{\prime}_{s-})\Big{\}}\bigg{|}{\mathcal{F}}\Bigg{]}.$
+For $J_{2}$, we see that
+$\big{[}$ $g^{(i)}(\bm{X}^{\prime}_{t}),g^{(j)}(\bm{X}^{\prime\prime}_{t})\big{]}_{t}=\big{[}(g^{(i)}(\bm{X}^{\prime}))^{c},(g^{j}(\bm{X}^{\prime\prime}))^{c}\big{]}_{t}+\sum_{0<s\leq t}\Delta g^{(i)}(\bm{X}^{\prime}_{s})\Delta g^{(j)}(\bm{X}^{\prime\prime}_{s})$
 $=\int_{0+}^{t}\Big{(}\nabla g^{(i)}(\bm{X}^{\prime}_{s-})^{\top}\nabla g^{(j)}(\bm{X}^{\prime\prime}_{s-})\Big{)}:{\mathrm{d}}[\bm{X}^{\prime},\bm{X}^{\prime\prime}]^{c}_{s}$
 $\qquad+\sum_{0<s\leq t}\Big{\{}\big{(}g^{(i)}(\bm{X}^{\prime}_{s})-g^{(i)}(\bm{X}^{\prime}_{s-})\big{)}\big{(}g^{(j)}(\bm{X}^{\prime\prime}_{s})-g^{(j)}(\bm{X}^{\prime\prime}_{s-})\big{)}\Big{\}},$
-and computing $[g^{(i)}(\bm{X}^{\prime}),Y^{(k)}]_{t}$ for $1\leq i\leq n$, $1\leq k\leq d$, we have
-[…]
+and
+$\big{[}$ $g^{(i)}(\bm{X}^{\prime}_{t}),Y^{(j)}\big{]}_{t}=\big{[}(g^{(i)}(\bm{X}^{\prime}))^{c},(Y^{j})^{c}\big{]}_{t}+\sum_{0<s\leq t}\Delta g^{(i)}(\bm{X}^{\prime}_{s})\Delta Y^{(j)}_{s}$
 $=\int_{0+}^{t}\nabla g^{(i)}(\bm{X}^{\prime}_{s-})\cdot{\mathrm{d}}[\bm{X}^{\prime},Y^{(j)}]^{c}_{s}+\sum_{0<s\leq t}\Big{\{}\big{(}g^{(i)}(\bm{X}^{\prime}_{s})-g^{(i)}(\bm{X}^{\prime}_{s-})\big{)}\big{(}Y^{(k)}_{s}-Y^{(k)}_{s-}\big{)}\Big{\}}$.
-Consequently, using (2d) of Theorem 3.3, we have
+Consequently, by (2) of Theorem 3.5,
 $J_{2}=\overline{{\mathbb{E}}}\Bigg{[}\int_{0+}^{t}$ $\bigg{(}\frac{1}{2}\sum_{i,j=1}^{n}\partial^{2}_{z_{i}z_{j}}f(\bm{Z}_{s-},\bm{Y}_{s-})\Big{(}\big{(}\nabla g^{(i)}(\bm{X}^{\prime}_{s-})\nabla g^{(j)}(\bm{X}^{\prime\prime}_{s-})^{\top}\big{)}:{\mathrm{d}}[\bm{X}^{\prime},\bm{X}^{\prime\prime}]^{c}_{s}\Big{)}$ (4.4)
@@
 $+\sum_{i=1}^{n}\sum_{k=1}^{d}\partial^{2}_{z_{i}y_{k}}f(\bm{Z}_{s-},\bm{Y}_{s-})\Big{\{}\big{(}g^{(i)}(\bm{X}^{\prime}_{s})-g^{(i)}(\bm{X}^{\prime}_{s-})\big{)}\big{(}Y^{(k)}_{s}-Y^{(k)}_{s-}\big{)}\Big{\}}\bigg{)}\bigg{|}{\mathcal{F}}\Bigg{]}$.
-Now, to deal with $J_{3}$, note that
+To deal with $J_{3}$, note that
 $\Delta Z^{(i)}_{t}$ $=\big{\langle}\mu_{t},g^{(i)}\big{\rangle}-\lim_{s\nearrow t}\big{\langle}\mu_{s},g^{(i)}\big{\rangle}=\overline{{\mathbb{E}}}\big{[}g^{(i)}(\bm{X}_{t})\big{|}{\mathcal{G}}_{T}\big{]}-\lim_{s\nearrow t}\overline{{\mathbb{E}}}\big{[}g^{(i)}(\bm{X}_{s})\big{|}{\mathcal{G}}_{T}\big{]}$
@@
 $=\overline{{\mathbb{E}}}\big{[}g^{(i)}(\bm{X}^{\prime}_{t})-g^{(i)}(\bm{X}^{\prime}_{t-})\big{|}{\mathcal{F}}\big{]},$
-where the third equation holds by the conditional dominated convergence theorem. Therefore, we can rewrite $J_{3}$ as
+where the third equality holds by the conditional dominated convergence theorem (Lemma 2.5). Therefore, we can rewrite $J_{3}$ as
 $J_{3}=\sum_{0<s\leq t}\bigg{\{}\overline{{\mathbb{E}}}\bigg{[}$ $\frac{1}{2}\sum_{i,j=1}^{n}\partial^{2}_{z_{i}z_{j}}f(\bm{Y}_{s-},\bm{Z}_{s-})\big{(}g^{(i)}(\bm{X}^{\prime}_{s})-g^{(i)}(\bm{X}^{\prime}_{s-})\big{)}\big{(}g^{(j)}(\bm{X}^{\prime\prime}_{s})-g^{(j)}(\bm{X}^{\prime\prime}_{s-})\big{)}$ (4.5)
@@
 $+\sum_{0<s\leq t}\overline{{\mathbb{E}}}\bigg{[}$ $\sum_{i=1}^{n}\sum_{k=1}^{d}\partial^{2}_{z_{i}y_{k}}f(\bm{Y}_{s-},\bm{Z}_{s-})\big{(}g^{(i)}(\bm{X}^{\prime}_{s})-g^{(i)}(\bm{X}^{\prime}_{s-})\big{)}\big{(}Y^{(k)}_{s}-Y^{(k)}_{s-}\big{)}\bigg{|}{\mathcal{F}}\Bigg{]}$,
-since $\partial^{2}_{z_{i}z_{j}}f(\bm{Y}_{s-},\bm{Z}_{s-})$, $\partial_{z_{i}}f(\bm{Y}_{s-},\bm{Z}_{s-})$, $\partial^{2}_{z_{i}y_{k}}f(\bm{Y}_{s-},\bm{Z}_{s-})$, $Y^{(k)}_{s}$ and $Y^{(k)}_{s-}$ are ${\mathcal{F}}$-measurable random variables for $s\in(0,t]$, $1\leq i,j\leq n$ and $1\leq k\leq d$.
-Recall that $\Phi(\mu,\bm{y})=f\big{(}\langle\mu,g^{(1)}\rangle,..\langle\mu,g^{(n)}\rangle,\bm{y}\big{)}$, simple calculation of linear derivatives gives us a version of the derivatives by
+since $\partial^{2}_{z_{i}z_{j}}f(\bm{Y}_{s-},\bm{Z}_{s-})$, $\partial_{z_{i}}f(\bm{Y}_{s-},\bm{Z}_{s-})$, $\partial^{2}_{z_{i}y_{k}}f(\bm{Y}_{s-},\bm{Z}_{s-})$, $Y^{(k)}_{s}$ and $Y^{(k)}_{s-}$ are $\mathcal{F}$}-measurable random variables for $s\in(0,t]$, $1\leq i,j\leq n$, and $1\leq k\leq d$.
+Recall that $\Phi(\mu,\bm{y})=f\big{(}\langle\mu,g^{(1)}\rangle,..\langle\mu,g^{(n)}\rangle,\bm{y}\big{)}$, simple calculation of linear derivatives yields a version of the derivatives by
 $\nabla_{\bm{y}}\Phi(\mu,\bm{y})=\nabla_{\bm{y}}f\big{(}\langle\mu,g^{(1)}\rangle,..\langle\mu,g^{(n)},\bm{y}\rangle\big{)},\quad\nabla^{2}_{0}\Phi(\mu,\bm{y})=\nabla^{2}_{\bm{y}\bm{y}}f\big{(}\langle\mu,g^{(1)}\rangle,..\langle\mu,g^{(n)}\rangle,\bm{y}\big{)},$
@@
 $\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}(\mu,\bm{y},\bm{x}_{1},\bm{x}_{2})=\sum_{i,j=1}^{n}\partial_{z_{i}z_{j}}f\big{(}\langle\mu,g^{(1)}\rangle,..\langle\mu,g^{(n)}\rangle,\bm{y}\big{)}g^{(i)}(\bm{x}_{1})g^{(j)}(\bm{x}_{2}).$
-Therefore, we can rewrite $J_{1}$ by
+Therefore, we can rewrite $J_{1}$ as
 $\;\overline{{\mathbb{E}}}\bigg{[}\int_{0+}^{t}\bigg{(}\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\cdot{\mathrm{d}}(\bm{X}^{\prime})^{c}_{s}+\frac{1}{2}\nabla^{2}_{\bm{x}_{1}\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s},\bm{X}^{\prime}_{s}\big{)}:{\mathrm{d}}[\bm{X}^{\prime},\bm{X}^{\prime}]^{c}_{t}\bigg{)}$ (4.6)
 $+\sum_{0<s\leq t}\bigg{(}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s}\big{)}-\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\bigg{)}\bigg{|}{\mathcal{F}}\bigg{]},$
-$J_{2}$ by
+$J_{2}$ as
 $\overline{{\mathbb{E}}}\bigg{[}$ $\int_{0+}^{t}\bigg{(}\frac{1}{2}\nabla^{2}_{\bm{x}_{1}\bm{x}_{2}}\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{Y}_{s},\bm{X}^{\prime}_{s},\bm{X}^{\prime\prime}_{s}\big{)}:{\mathrm{d}}[\bm{X}^{\prime},\bm{X}^{\prime\prime}]^{c}_{t}+\nabla^{2}_{\bm{x}_{1}\bm{y}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}:{\mathrm{d}}[\bm{X}^{\prime},\bm{Y}]^{c}_{s}\bigg{)}$ (4.7)
@@
 $\qquad\qquad\qquad+\bigg{(}\nabla_{\bm{y}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s}\big{)}-\nabla_{\bm{y}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\bigg{)}\cdot\Delta\bm{Y}_{s}\bigg{)}\bigg{|}{\mathcal{F}}\bigg{]},$
-and $J_{3}$ by
+and $J_{3}$ as
 $\sum_{0<s\leq t}\bigg{(}\frac{1}{2}\overline{{\mathbb{E}}}$ $\bigg{[}\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s},\bm{X}^{\prime\prime}_{s}\big{)}-\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-},\bm{X}^{\prime\prime}_{s}\big{)}$ (4.8)
@@
 $+$ $\overline{{\mathbb{E}}}\bigg{[}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s}\big{)}-\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\bigg{|}{\mathcal{F}}\bigg{]}\bigg{)}.$
-Using the argument at the end of [24, Section 3.2.1], one can cancel out redundant terms in $J_{2}$ and $J_{3}$ to get the first summation in the right hand side of (2.3).
-Therefore, combining (4.6), (4.7), and (4.8) and plugging them back into (4.2) complete the proof.
+Using the argument at the end of [31, Section 3.2.1], one can cancel out redundant terms in $J_{2}$ and $J_{3}$ to get the first summation in the right hand side of (2.3).
+Therefore, combining (4.6), (4.7), and (4.8) and plugging them back into (4.2) completes the proof.
 4.2. Step 2: localization argument.
-We claim that for a sequence of ${\mathbb{F}}$ stopping time $\{\tau_{n}\}_{n\in{\mathbb{N}}}$ such that $\tau_{n}\to T$, in order to prove that (2.3) holds for $\Phi\in C^{2,2}({\mathcal{P}}_{p}({\mathbb{R}}^{d})\times{\mathbb{R}}^{d})$ with the process $(\mu,\bm{Y})=(\mu_{t\wedge\tau_{n}},\bm{Y}_{t\wedge\tau_{n}})_{t\in[0,T]}$, it suffices to prove that (2.3) holds for $\Phi$ with the truncated process $(\mu^{(\tau_{n})},\bm{Y}^{(\tau_{n})}):=(\mu_{t\wedge\tau_{n}},\bm{Y}_{t\wedge\tau_{n}})_{t\in[0,T]}$.
+Take a sequence of $\mathbb{F}$}-stopping time $\{\tau_{n}\}_{n\in{\mathbb{N}}}$ such that $\tau_{n}\to T$, in order to prove (2.3) for $\Phi\in C^{2,2}({\mathcal{P}}_{p}({\mathbb{R}}^{d})\times{\mathbb{R}}^{d})$ with the process $(\mu,\bm{Y})=(\mu_{t\wedge\tau_{n}},\bm{Y}_{t\wedge\tau_{n}})_{t\in[0,T]}$, it suffices to prove (2.3) for $\Phi$ with the truncated process $(\mu^{(\tau_{n})},\bm{Y}^{(\tau_{n})}):=(\mu_{t\wedge\tau_{n}},\bm{Y}_{t\wedge\tau_{n}})_{t\in[0,T]}$.
 Indeed, suppose (2.3) holds for the truncated process, for $t<T$.
-Let $n\to\infty$, it is not hard to see that the left hand side of (2.3) converges and last three terms of the right hand side of (2.3) converge to the desired equation.
-Now, let us split the conditional expectation term of the right hand side of (2.3) into several terms.
-First, in order to prove that
+Let $n\to\infty$, it is not hard to see that the left hand side of (2.3) converges and the last three terms of the right hand side of (2.3) converge to the desired equation.
+Now, let us split the conditional expectation term of the right hand side of (2.3) into several terms. First, in order to prove that
 $\overline{{\mathbb{E}}}\bigg{[}\int_{0+}^{t\wedge\tau_{n}}\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\cdot{\mathrm{d}}(\bm{X}^{\prime})^{c}_{s}\bigg{|}{\mathcal{F}}\bigg{]}\to\overline{{\mathbb{E}}}\bigg{[}\int_{0+}^{t}\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\cdot{\mathrm{d}}(\bm{X}^{\prime})^{c}_{s}\bigg{|}{\mathcal{F}}\bigg{]},$ (4.9)
-almost surely as $n\to\infty$, we separate it by,
+a.s. as $n\to\infty$, we separate it by,
 $\int_{0+}^{t\wedge\tau_{n}}$ $\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\cdot{\mathrm{d}}(\bm{X}^{\prime})^{c}_{s}=$ (4.10)
-$\int_{0+}^{t\wedge\tau_{n}}\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\cdot{\mathrm{d}}\bm{X}^{\prime}_{s}-\sum_{0<s\leq t\wedge\tau_{n}}\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\cdot\Delta\bm{X}^{\prime}_{s}$
+$\int_{0+}^{t\wedge\tau_{n}}\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\cdot{\mathrm{d}}\bm{X}^{\prime}_{s}-\sum_{0<s\leq t\wedge\tau_{n}}\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\cdot\Delta\bm{X}^{\prime}_{s}.$
 Note that the first term is uniformly bounded by
@@
 $\leq c_{p}c_{1}C\big{(}\big{\|}\bm{X}\big{\|}^{p}_{{\mathcal{H}}_{p}}+\big{\|}\bm{Y}\big{\|}^{p}_{{\mathcal{H}}_{p}}+1\big{)}\big{\|}\bm{X}\big{\|}_{{\mathcal{H}}_{p}}<\infty,$
-where $p^{\prime}=\frac{p}{p-1}$, $c_{1},c_{p}$ are constants in Proposition 2.2 and $C$ is a generic constant that may vary line by line.
-The first inequality is due to Proposition 2.2, the second inequality holds because of Proposition 2.3, the third inequality holds by $\Phi\in C^{2,2}({\mathcal{P}}_{p}({\mathbb{R}}^{d})\times{\mathbb{R}}^{d})$, and the last one holds because $\frac{1}{p}+\frac{1}{p^{\prime}}=1$.
+where $p^{\prime}=\frac{p}{p-1}$, $c_{1}$ and $c_{p}$ are constants in Proposition 2.2, and $C$ is a generic constant that may vary line by line.
+The first inequality is due to Proposition 2.2, the second inequality is by Proposition 2.3, the third inequality is from $\Phi\in C^{2,2}({\mathcal{P}}_{p}({\mathbb{R}}^{d})\times{\mathbb{R}}^{d})$, and the last one holds because $\frac{1}{p}+\frac{1}{p^{\prime}}=1$.
 Meanwhile, the second term of the right hand side of (4.10) is bounded by
@@
 $\leq C\bigg{(}1+\sup_{0\leq s\leq T}|\bm{X}_{s}|^{p}+\sup_{0\leq s\leq T}|\bm{Y}_{s}|^{p}+\bigg{(}\sum_{0<s\leq T}|\Delta\bm{X}^{\prime}_{s}|\bigg{)}^{p}\bigg{)},$
-where $C$ is a generic constant that may vary line by line, and the right hand side of the above inequality is integrable due to the fact that $\|\bm{X}\|_{{\mathcal{S}}_{p}}\leq c_{p}\|\bm{X}\|_{{\mathcal{H}}_{p}}<\infty$, $\|\bm{Y}\|_{{\mathcal{S}}_{p}}\leq c_{p}\|\bm{Y}\|_{{\mathcal{H}}_{p}}<\infty$ and the assumption on $\bm{X}$.
-Therefore, using the conditional dominated theorem, we conclude (4.9).
+where $C$ is a generic constant, and the right hand side of the above inequality is integrable because $\|\bm{X}\|_{{\mathcal{S}}_{p}}\leq c_{p}\|\bm{X}\|_{{\mathcal{H}}_{p}}<\infty$, $\|\bm{Y}\|_{{\mathcal{S}}_{p}}\leq c_{p}\|\bm{Y}\|_{{\mathcal{H}}_{p}}<\infty$, and by the assumption of $\bm{X}$.
+Therefore, by the conditional dominated convergence theorem (see Lemma 2.5), we conclude (4.9).
 Next, to prove that
@@
 $\to\overline{{\mathbb{E}}}\bigg{[}\int_{0+}^{t}\nabla^{2}_{\bm{x}_{1}\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}:{\mathrm{d}}[\bm{X}^{\prime},\bm{X}^{\prime}]^{c}_{s}\bigg{|}{\mathcal{F}}\bigg{]},$
-almost surely as $n\to\infty$, we note that
+a.s. as $n\to\infty$, note that
 $\int_{0+}^{t\wedge\tau_{n}}\nabla^{2}_{\bm{x}_{1}\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}:{\mathrm{d}}[\bm{X}^{\prime},\bm{X}^{\prime}]^{c}_{s}$
 is bounded by
-$\sup_{0\leq s\leq T}\bigg{|}\int_{0+}^{s}\nabla^{2}_{\bm{x}_{1}\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{u-},\bm{Y}_{u-},\bm{X}^{\prime}_{u-}\big{)}:{\mathrm{d}}[\bm{X}^{\prime},\bm{X}^{\prime}]^{c}_{u}\bigg{|}\leq\sup_{0\leq s\leq T}\bigg{|}\nabla^{2}_{\bm{x}_{1}\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\bigg{|}\cdot[\bm{X}^{\prime},\bm{X}^{\prime}]^{c}_{T}$
+$\sup_{0\leq s\leq T}\bigg{|}\int_{0+}^{s}\nabla^{2}_{\bm{x}_{1}\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{u-},\bm{Y}_{u-},\bm{X}^{\prime}_{u-}\big{)}:{\mathrm{d}}[\bm{X}^{\prime},\bm{X}^{\prime}]^{c}_{u}\bigg{|}\leq\sup_{0\leq s\leq T}\bigg{|}\nabla^{2}_{\bm{x}_{1}\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\bigg{|}\cdot[\bm{X}^{\prime},\bm{X}^{\prime}]^{c}_{T},$
 and
@@
 where $c_{p^{\prime\prime}}$ is the constant in Proposition 2.2, $C$ is a generic constant that may vary from line to line and $p^{\prime\prime}=\frac{p-2}{2p}$.
-Thus, (4.12) holds due to the conditional dominated theorem, we conclude .
+Thus, (4.12) holds by Lemma 2.5.
 Similar arguments holds also for the convergence of $\int_{0+}^{t\wedge\tau_{n}}\nabla^{2}_{\bm{x}_{1}\bm{x}_{2}}\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-},\bm{X}^{\prime\prime}_{s-}\big{)}:{\mathrm{d}}[\bm{X}^{\prime},\bm{X}^{\prime\prime}]_{s}$ and $\int_{0+}^{t\wedge\tau_{n}}\nabla^{2}_{\bm{x}_{1}\bm{y}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}:{\mathrm{d}}[\bm{X}^{\prime},\bm{Y}]_{s}$.
@@
 $-\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s},\bm{X}^{\prime\prime}_{s-}\big{)}+\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-},\bm{X}^{\prime\prime}_{s-}\big{)}\bigg{)}\mathbbm{1}_{\{\mu_{s}=\mu_{s-}\}}.$
-With simple application of the fundamental theorem of calculus for the linear derivative, we have
+Applying the fundamental theorem of calculus for the linear derivative, we have
 $\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s},\bm{X}^{\prime\prime}_{s}\big{)}-\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-},\bm{X}^{\prime\prime}_{s}\big{)}$
@@
 $\leq C\bigg{(}\sup_{0\leq t\leq T}\Big{(}1+|\bm{X}^{\prime}_{t}|^{p}+|\bm{X}^{\prime\prime}_{t}|^{p}\Big{)}+\sum_{0<s\leq T}|\Delta\bm{X}^{\prime}_{s}|^{p}+\sum_{0<s\leq T}|\Delta\bm{X}^{\prime\prime}_{s}|^{p}\bigg{)},$
-where $C$ is a generic constant that may vary from line by line.
-Therefore, the right hand side of the above inequality is integrable and the convergence of the conditional expectation of (4.13) is ensured by conditional dominated convergence theorem.
-Similar arguments can be applied to deduce the convergence of the conditional expectation of $\big{(}\nabla_{\bm{y}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s}\big{)}-\nabla_{\bm{y}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\big{)}\cdot\Delta\bm{Y}_{s}$ and $\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s}\big{)}-\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}$, s and thus we conclude that (2.3) holds for the original process $(\mu,\bm{Y})$ for $0\leq t<T$.
-For the case of $t=T$, taking $n\to\infty$, left hand side equals $\Phi(\mu_{T-},\bm{Y}_{T-})-\Phi(\mu_{0},\bm{Y}_{0})$ while all terms at the right hand sides is evaluated at $T$ except the last term being
+where $C$ is a generic constant.
+The right hand side of the above inequality is integrable and the convergence of the conditional expectation of (4.13) is ensured by Lemma 2.5.
+Similar arguments can be applied to deduce the convergence of the conditional expectation of $\big{(}\nabla_{\bm{y}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s}\big{)}-\nabla_{\bm{y}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\big{)}\cdot\Delta\bm{Y}_{s}$ and $\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s}\big{)}-\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}$, and thus we conclude that (2.3) holds for the original process $(\mu,\bm{Y})$ for $0\leq t<T$.
+For the case of $t=T$, taking $n\to\infty$, then the left hand side equals $\Phi(\mu_{T-},\bm{Y}_{T-})-\Phi(\mu_{0},\bm{Y}_{0})$ while all terms on the right hand side are evaluated at $T$ except for the last term which is
 $\sum_{0<s<t}\Big{(}\Phi(\mu_{s},\bm{Y}_{s})-\Phi(\mu_{s-},\bm{Y}_{s-})\Big{)},$
@@
 4.3. Step 3: cylindrical functions are dense in $C^{2,2}({\mathbb{R}}^{d}\times\mathcal{P}_{p}({\mathbb{R}}^{d}))$.
-The following result slightly generalizes that in [16, Theorem 4.4, Theorem 4.10, Section 4] as it includes the convergences of the derivatives of the linear derivative.
+The following result slightly generalizes that in [21, Theorem 4.4, Theorem 4.10, Section 4] as it also includes the convergences of the associated linear derivatives.
 Proposition 4.2.
-Let $\Phi\in C^{2,2}({\mathcal{P}}_{p}({\mathbb{R}}^{d})\times{\mathbb{R}}^{d})$, then there exist a sequence of $C^{2,2}$ cylinder functions $\{f_{n}\}_{n\in{\mathbb{N}}^{+}}$ such that one has the point-wise convergence
+Let $\Phi\in C^{2,2}({\mathcal{P}}_{p}({\mathbb{R}}^{d})\times{\mathbb{R}}^{d})$, then there exists a sequence of $C^{2,2}$ cylinder functions $\{f_{n}\}_{n\in{\mathbb{N}}^{+}}$ such that one has the point-wise convergence
 $\bigg{(}f_{n},\nabla_{\bm{y}}f_{n}$ $,\nabla_{\bm{y}}^{2}f_{n},\frac{\delta f_{n}}{\delta\mu},\nabla_{\bm{x}_{1}}\frac{\delta f_{n}}{\delta\mu},\nabla^{2}_{\bm{x}_{1}}\frac{\delta f_{n}}{\delta\mu},\nabla_{\bm{x}_{1}\bm{y}}\frac{\delta f_{n}}{\delta\mu},\nabla_{\bm{x}_{1}\bm{x}_{2}}\frac{\delta^{2}f_{n}}{(\delta\mu)^{2}}\bigg{)}$ (4.14)
 $\to\bigg{(}\Phi,\nabla_{\bm{y}}\Phi,\nabla_{\bm{y}}^{2}\Phi,\frac{\delta\Phi}{\delta\mu},\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu},\nabla^{2}_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu},\nabla_{\bm{x}_{1}\bm{y}}\frac{\delta\Phi}{\delta\mu},\nabla_{\bm{x}_{1}\bm{x}_{2}}\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\bigg{)}$
-as $n\to\infty$, for all $\bm{x}_{1},\bm{x}_{2},\bm{y},\in{\mathbb{R}}^{d}$, $\mu\in{\mathcal{P}}_{p}({\mathbb{R}}^{d})$. Moreover, there exists a constant $c>0$ such that
+as $n\to\infty$, for all $\bm{x}_{1},\bm{x}_{2},\bm{y}\in{\mathbb{R}}^{d}$, $\mu\in{\mathcal{P}}_{p}({\mathbb{R}}^{d})$. Moreover, there exists a constant $c>0$ such that
 $|\nabla_{\bm{y}}f_{n}|(\mu,\bm{y})\leq c(1+|\bm{y}|^{p-1}),\qquad|\nabla^{2}_{\bm{y}}f_{n}|(\mu,\bm{y})$ $\leq c(1+|\bm{y}|^{p-2}),$ (4.15)
 $\bigg{|}\frac{\delta f_{n}}{\delta\mu}\bigg{|}\leq c\big{(}1+|\bm{x}_{1}|^{p}+|\bm{y}|^{p}\big{)},\qquad\bigg{|}\nabla_{\bm{x}_{1}}\frac{\delta f_{n}}{\delta\mu}\bigg{|}$ $\leq c\big{(}1+|\bm{x}_{1}|^{p-1}+|\bm{y}|^{p-1}\big{)},$
-$\bigg{|}\nabla^{2}_{\bm{x}_{1}}\frac{\delta f_{n}}{\delta\mu}\bigg{|}\leq c_{K}\big{(}1+|\bm{x}_{1}|^{p-2}+|\bm{y}|^{p-2}\big{)},\qquad\bigg{|}\nabla^{2}_{\bm{x}_{1}\bm{y}}\frac{\delta f_{n}}{\delta\mu}\bigg{|}$ $\leq c\big{(}1+|\bm{x}_{1}|^{{p-2}}+|\bm{y}|^{p-2}\big{)},$
+$\bigg{|}\nabla^{2}_{\bm{x}_{1}}\frac{\delta f_{n}}{\delta\mu}\bigg{|}\leq c\big{(}1+|\bm{x}_{1}|^{p-2}+|\bm{y}|^{p-2}\big{)},\qquad\bigg{|}\nabla^{2}_{\bm{x}_{1}\bm{y}}\frac{\delta f_{n}}{\delta\mu}\bigg{|}$ $\leq c\big{(}1+|\bm{x}_{1}|^{{p-2}}+|\bm{y}|^{p-2}\big{)},$
 $\bigg{|}\nabla^{2}_{\bm{x}_{1}\bm{x}_{2}}\frac{\delta^{2}f_{n}}{(\delta\mu)^{2}}\bigg{|}$ $\leq c\big{(}1+|\bm{x}_{1}|^{p-2}+|\bm{x}_{2}|^{p-2}+|\bm{y}|^{p-2}\big{)},$
 for all $\bm{x}_{1},\bm{x}_{2},\bm{y}\in{\mathbb{R}}^{d}$, $\mu\in{\mathcal{P}}_{p}({\mathbb{R}}^{d})$.
-The proof is similar to that of [16, Section 4].
-The main step of their proof consists of constructing an operator $T_{n}:C_{b}({\mathbb{R}}^{d})\to C_{b}({\mathbb{R}}^{d})$ and the corresponding adjoint $T_{n}^{*}:{\mathcal{P}}_{p}({\mathbb{R}}^{d})\to{\mathcal{P}}_{p}({\mathbb{R}}^{d})$ corresponding square $T_{n}^{\otimes 2}:C_{b}({\mathbb{R}}^{2d})\to C_{b}({\mathbb{R}}^{2d})$ such that $T_{n}\varphi_{n}\to\varphi$, when $\varphi_{n}\to\varphi$ and $T_{n}^{*}\mu\to\mu$.
-Define $f_{n}(\mu,\bm{y}):=\Phi(T_{n}^{*}\mu,\bm{y})$, then the key identities to be established are
-$\frac{\delta f_{n}}{\delta\mu}(\mu,\bm{y},\bm{x}_{1})=T_{n}\Big{(}\frac{\delta\Phi}{\delta\mu}(T_{n}^{*}\mu,\bm{y},\cdot)\Big{)}(\bm{x}_{1}),\quad\frac{\delta^{2}f_{n}}{(\delta\mu)^{2}}(\mu,\bm{y},\bm{x}_{1},\bm{x}_{2})=T_{n}^{\otimes 2}\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}(T_{n}^{*}\mu,\bm{y},\cdot,\cdot)(\bm{x}_{1},\bm{x}_{2}).$
-for all $\bm{x}_{1},\bm{x}_{2},\bm{y}\in{\mathbb{R}}^{d}$ and $\mu\in{\mathcal{P}}_{p}({\mathbb{R}}^{d})$. To this end, it suffices to note that the construction of $T_{n}$ ensures
+The proof is similar to that of [21, Section 4].
+Following their construction procedure, fix $n\in{\mathbb{N}}$ and cover the ball $B_{n}:=\{\bm{x}\in{\mathbb{R}}^{d}:|\bm{x}|\leq n\}$ by finitely many open sets of diameter at most $1/n$, denoted by $U^{n}_{i}$, with $i=1,...,N_{n}$.
+Set $U^{n}_{0}={\mathbb{R}}^{d}\setminus B_{n}$ to get an open cover in ${\mathbb{R}}^{d}$.
+Then, fix points ${\bm{x}}^{n}_{i}$ in $\overline{U^{n}_{i}}$ with the minimal norm for $i=0,1,..,N_{n}$.
+Next, define a family of functions $\{\psi^{n}_{i}\}$ as a partition of unity subordinate to $\{U^{n}_{i}\}$.
+(That is, each $\psi^{n}_{i}$ is a continuous function supported on $U^{n}_{i}$, such that $\sum_{i=0}^{N_{n}}\psi^{n}_{i}(\bm{x})=1$ for all $\bm{x}\in{\mathbb{R}}^{d}$).
+Finally, define the operator $T_{n}:C_{b}({\mathbb{R}}^{d})\to C_{b}({\mathbb{R}}^{d})$ such that for any function $\varphi$ on ${\mathbb{R}}^{d}$,
+$T_{n}\varphi(x)=\sum_{i=0}^{N_{n}}\varphi({\bm{x}}^{n}_{i})\psi^{n}_{i}(\bm{x}).$
+Its adjoint operator $T_{n}^{*}:{\mathcal{P}}_{p}({\mathbb{R}}^{d})\to{\mathcal{P}}_{p}({\mathbb{R}}^{d})$ and the tensor square operator $T_{n}^{\otimes 2}:C_{b}({\mathbb{R}}^{2d})\to C_{b}({\mathbb{R}}^{2d})$ are defined accordingly.
+Now, define $f_{n}(\mu,\bm{y}):=\Phi(T_{n}^{*}\mu,\bm{y})$, then $f_{n}$ are $C^{2,2}$ cylindrical functions and the key identities to be established are
+$\frac{\delta f_{n}}{\delta\mu}(\mu,\bm{y},\bm{x}_{1})=T_{n}\Big{(}\frac{\delta\Phi}{\delta\mu}(T_{n}^{*}\mu,\bm{y},\cdot)\Big{)}(\bm{x}_{1}),\quad\frac{\delta^{2}f_{n}}{(\delta\mu)^{2}}(\mu,\bm{y},\bm{x}_{1},\bm{x}_{2})=T_{n}^{\otimes 2}\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}(T_{n}^{*}\mu,\bm{y},\cdot,\cdot)(\bm{x}_{1},\bm{x}_{2}),$
+for all $\bm{x}_{1},\bm{x}_{2},\bm{y}\in{\mathbb{R}}^{d}$ and $\mu\in{\mathcal{P}}_{p}({\mathbb{R}}^{d})$ (see [21, Lemma 4.6, Lemma 4.12]). To this end, note that the construction of $T_{n}$ ensures
 $T_{n}\nabla_{\bm{x}_{1}}\varphi(\bm{x}_{1})=\nabla_{\bm{x}_{1}}T_{n}\varphi(\bm{x}_{1}),\quad T_{n}^{\otimes 2}\nabla^{2}_{\bm{x}_{1}\bm{x}_{2}}\psi(\bm{x}_{1},\bm{x}_{2})=\nabla^{2}_{\bm{x}_{1}\bm{x}_{2}}T_{n}\psi(\bm{x}_{1},\bm{x}_{2}),$
-for all $\varphi\in C^{1}_{b}({\mathbb{R}}^{d})$ and $\psi\in C^{2}_{b}({\mathbb{R}}^{2d}).$
+for all $\varphi\in C^{1}_{b}({\mathbb{R}}^{d})$ and $\psi\in C^{2}_{b}({\mathbb{R}}^{2d}).$ Thus repeating the proof of [21, Theorem 4.4, Theorem 4.10] finishes the proof of the proposition.
 4.4. Final step
-Define $\tau_{k}:=\inf\{t\in[0,T]:|\bm{Y}_{t}|>k\}$, then by Step 1, we know that Itô’s formula (2.3) holds for the truncated process $(\mu^{(\tau_{n})},\bm{Y}^{(\tau_{n})}):=(\mu_{t\wedge\tau_{n}},\bm{Y}_{t\wedge\tau_{n}})_{t\in[0,T]}$ with the sequences of cylindrical functions $\{f_{n}\}_{n\in{\mathbb{N}}^{+}}$.
+Define $\tau_{k}:=\inf\{t\in[0,T]:|\bm{Y}_{t}|>k\}$, then by Step 1, Itô’s formula (2.3) holds for the truncated process $(\mu^{(\tau_{n})},\bm{Y}^{(\tau_{n})}):=(\mu_{t\wedge\tau_{n}},\bm{Y}_{t\wedge\tau_{n}})_{t\in[0,T]}$ with the sequences of cylindrical functions $\{f_{n}\}_{n\in{\mathbb{N}}^{+}}$.
 By Step 2, the localization argument shows that Itô’s formula (2.3) holds for the sequences of cylindrical functions $\{f_{n}\}_{n\in{\mathbb{N}}^{+}}$.
@@
 $\int_{0+}^{t}\nabla_{\bm{x}_{1}}\frac{\delta f_{n}}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\cdot{\mathrm{d}}(\bm{X}^{\prime})^{c}_{s}\overset{L^{1}}{\to}\int_{0+}^{t}\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\cdot{\mathrm{d}}(\bm{X}^{\prime})^{c}_{s},$ (4.17)
-as $n\to\infty$. By Proposition 4.2, we have
+as $n\to\infty$. In fact, by Proposition 4.2, we have
 $\bigg{|}|\nabla_{\bm{x}_{1}}\frac{\delta f_{n}}{\delta\mu}\big{(}\mu_{s-},\bm{Y}_{s-},\bm{X}^{\prime}_{s-}\big{)}\bigg{|}\leq c\big{(}1+|\bm{X}_{s-}|^{p-1}+|\bm{Y}_{s-}|^{p-1}\big{)},$
-where the ${\mathcal{S}}_{p^{\prime}}$ norm of the right hand side is finite with $p^{\prime}=\frac{p}{p-1}$.
-Now (4.17) follows from $\bm{X}\in{\mathcal{H}}^{p}$ and the dominated convergence theorem of stochastic integral ([37, Page 273, Lemma]).
+where the ${\mathcal{S}}_{p^{\prime}}$ norm of the right hand side is finite with $p^{\prime}=\frac{p}{p-1}$. Now (4.17) follows from $\|\bm{X}\|_{{\mathcal{H}}^{p}}<\infty$ and Lemma 2.5.
 Next, since
@@
 and $L_{1}$ convergences implies convergence in subsequence, (4.16) holds. The remaining terms follow similarly due to the uniform bound (4.15), and thus the proof of Theorem 2.11 is complete.
-5. Proof of Section 3
-5.1. Construction of the conditional independent copy
-This subsection is devoted to the proof of the precise construction of conditional independent copy as in Theorem 3.1.
-For ease of exposition, we will show how to enlarge the probability space with two independent copies, and any finite number of copies can be easily adapted.
-Recall the definition in Section 3, let us consider the probability space $\overline{\Omega}$ and the $\sigma$-algrebra $\overline{\mathcal{F}}$ defined by
-$\overline{\Omega}=\Omega^{3}=\{(\omega_{0},\omega_{1},\omega_{2})|\omega_{0},\omega_{1},\omega_{2}\in\Omega\},\qquad\overline{\mathcal{F}}=\sigma\{A_{0}\times A_{1}\times A_{2}:A_{0},A_{1},A_{2}\in{\mathcal{F}}\}$ (5.1)
-and probability measure $\overline{\mathbb{P}}$ is defined as follows: for $A_{0},A_{1},A_{2}\in{\mathcal{F}}$,
-¯$\overline{\mathbb{P}}\big{(}A_{0}\times A_{1}\times A_{2}\big{)}:={\mathbb{E}}\Big{[}\mathbbm{1}_{A_{0}}{\mathbb{P}}(A_{1}|{\mathcal{G}}){\mathbb{P}}(A_{2}|{\mathcal{G}})\Big{]},$ (5.2)
-Since the set $\{A_{0}\times A_{1}\times A_{2}:A_{0},A_{1},A_{2}\in{\mathcal{F}}\}$ is a $\pi$-system, the measure on $(\overline{\Omega},\overline{\mathcal{F}},\overline{\mathbb{P}})$ is properly defined.
-Next, use the notation ~$\widetilde{\mathcal{F}},\widetilde{\mathcal{G}}$ for the nature extension into this enlarged probability space as
-~:=$\widetilde{\mathcal{F}}:=\{A\times\Omega\times\Omega:A\in{\mathcal{F}}\},\quad\widetilde{\mathcal{G}}:=\{A\times\Omega\times\Omega:A\in{\mathcal{G}}\},$ (5.3)
+5. Proof of Results in Section 3
+5.1. Proof of Theorem 3.1
+To prove the theorem, we first provide a constructive proof for the existence of conditionally independent copy claimed in Theorem 3.1.
+For ease of exposition, we will show how to enlarge the probability space with two independent copies, and the case with a finite number of copies can be easily adapted.
+Recall the discussions in Section 3, and consider the probability space $\overline{\Omega}$, the $\sigma$-algebra $\overline{{\mathcal{F}}}$ defined by
+$\overline{\Omega}=\Omega^{3}=\{(\omega_{0},\omega_{1},\omega_{2})|\omega_{0},\omega_{1},\omega_{2}\in\Omega\},\qquad\overline{{\mathcal{F}}}=\sigma\{A_{0}\times A_{1}\times A_{2}:A_{0},A_{1},A_{2}\in{\mathcal{F}}\},$ (5.1)
+and a probability measure $\overline{{\mathbb{P}}}$ defined as follows: for $A_{0},A_{1},A_{2}\in{\mathcal{F}}$,
+$\overline{{\mathbb{P}}}\big{(}A_{0}\times A_{1}\times A_{2}\big{)}:={\mathbb{E}}\Big{[}\mathbbm{1}_{A_{0}}{\mathbb{P}}(A_{1}|{\mathcal{G}}){\mathbb{P}}(A_{2}|{\mathcal{G}})\Big{]}.$ (5.2)
+Since the set $\{A_{0}\times A_{1}\times A_{2}:A_{0},A_{1},A_{2}\in{\mathcal{F}}\}$ is a $\pi$-system, the measure on $(\overline{\Omega},\overline{{\mathcal{F}}},\overline{{\mathbb{P}}})$ is well defined.
+Next, take $\widetilde{\mathcal{F}},\widetilde{\mathcal{G}}$ for the natural extension in this enlarged probability space as
+$\widetilde{\mathcal{F}}:=\{A\times\Omega\times\Omega:A\in{\mathcal{F}}\},\quad\widetilde{\mathcal{G}}:=\{A\times\Omega\times\Omega:A\in{\mathcal{G}}\},$ (5.3)
 and define $\mathcal{F}$$^{\prime}$, $\mathcal{F}$$^{\prime\prime}$ as
-:=${\mathcal{F}}^{\prime}:=\{\Omega\times A\times\Omega:A\in{\mathcal{F}}\},\quad{\mathcal{F}}^{\prime\prime}:=\{\Omega\times\Omega\times A:A\in{\mathcal{F}}\}.$ (5.4)
-With these constructions, we have the following propositions.
+${\mathcal{F}}^{\prime}:=\{\Omega\times A\times\Omega:A\in{\mathcal{F}}\},\quad{\mathcal{F}}^{\prime\prime}:=\{\Omega\times\Omega\times A:A\in{\mathcal{F}}\}.$ (5.4)
+In this framework, we have the following proposition.
 Proposition 5.1.
-Given $(\overline{\Omega},\overline{\mathcal{F}},\overline{\mathbb{P}})$ defined by (5.1) and (5.2) and the sub-algebras ~$\widetilde{\mathcal{F}},\widetilde{\mathcal{G}}$, ${\mathcal{F}}^{\prime}$, and ${\mathcal{F}}^{\prime\prime}$ defined by (5.3) and (5.4).
-Suppose $A,B\in{\mathcal{F}}$, let $\widetilde{A}=A\times\Omega\times\Omega$, $A^{\prime}=\Omega\times A\times\Omega$, $A^{\prime\prime}=\Omega\times\Omega\times A$, $B^{\prime}=\Omega\times B\times\Omega$ and $C^{\prime\prime}=\Omega\times\Omega\times C$.
+Consider $(\overline{\Omega},\overline{{\mathcal{F}}},\overline{{\mathbb{P}}})$ defined by (5.1) and (5.2) and the sub-algebras $\widetilde{\mathcal{F}},\widetilde{\mathcal{G}}$, $\mathcal{F}$$^{\prime}$, and $\mathcal{F}$$^{\prime\prime}$ defined by (5.3) and (5.4).
+Suppose $A,B\in{\mathcal{F}}$, and let $\widetilde{A}=A\times\Omega\times\Omega$, $A^{\prime}=\Omega\times A\times\Omega$, $A^{\prime\prime}=\Omega\times\Omega\times A$, $B^{\prime}=\Omega\times B\times\Omega$, and $C^{\prime\prime}=\Omega\times\Omega\times C$.
 Then
   1. (1)
-    For $\overline{\omega}=(\omega_{0},\omega_{1},\omega_{2})\in\overline{\Omega}$
-    ${\mathbb{P}}(A|\mathcal{G})(\omega_{0})=\overline{\mathbb{P}}(\widetilde{A}|\widetilde{\mathcal{G}})(\omega_{0},\omega_{1},\omega_{2}),\text{ a.s. under }\overline{\mathbb{P}}.$
+    For $\overline{\omega}=(\omega_{0},\omega_{1},\omega_{2})\in\overline{\Omega}$,
+[…]
   2. (2)
-    Given $\widetilde{\mathcal{G}}$, then $\widetilde{A},A^{\prime},$ and $A^{\prime\prime}$ have the same conditional law. That is,
-    ¯$\overline{\mathbb{P}}(\widetilde{A}|\widetilde{\mathcal{G}})=\overline{\mathbb{P}}(A^{\prime}|\widetilde{\mathcal{G}})=\overline{\mathbb{P}}(A^{\prime\prime}|\widetilde{\mathcal{G}}),\text{ a.s.
+    Given $\widetilde{\mathcal{G}}$, $\widetilde{A},A^{\prime},$ and $A^{\prime\prime}$ have the same conditional law. That is,
+[…]
+[…]
+[…]
+  3. (3)
+    Given $\widetilde{\mathcal{G}}$, $\widetilde{\mathcal{F}}$$\mathcal{F}$$^{\prime}$, and $\mathcal{F}$$^{\prime\prime}$ are conditionally independent. That is,
+    $\overline{{\mathbb{P}}}\big{(}\widetilde{A}\cap B^{\prime}\cap C^{\prime\prime}|\widetilde{\mathcal{G}}\big{)}=\overline{{\mathbb{P}}}\big{(}\widetilde{A}|\widetilde{\mathcal{G}}\big{)}\overline{{\mathbb{P}}}\big{(}B^{\prime}|\widetilde{\mathcal{G}}\big{)}\overline{{\mathbb{P}}}\big{(}C^{\prime\prime}|\widetilde{\mathcal{G}}\big{)},\text{ a.s.
 under }\overline{{\mathbb{P}}}.$
-    Moreover, this conditional probability ¯$\overline{\mathbb{P}}(\widetilde{A}|\widetilde{\mathcal{G}})$ coincides with the conditional probability of $A^{\prime}$ and $A^{\prime\prime}$ given $\widetilde{\mathcal{F}}$, that is
-    ¯$\overline{\mathbb{P}}(\widetilde{A}|\widetilde{\mathcal{G}})=\overline{\mathbb{P}}(A^{\prime}|\widetilde{\mathcal{F}})=\overline{\mathbb{P}}(A^{\prime\prime}|\widetilde{\mathcal{F}}),\text{ a.s.
-under }\overline{\mathbb{P}}.$
-  3. (3)
-    Given $\widetilde{\mathcal{G}}$, then $\widetilde{\mathcal{F}},{\mathcal{F}}^{\prime},$ and ${\mathcal{F}}^{\prime\prime}$ are conditionally independent. That is,
-    ¯$\overline{\mathbb{P}}\big{(}\widetilde{A}\cap B^{\prime}\cap C^{\prime\prime}|\widetilde{\mathcal{G}}\big{)}=\overline{\mathbb{P}}\big{(}\widetilde{A}|\widetilde{\mathcal{G}}\big{)}\overline{\mathbb{P}}\big{(}B^{\prime}|\widetilde{\mathcal{G}}\big{)}\overline{\mathbb{P}}\big{(}C^{\prime\prime}|\widetilde{\mathcal{G}}\big{)},\text{ a.s.
-under }\overline{\mathbb{P}}.$
-    Moreover, given $\widetilde{\mathcal{F}}$, then ${\mathcal{F}}^{\prime}$ and ${\mathcal{F}}^{\prime\prime}$ are conditionally independent, That is
-    ¯$\overline{\mathbb{P}}\big{(}B^{\prime}\cap C^{\prime\prime}|\widetilde{\mathcal{F}}\big{)}=\overline{\mathbb{P}}\big{(}B^{\prime}|\widetilde{\mathcal{F}}\big{)}\overline{\mathbb{P}}\big{(}C^{\prime\prime}|\widetilde{\mathcal{F}}\big{)},\text{ a.s.
+    Moreover, given $\widetilde{\mathcal{F}}$, then $\mathcal{F}$$^{\prime}$ and $\mathcal{F}$$^{\prime\prime}$ are conditionally independent. That is,
+    $\overline{{\mathbb{P}}}\big{(}B^{\prime}\cap C^{\prime\prime}|\widetilde{\mathcal{F}}\big{)}=\overline{{\mathbb{P}}}\big{(}B^{\prime}|\widetilde{\mathcal{F}}\big{)}\overline{{\mathbb{P}}}\big{(}C^{\prime\prime}|\widetilde{\mathcal{F}}\big{)},\text{ a.s.
 under }\overline{{\mathbb{P}}}.$
 Proof.
-To prove (1), note that ${\mathbb{P}}[A|{\mathcal{G}}]$ is ${\mathcal{G}}$ measurable in $(\Omega,{\mathcal{F}},{\mathbb{F}},{\mathbb{P}})$, then ${\mathbb{P}}[A|{\mathcal{G}}](\omega_{1})$ as a random variable in $(\overline{\Omega},\overline{\mathcal{F}},\overline{\mathbb{F}},\overline{\mathbb{P}})$ is $\widetilde{\mathcal{G}}$ measurable .
+(of Proposition 5.1).
+To prove (1), note that ${\mathbb{P}}[A|{\mathcal{G}}]$ is $\mathcal{G}$} measurable in $(\Omega,{\mathcal{F}},{\mathbb{F}},{\mathbb{P}})$, then ${\mathbb{P}}[A|{\mathcal{G}}](\omega_{1})$ as a random variable in $(\overline{\Omega},\overline{{\mathcal{F}}},\overline{{\mathbb{F}}},\overline{{\mathbb{P}}})$ is $\widetilde{\mathcal{G}}-$measurable.
 For any $\widetilde{G}\in\widetilde{\mathcal{G}}$, there exists a $G\in{\mathcal{G}}$ such that $\widetilde{G}=G\times\Omega\times\Omega\in\widetilde{\mathcal{G}}$.
-Recall the definition of $\overline{{\mathbb{P}}}$ at (5.2) and the property of conditional expectation, we have
-¯$\overline{\mathbb{P}}\big{[}\widetilde{A}\cap\widetilde{G}\big{]}=\overline{\mathbb{P}}\big{[}(A\cap G)\times\Omega\times\Omega\big{]}={\mathbb{P}}[A\cap G]={\mathbb{E}}\Big{[}{\mathbb{P}}[A|{\mathcal{G}}]\mathbbm{1}_{G}\Big{]}=\int_{\Omega}{\mathbb{P}}[A|{\mathcal{G}}](\omega)\mathbbm{1}_{G}(\omega){\mathbb{P}}({\mathrm{d}}\omega).$
-Again by (5.2), we obtain ¯$\overline{\mathbb{P}}({\mathrm{d}}\omega_{1}\times\Omega\times\Omega)={\mathbb{P}}({\mathrm{d}}\omega_{1})$, therefore,
-¯$\overline{\mathbb{P}}\big{[}\widetilde{A}\cap\widetilde{G}\big{]}$ $=\int_{\Omega^{3}}{\mathbb{P}}[A|{\mathcal{G}}](\omega_{0})\mathbbm{1}_{B}(\omega_{0})\overline{\mathbb{P}}({\mathrm{d}}\omega_{0}{\mathrm{d}}\omega_{1}{\mathrm{d}}\omega_{2})$ (5.5)
+By the definition of $\overline{{\mathbb{P}}}$ in ((5.2)) and the property of conditional expectation,
+$\overline{{\mathbb{P}}}\big{[}\widetilde{A}\cap\widetilde{G}\big{]}=\overline{{\mathbb{P}}}\big{[}(A\cap G)\times\Omega\times\Omega\big{]}={\mathbb{P}}[A\cap G]={\mathbb{E}}\Big{[}{\mathbb{P}}[A|{\mathcal{G}}]\mathbbm{1}_{G}\Big{]}=\int_{\Omega}{\mathbb{P}}[A|{\mathcal{G}}](\omega)\mathbbm{1}_{G}(\omega){\mathbb{P}}({\mathrm{d}}\omega).$
+Again by (5.2), $\overline{{\mathbb{P}}}({\mathrm{d}}\omega_{1}\times\Omega\times\Omega)={\mathbb{P}}({\mathrm{d}}\omega_{1})$, therefore,
+$\overline{{\mathbb{P}}}\big{[}\widetilde{A}\cap\widetilde{G}\big{]}$ $=\int_{\Omega^{3}}{\mathbb{P}}[A|{\mathcal{G}}](\omega_{0})\mathbbm{1}_{G}(\omega_{0})\overline{{\mathbb{P}}}({\mathrm{d}}\omega_{0}{\mathrm{d}}\omega_{1}{\mathrm{d}}\omega_{2})$ (5.5)
 $=\int_{\Omega^{3}}{\mathbb{P}}[A|{\mathcal{G}}](\omega_{0})\mathbbm{1}_{\widetilde{G}}(\omega_{0},\omega_{1},\omega_{2})\overline{{\mathbb{P}}}({\mathrm{d}}\omega_{0}{\mathrm{d}}\omega_{1}{\mathrm{d}}\omega_{2})=\overline{{\mathbb{E}}}\Big{[}{\mathbb{P}}[A|{\mathcal{G}}]\mathbbm{1}_{\widetilde{G}}\Big{]}$,
@@
 To prove (2), for any $\widetilde{G}\in\widetilde{\mathcal{G}}$, there exists a $G\in{\mathcal{G}}$ such that $\widetilde{G}=G\times\Omega\times\Omega\in\widetilde{\mathcal{G}}$. Then,
-¯$\overline{\mathbb{P}}\big{[}A^{\prime}\cap\widetilde{G}\big{]}=\overline{\mathbb{P}}\big{[}G\times A\times\Omega\big{]}={\mathbb{E}}\big{[}\mathbbm{1}_{G}{\mathbb{P}}(A|{\mathcal{G}})\big{]}=\overline{\mathbb{E}}\Big{[}{\mathbb{P}}[A|{\mathcal{G}}]\mathbbm{1}_{\widetilde{G}}\Big{]}=\overline{\mathbb{E}}\Big{[}\overline{\mathbb{P}}[\widetilde{A}|\widetilde{\mathcal{G}}]\mathbbm{1}_{\widetilde{G}}\Big{]},$
-where the second equation holds because of (5.2), the third equation holds due to the same calculation as (5.5) and the last equation holds by (1).
-This leads to ¯$\overline{\mathbb{P}}(A^{\prime}|\widetilde{\mathcal{G}})=\overline{\mathbb{P}}(\widetilde{A}|\widetilde{\mathcal{G}})$ and similarly, we have ¯$\overline{\mathbb{P}}(A^{\prime\prime}|\widetilde{\mathcal{G}})=\overline{\mathbb{P}}(\widetilde{A}|\widetilde{\mathcal{G}})$.
-As for the second statement of (2), note that ¯$\overline{\mathbb{P}}(A^{\prime}|\widetilde{\mathcal{G}})$ is $\widetilde{\mathcal{G}}$ measurable, thus $\widetilde{\mathcal{F}}$ measurable.
-For any $\widetilde{F}\in\widetilde{\mathcal{F}}$, there exists a $F\in{\mathcal{F}}$ such that $\widetilde{F}=F\times\Omega\times\Omega\in\widetilde{\mathcal{F}}$.
-Then,
-¯$\overline{\mathbb{P}}(A^{\prime}\cap\widetilde{F})=\overline{\mathbb{P}}\big{[}F\times A\times\Omega\big{]}={\mathbb{E}}\big{[}\mathbbm{1}_{F}{\mathbb{P}}(A|{\mathcal{G}})\big{]}=\overline{\mathbb{E}}\Big{[}{\mathbb{P}}[A|{\mathcal{G}}]\mathbbm{1}_{\widetilde{F}}\Big{]}=\overline{\mathbb{E}}\Big{[}\overline{\mathbb{P}}[\widetilde{A}|\widetilde{\mathcal{G}}]\mathbbm{1}_{\widetilde{F}}\Big{]},$
-where the second equation holds because of (5.2), the third equation holds due to the same calculation as (5.5) and the last equation holds by (1).
-This leads to ¯$\overline{\mathbb{P}}(A^{\prime}|\widetilde{\mathcal{F}})=\overline{\mathbb{P}}(\widetilde{A}|\widetilde{\mathcal{G}})$ and similarly, we have ¯$\overline{\mathbb{P}}(A^{\prime\prime}|\widetilde{\mathcal{F}})=\overline{\mathbb{P}}(\widetilde{A}|\widetilde{\mathcal{G}})$.
-To prove (3), for any $\widetilde{G}\in\widetilde{\mathcal{G}}$, there exists a $G\in{\mathcal{G}}$ such that $\widetilde{G}=G\times\Omega\times\Omega\in\widetilde{\mathcal{G}}$. Then,
-¯$\overline{{\mathbb{P}}}(\widetilde{A}\cap B^{\prime}\cap C^{\prime\prime}\cap\widetilde{G})$ =¯$=\overline{{\mathbb{P}}}((A\cap G)\times B\times C)={\mathbb{E}}\Big{[}\mathbbm{1}_{A\cap G}{\mathbb{P}}(B|{\mathcal{G}}){\mathbb{P}}(C|{\mathcal{G}})\Big{]}$
-$={\mathbb{E}}\Big{[}{\mathbb{E}}[\mathbbm{1}_{A\cap G}|{\mathcal{G}}]{\mathbb{P}}(B|{\mathcal{G}}){\mathbb{P}}(C|{\mathcal{G}})\Big{]}={\mathbb{E}}\Big{[}\mathbbm{1}_{G}{\mathbb{P}}(A|{\mathcal{G}}){\mathbb{P}}(B|{\mathcal{G}}){\mathbb{P}}(C|{\mathcal{G}})\Big{]}$
-Using the same calculation as in (5.5) and note that ${\mathbb{P}}(A|{\mathcal{G}})=\overline{\mathbb{P}}(\widetilde{A}|\widetilde{\mathcal{G}})$, ${\mathbb{P}}(B|{\mathcal{G}})=\overline{\mathbb{P}}(B^{\prime}|\widetilde{\mathcal{G}})$ and ${\mathbb{P}}(C|{\mathcal{G}})=\overline{\mathbb{P}}(C^{\prime\prime}|\widetilde{\mathcal{G}})$, we have
-¯$\overline{{\mathbb{P}}}(\widetilde{A}\cap B^{\prime}\cap C^{\prime\prime}\cap\widetilde{G})$ $=\overline{\mathbb{E}}\Big{[}\mathbbm{1}_{\widetilde{G}}{\mathbb{P}}(A|{\mathcal{G}}){\mathbb{P}}(B|{\mathcal{G}}){\mathbb{P}}(C|{\mathcal{G}})\Big{]}=\overline{\mathbb{E}}\Big{[}\mathbbm{1}_{\widetilde{G}}\overline{\mathbb{P}}(\widetilde{A}|\widetilde{\mathcal{G}})\overline{\mathbb{P}}(B^{\prime}|\widetilde{\mathcal{G}})\overline{\mathbb{P}}(C^{\prime\prime}|\widetilde{\mathcal{G}})\Big{]}.$
-This leads to
-¯$\overline{\mathbb{P}}\big{(}\widetilde{A}\cap B^{\prime}\cap C^{\prime\prime}|\widetilde{\mathcal{G}}\big{)}=\overline{\mathbb{P}}(\widetilde{A}|\widetilde{\mathcal{G}})\overline{\mathbb{P}}(B^{\prime}|\widetilde{\mathcal{G}})\overline{\mathbb{P}}(C^{\prime\prime}|\widetilde{\mathcal{G}}),\text{ a.s.
+$\overline{{\mathbb{P}}}\big{[}A^{\prime}\cap\widetilde{G}\big{]}=\overline{{\mathbb{P}}}\big{[}G\times A\times\Omega\big{]}={\mathbb{E}}\big{[}\mathbbm{1}_{G}{\mathbb{P}}(A|{\mathcal{G}})\big{]}=\overline{{\mathbb{E}}}\Big{[}{\mathbb{P}}[A|{\mathcal{G}}]\mathbbm{1}_{\widetilde{G}}\Big{]}=\overline{{\mathbb{E}}}\Big{[}\overline{{\mathbb{P}}}[\widetilde{A}|\widetilde{\mathcal{G}}]\mathbbm{1}_{\widetilde{G}}\Big{]}$,
+[…]
+[…]
+$\overline{{\mathbb{P}}}(A^{\prime}\cap\widetilde{F})=\overline{{\mathbb{P}}}\big{[}F\times A\times\Omega\big{]}={\mathbb{E}}\big{[}\mathbbm{1}_{F}{\mathbb{P}}(A|{\mathcal{G}})\big{]}=\overline{{\mathbb{E}}}\Big{[}{\mathbb{P}}[A|{\mathcal{G}}]\mathbbm{1}_{\widetilde{F}}\Big{]}=\overline{{\mathbb{E}}}\Big{[}\overline{{\mathbb{P}}}[\widetilde{A}|\widetilde{\mathcal{G}}]\mathbbm{1}_{\widetilde{F}}\Big{]}$,
+[…]
+To prove (3), for any $\widetilde{G}\in\widetilde{\mathcal{G}}$, there exists a $G\in{\mathcal{G}}$ such that $\widetilde{G}=G\times\Omega\times\Omega\in\widetilde{\mathcal{G}}$, and
+[…]
+$={\mathbb{E}}\Big{[}{\mathbb{E}}[\mathbbm{1}_{A\cap G}|{\mathcal{G}}]{\mathbb{P}}(B|{\mathcal{G}}){\mathbb{P}}(C|{\mathcal{G}})\Big{]}={\mathbb{E}}\Big{[}\mathbbm{1}_{G}{\mathbb{P}}(A|{\mathcal{G}}){\mathbb{P}}(B|{\mathcal{G}}){\mathbb{P}}(C|{\mathcal{G}})\Big{]}.$
+[…]
+[…]
+Therefore,
+$\overline{{\mathbb{P}}}\big{(}\widetilde{A}\cap B^{\prime}\cap C^{\prime\prime}|\widetilde{\mathcal{G}}\big{)}=\overline{{\mathbb{P}}}(\widetilde{A}|\widetilde{\mathcal{G}})\overline{{\mathbb{P}}}(B^{\prime}|\widetilde{\mathcal{G}})\overline{{\mathbb{P}}}(C^{\prime\prime}|\widetilde{\mathcal{G}}),\text{ a.s.
 under }\overline{{\mathbb{P}}}.$
-As for the second statement of (3), for any $\widetilde{F}\in\widetilde{\mathcal{F}}$, there exists a $F\in{\mathcal{F}}$ such that $\widetilde{F}=F\times\Omega\times\Omega\in\widetilde{\mathcal{F}}$.
-Then,
-¯$\overline{\mathbb{P}}(B^{\prime}\cap C^{\prime\prime}\cap\widetilde{F})$ =¯$=\overline{\mathbb{P}}(F\times B\times C)={\mathbb{E}}\big{[}\mathbbm{1}_{F}{\mathbb{P}}(B|{\mathcal{G}}){\mathbb{P}}(C|{\mathcal{G}})\big{]}$
+As for the second statement of (3), for any $\widetilde{F}\in\widetilde{\mathcal{F}}$, there exists an $F\in{\mathcal{F}}$ such that $\widetilde{F}=F\times\Omega\times\Omega\in\widetilde{\mathcal{F}}$, and by (5.2),
+[…]
 $=\overline{{\mathbb{E}}}\big{[}\mathbbm{1}_{\widetilde{F}}{\mathbb{P}}(B|{\mathcal{G}}){\mathbb{P}}(C|{\mathcal{G}})\big{]}=\overline{{\mathbb{E}}}\big{[}\mathbbm{1}_{\widetilde{F}}\overline{{\mathbb{P}}}(B^{\prime}|\widetilde{\mathcal{F}})\overline{{\mathbb{P}}}(C^{\prime\prime}|\overline{{\mathcal{F}}})\big{]},$
-where the second equation holds because of (5.2), the third equation holds due to the same calculation as (5.5) and the last equation holds by (1) and (2). Therefore,
-¯$\overline{\mathbb{P}}\big{(}B^{\prime}\cap C^{\prime\prime}|\widetilde{\mathcal{F}}\big{)}=\overline{\mathbb{P}}\big{(}B^{\prime}|\widetilde{\mathcal{F}}\big{)}\overline{\mathbb{P}}\big{(}C^{\prime\prime}|\widetilde{\mathcal{F}}\big{)},\text{ a.s.
-under }\overline{\mathbb{P}},$
-which completes the whole proof. ∎
-Proof of Theorem 3.1.
-For any random variable $R$ in $(\Omega,{\mathcal{F}},{\mathbb{P}})$, recall that the natural extension $\widetilde{R}$ and the two conditional independent copies $R^{\prime},R^{\prime\prime}$ of $R$ are defined by
+where the the third equality follows from the same calculation as (5.5) and the last equation holds by (1) and (2). Therefore,
+$\overline{{\mathbb{P}}}\big{(}B^{\prime}\cap C^{\prime\prime}|\widetilde{\mathcal{F}}\big{)}=\overline{{\mathbb{P}}}\big{(}B^{\prime}|\widetilde{\mathcal{F}}\big{)}\overline{{\mathbb{P}}}\big{(}C^{\prime\prime}|\widetilde{\mathcal{F}}\big{)},\text{ a.s.
+under }\overline{{\mathbb{P}}}.$
+Proof.
+(of Theorem 3.1).
+For any random variable $R$ taking values in a standard Borel space $(E,\mathcal{E})$, recall that the natural extension $\widetilde{R}$ and the two conditionally independent copies $R^{\prime}$ and $R^{\prime\prime}$ of $R$ are defined by
 $\widetilde{R}(\omega_{0},\omega_{1},\omega_{2})=R(\omega_{0}),\;R^{\prime}(\omega_{0},\omega_{1},\omega_{2})=R(\omega_{1}),\;R^{\prime\prime}(\omega_{0},\omega_{1},\omega_{2})=R(\omega_{2}).$
 Then
-$\text{Law}\big{(}R\big{|}{\mathcal{G}}\big{)}(\omega_{0})=\text{Law}\big{(}\widetilde{R}\big{|}\widetilde{\mathcal{G}}\big{)}(\overline{\omega})=\text{Law}\big{(}R^{\prime}\big{|}{\widetilde{\mathcal{G}}}\big{)}(\overline{\omega})=\text{Law}\big{(}R^{\prime\prime}\big{|}{\widetilde{\mathcal{F}}}\big{)}(\overline{\omega})\text{ a.s.
-under }\overline{\mathbb{P}},$
 […]
-$R,R^{\prime}\text{ and }R^{\prime\prime}\text{ are independent given }\widetilde{\mathcal{G}}\text{ and }R,R^{\prime}\text{ and }R^{\prime\prime}\text{ are independent given }{\mathcal{F}},$
-which completes the proof of Theorem 3.1.∎
-Proof of Corollary 3.2.
-It suffices to show that extension of the sub-filtration $\widetilde{\mathbb{G}}\subset\widetilde{\mathbb{F}}$ satisfies the compatibility assumption or the conditional independence condition, that is
-~${\widetilde{\mathcal{F}}}_{t}\perp\!\!\!\perp{\widetilde{\mathcal{G}}}_{T}\big{|}{\widetilde{\mathcal{G}}}_{t},\text{ for all }0\leq t\leq T.$
-For any $\widetilde{F}_{t}\in\widetilde{\mathcal{F}}_{t}$, $\widetilde{G}_{t}\in{\widetilde{\mathcal{G}}}_{t}$ and $\widetilde{G}_{T}\in{\widetilde{\mathcal{G}}}_{T}$, there exists $F_{t}\in{\mathcal{F}}_{t}$, $G_{t}\in{\mathcal{G}}_{t}$ and $G_{T}\in{\mathcal{G}}_{T}$ such that $\widetilde{F}_{t}=F_{t}\times\Omega\times\Omega$, $\widetilde{G}_{t}=G_{t}\times\Omega\times\Omega$ and $\widetilde{G}_{T}=G_{T}\times\Omega\times\Omega$.
+[…]
+Remark 5.2.
+Our construction is inspired by the proof of Yamada-Watanabe theorem (see for instance [39, 42]) and the theory of weak and strong solutions of stochastic differential equations (see for instance [37, 43]).
+Note that in these existing approaches, the common information $\mathcal{G}$} is modeled as part of the stochastic input.
+Moreover, in the context of weak and strong solutions of SDEs, one often uses the existence of regular conditional distributions to construct an extended probability space by embedding different weak solutions into a common space.
+To illustrate this, take the example of $(X^{(1)},W^{(1)})$ and $(X^{(2)},W^{(2)})$, let $\nu({\mathrm{d}}w)$ denote the common law of a Brownian motion, and define the regular conditional law $\eta^{(i)}({\mathrm{d}}x;w):=\text{Law}(X^{(i)}|W^{(i)}=w)$ for $i=1,2$.
+In such a setting, an extended probability measure is often defined (informally) by ${\mathbb{P}}({\mathrm{d}}x^{(1)}{\mathrm{d}}x^{(2)}{\mathrm{d}}w):=\eta^{(1)}({\mathrm{d}}x^{(1)};w)\cdot\eta^{(2)}({\mathrm{d}}x^{(2)};w)\nu({\mathrm{d}}w).$
+Motivated by these constructions, we consider a more general setting in which the common information $\mathcal{G}$} may be abstract and regular conditional distributions may not be available.
+We define conditionally independent copies of some random variable $X$ directly by specifying the measure on the extended space supporting $X,X^{\prime}$ such that
+${\mathbb{P}}\big{(}(X,X^{\prime})\in A\times B\big{)}:={\mathbb{E}}[\mathbbm{1}_{X\in A}\cdot{\mathbb{P}}(X\in B|{\mathcal{G}})],$
+which remains consistent with the spirit of the classical constructions while avoiding the explicit use of regular conditional distributions.
+Proof of Corollary 3.2. It suffices to show that the extension of the sub-filtration $\widetilde{\mathbb{G}}\subset\widetilde{\mathbb{F}}$ satisfies
+${\widetilde{\mathcal{F}}}_{t}\perp\!\!\!\perp{\widetilde{\mathcal{G}}}_{T}\big{|}{\widetilde{\mathcal{G}}}_{t},\text{ for all }0\leq t\leq T.$
+For any $\widetilde{F}_{t}\in\widetilde{\mathcal{F}}_{t}$, $\widetilde{G}_{t}\in{\widetilde{\mathcal{G}}}_{t}$ and $\widetilde{G}_{T}\in{\widetilde{\mathcal{G}}}_{T}$, there exist $F_{t}\in{\mathcal{F}}_{t}$, $G_{t}\in{\mathcal{G}}_{t}$, and $G_{T}\in{\mathcal{G}}_{T}$ such that $\widetilde{F}_{t}=F_{t}\times\Omega\times\Omega$, $\widetilde{G}_{t}=G_{t}\times\Omega\times\Omega$ and $\widetilde{G}_{T}=G_{T}\times\Omega\times\Omega$.
 Then,
-¯$\overline{{\mathbb{P}}}(\widetilde{F}_{t}\cap\widetilde{G}_{t}\cap\widetilde{G}_{T})$ =¯$=\overline{\mathbb{P}}((F_{t}\cap G_{t}\cap G_{T})\times\Omega\times\Omega)={\mathbb{P}}(F_{t}\cap G_{t}\cap G_{T})$
+$\overline{{\mathbb{P}}}(\widetilde{F}_{t}\cap\widetilde{G}_{t}\cap\widetilde{G}_{T})$ $=\overline{{\mathbb{P}}}((F_{t}\cap G_{t}\cap G_{T})\times\Omega\times\Omega)={\mathbb{P}}(F_{t}\cap G_{t}\cap G_{T})$
 $={\mathbb{E}}\big{[}{\mathbb{P}}(F_{t}\cap G_{T}|{\mathcal{G}}_{t})\mathbbm{1}_{G_{t}}\big{]}={\mathbb{E}}\big{[}{\mathbb{P}}(F_{t}|{\mathcal{G}}_{t}){\mathbb{P}}(G_{T}|{\mathcal{G}}_{t})\mathbbm{1}_{G_{t}}\big{]}$
 $=\overline{{\mathbb{E}}}\big{[}{\mathbb{P}}(F_{t}|{\mathcal{G}}_{t}){\mathbb{P}}(G_{T}|{\mathcal{G}}_{t})\mathbbm{1}_{\widetilde{G}_{t}}\big{]}=\overline{{\mathbb{E}}}\big{[}\overline{{\mathbb{P}}}(\widetilde{F}_{t}|\widetilde{\mathcal{G}}_{t})\overline{{\mathbb{P}}}(\widetilde{G}_{T}|\widetilde{\mathcal{G}}_{t})\mathbbm{1}_{\widetilde{G}_{t}}\big{]},$
-where the fourth equation holds since ${\mathcal{F}}_{t}\perp\!\!\!\perp{\mathcal{G}}_{T}|{\mathcal{G}}_{t}$, the fifth equation holds because of the same calculation as (5.5) and the last equation holds because of (2) in Proposition 5.1.
+where the fourth equality holds by ${\mathcal{F}}_{t}\perp\!\!\!\perp{\mathcal{G}}_{T}|{\mathcal{G}}_{t}$, the fifth equality follows from the same calculation as (5.5), and the last equality is by (2) in Proposition 5.1.
 Therefore,
-¯$\overline{\mathbb{P}}(\widetilde{F}_{t}\cap\widetilde{G}_{T}|\widetilde{\mathcal{G}}_{t})=\overline{\mathbb{P}}(\widetilde{F}_{t}|\widetilde{\mathcal{G}}_{t})\overline{\mathbb{P}}(\widetilde{G}_{T}|\widetilde{\mathcal{G}}_{t}),$
-and thus completes the proof.∎
-5.2. Proof of Theorem 3.3
-For (1), due to the decomposablity (see [37, Theorem 1, Chapter 4]) let us assume that
+$\overline{{\mathbb{P}}}(\widetilde{F}_{t}\cap\widetilde{G}_{T}|\widetilde{\mathcal{G}}_{t})=\overline{{\mathbb{P}}}(\widetilde{F}_{t}|\widetilde{\mathcal{G}}_{t})\overline{{\mathbb{P}}}(\widetilde{G}_{T}|\widetilde{\mathcal{G}}_{t}).$
+5.2. Proof of Proposition 3.3
+For (1), let us assume that ([51, Theorem 1, Chapter 4])
 $X_{t}=M_{t}+V_{t},$
-where $M_{t}$ is an ${\mathbb{F}}$-local martingale and $V_{t}$ is finite variation process.
-To prove that $X$ is a $\overline{\mathbb{F}}$-semimartingale, it suffices to show that $M_{t}$ is a $\overline{\mathbb{F}}$-local martingale.
-Let $\{\tau_{k}\}_{k\in{\mathbb{N}}^{+}}$ be a sequence of ${\mathbb{F}}$-stopping times such that $M^{\tau_{k}}:=(M_{t\wedge\tau_{k}})_{t\in[0,T]}$ is an ${\mathbb{F}}$-martingale for all $k\in{\mathbb{N}}^{+}$ and as $k\to\infty$, we have $\tau_{k}\to\infty$.
+where $M_{t}$ is an $\mathbb{F}$}-local martingale and $V_{t}$ is a finite variation process.
+To prove that $X$ is an $\overline{{\mathbb{F}}}$-semimartingale, it suffices to show that $M_{t}$ is an $\overline{{\mathbb{F}}}$-local martingale.
+Let $\{\tau_{k}\}_{k\in{\mathbb{N}}^{+}}$ be a sequence of $\mathbb{F}$}-stopping times such that $M^{\tau_{k}}:=(M_{t\wedge\tau_{k}})_{t\in[0,T]}$ is an $\mathbb{F}$}-martingale for all $k\in{\mathbb{N}}^{+}$, and $\tau_{k}\to\infty$ as $k\to\infty$.
 We claim that
@@
 Recall that
-¯$\overline{\mathcal{F}}_{s}=\sigma\{A_{0}\times A_{1}\times A_{2}:A_{0},A_{1},A_{2}\in{\mathcal{F}}_{s}\},$
+$\overline{{\mathcal{F}}}_{s}=\sigma\{A_{0}\times A_{1}\times A_{2}:A_{0},A_{1},A_{2}\in{\mathcal{F}}_{s}\},$
 it suffices to show that
@@
 for all $A_{0},A_{1},A_{2}\in{\mathcal{F}}_{s}$.
-Note that $M_{t\wedge\tau_{k}}\mathbbm{1}_{A_{0}\times\Omega\times\Omega}$, $\mathbbm{1}_{\Omega\times A_{1}\times\Omega}$ and $\mathbbm{1}_{\Omega\times\Omega\times A_{2}}$ are conditional independent of given ${\mathcal{G}}_{T}$ by (3) in Proposition 5.1, we have
+Note that $M_{t\wedge\tau_{k}}\mathbbm{1}_{A_{0}\times\Omega\times\Omega}$, $\mathbbm{1}_{\Omega\times A_{1}\times\Omega}$ and $\mathbbm{1}_{\Omega\times\Omega\times A_{2}}$ are conditionally independent given ${\mathcal{G}}_{T}$ by (3) in Proposition 5.1, we have
 $\overline{{\mathbb{E}}}\big{[}M_{t\wedge\tau_{k}}\mathbbm{1}_{A_{0}\times A_{1}\times A_{2}}\big{]}=\overline{{\mathbb{E}}}\Big{[}\overline{{\mathbb{E}}}\big{[}M_{t\wedge\tau_{k}}\mathbbm{1}_{A_{0}\times\Omega\times\Omega}\big{|}{\mathcal{G}}_{T}\big{]}\overline{{\mathbb{E}}}\big{[}\mathbbm{1}_{\Omega\times A_{1}\times\Omega}\big{|}{\mathcal{G}}_{T}\big{]}\overline{{\mathbb{E}}}\big{[}\mathbbm{1}_{\Omega\times\Omega\times A_{2}}\big{|}{\mathcal{G}}_{T}\big{]}\Big{]}.$ (5.7)
-By (2) in Proposition 5.1 and ${\mathcal{F}}_{s}\perp\!\!\!\perp{\mathcal{G}}_{T}|{\mathcal{G}}_{s}$, we know that
+By (2) in Proposition 5.1 and ${\mathcal{F}}_{s}\perp\!\!\!\perp{\mathcal{G}}_{T}|{\mathcal{G}}_{s}$,
 $\overline{{\mathbb{E}}}\big{[}\mathbbm{1}_{\Omega\times A_{1}\times\Omega}\big{|}{\mathcal{G}}_{T}\big{]}=\overline{{\mathbb{P}}}(A_{1}|{\mathcal{G}}_{T})=\overline{{\mathbb{P}}}(A_{1}|{\mathcal{G}}_{s}),\text{ and similarly }\overline{{\mathbb{E}}}\big{[}\mathbbm{1}_{\Omega\times\Omega\times A_{2}}\big{|}{\mathcal{G}}_{T}\big{]}=\overline{{\mathbb{P}}}(A_{2}|{\mathcal{G}}_{s}).$ (5.8)
-Plugging back in (5.7), we have
+Therefore (5.7) becomes
 $\overline{{\mathbb{E}}}\big{[}M_{t\wedge\tau_{k}}\mathbbm{1}_{A_{0}\times A_{1}\times A_{2}}\big{]}$ $=\overline{{\mathbb{E}}}\Big{[}\overline{{\mathbb{E}}}\big{[}M_{t\wedge\tau_{k}}\mathbbm{1}_{A_{0}\times\Omega\times\Omega}\big{|}{\mathcal{G}}_{T}\big{]}\overline{{\mathbb{P}}}\big{(}A_{1}|{\mathcal{G}}_{s}\big{)}\overline{{\mathbb{P}}}\big{(}A_{2}|{\mathcal{G}}_{s}\big{)}\Big{]}$ (5.9)
-$=\overline{\mathbb{E}}\Big{[}\overline{\mathbb{E}}\big{[}M_{t\wedge\tau_{k}}\mathbbm{1}_{A_{0}\times\Omega\times\Omega}\big{|}{\mathcal{G}}_{s}\big{]}\overline{\mathbb{P}}\big{(}A_{1}|{\mathcal{G}}_{s}\big{)}\overline{\mathbb{P}}\big{(}A_{2}|{\mathcal{G}}_{s}\big{)}\Big{]}$
-where the second inequality holds by the tower rule. Since $M^{\tau_{k}}$ is a ${\mathcal{F}}$-martingale, we get
+$=\overline{{\mathbb{E}}}\Big{[}\overline{{\mathbb{E}}}\big{[}M_{t\wedge\tau_{k}}\mathbbm{1}_{A_{0}\times\Omega\times\Omega}\big{|}{\mathcal{G}}_{s}\big{]}\overline{{\mathbb{P}}}\big{(}A_{1}|{\mathcal{G}}_{s}\big{)}\overline{{\mathbb{P}}}\big{(}A_{2}|{\mathcal{G}}_{s}\big{)}\Big{]},$
+where the second equality holds by the tower rule. Since $M^{\tau_{k}}$ is an $\mathbb{F}$}-martingale,
 $\overline{{\mathbb{E}}}\big{[}M_{t\wedge\tau_{k}}\mathbbm{1}_{A_{0}\times\Omega\times\Omega}\big{|}{\mathcal{G}}_{s}\big{]}=\overline{{\mathbb{E}}}\big{[}\mathbbm{1}_{A_{0}\times\Omega\times\Omega}{\mathbb{E}}[M_{t\wedge\tau_{k}}|{\mathcal{F}}_{s}]\big{|}{\mathcal{G}}_{s}\big{]}=\overline{{\mathbb{E}}}\big{[}M_{s\wedge\tau_{k}}\mathbbm{1}_{A_{0}\times\Omega\times\Omega}\big{|}{\mathcal{G}}_{s}\big{]}.$
@@
 $\overline{{\mathbb{E}}}\big{[}M_{t\wedge\tau_{k}}\mathbbm{1}_{A_{0}\times A_{1}\times A_{2}}\big{]}$ $=\overline{{\mathbb{E}}}\Big{[}\overline{{\mathbb{E}}}\big{[}M_{s\wedge\tau_{k}}\mathbbm{1}_{A_{0}\times\Omega\times\Omega}\big{|}{\mathcal{G}}_{s}\big{]}\overline{{\mathbb{P}}}\big{(}A_{1}|{\mathcal{G}}_{s}\big{)}\overline{{\mathbb{P}}}\big{(}A_{2}|{\mathcal{G}}_{s}\big{)}\Big{]}.$
-Repeating the calculation as in (5.7), (5.8) and (5.9) we achieve
+Repeating the calculation as in (5.7), (5.8), and (5.9) yields
 $\overline{{\mathbb{E}}}\big{[}M_{s\wedge\tau_{k}}\mathbbm{1}_{A_{0}\times A_{1}\times A_{2}}\big{]}$ $=\overline{{\mathbb{E}}}\Big{[}\overline{{\mathbb{E}}}\big{[}M_{s\wedge\tau_{k}}\mathbbm{1}_{A_{0}\times\Omega\times\Omega}\big{|}{\mathcal{G}}_{s}\big{]}\overline{{\mathbb{P}}}\big{(}A_{1}|{\mathcal{G}}_{s}\big{)}\overline{{\mathbb{P}}}\big{(}A_{2}|{\mathcal{G}}_{s}\big{)}\Big{]},$
 which leads to (5.6).
-To show that $X^{\prime}$ is an $\overline{\mathbb{F}}$-semimartingale, by the construction of the extended probability space, we can define $M^{\prime},V^{\prime}$, and $\tau_{k}^{\prime}$ to be respectively the conditional independent copy of $M,V$, and $\tau_{k}$.
-Then it suffices to show that it suffices to show that
+To show that $X^{\prime}$ is an $\overline{{\mathbb{F}}}$-semimartingale, by the construction of the extended probability space, we can define $M^{\prime},V^{\prime}$, and $\tau_{k}^{\prime}$ to be respectively the conditionally independent copy of $M,V$, and $\tau_{k}$.
+Then it suffices to show that
 ${\mathbb{E}}\big{[}M^{\prime}_{t\wedge\tau^{\prime}_{k}}\mathbbm{1}_{A_{0}\times A_{1}\times A_{2}}\big{]}={\mathbb{E}}\big{[}M^{\prime}_{s\wedge\tau^{\prime}_{k}}\mathbbm{1}_{A_{0}\times A_{1}\times A_{2}}\big{]}.$
-Similarly as the procedure of showing (5.6), we have
+Note that
 $\overline{{\mathbb{E}}}\big{[}M^{\prime}_{t\wedge\tau_{k}^{\prime}}\mathbbm{1}_{A_{0}\times A_{1}\times A_{2}}\big{]}=\overline{{\mathbb{E}}}\Big{[}\overline{{\mathbb{E}}}\big{[}M^{\prime}_{t\wedge\tau^{\prime}_{k}}\mathbbm{1}_{\Omega\times A_{2}\times\Omega}\big{|}{\mathcal{G}}_{T}\big{]}\overline{{\mathbb{P}}}(A_{0}|{\mathcal{G}}_{s})\overline{{\mathbb{P}}}(A_{2}|{\mathcal{G}}_{s})\Big{]}$
@@
 $=\overline{{\mathbb{E}}}\Big{[}\overline{{\mathbb{E}}}\big{[}M^{\prime}_{s\wedge\tau^{\prime}_{k}}\mathbbm{1}_{\Omega\times A_{2}\times\Omega}\big{|}{\mathcal{G}}_{T}\big{]}\overline{{\mathbb{P}}}(A_{0}|{\mathcal{G}}_{s})\overline{{\mathbb{P}}}(A_{2}|{\mathcal{G}}_{s})\Big{]}={\mathbb{E}}\big{[}M^{\prime}_{s\wedge\tau^{\prime}_{k}}\mathbbm{1}_{A_{0}\times A_{1}\times A_{2}}\big{]},$
-where the first and last equation mimic the argument in (5.7), (5.8) and (5.9), the second and the sixth equation hold since (2) in Proposition 5.1, the third and the fifth equation is because of the tower rule and the fourth equation follows from the martingale property.
+where the first and last equalities mimic the argument in (5.7), (5.8) and (5.9), the second and the sixth equalities hold from (2) in Proposition 5.1, the third and the fifth equalities follow from the tower rule, and the fourth equality is the martingale property.
 This completes the proof of (1).
@@
 which leads to (2a).
-For (2b), due to (2a) and the decomposablity (see [37, Theorem 1, Chapter 4]), it suffices to show that if $X$ is a ${\mathbb{F}}$-local martingale, then $X^{\mathcal{G}}$ is a local martingale with respect to $\overline{\mathbb{F}}$, ${\mathbb{F}}$ and ${\mathbb{G}}$ and thanks to (1), it can be further reduced to prove that $X^{\mathcal{G}}$ is a local martingale with respect to ${\mathbb{F}}$ and ${\mathbb{G}}$.
+For (2b), due to (2a) and the decomposability (see [51, Theorem 1, Chapter 4]), it suffices to show that if $X$ is an $\mathbb{F}$}-local martingale, then $X^{\mathcal{G}}$ is a local martingale with respect to $\overline{{\mathbb{F}}}$, $\mathbb{F}$} and $\mathbb{G}$}, which thanks to (1) can be further reduced to proving that $X^{\mathcal{G}}$ is a local martingale with respect to $\mathbb{F}$} and $\mathbb{G}$}.
 Let $\tau_{k}$ be the stopping time $\tau_{k}:=\inf\{t\in[0,T]:[X,X]_{t}>k\}$ and consider the truncated process $X^{\tau_{k}}:=\{X_{t\wedge\tau_{k}}\}_{t\in[0,T]}$.
@@
 $\overline{{\mathbb{E}}}\big{[}(X^{\tau_{k}})^{*}_{t}\big{]}^{2}\leq\overline{{\mathbb{E}}}\big{[}\big{(}(X^{\tau_{k}}\big{)}^{*}_{t})^{2}\big{]}\leq C\overline{{\mathbb{E}}}\big{[}\big{[}X^{\tau_{k}},X^{\tau_{k}}\big{]}_{t}\big{]}\leq k,$ (5.10)
-for some constant $C$ and $(X^{\tau_{k}})^{*}_{t}=\sup_{0\leq s\leq t}|X_{s\wedge\tau_{k}}|.$ Let $\{\widetilde{\tau}_{\ell}\}_{\ell\in{\mathbb{N}}^{+}}$ be the sequence of the ${\mathbb{F}}$-stopping times such that $\tau_{\ell}\to T$ and $X^{\tilde{\tau}_{\ell}}=\{X_{t\wedge\tilde{\tau}_{\ell}}\}_{t\in[0,T]}$ are ${\mathbb{F}}$-martingales for all $\ell\in{\mathbb{N}}^{+}$.
-Then $X^{\tilde{\tau}_{\ell}\wedge\tau_{k}}=\{X_{t\wedge\tilde{\tau}_{\ell}\wedge\tau_{k}}\}_{t\in[0,T]}$ are ${\mathbb{F}}$-martingales for all $\ell\in{\mathbb{N}}^{+}$.
-Using conditional dominate convergence theorem, $|X_{t\wedge\tau_{k}\wedge\tilde{\tau}_{\ell}}|\leq(X^{\tau_{k}})^{*}_{t}$ and (5.10), we have
+for some constant $C$ and $(X^{\tau_{k}})^{*}_{t}=\sup_{0\leq s\leq t}|X_{s\wedge\tau_{k}}|.$ Let $\{\widetilde{\tau}_{\ell}\}_{\ell\in{\mathbb{N}}^{+}}$ be the sequence of the $\mathbb{F}$}-stopping times such that $\tau_{\ell}\to T$ and $X^{\tilde{\tau}_{\ell}}=\{X_{t\wedge\tilde{\tau}_{\ell}}\}_{t\in[0,T]}$ is an $\mathbb{F}$}-martingale for any $\ell\in{\mathbb{N}}^{+}$.
+Then $X^{\tilde{\tau}_{\ell}\wedge\tau_{k}}=\{X_{t\wedge\tilde{\tau}_{\ell}\wedge\tau_{k}}\}_{t\in[0,T]}$ is an $\mathbb{F}$}-martingale for any $\ell\in{\mathbb{N}}^{+}$.
+By Lemma 2.5, $|X_{t\wedge\tau_{k}\wedge\tilde{\tau}_{\ell}}|\leq(X^{\tau_{k}})^{*}_{t}$, and from (5.10),
 $X_{s\wedge\tau_{k}}=\lim_{\ell\to\infty}X_{s\wedge\tau_{k}\wedge\tilde{\tau}_{\ell}}=\lim_{\ell\to\infty}\overline{{\mathbb{E}}}\big{[}X_{t\wedge\tau_{k}\wedge\tilde{\tau}_{\ell}}|{\mathcal{F}}_{s}\big{]}=\overline{{\mathbb{E}}}\Big{[}\lim_{\ell\to\infty}X_{t\wedge\tau_{k}\wedge\tilde{\tau}_{\ell}}|{\mathcal{F}}_{s}\Big{]}=\overline{{\mathbb{E}}}\big{[}X_{t\wedge\tau_{k}}|{\mathcal{F}}_{s}\big{]}.$
-Combining with (5.10), we have that $X^{\tau_{k}}=\{X_{t\wedge\tau_{k}}\}_{t\in[0,T]}$ is indeed a square integrable ${\mathbb{F}}$-martingale.
-Now we claim that $(X^{\tau_{k}})^{\mathcal{G}}$ is a ${\mathbb{G}}$-martingale, and using the property of H hyphothesis, and the compatiability condition of ${\mathbb{F}}$ and ${\mathbb{G}}$ (i.e.
-${\mathcal{F}}_{t}\perp\!\!\!\perp{\mathcal{G}}_{T}|{\mathcal{G}}_{t}$ for all $t\in[0,T]$), it is also an ${\mathbb{F}}$-martingale (see [20]).
-The integrability holds trivially since $X_{t}$ is integrable.
-And the martingale property holds since
+Combined with (5.10), we have that $X^{\tau_{k}}=\{X_{t\wedge\tau_{k}}\}_{t\in[0,T]}$ is indeed a square integrable $\mathbb{F}$}-martingale.
+Now we claim that $(X^{\tau_{k}})^{\mathcal{G}}$ is a square integrable $\mathbb{G}$}-martingale.
+(Hence, by the H hypothesis and Remark 2.7 also an $\mathbb{F}$}-martingale).
+The integrability holds trivially since $X_{t}$ is integrable, and the martingale property holds since
 $\overline{{\mathbb{E}}}\Big{[}X_{t\wedge\tau_{k}}^{\mathcal{G}}\big{|}{\mathcal{G}}_{s}\Big{]}$ $=\overline{{\mathbb{E}}}\Big{[}\overline{{\mathbb{E}}}\big{[}X_{t\wedge\tau_{k}}|{\mathcal{G}}_{T}\big{]}\big{|}{\mathcal{G}}_{s}\Big{]}=\overline{{\mathbb{E}}}\big{[}X_{t\wedge\tau_{k}}|{\mathcal{G}}_{s}\big{]}=\overline{{\mathbb{E}}}\Big{[}\overline{{\mathbb{E}}}\big{[}X_{t\wedge\tau_{k}}|{\mathcal{F}}_{s}\big{]}\big{|}{\mathcal{G}}_{s}\Big{]}$
 $=\overline{{\mathbb{E}}}\big{[}X_{s\wedge\tau_{k}}|{\mathcal{G}}_{s}\big{]}=\overline{{\mathbb{E}}}\big{[}X_{s\wedge\tau_{k}}|{\mathcal{G}}_{T}\big{]}=X_{s\wedge\tau_{k}}^{\mathcal{G}}.$
-This completes the proof of the claim that $X^{\mathcal{G}}$ is an ${\mathbb{G}}$-martingale and thus leads to(2b).
-For (2c), let $\sigma_{m}:0=t_{0}^{m}\leq t_{1}^{m}\leq...\leq t_{k_{m}}^{m}=t$ be a sequence of partitions tend to identity (the definition of a sequence tending to identity can be found in the [37, Page 64, defintion] with a slight modification of the terminal time), such that
+This completes the proof that $X^{\mathcal{G}}$ is an $\mathbb{G}$}-local martingale and (2b).
+5.3. Proof of Theorem 3.5
+For (1), let $\sigma_{m}:0=t_{0}^{m}\leq t_{1}^{m}\leq...\leq t_{k_{m}}^{m}=t$ be a sequence of partitions which tends to identity in the sense of [51, Page 64, definition] and satisfies
 $\int_{0}^{t}Z_{s-}{\mathrm{d}}X^{\mathcal{G}}_{s}=\lim_{m\to\infty}\sum_{k=0}^{k_{m}-1}Z_{t_{k}^{m}}(X^{\mathcal{G}}_{t_{k+1}^{m}}-X^{\mathcal{G}}_{t_{k}^{m}}),\;\;\int_{0}^{t}Z_{s-}{\mathrm{d}}X^{\prime}_{s}=\lim_{m\to\infty}\sum_{k=0}^{k_{m}-1}Z_{t_{k}^{m}}(X^{\prime}_{t_{k+1}^{m}}-X^{\prime}_{t_{k}^{m}}),$
-where the convergence is understood as almost surely convergence. Such partition can be found thanks to [37, Theorem 21, Chapter II]. Then
+where the convergence is understood as a.s. convergence. Then
 $\int_{0}^{t}Z_{s-}{\mathrm{d}}X^{\mathcal{G}}_{s}$ $=\lim_{m\to\infty}\sum_{k=0}^{k_{m}-1}Z_{t_{k}^{m}}(X^{\mathcal{G}}_{t_{k+1}^{m}}-X^{\mathcal{G}}_{t_{k}^{m}})=\lim_{m\to\infty}\sum_{k=0}^{k_{m}-1}Z_{t_{k}^{m}}{\mathbb{E}}\Big{[}X_{t_{k+1}^{m}}-X_{t_{k}^{m}}\Big{|}{\mathcal{G}}_{T}\Big{]}$
 $=\lim_{m\to\infty}\sum_{k=0}^{k_{m}-1}Z_{t_{k}^{m}}\overline{{\mathbb{E}}}\Big{[}X^{\prime}_{t_{k+1}^{m}}-X^{\prime}_{t_{k}^{m}}\Big{|}{\mathcal{F}}\Big{]}=\lim_{m\to\infty}\sum_{k=0}^{k_{m}-1}{\mathbb{E}}\Big{[}Z_{t_{k}^{m}}(X^{\prime}_{t_{k+1}^{m}}-X^{\prime}_{t_{k}^{m}})\Big{|}{\mathcal{F}}\Big{]}.$
-Define ${Z}^{m}$ as the process
+Define ${Z}^{m}$ as
 $Z^{m}_{t}=Z_{0}\mathbbm{1}_{\{0\}}(t)+\sum_{k=0}^{k_{m}-1}Z_{t_{k}^{m}}\mathbbm{1}_{(t_{k}^{m},t_{k+1}^{m}]}(t),$
@@
 $\sum_{k=0}^{k_{m}-1}\overline{{\mathbb{E}}}\Big{[}Z_{t_{k}^{m}}\big{(}X^{\prime}_{t_{k+1}^{m}}-X^{\prime}_{t_{k}^{m}}\big{)}\Big{]}-\overline{{\mathbb{E}}}\bigg{[}\int_{0}^{t}Z_{s-}{\mathrm{d}}X^{\prime}_{s}\bigg{]}=\overline{{\mathbb{E}}}\bigg{[}\int_{0}^{t}(Z_{s}^{m}-Z_{s-}){\mathrm{d}}X^{\prime}_{s}\bigg{]}.$
-Recall that $X^{\prime}$ has the decomposition $X^{\prime}=M^{\prime}+V^{\prime}$ where $M^{\prime}$ is a local martingale and $V^{\prime}$ being a adapted finite variation process and
-$\bigg{\|}\sqrt{[M^{\prime},M^{\prime}]_{T}}+\int_{0}^{T}\big{|}{\mathrm{d}}V^{\prime}_{s}\big{|}\bigg{\|}_{L^{p}}<\infty.$
-Therefore,
+Recall that $X^{\prime}$ has the decomposition $X^{\prime}=M^{\prime}+V^{\prime}$, with $M^{\prime}$ a local martingale, $V^{\prime}$ an adapted finite variation process, and $\bigg{\|}\sqrt{[M^{\prime},M^{\prime}]_{T}}+\int_{0}^{T}\big{|}{\mathrm{d}}V^{\prime}_{s}\big{|}\bigg{\|}_{L^{p}}<\infty.$ Therefore,
 $\overline{{\mathbb{E}}}\bigg{[}\bigg{|}\int_{0}^{t}(Z_{s}^{m}-Z^{\prime\prime}_{s-}){\mathrm{d}}X^{\prime}_{s}\bigg{|}\bigg{]}\leq\overline{{\mathbb{E}}}\bigg{[}\bigg{|}\int_{0}^{t}(Z_{s}^{m}-Z_{s-}){\mathrm{d}}M^{\prime}_{s}\bigg{|}\bigg{]}+\overline{{\mathbb{E}}}\bigg{[}\bigg{|}\int_{0}^{t}(Z_{s}^{m}-Z_{s-}){\mathrm{d}}V^{\prime}_{s}\bigg{|}\bigg{]}.$ (5.11)
-For the first term in (5.11), we use Burkholder-Davis-Gundy inequality so that
+For the first term in (5.11), applying Burkholder-Davis-Gundy inequality to get
 $\overline{{\mathbb{E}}}\bigg{[}\bigg{|}\int_{0}^{t}(Z_{s}^{m}-Z_{s-}){\mathrm{d}}M^{\prime}_{s}\bigg{|}\bigg{]}$ $\leq\overline{{\mathbb{E}}}\bigg{[}\sup_{0\leq s\leq t}\bigg{|}\int_{0}^{s}(Z_{u}^{m}-Z_{u-}){\mathrm{d}}M^{\prime}_{u}\bigg{|}^{p}\bigg{]}$
@@
 for some constant $c_{1}$. Note that $\int_{0}^{t}(Z_{u}^{m}-Z_{u-})^{2}{\mathrm{d}}[M^{\prime},M^{\prime}]_{s}\leq 4\sup_{0\leq s\leq t}(Z_{s})^{2}[M^{\prime},M^{\prime}]_{t}$, and
-$\overline{\mathbb{E}}\Big{[}\Big{(}\sup_{0\leq s\leq t}(Z_{s})^{2}[M^{\prime},M^{\prime}]_{t}\Big{)}^{1/2}\Big{]}\leq\bigg{(}\overline{\mathbb{E}}\Big{[}\sup_{0\leq s\leq t}(Z_{s})^{q}\Big{]}\overline{\mathbb{E}}\big{[}[M^{\prime},M^{\prime}]^{p/2}_{t}\big{]}\bigg{)}^{1/2}\leq C\sqrt{\|Z\|^{q}_{{\mathcal{H}}^{q}}\overline{\mathbb{E}}\big{[}[M^{\prime},M^{\prime}]_{t}^{p/2}\big{]}}<\infty$
-Since $Z$ is a càdlàg function, we know that $Z^{m}_{s}\to Z_{s-},a.s.$ and $|Z^{m}_{s}-Z_{s}|\leq\sup_{0\leq u\leq s}|Z_{u}|\in S_{q}$, and we can use the dominated convergence theorem (the classical one and the one in ${\mathcal{H}}_{p}$, see [37, Page 273, Lemma]) to get
+$\overline{{\mathbb{E}}}\Big{[}\Big{(}\sup_{0\leq s\leq t}(Z_{s})^{2}[M^{\prime},M^{\prime}]_{t}\Big{)}^{1/2}\Big{]}\leq\bigg{(}\overline{{\mathbb{E}}}\Big{[}\sup_{0\leq s\leq t}(Z_{s})^{q}\Big{]}\overline{{\mathbb{E}}}\big{[}[M^{\prime},M^{\prime}]^{p/2}_{t}\big{]}\bigg{)}^{1/2}\leq C\sqrt{\|Z\|^{q}_{{\mathcal{H}}^{q}}\overline{{\mathbb{E}}}\big{[}[M^{\prime},M^{\prime}]_{t}^{p/2}\big{]}}<\infty.$
+Since $Z$ is càdlàg, $Z^{m}_{s}\to Z_{s-},a.s.$ and $|Z^{m}_{s}-Z_{s}|\leq\sup_{0\leq u\leq s}|Z_{u}|\in S_{q}$; and Lemma 2.5 leads to
 $\overline{{\mathbb{E}}}\bigg{[}\int_{0}^{t}(Z_{s}^{m}-Z_{s-}){\mathrm{d}}M^{\prime}_{s}\bigg{]}\to 0.$
@@
 $\overline{{\mathbb{E}}}\bigg{[}\bigg{|}\int_{0}^{t}(Z_{s}^{m}-Z_{s-}){\mathrm{d}}V^{\prime}_{s}\bigg{|}\bigg{]}\leq\overline{{\mathbb{E}}}\bigg{[}\int_{0}^{t}|Z_{s}^{m}-Z_{s-}|\big{|}{\mathrm{d}}V^{\prime}_{s}\big{|}\bigg{]}.$
-Noting that $\int_{0}^{t}|Z_{s}^{m}-Z_{s-}|\big{|}{\mathrm{d}}V^{\prime}_{s}\big{|}\leq 2\sup_{0\leq s\leq t}|Z_{s}|\int_{0}^{t}|{\mathrm{d}}V_{t}|$, and similarly, we have
+Note that $\int_{0}^{t}|Z_{s}^{m}-Z_{s-}|\big{|}{\mathrm{d}}V^{\prime}_{s}\big{|}\leq 2\sup_{0\leq s\leq t}|Z_{s}|\int_{0}^{t}|{\mathrm{d}}V_{t}|$, and similarly, we have
 $\overline{{\mathbb{E}}}\bigg{[}\bigg{|}\int_{0}^{t}(Z_{s}^{m}-Z_{s-}){\mathrm{d}}V^{\prime}_{s}\bigg{|}\bigg{]}\to 0,$
-and therefore,
+and
 $\lim_{m\to\infty}\overline{{\mathbb{E}}}\bigg{[}\bigg{|}\int_{0}^{t}Z_{s}^{m}{\mathrm{d}}X^{\prime}_{s}-\int_{0}^{t}Z_{s-}{\mathrm{d}}X^{\prime}_{s}\bigg{|}\bigg{]}=0.$
@@
 Let $A=\{R-R_{m}>\epsilon\}$, we have
-¯$\overline{\mathbb{P}}\big{(}R_{m}-R>\epsilon\big{)}\leq\frac{\overline{\mathbb{E}}[(R_{m}-R)\mathbbm{1}_{A}]}{\epsilon}\to 0,$
+$\overline{{\mathbb{P}}}\big{(}R_{m}-R>\epsilon\big{)}\leq\frac{\overline{{\mathbb{E}}}[(R_{m}-R)\mathbbm{1}_{A}]}{\epsilon}\to 0,$
 and similarly
-¯$\overline{\mathbb{P}}\big{(}R_{m}-R<\epsilon\big{)}\to 0.$
-Therefore $R_{m}\to R$ in probability. Meanwhile $R_{m}\to\int_{0}^{t}Z_{s-}{\mathrm{d}}X^{\mathcal{G}}_{s}$ almost surely. Therefore,
+$\overline{{\mathbb{P}}}\big{(}R_{m}-R<\epsilon\big{)}\to 0.$
+Therefore, $R_{m}\to R$ in probability. Meanwhile, $R_{m}\to\int_{0}^{t}Z_{s-}{\mathrm{d}}X^{\mathcal{G}}_{s}$ a.s., hence
 $\int_{0}^{t}Z_{s-}{\mathrm{d}}X^{\mathcal{G}}_{s}=\overline{{\mathbb{E}}}\bigg{[}\int_{0}^{t}Z^{m}_{s-}{\mathrm{d}}X^{\prime}_{s}\Big{|}{\mathcal{F}}\bigg{]}.$
-For (2d), let us first prove that
-$[X^{\mathcal{G}},Y^{\mathcal{G}}]_{t}=\overline{\mathbb{E}}\Big{[}[X^{\prime\prime},Y^{\prime}]_{t}\Big{|}{\mathcal{G}}\Big{]}.$
+For (2), let us first prove that
+$X^{\mathcal{G}},Y^{\mathcal{G}}]_{t}=\overline{{\mathbb{E}}}\Big{[}[X^{\prime\prime},Y^{\prime}]_{t}\Big{|}{\mathcal{G}}\Big{]}.$
 By the definition of the quadratic variation,
 $\big{[}X^{\mathcal{G}},Y^{\mathcal{G}}\big{]}_{t}=X^{\mathcal{G}}_{t}Y^{\mathcal{G}}_{t}-\int_{0}^{t}X^{\mathcal{G}}_{s-}{\mathrm{d}}Y^{\mathcal{G}}_{s}-\int_{0}^{t}Y^{\mathcal{G}}_{s-}{\mathrm{d}}X^{\mathcal{G}}_{s},$
-where the last two terms are well defined because of (2b). The first term $X^{\mathcal{G}}_{t}Y^{\mathcal{G}}_{t}$ can be calculated by
+where the last two terms are well defined because of (2b). The first term $X^{\mathcal{G}}_{t}Y^{\mathcal{G}}_{t}$ is
 $X^{\mathcal{G}}_{t}Y^{\mathcal{G}}_{t}=\overline{{\mathbb{E}}}\big{[}X_{t}|{\mathcal{G}}_{T}\big{]}\overline{{\mathbb{E}}}\big{[}Y_{t}|{\mathcal{G}}_{T}\big{]}=\overline{{\mathbb{E}}}\big{[}X^{\prime\prime}_{t}|{\mathcal{F}}\big{]}\overline{{\mathbb{E}}}\big{[}Y^{\prime}_{t}|{\mathcal{F}}\big{]}=\overline{{\mathbb{E}}}\big{[}X^{\prime\prime}_{t}Y^{\prime}_{t}|{\mathcal{F}}\big{]},$
 where the last equation holds due to (3) in Proposition 5.1.
-For the second term $\int_{0}^{t}X^{\mathcal{G}}_{s-}{\mathrm{d}}Y^{\mathcal{G}}_{s}$, let $\sigma_{m}:0=t_{0}^{m}\leq t_{1}^{m}\leq...\leq t_{k_{m}}^{m}=t$ be a sequence of partitions tend to identity in the sense of [37, Page 64, defintion], with a slight modification of the terminal time, such that
+For the second term $\int_{0}^{t}X^{\mathcal{G}}_{s-}{\mathrm{d}}Y^{\mathcal{G}}_{s}$, let $\sigma_{m}:0=t_{0}^{m}\leq t_{1}^{m}\leq...\leq t_{k_{m}}^{m}=t$ be a sequence of partitions tending to identity in the sense of [51, Page 64, definition] and satisfying
 $\int_{0}^{t}X^{\mathcal{G}}_{s-}{\mathrm{d}}Y^{\mathcal{G}}_{s}=\lim_{m\to\infty}\sum_{k=0}^{k_{m}-1}X^{\mathcal{G}}_{t_{k}^{m}}(Y^{G}_{t_{k+1}^{m}}-Y^{\mathcal{G}}_{t_{k}^{m}}),\;\;\int_{0}^{t}X^{\prime\prime}_{s-}{\mathrm{d}}Y^{\prime}_{s}=\lim_{m\to\infty}\sum_{k=0}^{k_{m}-1}X^{\prime\prime}_{t_{k}^{m}}(Y^{\prime}_{t_{k+1}^{m}}-Y^{\prime}_{t_{k}^{m}}),$
-where the convergence is understood as almost surely convergence. Therefore,
+where the convergence is understood as a.s. convergence. Therefore,
 $\int_{0}^{t}X^{\mathcal{G}}_{s-}{\mathrm{d}}Y^{\mathcal{G}}_{s}$ $=\lim_{m\to\infty}\sum_{k=0}^{k_{m}-1}X^{\mathcal{G}}_{t_{k}^{m}}(Y^{\mathcal{G}}_{t_{k+1}^{m}}-Y^{\mathcal{G}}_{t_{k}^{m}})=\lim_{m\to\infty}\sum_{k=0}^{k_{m}-1}{\mathbb{E}}\big{[}X_{t_{k}^{m}}|{\mathcal{G}}_{T}\big{]}\overline{{\mathbb{E}}}\big{[}Y_{t_{k+1}^{m}}-Y_{t_{k}^{m}}|{\mathcal{G}}_{T}\big{]}$
 $=\lim_{m\to\infty}\sum_{k=0}^{k_{m}-1}{\mathbb{E}}\big{[}X^{\prime\prime}_{t_{k}^{m}}|{\mathcal{F}}\big{]}{\mathbb{E}}\big{[}Y^{\prime}_{t_{k+1}^{m}}-Y^{\prime}_{t_{k}^{m}}|{\mathcal{F}}\big{]}=\lim_{m\to\infty}\sum_{k=0}^{k_{m}-1}\overline{{\mathbb{E}}}\big{[}X^{\prime\prime}_{t_{k}^{m}}\big{(}Y^{\prime}_{t_{k+1}^{m}}-Y^{\prime}_{t_{k}^{m}}\big{)}|{\mathcal{F}}\big{]},$
-where the last equation holds due to (3) in Proposition 5.1. Similar as the proof of (2c), we can conclude that
+where the last equation holds due to (3) in Proposition 5.1. Similar to the proof of (1), we conclude that
 $\int_{0}^{t}X^{\mathcal{G}}_{s-}{\mathrm{d}}Y^{\mathcal{G}}_{s}=\overline{{\mathbb{E}}}\bigg{[}\int_{0}^{t}X^{\prime\prime}_{s-}{\mathrm{d}}Y^{\prime}_{s}\bigg{|}{\mathcal{F}}\bigg{]},$
@@
 Therefore
-$[X^{\mathcal{G}},Y^{\mathcal{G}}]_{t}=\overline{\mathbb{E}}\Big{[}[X^{\prime\prime},Y^{\prime}]_{t}\Big{|}{\mathcal{F}}\Big{]}.$
-Now let $\sigma_{m}:0=t_{0}^{m}\leq t_{1}^{m}\leq...\leq t_{k_{m}}^{m}=t$ being a sequence of partitions tend to identity in the sense of [37, Page 64, defintion] with a slight modification of the terminal time, such that
+$X^{\mathcal{G}},Y^{\mathcal{G}}]_{t}=\overline{{\mathbb{E}}}\Big{[}[X^{\prime\prime},Y^{\prime}]_{t}\Big{|}{\mathcal{F}}\Big{]}.$
+Meanwhile, we can further assume that the partition $\sigma_{m}:0=t_{0}^{m}\leq t_{1}^{m}\leq...\leq t_{k_{m}}^{m}=t$ satisfies
 $\int_{0}^{t}Z_{s-}{\mathrm{d}}[X^{\mathcal{G}},Y^{\mathcal{G}}]_{s}=\lim_{m\to\infty}\sum_{k=0}^{k_{m}-1}Z_{t_{k}^{m}}\big{(}[X^{\mathcal{G}},Y^{\mathcal{G}}]_{t_{k+1}^{m}}-[X^{\mathcal{G}},Y^{\mathcal{G}}]_{t_{k}^{m}}\big{)},$
 and
-$\int_{0}^{t}Z_{s-}{\mathrm{d}}[X^{\prime\prime},Y^{\prime}]_{s}=\lim_{m\to\infty}\sum_{k=0}^{k_{m}-1}Z_{t_{k}^{m}}([X^{\prime\prime},Y^{\prime}]_{t_{k+1}^{m}}-[X^{\prime\prime},Y^{\prime}]_{t_{k}^{m}}),$
-To repeat the same argument as before, we need to notice that $\big{|}\int_{0}^{t}(Z_{s}^{m}-Z_{s-}){\mathrm{d}}[X^{\prime\prime},Y^{\prime}]_{s}\big{|}\leq 2\sup_{0\leq s\leq t}|Z_{s}|[X^{\prime\prime},Y^{\prime}]_{t}$ and
-[…]
-where the first inequality is due to Cauchy inequality and Kunita-Watanabe Inequality and $C$ is a generic constant. Now repeating the arguments in (2c), we achieve
+$\int_{0}^{t}Z_{s-}{\mathrm{d}}[X^{\prime\prime},Y^{\prime}]_{s}=\lim_{m\to\infty}\sum_{k=0}^{k_{m}-1}Z_{t_{k}^{m}}([X^{\prime\prime},Y^{\prime}]_{t_{k+1}^{m}}-[X^{\prime\prime},Y^{\prime}]_{t_{k}^{m}}).$
+Repeat the same argument as before and notice that
+$\big{|}\int_{0}^{t}(Z_{s}^{m}-Z_{s-}){\mathrm{d}}[X^{\prime\prime},Y^{\prime}]_{s}\big{|}\leq 2\sup_{0\leq s\leq t}|Z_{s}|[X^{\prime\prime},Y^{\prime}]_{t}.$
+Then we have
+$\overline{{\mathbb{E}}}\bigg{[}\sup_{0\leq s\leq t}|Z_{s}|[X^{\prime\prime},Y^{\prime}]_{s}\bigg{]}$ $\leq\overline{{\mathbb{E}}}[(\sup_{0\leq s\leq t}|Z_{s}|)^{r}]^{1/r}\overline{{\mathbb{E}}}\Big{[}[X^{\prime\prime},X^{\prime\prime}]_{t}^{p}\Big{]}^{1/p}\overline{{\mathbb{E}}}\Big{[}[Y^{\prime},Y^{\prime}]_{t}^{q}\Big{]}^{1/q}$
+$\leq C\|Z\|_{H_{r}}\|X\|_{H_{p}}\|Y\|_{H_{q}},$
+where the first inequality is by Cauchy inequality and Kunita-Watanabe inequality with $C$ a generic constant. Now repeating the arguments in (1), we achieve
 $\int_{0}^{t}Z_{s-}{\mathrm{d}}\big{[}X^{\mathcal{G}},Y^{\mathcal{G}}\big{]}_{t}=\overline{{\mathbb{E}}}\bigg{[}\int_{0}^{t}Z_{s-}{\mathrm{d}}\big{[}X^{\prime\prime},Y^{\prime}\big{]}_{s}\bigg{|}{\mathcal{F}}\bigg{]},$
-and similarly, we can get
+and similarly,
 $\int_{0}^{t}Z_{s-}\big{[}X^{\mathcal{G}},Y\big{]}_{s}=\overline{{\mathbb{E}}}\bigg{[}\int_{0}^{t}Z_{s-}\big{[}X^{\prime},Y\big{]}_{s}\bigg{|}{\mathcal{F}}\bigg{]},$
 which completes the proof.∎
+6. Application to control problems with partial observed jump-diffusions
+We now apply our result (Theorem 2.11) to analyze control problems with partially observed jump-diffusions.
+We will derive an appropriate form of the verification theorem and recover the results in [4] as special cases.
+Consider a complete filtered space $(\Omega,{\mathcal{F}},{\mathbb{F}}=({\mathcal{F}}_{t})_{t\in[0,T]},{\mathbb{P}})$ with a sub-filtration ${\mathbb{G}}=({\mathcal{G}}_{t})_{t\in[0,T]}\subset{\mathbb{F}}$ satisfying the compatibility condition as in Assumption 2.6.
+We define as an admissible control process any $\mathbb{G}$}-progressive process $\alpha$ with values in some given Borel space $A$.
+The set of admissible control processes is denoted by ${\mathcal{A}}^{\mathbb{G}}$.
+Given a square-integrable initial condition $\bm{\xi}\in{\mathbb{R}}^{n}$, $d$-dimensional $\mathbb{F}$}-Brownian motion $\bm{B}$, an independent Poisson random measure $N$ with a Lévy measure $\nu$, with $\widetilde{N}(\ d\bm{\theta},{\mathrm{d}}t):=N({\mathrm{d}}\bm{\theta},{\mathrm{d}}t)-\nu({\mathrm{d}}\bm{\theta}){\mathrm{d}}t$ the compensated Poisson random measure and $\alpha\in{\mathcal{A}}^{\mathbb{G}}$, the $n$-dimensional controlled state equation is defined by
+${\mathrm{d}}\bm{X}_{t}^{\alpha}=\bm{b}\big{(}\bm{X}_{t}^{\alpha},\alpha_{t}\big{)}{\mathrm{d}}t+\bm{\sigma}\big{(}\bm{X}_{t}^{\alpha},\alpha_{t}\big{)}{\mathrm{d}}\bm{B}_{t}+\int_{{\mathbb{R}}^{q}}\bm{\beta}(\bm{X}^{\alpha}_{t-},\alpha_{t-},\bm{\theta})\widetilde{N}({\mathrm{d}}\bm{\theta},{\mathrm{d}}t),\text{ for }0\leq t\leq T,\;\bm{X}^{\alpha}_{0}=\bm{\xi}.$ (6.1)
+The coefficients $\bm{b}$ and $\bm{\sigma}$ are deterministic measurable functions from ${\mathbb{R}}^{n}\times A$ into ${\mathbb{R}}^{n}$ and ${\mathbb{R}}^{n\times d}$, respectively, and the coefficient $\bm{\beta}$ are deterministic measurable functions from ${\mathbb{R}}^{n}\times A\times{\mathbb{R}}^{q}$ into ${\mathbb{R}}^{n}$ and ${\mathbb{R}}^{n\times d}$ assumed to satisfy the following standing assumptions: there exists some positive constant $C_{1}$ such that for all $\bm{x},\bm{x}^{\prime}\in{\mathbb{R}}^{n}$, $a\in A$,
+$\big{|}\bm{b}(\bm{x},a)-\bm{b}(\bm{x}^{\prime},a)\big{|}+\big{|}\bm{\sigma}(\bm{x},a)-\bm{\sigma}(\bm{x}^{\prime},a)\big{|}+\int_{{\mathbb{R}}^{q}}\big{|}\bm{\beta}(\bm{x},a,\bm{\theta})-\bm{\beta}(\bm{x}^{\prime},a,\bm{\theta})\big{|}\nu({\mathrm{d}}\bm{\theta})$ $\leq C_{1}|\bm{x}-\bm{x}^{\prime}|;$
+$\big{|}\bm{b}(\bm{0},a)\big{|}+\big{|}\bm{\sigma}(\bm{0},a)\big{|}+\int_{{\mathbb{R}}^{q}}\big{|}\bm{\beta}(\bm{0},a,\bm{\theta})\big{|}\nu({\mathrm{d}}\bm{\theta})$ $\leq C_{1}.$
+It is clear from standard arguments that there exists a unique strong solution $\{\bm{X}_{t}^{\alpha}\}_{t\in[0,T]}$ to (6.1), which is $\mathbb{F}$}-adapted and satisfies
+$\sup_{\alpha\in{\mathcal{A}}^{\mathbb{G}}}{\mathbb{E}}\bigg{[}\sup_{0\leq t\leq T}\big{|}\bm{X}_{t}^{\alpha}\big{|}^{2}\bigg{]}\leq C\Big{(}1+\textasciimacron{\mathbb{E}}\big{[}|\bm{\xi}|^{2}\big{]}\Big{)}<\infty,$
+for some positive constant $C$.
+The goal is to maximize, over all admissible control processes $\alpha\in{\mathcal{A}}^{\mathbb{G}}$, the following objective:
+${\mathcal{J}}(\alpha):={\mathbb{E}}\bigg{[}\int_{0}^{T}f(\bm{X}_{t}^{\alpha},\alpha_{t}){\mathrm{d}}t+g(\bm{X}_{T}^{\alpha})\bigg{]}.$
+Here $f:{\mathbb{R}}^{n}\times A\to{\mathbb{R}}$, and $g:{\mathbb{R}}^{n}\to{\mathbb{R}}$ are continuous functions satisfying the quadratic growth condition: there exists some positive constant $C_{2}$ such that for all $(\bm{x},a)\in{\mathbb{R}}^{n}\times A$,
+$\big{|}f(\bm{x},a)\big{|}+\big{|}g(\bm{x})\big{|}\leq C_{2}\big{(}1+|\bm{x}|^{2}\big{)}.$
+We can now define the objective function starting from time $t$, and for $\mu\in{\mathcal{P}}_{2}({\mathbb{R}}^{d})$ by
+$J(t,\alpha,\mu):={\mathbb{E}}\bigg{[}\int_{t}^{T}f(\bm{X}_{s}^{\alpha},\alpha_{s}){\mathrm{d}}s+g(\bm{X}_{T}^{\alpha})\Big{|}\text{Law}(\bm{X}^{\alpha}_{t}|{\mathcal{G}}_{T})=\text{Law}(\bm{X}^{\alpha}_{t}|{\mathcal{G}}_{t})=\mu\bigg{]},$
+and the corresponding value function by
+$V(t,\mu):=\sup_{\alpha}J(t,\alpha,\mu).$
+It is not hard to see that both $J$ and $V$ satisfy the quadratic growth estimate:
+$|J(t,\alpha,\mu)|\leq|V(t,\mu)|\leq C\Big{(}1+\textasciimacron{\mathbb{E}}\big{[}|\bm{\xi}|^{2}\big{]}\Big{)}$
+for all $\mu\in{\mathcal{P}}_{2}({\mathbb{R}}^{d}),\alpha\in{\mathcal{A}}^{\mathcal{G}}$ for some constant $C$.
+Note that [4] considers the case where the information $\mathbb{G}$} is a filtration generated by a lower-dimensional Brownian motion $\bm{W}$ such that $\bm{B}=(\bm{V},\bm{W})$ and $\bm{V}$ independent of $\bm{W}$.
+This corresponds to the special case of (6.1) with $\bm{\beta}=0$.
+In this case, a form of Itô’s formula without proof is presented (see (5.7) of [4]), from which a verification theorem for the control problem with partial observation is derived.
+Here we will derive rigorously, as a corollary of Theorem 2.11, a generalization of (5.7) in [4] with a general form of $\mathbb{G}$} satisfying the compatibility condition.
+Corollary 6.1 (Itô’s formula for flows of partial observed jump-diffusion).
+[…]
+$\Phi(\mu_{t})-\Phi(\mu_{0})$ (6.2)
+$=$ $\;\overline{{\mathbb{E}}}\Bigg{[}\int_{0}^{t}\bigg{(}\Big{(}\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s},\bm{X}^{\alpha,1}_{s}\big{)}\cdot\bm{b}(\bm{X}^{\alpha,1}_{s},\alpha_{s})+\frac{1}{2}\nabla^{2}_{\bm{x}_{1}\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s},\bm{X}^{\alpha,1}_{s}\big{)}:\bm{\sigma}\bm{\sigma}^{\top}\big{(}\bm{X}^{\alpha,1}_{s},\alpha_{s}\big{)}\Big{)}{\mathrm{d}}s$
+$\qquad\qquad+\frac{1}{2}\nabla^{2}_{\bm{x}_{1}\bm{x}_{2}}\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s},\bm{X}^{\alpha,1}_{s},\bm{X}^{\alpha,2}_{s}\big{)}:\bm{\sigma}\big{(}\bm{X}^{\alpha,1}_{s},\alpha_{s}\big{)}\bm{\sigma}^{\top}\big{(}\bm{X}^{\alpha,2}_{s},\alpha_{s}\big{)}{\mathrm{d}}[\bm{B}^{1},\bm{B}^{2}]_{s}$
+$\qquad\qquad+\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s},\bm{X}^{\alpha,1}_{s}\big{)}\cdot\bm{\sigma}(\bm{X}^{\alpha,1}_{s},\alpha_{s}){\mathrm{d}}\bm{B}^{1}_{s}\bigg{)}$
+$+\sum_{0<s\leq t}\bigg{(}\frac{1}{2}\bigg{(}\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{X}^{\alpha,1}_{s},\bm{X}^{\alpha,2}_{s}\big{)}-\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{X}^{\alpha,1}_{s-},\bm{X}^{\alpha,2}_{s}\big{)}$
+$\qquad\qquad\qquad-\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{X}^{\alpha,1}_{s},\bm{X}^{\alpha,2}_{s-}\big{)}+\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s-},\bm{X}^{\alpha,1}_{s-},\bm{X}^{\alpha,2}_{s-}\big{)}\bigg{)}$
+$\quad\qquad\qquad+\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{X}^{\alpha,1}_{s}\big{)}-\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{X}^{\alpha,1}_{s-}\big{)}\bigg{)}\mathbbm{1}_{\{\mu_{s}=\mu_{s-}\}}\bigg{|}{\mathcal{F}}\bigg{]}$
+$+\sum_{0<s\leq t}\Big{(}\Phi(\mu_{s})-\Phi(\mu_{s-})\Big{)},$
+for all $t\in[0,T]$, where $(\bm{X}^{\alpha,1},\bm{B}^{1},N^{1}),(\bm{X}^{\alpha,1},\bm{B}^{2},N^{2})$ satisfying for $0\leq t\leq T$,
+${\mathrm{d}}\bm{X}_{t}^{\alpha,1}=\bm{b}\big{(}\bm{X}_{t}^{\alpha,1},\alpha_{t}\big{)}{\mathrm{d}}t+\bm{\sigma}\big{(}\bm{X}_{t}^{\alpha,1},\alpha_{t}\big{)}{\mathrm{d}}\bm{B}^{1}_{t}+\int_{{\mathbb{R}}^{q}}\bm{\beta}(\bm{X}^{\alpha,1}_{t-},\alpha_{t-},\bm{\theta})N^{1}({\mathrm{d}}\bm{\theta},{\mathrm{d}}t),$
+${\mathrm{d}}\bm{X}_{t}^{\alpha,2}=\bm{b}\big{(}\bm{X}_{t}^{\alpha,2},\alpha_{t}\big{)}{\mathrm{d}}t+\bm{\sigma}\big{(}\bm{X}_{t}^{\alpha,2},\alpha_{t}\big{)}{\mathrm{d}}\bm{B}^{2}_{t}+\int_{{\mathbb{R}}^{q}}\bm{\beta}(\bm{X}^{\alpha,2}_{t-},\alpha_{t-},\bm{\theta})N^{2}({\mathrm{d}}\bm{\theta},{\mathrm{d}}t),$
+are two conditionally independent copies of $(\bm{X}^{\alpha},\bm{B},N)$ given the sub $\sigma$-algebra ${\mathcal{G}}_{T}$ defined in an enlarged probability space $(\overline{\Omega},\overline{{\mathcal{F}}},\overline{{\mathbb{P}}})$, and $\overline{{\mathbb{E}}}[\cdot|{\mathcal{F}}]$ is the conditional expectation given $\mathcal{F}$} in the enlarged probability space.
+HJB equations and verification theorem.
+In the case where $\mathbb{G}$} is the filtration generated by Brownian motion $\bm{W}$, with $\bm{B}=(\bm{V},\bm{W})$ and $\bm{V}\perp\!\!\!\perp\bm{W}$, we can rewrite the dynamic by
+${\mathrm{d}}\bm{X}_{t}^{\alpha}=\bm{b}\big{(}\bm{X}_{t}^{\alpha},\alpha_{t}\big{)}{\mathrm{d}}t+\bm{\sigma}^{\bm{V}}\big{(}\bm{X}_{t}^{\alpha},\alpha_{t}\big{)}{\mathrm{d}}\bm{V}_{t}+\bm{\sigma}^{\bm{W}}\big{(}\bm{X}_{t}^{\alpha},\alpha_{t}\big{)}{\mathrm{d}}\bm{W}_{t}+\int_{{\mathbb{R}}^{q}}\bm{\beta}(\bm{X}^{\alpha}_{t-},\alpha_{t-},\bm{\theta})\widetilde{N}({\mathrm{d}}\bm{\theta},{\mathrm{d}}t),$
+and the conditionally independent copies
+${\mathrm{d}}\bm{X}_{t}^{\alpha,1}=\bm{b}\big{(}\bm{X}_{t}^{\alpha,1},\alpha_{t}\big{)}{\mathrm{d}}t+\bm{\sigma}^{\bm{V}}\big{(}\bm{X}_{t}^{\alpha,1},\alpha_{t}\big{)}{\mathrm{d}}\bm{V}^{1}_{t}+\bm{\sigma}^{\bm{W}}\big{(}\bm{X}_{t}^{\alpha,1},\alpha_{t}\big{)}{\mathrm{d}}\bm{W}_{t}+\int_{{\mathbb{R}}^{q}}\bm{\beta}(\bm{X}^{\alpha,1}_{t-},\alpha_{t-},\bm{\theta})\widetilde{N}^{1}({\mathrm{d}}\bm{\theta},{\mathrm{d}}t),$
+${\mathrm{d}}\bm{X}_{t}^{\alpha,2}=\bm{b}\big{(}\bm{X}_{t}^{\alpha,2},\alpha_{t}\big{)}{\mathrm{d}}t+\bm{\sigma}^{\bm{V}}\big{(}\bm{X}_{t}^{\alpha,2},\alpha_{t}\big{)}{\mathrm{d}}\bm{V}^{2}_{t}+\bm{\sigma}^{\bm{W}}\big{(}\bm{X}_{t}^{\alpha,2},\alpha_{t}\big{)}{\mathrm{d}}\bm{W}_{t}+\int_{{\mathbb{R}}^{q}}\bm{\beta}(\bm{X}^{\alpha,2}_{t-},\alpha_{t-},\bm{\theta})\widetilde{N}^{2}({\mathrm{d}}\bm{\theta},{\mathrm{d}}t).$
+Note that under this setting, at time $s$, $\bm{X}^{\alpha,1}_{s-}$ and $\bm{X}^{\alpha,2}_{s-}$ has a jump $\bm{\beta}(\bm{X}^{\alpha,1}_{s-},\alpha_{s-},\theta^{1}),\bm{\beta}(\bm{X}^{\alpha,2}_{s-},\alpha_{s-},\theta^{2})$ with Poisson random measure $N^{1}({\mathrm{d}}\theta^{1},{\mathrm{d}}s)\times N^{2}({\mathrm{d}}\theta^{2},{\mathrm{d}}s)$, we can simplify the Itô’s formula by
+$\Phi(\mu_{t})-\Phi(\mu_{0})$ (6.3)
+$=$ $\;\int_{0}^{t}\overline{{\mathbb{E}}}\Big{[}\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s},\bm{X}^{\alpha,1}_{s}\big{)}\cdot\bm{b}(\bm{X}^{\alpha,1}_{s},\alpha_{s})+\frac{1}{2}\nabla^{2}_{\bm{x}_{1}\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s},\bm{X}^{\alpha,1}_{s}\big{)}:\bm{\sigma}\bm{\sigma}^{\top}\big{(}\bm{X}^{\alpha,1}_{s},\alpha_{s}\big{)}$
+$\qquad\qquad+\frac{1}{2}\nabla^{2}_{\bm{x}_{1}\bm{x}_{2}}\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s},\bm{X}^{\alpha,1}_{s},\bm{X}^{\alpha,2}_{s}\big{)}:\bm{\sigma}^{\bm{W}}\big{(}\bm{X}^{\alpha,1}_{s},\alpha_{s}\big{)}{\bm{\sigma}^{\bm{W}}}^{\top}\big{(}\bm{X}^{\alpha,2}_{s},\alpha_{s}\big{)}\Big{|}{\mathcal{F}}\Big{]}{\mathrm{d}}s$
+$\quad+\frac{1}{2}\int_{0}^{t}\int_{{\mathbb{R}}^{2q}}\overline{{\mathbb{E}}}\Big{[}\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s},\bm{X}^{\alpha,1}_{s}+\bm{\beta}(\bm{X}^{\alpha,1}_{s},\alpha_{s},\bm{\theta}^{1}),\bm{X}^{\alpha,2}_{s}+\bm{\beta}(\bm{X}^{\alpha,2}_{s},\alpha_{s},\bm{\theta}^{2})\big{)}$
+$-\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s},\bm{X}^{\alpha,1}_{s},\bm{X}^{\alpha,2}_{s}+\bm{\beta}(\bm{X}^{\alpha,2}_{s},\alpha_{s},\bm{\theta}^{2})\big{)}-\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s},\bm{X}^{\alpha,1}_{s}+\bm{\beta}(\bm{X}^{\alpha,1}_{s},\alpha_{s},\bm{\theta}^{1}),\bm{X}^{\alpha,2}_{s}\big{)}$
+$\qquad\qquad\qquad\qquad+\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu_{s},\bm{X}^{\alpha,1}_{s},\bm{X}^{\alpha,2}_{s}\big{)}\Big{|}{\mathcal{F}}\Big{]}\nu({\mathrm{d}}\bm{\theta}^{1})\nu({\mathrm{d}}\bm{\theta}^{2}){\mathrm{d}}s$
+$\quad\qquad+\int_{0}^{t}\int_{{\mathbb{R}}^{q}}\overline{{\mathbb{E}}}\Big{[}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{X}^{\alpha,1}_{s}+\bm{\beta}(\bm{X}^{\alpha,1}_{s},\alpha_{s},\bm{\theta})\big{)}-\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{X}^{\alpha,1}_{s}\big{)}$
+$\qquad\qquad\qquad\qquad\quad-\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s-},\bm{X}^{\alpha,1}_{s}\big{)}\cdot\bm{\beta}(\bm{X}^{\alpha,1}_{s},\alpha_{s},\bm{\theta})\Big{|}{\mathcal{F}}\Big{]}\nu({\mathrm{d}}\bm{\theta}){\mathrm{d}}s$
+$+\int_{0}^{t}\overline{{\mathbb{E}}}\Big{[}\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu_{s},\bm{X}^{\alpha,1}_{s}\big{)}\cdot\bm{\sigma}^{\bm{W}}(\bm{X}^{\alpha,1}_{s},\alpha_{s})\Big{|}{\mathcal{F}}\Big{]}{\mathrm{d}}\bm{W}_{s},$
+Proceeding by the same argument as in [20], we have the dynamic programming principle for the control problem with partial observation. In particular, for all $0\leq t<s\leq T$,
+$V(t,\mu)=\inf_{\alpha\in{\mathcal{A}}^{G}}{\mathbb{E}}\bigg{[}\int_{t}^{s}f(\bm{X}^{\alpha}_{u},\alpha_{u}){\mathrm{d}}u+V\big{(}s,\text{Law}(\bm{X}_{s}^{\alpha}|{\mathcal{G}}_{T})\big{)}\bigg{|}\text{Law}(\bm{X}_{t}^{\alpha}|{\mathcal{G}}_{T})=\mu\bigg{]}.$
+We have now the ingredients for deriving the Bellman equation.
+For $\Phi\in C^{2,2}({\mathcal{P}}_{2}({\mathbb{R}}^{n}))$, $\mu\in{\mathcal{P}}_{2}({\mathbb{R}}^{n})$ and $a\in A$, the operator $\mathcal{L}^{\mu,a}$ maps from $C^{2}({\mathcal{P}}_{2}({\mathbb{R}}^{n}))$ to a function ${\mathbb{R}}^{n}\to{\mathbb{R}}$ defined by
+$\big{(}\mathcal{L}^{\mu,a}\Phi\big{)}(\bm{x}):=$ $\;\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu,\bm{x}\big{)}\cdot\bm{b}(\bm{x},a)+\frac{1}{2}\nabla^{2}_{\bm{x}_{1}\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu,\bm{x}\big{)}:\bm{\sigma}\bm{\sigma}^{\top}\big{(}\bm{x},a\big{)}$
+$+\int_{{\mathbb{R}}^{q}}\bigg{(}\frac{\delta\Phi}{\delta\mu}\big{(}\mu,\bm{x}+\bm{\beta}(\bm{x},a,\bm{\theta})\big{)}-\frac{\delta\Phi}{\delta\mu}\big{(}\mu,\bm{x}\big{)}-\nabla_{\bm{x}_{1}}\frac{\delta\Phi}{\delta\mu}\big{(}\mu,\bm{x}\big{)}\cdot\bm{\beta}(\bm{x},a,\bm{\theta})\bigg{)}\nu({\mathrm{d}}\bm{\theta}).$
+Furthermore, for $\Phi\in C^{2,2}({\mathcal{P}}_{2}({\mathbb{R}}^{n}))$, $\mu\in{\mathcal{P}}_{2}({\mathbb{R}}^{n})$, and $a\in A$, the operator $\mathcal{M}^{\mu,a}$ maps from $C^{2,2}({\mathcal{P}}_{2}({\mathbb{R}}^{n}))$ to a function ${\mathbb{R}}^{2n}\to{\mathbb{R}}$ defined by
+$\big{(}\mathcal{M}^{\mu,a}\Phi\big{)}(\bm{x},\bm{x}^{\prime}):=$ $\;\frac{1}{2}\nabla^{2}_{\bm{x}_{1}\bm{x}_{2}}\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu,\bm{x},\bm{x}^{\prime}\big{)}:\bm{\sigma}^{\bm{W}}\big{(}\bm{x},a\big{)}{\bm{\sigma}^{\bm{W}}}^{\top}\big{(}\bm{x}^{\prime},a\big{)}\Big{)}$
+$+\int_{{\mathbb{R}}^{2q}}\bigg{(}\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu,\bm{x}+\bm{\beta}(\bm{x},a,\bm{\theta}^{1}),\bm{x}^{\prime}+\bm{\beta}(\bm{x}^{\prime},a,\bm{\theta}^{2})\big{)}$
+$-\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu,\bm{x},\bm{x}^{\prime}+\bm{\beta}(\bm{x}^{\prime},a,\bm{\theta}^{2})\big{)}-\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu,\bm{x}+\bm{\beta}(\bm{x},a,\bm{\theta}^{1}),\bm{x}^{\prime}\big{)}$
+$\qquad\qquad+\frac{\delta^{2}\Phi}{(\delta\mu)^{2}}\big{(}\mu,\bm{x},\bm{x}^{\prime}\big{)}\bigg{)}\nu({\mathrm{d}}\bm{\theta}^{1})\nu({\mathrm{d}}\bm{\theta}^{2}).$
+Now, the Bellman equation takes the following form:
+$\begin{cases}\partial_{t}v(t,\mu)+\sup_{a\in A}\Big{(}\big{\langle}\mu,f(\cdot,a)+\mathcal{L}^{\mu,a}v(t,\mu)(\cdot)\big{\rangle}+\big{\langle}\mu\otimes\mu,\mathcal{M}^{\mu,a}v(t,\mu)(\cdot,\cdot)\big{\rangle}\Big{)}=0,\\ v(T,\mu)=\big{\langle}\mu,g\big{\rangle}.\end{cases}$ (6.4)
+The Bellman equation (6.4) is validated by the following verification theorem, which can be proved by following the similar argument as in [4], with the Itô’s formula (6.3).
+Theorem 6.2.
+Let $v$ be a real-valued function in $C^{1,2}([0,T]\times{\mathcal{P}}_{2}({\mathbb{R}}^{n}))$, i.e., $v$ is continuous on $0,T]\times{\mathcal{P}}_{2}({\mathbb{R}}^{n})$, $v(t,\cdot)\in C^{2}({\mathcal{P}}_{2}({\mathbb{R}}^{n}))$ for all $t\in[0,T]$, and $v(\cdot,\mu)\in C^{1}([0,T))$ for all $\mu\in{\mathcal{P}}_{2}({\mathbb{R}}^{n})$.
+Suppose that $v$ is a solution to the Bellman equation (6.4), and there exists for all $(t,\mu)\in[0,T)\times{\mathcal{P}}_{2}({\mathbb{R}}^{n})$ an element $\hat{a}(t,\mu)$ attaining the supremum in (6.4) such that for a given square integrable initial condition, the stochastic McKean-Vlasov SDE
+${\mathrm{d}}\hat{\bm{X}}_{t}=$ $\bm{b}\Big{(}\hat{\bm{X}}_{t},\hat{a}\big{(}t,\text{Law}(\hat{\bm{X}_{t}}|\bm{W})\big{)}\Big{)}{\mathrm{d}}t+\bm{\sigma}^{\bm{V}}\Big{(}\hat{\bm{X}}_{t},\hat{a}\big{(}t,\text{Law}(\hat{\bm{X}_{t}}|\bm{W})\big{)}\Big{)}{\mathrm{d}}\bm{V}_{t}$
+$+\bm{\sigma}^{\bm{W}}\Big{(}\hat{\bm{X}}_{t},\hat{a}\big{(}t,\text{Law}(\hat{\bm{X}_{t}}|\bm{W})\big{)}\Big{)}{\mathrm{d}}\bm{W}_{t}+\int_{{\mathbb{R}}^{q}}\bm{\beta}\Big{(}\hat{\bm{X}}_{t},\hat{a}\big{(}t,\text{Law}(\hat{\bm{X}_{t}}|\bm{W})\big{)},\bm{\theta}\Big{)}N({\mathrm{d}}\bm{\theta},{\mathrm{d}}t)$
+admits a unique solution $\hat{\bm{X}}$. Then, $v=V$ and the feedback control in ${\mathcal{A}}^{\mathcal{G}}$ defined by
+$\alpha^{\star}_{s}:=\hat{a}\big{(}s,\text{Law}(\hat{\bm{X}_{s}}|\bm{W})\big{)},\text{ for }t\leq s\leq T,$
+is an optimal control for $V(t,\mu)$. That is, $V(t,\mu)=J(t,\alpha^{\star},\mu)$.
+Note that this generalizes the verification results in [4] to the case of jump-diffusions.